### v11

# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

Abstract
As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions, to compute means and standard deviations of the results. Statistical Taylor expansion traces the dependency of the input uncertainties in the intermediate steps, so that the variables in the intermediate analytic expressions can no longer be regarded as independent of each other, and the result of the analytic expression is path independent. Thus, it differs fundamentally from the conventional common approaches in applied mathematics which optimize execution path for each calculation. In fact, statistical Taylor expansion may standardize numerical calculations for analytic expressions. Its statistical nature allows religious testing of its result when the sample size is large enough. This paper also introduces an implementation of statistical Taylor expansion called variance arithmetic and presents corresponding test results in a very wide range of mathematical applications.
Another important conclusion of this paper is that the numerical errors in the library function can have significant effects on the result. For example, the periodic numerical errors in the trigonometric library functions can resonate with periodic signals, producing large numerical errors in the results.
Keywords: computer arithmetic, error analysis, interval arithmetic, uncertainty, numerical algorithms.
AMS subject classifications: G.1.0
Copyright ©2024
1 Introduction
1.1 Measurement Uncertainty
Except for the most basic counting, scientific and engineering measurements never yield completely precise results [2][3]. In such measurements, the uncertainty of a quantity $x$ is typically expressed by either the sample deviation $\delta x$ or the uncertainty range $\Delta x$ [2][3].
  • If $\delta x=0$ or $\Delta x=0$, $x$ is a precise value.
  • Otherwise, $x$ is an imprecise value.
$P(x)\equiv\delta x/|x|$ is defined as the statistical precision (hereafter referred to simply as precision) of the measurement, where $x$ denotes the value, and $\delta x$ represents the uncertainty deviation. A larger precision indicates a coarser measurement whereas a smaller precision indicates a finer measurement. The precision of measured values can range from order-of-magnitude estimates to $10^{-2}$ to $10^{-4}$ in common measurements, and to $10^{-15}$ in state-of-the-art determinations of fundamental physical constants [4].
The focus of this paper is on determining the result value and uncertainty of a general analytic expression with imprecise input values.
1.2 Extension to Existing Statistics
Rather than focusing on the result uncertainty distribution when applying a function to a random variable with known uncertainty distribution [5], statistical Taylor expansion provides the result mean and variance for general analytic expressions without the need for the result uncertainty distribution.
  • Previous work has addressed the effect of input uncertainties to output values for special cases [6]. Statistical Taylor expansion generalizes this as the uncertainty bias as shown in Formula (2.6), and (2.9) in this paper.
  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [5][6], whereas statistical Taylor expansion extends this to higher-order interactions as expressed in Formula (2.3) in this paper.
1.3 Problem of the Related Numerical Arithmetics
Variance arithmetic implements statistical Taylor expansion and surpasses all existing related numerical arithmetic methods.
1.3.1 Conventional Floating-Point Arithmetic
Conventional floating-point arithmetic [7][8][9] assumes that every bit of its significand is valid at all times, and continuously either generates artificial information or erase useful information to maintain this assumption during its normalization process [10]. For example, the following calculation is performed exactly in integer format:
$64919224\times 205117922-159018721\times 83739041=\\ 13316075197586562-13316075197586561=1;$ (1.1)
If Formula (1.1) is carried out using conventional 64-bit floating-point arithmetic:
$64919224\times 205117922-159018721\times 83739041=\\ 64919224.000000000\times 205117922.000000000-159018721.000000000\times 83739041.000000000=\\ 13316075197586562.-13316075197586560.=2.=2.0000000000000000;$ (1.2)
  1. 1.
    The normalization of the input values may pad zero bits artificially in the least significant positions until every bit in the 53-bit significand are used, for example, converting $64919224$ to $64919224.000000000$.
  2. 2.
    The addition or multiplication results may exceed the bit range of the 53-bit significand, which is round off from the least significant bits, to fit into the 53-bit significand. This process may generate rounding errors, for example, from $13316075197586561$ to $13316075197586560$, to generate a rounding error of $1$ at the least significant bit.
  3. 3.
    The subtraction or division results can cancel values from high significant bits, exposing the rounding errors in the low significant bits, for example, 13316075197586562.13316075197586560.=2.13316075197586562.-13316075197586560.=2..
  4. 4.
    The normalization of the subtraction or division results can amplify the rounding error toward the most significant bit by padding zeros, producing catastrophic cancellation [11][12] in this case.
Because rounding errors from lower digits can propagate to higher digits, the $10^{-7}$ significance of the 32-bit IEEE floating-point format [7][8][9] is usually insufficient for calculations involving input data with a precision of $10^{-2}$ to $10^{-4}$. For more complex calculations, even the $10^{-16}$ significance of the 64-bit IEEE floating-point format [7][8][9] may not enough for inputs with $10^{-2}$ to $10^{-4}$ precision. This is a fundamental controversy of the conventional floating-point arithmetic.
Because rounding-error is path-dependent, a major effort of conventional numerical methods is to identify optimal computation strategies to minimize rounding errors, such as in Gaussian elimination [12][13][14], even though all other paths of root finding methods in linear algebra are equivalent mathematically [37].
Self-censoring rules have been developed to limit such rounding error propagation [13][14], such as avoiding the subtraction of results from large multiplications, as in Formula (1.2). However, these rules are neither enforceable nor easily adoptable, and they are even more difficult to quantify. To date, research on rounding error propagation has focused mainly either on linear calculations [11][12][15], or special cases [13][16][17], whereas in practice, rounding errors often appear as the pervasive and mysterious numerical instability [18].
The forward rounding error study [12] compares 1) the result with rounding error and 2) the ideal result without rounding error, such as comparing the result using 32-bit IEEE floating-point arithmetic with the corresponding result using 64-bit IEEE floating-point arithmetic [19]. The most recent study of this type presents an extremely optimistic view of numerical library errors, reporting them as a fraction of the least significant bit of the floating-point significands [19]. However, such optimism contradicts statistical tests on numerical library functions as presented in this paper.
The backward rounding error study [11][12][15] only estimates the result uncertainty due to rounding errors, thereby overlooking the bias caused by rounding errors on the result value. This analysis is typically restricted to very small uncertainty uncertainties because it relies on perturbation theory, and it tailored to each specific algorithm [11][12][15].
In contrast, variance arithmetic traces the rounding error directly as part of the uncertainty. Statistical Taylor expansion applies generically to any analytic function, for both result mean and deviation, and accommodates input uncertainties of any magnitude. By demonstrating that the analytic result should be path-independent, statistical Taylor expansion fundamentally challenges the conventional methodology for seeking optimal execution strategy for a given analytic expression.
1.3.2 Interval Arithmetic
Interval arithmetic [14][20][21][22][23][24] is currently a standard method to track calculation uncertainty. Its goal is to ensure that a value x remains absolutely bound within its bounding range throughout the computation.
However, the bounding range in interval arithmetic is not compatible with the approach commonly employed in scientific and engineering measurements, which instead characterizes uncertainty using the statistical mean and deviation [2][3]  [^2] There is one attempt to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics [25]. Because this attempt seems to rely heavily on 1) specific properties of the uncertainty distribution within the interval and/or 2) specific properties of the functions upon which the interval arithmetic is used, this attempt does not seem to be generic. If probability model is introduced to interval arithmetic to allow tiny bounding leakage, the bounding range is much less than the corresponding pure bounding range [15]. Anyway, these attempts seem to be outside the main course of interval arithmetic. .
Interval arithmetic represents only the worst-case scenario of uncertainty propagation. For example, in addition, it assumes that the two input variables are perfectly positively correlated [26], thereby yielding the widest possible bounding range. In contrast, if the variables were perfectly negatively correlated, the bounding range after addition would reduce [27]  [^3] Such case is called the best case in random interval arithmetic. The vast overestimation of bounding ranges in these two worst cases prompts the development of affine arithmetic [26][28], which traces error sources using a first-order model. Being expensive in execution and depending on approximate modeling even for such basic operations as multiplication and division, affine arithmetic has not been widely used. In another approach, random interval arithmetic [27] randomly chooses between the best-case and the worst-case intervals, so that it can no longer guarantee bounding without leakage. Anyway, these attempts seem to be outside the main course of interval arithmetic. . This worst-case assumption can lead to order-of-magnitude overestimates [1].
The results of interval arithmetic can depend strongly on the specific algebraic form of an analytic function $f(x)$, a phenomenon known as the dependency problem. This issue is amplified in interval arithmetic [23] but also presents in conventional floating-point arithmetic [13].
Furthermore, interval arithmetic lacks a mechanism to reject invalid calculations, even though every mathematical operation has a valid input range. For example, it produces branched results for $1/(x\pm\Delta x)$ or $\sqrt{x\pm\Delta x}$ when $0\in[x-\Delta x,x+\Delta x]$, whereas a context-sensitive uncertainty bearing arithmetic should reject such calculation naturally.
In contrast, variance arithmetic specifies a value with mean and deviation. It has no dependency problem. Its statistical context rejects certain input intervals mathematically, such as inversion and square root when the statistical bounding range contains zero.
1.3.3 Statistical Propagation of Uncertainty
Statistical propagation of uncertainty treats the uncertainty of a value as the random part of each imprecise value. When the statistical correlation between any two input variables is known, statistical propagation can be used to determine result mean and variance [29][30].
In contrast, statistical Taylor expansion treats the uncertainty of a value as the limitation of obtaining the precise value. This assumption aligns with most error analyses in the literature [2][3]. Its statistical foundation is the uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough precision, so that their uncertainties are independent of each other, even though the input variables could be highly correlated. This statistical foundation can be turned into a quantitative test on the input variables on the precision vs the correlation [1].
Statistical propagation of uncertainty seems to have applied a wrong statistical context for uncertainty. For example, a time series is a random variable while each value in the time series is just an imprecise value. If two such imprecise values have correlated uncertainty, they have systematic errors [2]. On the other hand, the uncorrelated uncertainty assumption can hold for each value in the time series, while the time series can be self-correlated [5] such as being periodic.
1.3.4 Significance Arithmetic
Significance arithmetic [31] seeks to track the number of reliable bits in an imprecise value through calculation. In two early implementations [32][33], significance arithmetic was based on simple operational rules applied to reliable bit counts, rather than on formal statistical approaches. In these methods, the reliable bit count is treated as an integer, even though in practice it can take a fractional value [34]. This constraint can cause an artificial step-wise reduction in significance. The implementation of significance arithmetic in Mathematica [34] employs a linear error model consistent with the first-order approximation of interval arithmetic [14][22][23].
One limitation of significance arithmetic is its inability to specify uncertainty accurately [1]. For example, if the least significant bit of significand is used to represent uncertainty, the result precision can be very coarse as in $1\pm 10^{-3}=1024\times 2^{-10}$ [1]. Introducing a limited number of bits calculated inside uncertainty does not fully resolve this issue. Therefore, various attempts for floating-point arithmetic without normalization [33] are not widely adopted, and the conventional floating-point arithmetic dominates the numerical world. For this reason, variance arithmetic has abandoned the significance arithmetic principle of its predecessor [1].
1.3.5 Stochastic Arithmetic
Stochastic arithmetic [35][36] randomizes the least significant bits (LSB) of each input floating-point value, repeats the same calculation multiple times, and then uses statistical analysis to identify invariant digits among the calculation results as significant digits. However, this approach can require excessive computation as the number of repetitions needed for each input is algorithm-independent, particularly when the algorithm contains branches.
In contrast, statistical Taylor expansion provides a direct characterization of the result’s mean and deviation without sampling.
1.4 An Overview of This Paper
This paper presents the theory of statistical Taylor expansion, its implementation as variance arithmetic, and the corresponding validation tests. Section 1 compares statistical Taylor expansion and variance arithmetic with other established uncertainty-bearing arithmetic. Section 2 develops the theoretical foundation of statistical Taylor expansion. Section 3 describes the variance arithmetic as an implementation of the statistical Taylor expansion. Section 4 discusses the standards and methodologies used to validate variance arithmetic. Section 5 evaluates variance arithmetic for common mathematical library functions. Section 6 applies variance arithmetic to adjugate matrix and matrix inversion. Section 7 applies variance arithmetic to process time-series data. Section 8 examines the impact of numerical library errors and shows that these errors can be significant. Section 9 applies variance arithmetic to regression analysis. Section 10 concludes with a summary and discussion.
2 Statistical Taylor Expansion
2.1 The Uncorrelated Uncertainty Assumption
The uncorrelated uncertainty assumption [1] states that the uncertainty of any input is statistically uncorrelated with the uncertainty of any other input. This condition is satisfied when uncorrelated noise is the primary contributor to the uncertainty of the signals. It is consistent with standard methods for processing experimental data [2][3].
The uncorrelated uncertainty assumption permits the signals themselves to be highly correlated. Suppose two signals have a correlation coefficient $\gamma$, and measured precision $P_{1}$ and $P_{2}$, respectively. Let $P$ be the coarser of $P_{1}$ and $P_{2}$. At the level of $P$, the correlation is reduced to $\gamma_{P}$ according to Formula (2.1) [1]. The value of $\gamma_{P}$ decreases rapidly as $P$ becomes finer, so that when $P$ is sufficiently fine, the correlation $\gamma_{P}$ of the uncertainties of two signals are effectively zero [1].
$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}};$ (2.1)
The uncorrelated uncertainty assumption ensures no systematic error [2].
2.2 Distributional Zero and Distributional Pole
Figure 1: Probability density function of $\tilde{y}=\tilde{x}^{2}$, for various values of $\mu$ as indicated in the legend. The variable $\tilde{x}$ follows a Gaussian distribution with mean $\mu$ and deviation $1$. The horizontal axis is scaled as $\sqrt{\tilde{y}}$.
Figure 2: Probability density function for $\tilde{y}=\sqrt{\tilde{x}}$, various values of $\mu$ as indicated in the legend. The variable $\tilde{x}$ follows a Gaussian distribution with the distributional mean $\mu$ and deviation $1$. The horizontal axis is scaled as $\tilde{y}^{2}$.
[…]
$\rho(\tilde{z})d\tilde{z}$= $\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y}$; (2.2)
$y=x^{c}:\;\;\;\rho(\tilde{y},\mu_{y},\sigma_{y})=c\tilde{y}^{\frac{1}{c}-1}\frac{1}{\sigma}N(\frac{\tilde{y}^{\frac{1}{c}}-\mu}{\sigma});$ (2.3)
[…]
  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, sine $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [5]. A linear transformation generates neither a distributional zero nor a distributional pole, in accordance with to Formula (2.2).
  • Figure 1 illustrates the probability density function for $(x\pm 1)^{2}$ according to Formula (2.3), which exhibits a distributional pole at $x=0$. The distribution $(0\pm 1)^{2}$ corresponds to the $\chi^{2}$ distribution [2]. At the distributional pole, the probability density function resembles a Delta distribution.
  • Figure 2 illustrates the probability density function for $\sqrt{x\pm 1}$ according to Formula (2.3), which has a distributional zero at $x=0$. At the distributional zero, the probability density function is zero.
In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ more when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero. This resemblance allows for a generic characterization of the output.
2.3 Statistical Taylor Expansion
Formula (2.2) provides the uncertainty distribution of an analytic function. However, in most scientific and engineering calculations, the primary interest is not the full result distribution, but rather just few summary statistics of the result, such as the mean and deviation [2][3]. These simplified statistics can be obtained through a statistical Taylor expansion.
Let $\overline{x}$ and $\delta x$ denote the sample mean and deviation of a random variable. Let $\mu$ and $\sigma$ denote the distribution mean and deviation of the same variable. An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.5). Formula (2.4) defines the bound moment $\zeta(n,\kappa)$, where $0<\varrho,\kappa$ specify the bounding ranges. Using Formula (2.4), Formula (2.6) and Formula (2.3) yields the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively. The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the result value.
$\zeta(n)\equiv$ $\int_{\mu-\varrho\sigma}^{\mu+\kappa\sigma}\tilde{x}^{n}\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.4)
[…]
$\overline{f(x)}=$ $\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\sum_{n=0}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.6)
$\delta^{2}f(x)=$ $\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$
$=$ $\sum_{n=1}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.7)
Formula (2.4) defines a bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ for the underlying distribution, where $\mu$ and $\sigma$ are the distribution mean and distribution deviation, respectively. As discussed later in this paper, $\kappa$ statistically determines $\varrho$, while $\kappa$ itself is determined by both the sample size and the underlying distribution of the input. Certain $f(x)$ such as $\log(x)$ and $x^{c}$ also require an upper bound on $\kappa$ to ensure the convergence of Formula (2.6) and (2.3). The probability $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
Under the uncorrelated uncertainty assumption, Formula (2.9) and (2.3) compute the mean and variance of the Taylor expansion in Formula (2.8), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ are the variance moments for $x$ and $y$, respectively:
$f(x+\tilde{x},y+\tilde{y})=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\frac{f^{(m,n)}_{(x,y)}}{m!n!}\tilde{x}^{m}\tilde{y}^{n};$ (2.8)
$\overline{f(x,y)}=$ $\int\int f(x+\tilde{x},y+\tilde{y})\rho(\tilde{x},\mu_{x},\sigma_{x})\rho(\tilde{y},\mu_{y},\sigma_{y})\;d\tilde{x}d\tilde{y}$ (2.9)
$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$
$\delta^{2}f(x,y)=$ $\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\big(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\big);$ (2.10)
Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
2.4 Bounding Symmetry
A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbound distribution. To achieve mean preserving, $\kappa$ determines $\varrho$, so that $\zeta(n)$ becomes $\zeta(n,\kappa)$. Under a mean-preserving bounding, Formula (2.11) and (2.12) provide the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
$\overline{x\pm y}$ $=\zeta(0,\kappa_{x})x\pm\zeta(0,\kappa_{x})y;$ (2.11)
$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}+\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.12)
$\overline{xy}$ $=\zeta(0,\kappa_{x})x\;\zeta(0,\kappa_{y})y;$ (2.13)
$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\sigma_{y})^{2}+\zeta(2,\kappa_{x})(\sigma_{x})^{2}\;\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.14)
When $4\leq\kappa$, $\zeta(0,\kappa)\simeq 1$ and $\zeta(2,\kappa)\simeq 1$, making Formula (2.11) and (2.12) the convolution results for $x\pm y$ [5], and Formula (2.13) and (2.14) the corresponding results of the product distribution for $xy$ [5].
For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
2.5 Bounding Asymptote
Empirically, Formula (2.15), (2.19), and (2.23) demonstrate that as $2n\rightarrow+\infty$, $\zeta(2n)\rightarrow\kappa^{2n}/(2n)$.
2.5.1 Uniform Input Uncertainty
For uniform distribution, the bounding range is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. Formula (2.15) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
$\eta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.15)
2.5.2 Gaussian Input Uncertainty
The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [5]. This convergence occurs rapidly [1]. In a digital computer, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sum of expansion terms [8][9]. Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [2][3][5].
For Gaussian input uncertainty, $\zeta(2n+1)=0$.
If $\kappa=\infty$, then $\zeta(2n)=(2n-1)!!$, which causes Formula (2.29) for $\log(x\pm\delta x)$ and Formula (2.35) for $(x\pm\delta x)^{c}$ to diverge. As discussed, $\kappa$ is determine by the sample count $N$. When $\kappa$ has an upper bound, Formula (2.4) reduces to Formula (2.16):
$\zeta(2n,\kappa)=$ $\;(2n-1)!!\left(\xi(\frac{\kappa}{\sqrt{2}})-2N(\kappa)\sum_{j=0}^{n-1}\frac{\kappa^{2j+1}}{(2j+1)!!}\right);$ (2.16)
$=$  $\;2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.17)
$=$ $\;(2n-1)\zeta(2n-2,\kappa)-2N(\kappa)\kappa^{2n-1};$ (2.18)
$\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)\simeq 2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.19)
  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$ according to Formula (2.18). When $\kappa=5$, and $n<5$, the relative error satisfies $\left|\eta(2n)/(2n-1)!!-1\right|<10^{-3}$.
  • For large $2n$, Formula (2.16) reduces to Formula (2.19), showing that $\zeta(2n)$ increases more slowly than $\kappa^{2n}$ as $2n$ grows.
2.5.3 An Input Uncertainty with Limited Range
$\rho(\tilde{x},\mu,\sigma)$ $=\frac{\tilde{x}}{\lambda^{2}}e^{-\frac{\tilde{x}}{\lambda}};\;\;\;\mu=2\lambda;\;\;\;\sigma=\sqrt{2}\lambda;$ (2.20)
$e^{-\varrho}\varrho^{2}=e^{-\kappa}\kappa^{2};$ (2.21)
$\zeta(n)$ $=(n+1)!\left(e^{-\varrho}\sum_{j=0}^{n+1}\frac{\varrho^{j}}{j!}-e^{-\kappa}\sum_{j=0}^{n+1}\frac{\kappa^{j}}{j!}\right);$ (2.22)
$\lim_{n\rightarrow+\infty}\zeta(n)=\frac{\kappa^{n+2}}{n+2}e^{-\kappa};$ (2.23)
Formula (2.20) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<1<\kappa$. Formula (2.21) gives its mean-preserving equation. Formula (2.22) expresses its bound moment. And Formula (2.23) describes its asymptotic behavior.
2.6 One-Dimensional Examples
Formula (2.25) and (2.26) give the mean and variance for $e^{x}$, respectively:
$e^{x+\tilde{x}}=e^{x}\sum_{n=0}^{\infty}\frac{\tilde{x}^{n}}{n!};$ (2.24)
$\frac{\overline{e^{x}}}{e^{x}}$ $=\sum_{n=0}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.25)
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.26)
Formula (2.28) and (2.29) give the mean and variance for $\log(x)$, respectively:
$\log(x+\tilde{x})-\log(x)=\log(1+\frac{\tilde{x}}{x})=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j}\frac{\tilde{x}^{j}}{x^{j}};$ (2.27)
$\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n)}{n};$ (2.28)
$\delta^{2}\log(x)$ $=\sum_{n=2}^{+\infty}P(x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j(n-j)};$ (2.29)
Formula (2.31) and (2.32) give the mean and variance for $\sin(x)$, respectively:
[…]
$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}\sigma^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.31)
$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.32)
Formula (2.34) and (2.35) give the mean and variance for $x^{c}$, respectively:
$(x+\tilde{x})^{c}=x^{c}(1+\frac{\tilde{x}}{x})^{c}=x^{c}+x^{c}\sum_{n=1}^{\infty}\frac{\tilde{x}^{n}}{x^{n}}\begin{pmatrix}c\\ n\end{pmatrix};\;\;\;\begin{pmatrix}c\\ n\end{pmatrix}\equiv\frac{\prod_{j=0}^{n-1}(c-j)}{n!};$ (2.33)
$\frac{\overline{x^{c}}}{x^{c}}$ $=1+1\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.34)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.35)
The result variance in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
2.7 Low-Order Approximation
When $n<5\leq\kappa$, $\eta(n)\simeq n!!$. Under these conditions, Formula (2.26), (2.29), (2.32), and (2.35) can be simplified as Formula (2.36), (2.38), (2.7), and (2.7), respectively.
$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}\simeq$ $\;\sigma^{2}+\frac{3}{2}\sigma^{4}+\frac{7}{6}\sigma^{6}+\frac{5}{8}\sigma^{8}+o((\delta x)^{10});$ (2.36)
$\delta^{2}\sin(x)\simeq$ $\;\sigma^{2}\cos(x)^{2}-(\delta x)^{4}(\cos(x)^{2}\frac{3}{2}-\frac{1}{2})$
$+\sigma^{6}(\cos(x)^{2}\frac{7}{6}-\frac{1}{2})-\sigma^{8}(\cos(x)^{2}\frac{5}{8}-\frac{7}{24})+o((\delta x)^{10});$ (2.37)
$\delta^{2}\log(x)\simeq$ $\;P(x)^{2}+P(x)^{4}\frac{9}{8}+P(x)^{6}\frac{119}{24}+P(x)^{8}\frac{991}{32}+o(P(x)^{10});$ (2.38)
$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}\simeq$ $\;c^{2}P(x)^{2}+\frac{3}{2}c^{2}(c-1)(c-\frac{5}{3})P(x)^{4}$
$+\frac{7}{6}c^{2}(c-1)(c-2)^{2}(c-\frac{16}{7})P(x)^{6}+o(P(x)^{8});$ (2.39)
Formula (2.40), (2.41), and (2.42) are special cases of Formula (2.7).
$\delta^{2}x^{2}\simeq$  $\;4x^{2}(\delta x)^{2}+2(\delta x)^{4};$ (2.40)
$\frac{\delta^{2}\sqrt{x}}{(\sqrt{x})^{2}}\simeq$ $\;\frac{1}{4}P(x)^{2}+\frac{7}{32}P(x)^{4}+\frac{75}{128}P(x)^{6}+o(P(x)^{8});$ (2.41)
$\frac{\delta^{2}1/x}{(1/x)^{2}}\simeq$ $\;P(x)^{2}+8P(x)^{4}+69P(x)^{6}+o(P(x)^{8});$ (2.42)
2.8 Convergence
Formula (2.26) for $e^{x\pm\delta x}$ and Formula (2.32) for $\sin(x\pm\delta x)$ both converge unconditionally. However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can take negative values for large $\delta x$, which imposes a constraint upper bound on the input $\delta x$.
Formula (2.29) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) when $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$ for a Gaussian distribution, or $P(x)<1/\sqrt{3}$ for a uniform distribution.
$\delta^{2}\log(x\pm\delta x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}$
$\simeq$  $\;2\nu(\kappa)\log(2)\sum_{n=1}^{+\infty}\frac{(P(x)\kappa)^{2n}}{(2n)^{2}},\begin{cases}\text{Gaussian}:\nu(\kappa)=N(\kappa)\kappa\\ \text{Uniform}:\nu(\kappa)=1,\;\;\;\kappa=\sqrt{3}\end{cases};$ (2.43)
Formula (2.35) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.44) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ . This expression converges when $P(x)\lesssim 1/\kappa$ although the precisse upper bound for $P(x)$ varies with $c$.
$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.44)
2.9 Bounding Range
Figure 3: The result variance ratio $\delta^{2}f/\widehat{\delta^{2}f}$ (left y-axis) and the bounding leakage (right y-axis) for varying bounding range $\kappa$ (x-axis) for the selected function $f(x)$ (legend), where $\widehat{\delta^{2}f}$ denotes the corresponding variance for $\kappa=5$.
Figure 4: The ratio of adjusted $\widehat{\delta^{2}f}$ to the original $\widehat{\delta^{2}f}$ (as shown by the y-axis) for varying bounding range $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), where $\widehat{\delta^{2}f}$ is approximated by normalizing the input variance with $\zeta(2,\kappa)$ such that $\delta^{2}x=(\delta x)^{2}$.
The bounding range $\kappa$ influences the results of statistical Taylor expansion unless it is sufficiently large. Figure 3 shows that $\epsilon(\kappa)$ decreases as $\kappa$ increases, while $\delta^{2}f$ for the selected functions rises with $\kappa$ until reaching stable value when $\kappa\geq 4$. When $\kappa\geq 5$, the results show negligible difference for varying $\kappa$. In practice, based on the 5-$\sigma$ rule [3] for reliably distinguishing two random variables, let $\widehat{\delta^{2}f}$ denote the ideal variance obtained once the result variance stabilizes for $\kappa=5$. Figure 3 indicates that $\delta^{2}f\leq\widehat{\delta^{2}f}$ due to $\epsilon(\kappa)$, and when $\epsilon(\kappa)\simeq 0$, $\delta^{2}f\simeq\widehat{\delta^{2}f}$. The variance ratio is defined as $\delta^{2}f/\widehat{\delta^{2}f}$. Figure 3 also demonstrates that variance ratio remains approximately a constant for a $\kappa$ value.
To approach $\widehat{\delta^{2}f}$ when $\kappa<5$, instead of $(\delta x)^{2}$, the adjusted input uncertainty variance $(\delta x)^{2}/\zeta(2,\kappa)$ is used in Formula (2.3) to obtain the adjusted variance whose $(\delta x)^{2}$-term has the value as if $\zeta(2,\kappa)=1$, which is same as that of the ideal variance $\widehat{\delta^{2}f}$. Figure 4 shows that the adjusted variances are close to the corresponding ideal variance, however as a first order approximation, their ratios depend on the specific form of $f(x)$. Figure 4 also indicates a universal worst $\kappa\simeq 1.75$, below which the tail effect of Gaussian distributions vanishes.
$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.45)
$\zeta(2,\kappa)\equiv$ $\frac{\delta^{2}f}{\widehat{\delta^{2}f}}$; (2.46)
$x\pm y:$ $\;\;\;\zeta(2,\kappa)=\frac{\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.47)
Formula (2.45) is a special case for Formula (2.3), which can be reinterpreted as obtaining the bounding range $\kappa$ from a variance ratio. This relation between variance ratio and bounding range can be generalized as in Formula (2.46). For example, Formula (2.47) shows that the result $\kappa$ for $x\pm y$ is a variance-weighted average of the input $\kappa_{x}$ and $\kappa_{y}$ via $\zeta(2,\kappa)$.
2.10 Statistical Bounding
Figure 5: Measured bounding leakage $\epsilon(\kappa_{s},N)$ (y-axis) for varying measuring bounding range $\kappa_{s}$ (x-axis) and sample count NN (legend).
Figure 6: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range for the normal distribution.
When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [5]. This yields the sample bounding leakage $\epsilon(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional bounding leakage $\epsilon(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$. Because $\epsilon(\kappa)\neq\epsilon(\kappa,N)$ for finite $N$, let $\epsilon(\kappa)=\epsilon(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bonding range, and $\kappa(\kappa_{s},N)$ is the measured bounding range.
$\epsilon(\kappa)=$  $\;1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.48)
$\epsilon(\kappa_{s},N)=$  $\;1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.49)
[…]
When the underlying distribution is uniform, $\kappa_{s}=\sqrt{3}$ and $\epsilon=0$. Figure 6 shows that $0<\epsilon(N)\sim N^{-0.564}$, resulting in a measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.4).
In Figure 6, when the underlying distribution is Gaussian, the bounding range $\kappa(5,N)$ start to reach it stable value at sample count $N=10^{3}$. In contrast, when the underlying distribution is uniform, such sample count $N$ is $10^{2}$. This difference of the sample count upper bounds necessary for computing adjusted variances is caused by the underlying distributions.
2.11 Dependency Tracing
When all inputs satisfy the uncorrelated uncertainty assumption, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
  • Formula (2.50) expresses $\delta^{2}(f+g)$, whose dependency tracing is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.12). Formula (2.54) shows that Formula (2.3) applies Formula (2.50) between any two terms in the Taylor expansion in Formula (2.5).
  • Formula (2.51) expresses $\delta^{2}(fg)$, illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.14).
  • Formula (2.52) shows $\delta^{2}f(g(x))$, whose dependency tracing is demonstrated by $\delta^{2}(f^{-1}(f))=(\delta x)^{2}$.
  • Formula (2.53) gives the variance of the linear transformation of a function, which can be applied to Formula (2.50) and (2.51) for more general dependency tracing.
$\delta^{2}(f+g)=$ $\;\delta^{2}f+\delta^{2}g+2(\overline{fg}-\overline{f}\overline{g});$ (2.50)
$\delta^{2}(fg)=$ $\;\overline{f^{2}g^{2}}-\overline{fg}^{2};$ (2.51)
$\delta^{2}f(g)=$ $\;\overline{f(g)^{2}}-\overline{f(g)}^{2};$ (2.52)
$\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.53)
$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;\sigma^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+\sigma^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.54)
$+2\sigma^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance fit statistics rigorously. However, dependency tracing comes with a price: variance calculations are generally more complex than value calculations, exhibiting narrower convergence range for input variables. Dependency tracing also implies that the results of statistical Taylor expansion should remain path independent.
2.12 Traditional Execution and Dependency Problem
Dependency tracing requires an analytic form of the function to apply statistical Taylor expansion for the result mean and variance, as in Formula (2.6), (2.3), (2.9), and (2.3). This requirement often conflicts with conventional numerical methods for on analytic functions:
  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships between original input variables.
  • Similarly, conditional executions are commonly employed to optimize performance and minimize rounding errors, such as using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [37]. For dependency tracing, such conditional executions should be replaced by direct matrix inversion as described in Section 6.
  • Furthermore, traditional approaches frequently apply approximations to the result values during executions. Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than the value in Taylor expansion. Consequently, approximation strategies should prioritize variances than values. Section 6 illustrates this with a first-order approximation for computing a matrix determinant.
  • Traditionally, results from mathematical library functions are accepted without question, with accuracy assumed down to the last bit. As demonstrated in Section 8, statistical Taylor expansion enables the detection of numerical errors within these functions and necessitates recalculating them with uncertainty explicitly included in the output.
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, and division  [^5] Sometimes square root is also regarded as an arithmetic operation.. However, this decomposition introduces dependency problems in floating-point arithmetic, interval arithmetic, and statistical Taylor expansion. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
  • Similarly, large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$. This approach fails in statistical Taylor expansion because dependency tracing within $g(x)$ affects $f(g(x))$. In this context, $\overline{f(g(x))}\neq\overline{f(y)|_{y=g(x)}}$ and $\delta^{2}f(g(x))\neq\delta^{2}f(y)|_{y=g(x)}$. The path dependence of $f(y)|_{y=g(x)}$ and $\delta^{2}f(y)|_{y=g(x)}$ are evident in cases such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
Dependency tracing thus removes nearly all flexibility in traditional numerical executions, eliminating associated dependency problems. Consequently, all conventional numerical algorithms require reevaluation or redesign to align with the principles of statistical Taylor expansion.
3 Variance Arithmetic
This paper confines the calculating of variances to ideal variances $\widehat{\delta^{2}f}$ when $\kappa=5$, and assumes that that the input distribution is Gaussian, although variance arithmetic also implements statistical Taylor expansion for other cases. Furthermore, for discussion simplicity, Taylor coefficients in Formula (2.5) and (2.8) are all assumed to be precise.
3.1 Numerical Representation
Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. Other conventional numerical numbers need to convert to this format.
If the least $20$ significant bits of its significand are all 0 in a standard floating-point number, the value is considered precise, representing a 2’s fractional with a possibility no less than $1-2^{-20}=1-2.384\;10^{-7}$. Otherwise, a standard floating-point value is considered imprecise with the uncertainty given as $1/\sqrt{3}$-fold of the ULP of the value, where ULP stands for Unit in the Last Place for a conventional floating-point number [9]. This follows from the fact that the pure rounding error of round-to-nearest is uniformly distributed within half bit of the significand of the floating-point value [1].
If an integer number is within the range $2^{53}+1,+2^{53}-1]$ of the significand of a 64-bit standard floating-point number, its uncertainty is zero. Otherwise, it is first converted to a conventional floating-point value before being converted to an imprecise value.
3.2 Calculation Rules
Variance arithmetic implements statistical Taylor expansion with several practical assumptions. Due to finite precision and range of conventional floating-point representation, $\zeta(n)$ can be computed only for $n<448$ before becoming infinite under the IEEE 64-bit floating-point format, Consequently, the following numerical rules are introduced:
  • finite: The result must remain finite.
  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion to an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$. Unless all the remaining terms in the expansion are known to be precise zeros, an expansion is always executed to the full 448 terms for the monotonic check.
  • positive: At any order, the expansion variance must be positive.
  • stable: For sufficiently fast convergence, the value of the last expansion term must be less than the $\epsilon(5)=5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value.
  • reliable: At any order, the value of the variance must be less than $1/\kappa$-fold of the uncertainty of the variance.
The validity of these numerical rules of variance arithmetic will be demonstrated in subsequent sections.
3.2.1 Finite
For $(1\pm\delta x)^{-2}$, the Taylor coefficient increases with the expansion order $n$ as $(-1)^{n}(n+1)$. When $\delta x=0.5$, this growth causes the result variance to diverge to infinity.
3.2.2 Monotonic
Figure 7: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$. When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
Figure 8: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$. When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
From Formula (2.8), the convergence condition applied to $\log(x\pm\delta x)$ is $P(x)\leq 1/\kappa=1/5$, which is numerically confirmed as $P(x)\lesssim 0.20086$. Beyond this upper bound, the expansion is no longer monotonic. Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.29) mathematically, with $\zeta(2n)$ as the connection of these two perspectives.
For $e^{x\pm\delta x}$ the convergence holds for $\delta x\lesssim 19.864$ regardless of $x$, while the result $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$. These limits follow directly from the relationship $\delta x\rightarrow P(e^{x})$ and $P(x)\rightarrow\delta\log(x)$, as indicated in Formula (2.26) and (2.29).
From Formula (2.44), and except when $c$ is a natural number, Formula (2.35) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$. This trend is approximately confirmed in Figure 7, and beyond the upper bound $P(x)$, the expansion is no longer monotonic. Qualitatively, $\delta^{2}1/x$ converges more slowly than $\delta^{2}\sqrt{x}$, consistent with Formula (2.42) and (2.41).
When the input uncertainty is Uniformly distributed, Figure 8 illustrates that the input uncertainty upper bound is close to $1/\sqrt{3}$ according to Formula (2.44). The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such the expansion in Formula (2.44) contains $652$ terms instead of $448$ terms. The upper bounds of input uncertainty in Figure 7 and 8 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 7 are $3$ order-of-magnitude smaller than those in Figure 8 when the exponent is less than 0. This suggests that convergence is not the only consideration in statistical Taylor expansion because uncertainty bias or uncertainty can also invalidate a result in practice.
3.2.3 Positive
Figure 9: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
Figure 10: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$.
In some cases, the variance expansion may produce negative results, as in Formula (2.32) for $sin(x\pm\delta x)$. Figure 9 demonstrates that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$, which is less than $\pi$ of $\sin(x)$. Beyond this upper bound, the expansion is no longer positive in Figure 9.
In Figure 10, when the bounded momentum $\zeta(2n)$ for Uniform distribution is used instead of that for Normal distribution, the input uncertainty upper bound is $4$ times larger than that in Figure 9 but is still periodic. The resulting uncertainty increases not much and still remains less than $1$.
3.2.4 Stable
The unstable condition happens rarely.
3.2.5 Reliable
The condition of not being reliable seldom happens.
3.3 Continuity
Figure 11: Histograms of normalized errors for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space. For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remains continuous around $c=n$. Table 1 shows that the result of $(1\pm\delta x)^{n\pm d},0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former had upper bound for $\delta x$, while the latter does not.
A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity. The presence of such poles does not disrupt continuity in of the result mean, variance, or histogram. Figure 12 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
  • When the second derivative is zero, the resulting distribution is symmetric two-sided and Delta-like, such as when $n=3,x=0$.
  • When the second derivative is positive, the resulting distribution is right-sided Delta-like, such as the distribution when $n=2,x=0$, or when $n=2,x=\pm 0.2$, or when $n=3,x=0.2$.
  • When the second derivative is negative, the resulted distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
  • In each case, the transition from $x=0$ to $x=0.2$ is continuous.
A statistical bounding range inf variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative. An illustrative example is $\sin(x)$ in Figure 9.
A statistical bounding range in variance arithmetic cannot include any distributional zero because the result will diverge, such as at $x=0$ for both $(x\pm\delta x)^{c},c<1$ and $\log(x\pm\delta x)$.
$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
Value $1\mp 2.155\;10^{-8}$ $1\pm 2.073\;10^{-8}$ $041\pm 6.065\;10^{-8} 1.1$$122\pm 1.328\;10^{-7}$
Uncertainty $0+2.127\;10^{-7}$ $0.201-1.358\;10^{-6}$ $0.407\pm 2.201\;10^{-7}$ $0.654\pm 2.784\;10^{-7}$
Table 1: The result value and uncertainty of $(1\pm\delta x)^{n\pm d}$ vs $(1\pm\delta x)^{n}$, in which $n$ is a natural number, $0<d\ll 1$, and $\delta x$ is the upper bound for $(1\pm\delta x)^{n\pm d}$. The value and the uncertainty are expressed as the difference with the corresponding value and uncertainty of those of $(1\pm\delta x)^{n}$.
3.4 Tracking Rounding Error
Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x axis). The y-axis to the left shows both the value and the uncertainty of the residual errors. The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
$\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}=\sum_{j=0}^{N}\tilde{x}^{j}\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (3.1)
Variance arithmetic can track rounding errors effectively without the need for additional rules.
Figure 12 illustrates the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (3.1), $\frac{1}{1-x}$ is computed using Formula (2.33), and $x$ is initiated as a floating-point value. Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (3.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$. Figure 12 shows:
  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, meaning that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$. The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
  • When $x\not\in[-0.74,+0.75]$, for example when x=0.75,0.76x=-0.75,0.76, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$. The residual error magnitude increases as $|x|\rightarrow 1$, reaching about $50$ when $x=0.98$.
3.5 Comparison
Two imprecise values can be compared statistically for their difference.
When the value difference is zero, the two imprecise values are equal. In statistics, such two imprecise values have $50\%$ possibility to be less than or greater to each other but no chance to be equal to each other [5]. In variance arithmetic, however, they are treated as neither less nor greater than each other, thus they are equal.
Otherwise, the standard z-statistic method [5] is applied to determine whether they are equal, less than, or more than each other. For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$, and the probability for them to be not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5]. If the threshold probability for inequality is $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$. Alternatively, an equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for an equal probability threshold of $50\%$.
4 Verification of Statistical Taylor Expansion
4.1 Verification Methods and Standards
Statistical Taylor expansion is verified through the validation of variance arithmetic. Analytic functions or algorithms with precisely known results are used to compare the outputs of variance arithmetic using the following statistical properties:
  • Value error: the difference between the numerical result and the corresponding known precise analytic result.
  • Value deviation: the standard deviation of the value errors.
  • Normalized error: the ratio of a value error to the corresponding uncertainty.
  • Error deviation: the standard deviation of normalized errors.
  • Error distribution: the histogram of the normalized errors.
  • Uncertainty mean: the sample mean of the result uncertainties.
  • Uncertainty response: the relationship between input uncertainties and output uncertainties.
  • Calculation response: the relationship between the amount of calculation and output uncertainties.
One objective of uncertainty-based calculation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
  1. 1.
    The error deviation is exactly $1$.
  2. 2.
    The error distribution should follow Normal distribution when an imprecise value is expected, or Delta distribution when a precise value is expected.
  3. 3.
    The uncertainty response should match the expected functional form. For a linear function, the output uncertainties should increase linearly with input uncertainties, with the ratio of the output to input uncertainty means defined as uncertainty response ratio.
  4. 4.
    The calculation response should follow the expected trend; for instance, increasing the amount of calculations should result in larger uncertainties.
If the precise result is unknown, as outlined in Section 6, the result error distribution can be used to determine whether ideal coverage is achieved.
If however, the input uncertainty is only accurate to the order of magnitude, the proper coverage is obtained when the error deviations fall within the range of $[0.1,10]$.
When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained. The minimal noise deviation required provides a good estimation of the magnitude of unspecified input uncertainties. The presence of ideal coverage is a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly in the given context. The input noise range yielding ideal coverage defines the ideal application range for input uncertainties.
Gaussian noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
4.2 Types of Uncertainties
There are five primary sources of result uncertainty in a calculation [1][2][13]:
  • Input uncertainties: The examples in this paper demonstrate that when input uncertainty precision is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
  • Rounding errors: Empirical results indicate that variance arithmetic can provide proper coverage for rounding errors.
  • Truncation errors: Variance arithmetic avoids truncation errors with its stable rule.
  • External errors: External errors are the value errors which are not specified in input uncertainties, such as numerical errors in library functions. Section 8 examines the effects of numerical errors in $\sin$ and $\tan$, showing that when these external errors are sufficiently large enough, neither ideal coverage nor proper coverage can be achieved. This finding indicates that library functions must be recalculated to include the corresponding uncertainty of each value.
  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to obtain a solution. For example, the predecessor of statistical Taylor expansion [1] demonstrated that the discrete Fourier transformation is only an approximation for the mathematically defined continuous Fourier transformation, containing modeling errors. Conceptually, modeling errors originate in mathematics and are therefore outside the domain of statistical Taylor expansion.
4.3 Types of Calculations to Verify
Algorithms of distinct nature each representative of its category are required to test the broad applicability of variance arithmetic [1]. An algorithm can be categorized by comparing the amount of its input and output data as [1]:
  • Application,
  • Transformation,
  • Generation,
  • Reduction.
An application algorithm computes numerical values from an analytic formula. Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems.
A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information contained remaining largely unchanged. For reversible transformations, a unique requirement is to recover the original input uncertainties after a round-trip transformation: performing a forward transformation followed by its reverse transformation. DFT (discrete Fourier transformation) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back to the input using essentially the same process. A test of variance arithmetic using FFT (fast Fourier transformation, which is an implementation of DFT) algorithms is presented in Section 8.
A generation algorithm produces substantially more output data than input data. Such algorithms encode mathematical knowledge into data. Some generation algorithms are theoretical calculations which involve no imprecise input, so all resulting uncertainty arises from rounding errors. Section 9 presents a generation algorithm, which generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or statistical characterization of a data set. In this process, certain information is lost while other information is extracted. Conventional wisdom is that a reduction algorithm generally benefits from a larger input data set [5]. A test of statistical characterization through sampling is described in Section 5.
5 Mathematical Library Functions
Figure 13: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $e^{x\pm\delta x}$, for different $x$ (x-axis), and different $\delta x$ (legend).
Figure 14: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $\log(x\pm\delta x)$, for different $x$ (x-axis), and different $\delta x$ (legend).
Figure 15: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $(1\pm\delta x)^{c}$, for different $c$ (x-axis), and different $\delta x$ (legend).
Figure 16: Error deviation for $(1\pm\delta x)^{c}$ as a function of $c$ and $\delta x$. The x-axis represents $c$ value between 2-2 and +3+3. The y-axis represents $\delta x$ value between $10^{-16}$ and $1$. The z-axis shows the corresponding error deviation.
Figure 17: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $\sin(x\pm\delta x)$, for different $x$ (x-axis), and different $\delta x$ (legend).
Figure 18: Error deviation for $\sin(x\pm\delta x)$ as a function of $x$ and $\delta x$. The x-axis represents $x$ value between $\pi$ and $\pi$. The y-axis represents $\delta x$ value between $10^{-16}$ and $1$. The z-axis shows the corresponding error deviation.
Figure 19: Values and uncertainties of $\log(e^{x})-x$ and $e^{\log(x)}-x$ as functions of $x$, evaluated at $0.1$ increment. When $x$ is 2’s fractional such as $1/2$ or $1$, the result uncertainties are significantly smaller.
Figure 20: Normalized errors of $(x^{p})^{\frac{1}{p}}-x$ as functions of $x$ and $p$.
Formula (2.26), (2.29), (2.32), and (2.35) are evaluated using the corresponding mathematical library functions exp, log, sin, and pow, respectively.
At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as follows:
  1. 1.
    Generate $10000$ samples from a Gaussian noise distribution, each with $\delta x$ as the distributional deviation, and construct $\tilde{x}$ by adding the sampled noise to $x$.
  2. 2.
    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between using $\tilde{x}$ and using $x$ as the input.
  3. 3.
    Divide the value error by the result uncertainty to obtain the normalized error.
  4. 4.
    Calculate the standard deviation of the normalized errors to determine the error deviation.
5.1 Exponential
Figure 13 demonstrates that the calculated uncertainties using Formula (2.26) align closely with the measured value deviations for $e^{x+\delta x}$. Consequently, all error deviations remain very close to $1$, even though both the uncertainty and the value deviations increase exponentially with $x$ and $\delta x$.
5.2 Logarithm
Because $\log(x)$ has a distributional zero at $x=0$ all $log(x\pm\delta x)$ values are rejected when $P(x)>1/5$. Figure 14 demonstrates that the uncertainties calculated using Formula (2.29) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
5.3 Power
Figure 15 demonstrates that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.35) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11. Figure 16 reveals that the error deviations of $\sin(x+\delta x)$ align close to 11 for all the input exponent $c$ and the input uncertainty $\delta x>10^{15}$.
5.4 Sine
Figure 17 demonstrate that the uncertainties calculated using Formula (2.32) correspond closely to the measured value deviations for $\sin(x+\delta x)$. It also reveals that $\delta^{2}\sin(x)$ has the same periodicity as $\sin(x)$:
  • When $x=0$, $\sin(x)\simeq x$, so that $\delta^{2}\sin(x)\simeq(\delta x)^{2}$.
  • When $x=\pi/2$, $\sin(x)\simeq 1$, so that $\delta^{2}\sin(x)\simeq 0$.
Since $\sin(x)$ has distributional poles at $x=\pm\pi/2$, Figure 18 shows that the error deviation for $\sin(x+\delta x)$ equals $1$ except when $x=\pm\pi/2$ and $\delta x<10^{-8}$, matching the expected Delta distribution these poles. Elsewhere, the error deviations remain close to $1$.
5.5 Numerical Errors for Library Functions
The combined numerical error of the library functions $e^{x}$ and $\log(x)$ is evaluated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$. Using either variance arithmetic or conventional floating-point library functions produce identical value errors. Figure 19 demonstrates the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$. Figure 19 also shows that $e^{\log(x)}$ has much less error than $\log(e^{x})$. For $\log(e^{x})-x$, the error deviation is $0.409$ when $|x|\leq 1$ or approaches zero otherwise. The reason for surprising small value errors for $e^{\log(x)}-x$ is not clear at the moment.
The numerical error of the library function $x^{p}$ is computed as $(x^{p})^{1/p}-x$. Figure 20 shows that the normalized errors do not depend on either $x$ or $p$, resulting in an error deviation $0.548$.
The numerical errors of the library functions $\sin(x)$, $\cos(x)$, and $\tan(x)$ will be examined in greater detail in Section 8.
5.6 Summary
Formula (2.26), (2.29), (2.35), and (2.32) provide effective estimates for their respective uncertainties for library functions. With added noise larger than $10^{-15}$, ideal coverage is achieved unless near a distributional pole where the error deviation approaches 0. In other cases, proper coverage is attainable.
6 Matrix Calculations
6.1 Matrix Determinant
Let vector $p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [37]. Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$37]. Formula (6.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [37]. The sub matrix $\mathbf{M_{i,j}}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (6.2) [37]. For discussion simplicity, sub determinant $|\mathbf{M}|_{i,j}$ in Formula (6.2) contains the permutation sign, which is different from the determinant of the sub matrix $|\mathbf{M}_{i,j}|$ [37] that treats the sub matrix $\mathbf{M_{i,j}}$ as an independent matrix. Formula (6.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [37].
$|\mathbf{M}|$ $\equiv\sum_{[p_{1}\dots p_{n}]_{n}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}x_{k,p_{k}};$ (6.1)
$|\mathbf{M}|_{i,j}$ $\equiv\sum_{[p_{1}\dots p_{n}]_{n}}^{p_{i}=j}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}^{k\neq i}x_{k,p_{k}};$ (6.2)
$|\mathbf{M}|$ $=\sum_{j=1\dots n}|\mathbf{M}_{i,j}|x_{i,j}=\sum_{i=1\dots n}|\mathbf{M}|_{i,j}x_{i,j};$ (6.3)
Let $i_{1},i_{2},\dots>$} denote an ordered permutation of a subset from $1\dots n$, and $i_{1},i_{2},\dots]$} an unordered permutation [37]. Apply Formula (6.3) progressively to $M_{i,j}$, to expand Formula (6.2) as (6.4), and Formula (6.1) as (6.5). The (nm)(n-m)-by-(nm)(n-m) sub matrix in $|\mathbf{M}_{<i_{1}\dots i_{m}>_{n}[j_{1}\dots j_{m}]_{n}}|$ is obtained by deleting the rows in $\{i_{1}\dots i_{m}\}$ and the columns in $\{j_{1}\dots j_{m}\}$. This leads to Formula (6.6).
$|\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\equiv\sum_{[p_{1}\dots p_{n}]_{n}}^{k\in\{i_{1}\dots i_{m}\}:p_{k}=j_{k}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}^{k\not\in\{i_{1}\dots i_{m}\}}x_{k,p_{k}};$ (6.4)
$|\mathbf{M}|$ $=\sum_{{[j_{1}\dots j_{m}]}_{n}}|\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{k=1}^{m}x_{i_{k},j_{k}};$ (6.5)
$\left||\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\right|=\left||\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<j_{1}\dots j_{m}>_{n}|}\right|;$ (6.6)
[…]
$=\sum_{m=0\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$
$\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (6.8)
$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<j_{1}\dots j_{m}>_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (6.9)
Formula (6.8) and (6.9) assume that the uncertainties of matrix elements are all independent of each other. This assumption maximized the result uncertainties. For discussion simplicity, other uncertainty assumptions are ignored in this paper.
6.2 Adjugate Matrix
Figure 21: Uncertainty means (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
Figure 22: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
Figure 23: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (6.10).
Figure 24: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (6.11).
The square matrix whose element is $a_{i,j}=(-1)^{i+j}|\mathbf{M}_{j,i}|$ is defined as the adjugate matrix [37] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$. Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [37]. Formula (6.10) and (6.11) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ [37].
$\mathbf{M}\times\mathbf{M}^{A}$ $=\mathbf{M}^{A}\times\mathbf{M}=|\mathbf{M}|\mathbf{I};$ (6.10)
$|\mathbf{M}|\;\mathbf{M}^{A}$ $=|\mathbf{M}^{A}|\;\mathbf{M};$ (6.11)
To test Formula (6.9):
  1. 1.
    A matrix $\mathbf{M}$ is constructed using random integers uniformly distributed in the range of $2^{8},+2^{8}]$, which has a distribution deviation of $2^{8}/\sqrt{3}$. The integer arithmetic ensures that $|\mathbf{M}|$, $\mathbf{M}^{A}$, and $|\mathbf{M}^{A}|$ are precise.
  2. 2.
[…]
The difference between $\widetilde{\mathbf{M}}^{A}$ and $\mathbf{M}^{A}$ defines the Adjugate Test. Figure 21 shows that the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$. Figure 22 shows that the ideal coverage is achieved for the input precision except at matrix size $8$ and input precision $10^{-17}$. The existence of ideal coverage validates Formula (6.9).
Formula (6.10) defines the Forward Test. Figure 23 shows that the difference of the two sides of Formula (6.10) is precise zero, whether it is $\widetilde{\mathbf{M}}\times\widetilde{\mathbf{M}}^{A}-|\widetilde{\mathbf{M}}|\mathbf{I}$, or $\widetilde{\mathbf{M}}^{A}\times\widetilde{\mathbf{M}}-|\widetilde{\mathbf{M}}|\mathbf{I}$. The difference in Formula (6.10) is closer to Delta distribution for larger input precision, because without input noise, rounding errors dominate value errors. The validation of Formula (6.9) leads naturally to the validation of Formula (6.10).
Formula (6.11) defines the Roundtrip Test. Similarly, Figure 24 validates Formula (6.11). Because roundtrip test is no longer linear, error deviation in Figure 24 no longer increases linearly with input precision.
6.3 Floating Point Rounding Errors
Figure 25: Histograms of normalized errors of the adjugate matrix as a function of matrix size without input noise (legend).
Figure 26: Histograms of normalized errors of the adjugate matrix as a function of matrix size with $10^{-11}$ input noise (legend).
The significand of the conventional floating-point representation [9] has $53$-bit resolution. Figure 25 shows that the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$. When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs, which results in non-Delta distribution in Figure 25. The rounding error is also the reason why only at matrix size 8 and input noise precision $10^{-10}$, error deviation is no longer $1$ in Figure 22.
With $10^{-11}$ noises added to the input, Figure 26, the distribution becomes Gaussian with a hint of Delta distribution. Such Delta-like distribution persists until the input noise precision reaches $10^{-10}$, which is also the transition of the two trends in Figure 21. Figure 26 shows the distribution when the input noise precision is $10^{-11}$, where the distinction due to matrix size vanishes.
6.4 First Order Approximation
[…]
$|\widetilde{\mathbf{M}}|\simeq\sum_{[p_{1}\dots p_{n}]_{n}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}(x_{k,p_{k}}+\tilde{x}_{k,p_{k}});\;\;\;\Rightarrow\;\;\;\delta^{2}|\mathbf{M}|\simeq\sum_{i}^{n}\sum_{j}^{n}M_{i,j}(\delta x_{i,j})^{2};$ (6.12)
Figure 27 contains the result of applying Formula (6.12). It is very similar to Figure 22, validating Formula (6.12).
6.5 Matrix Inversion
Figure 28: The linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis). The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
$\mathbf{M}^{-1}\equiv$ $\;\mathbf{M}^{A}/|\mathbf{M};$ (6.13)
$\mathbf{M}^{-1}\times\mathbf{M}=\mathbf{M}\times\mathbf{M}^{-1}=\mathbf{I};$ (6.14)
$(\mathbf{M}^{-1})^{-1}=\mathbf{M};$ (6.15)
An inverse matrix is defined by Formula (6.13), which satisfies Formula (6.14) and (6.15) [37]. However, this definition is seldom used conventionally to calculate inverse matrix due to large uncertainty response ratio which appears as small input uncertainties causing large output uncertainties [13][37]. Traditionally, matrix condition number [37] is a proxy for uncertainty response ratio of matrix inversion. In Formula (6.13), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$. Figure 28 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices. As a reference, Figure 28 presents the Hilbert matrix [37] for each matrix size, and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number. Thus, determinant precision can replace condition number to estimate uncertainty response ratio of matrix inversion.
$\mathbf{M}=$ $\;\left(\begin{matrix}w,x\\ y,z\end{matrix}\right);\;\;\;\mathbf{M}^{-1}=\frac{\left(\begin{matrix}z,-x\\ -y,w\end{matrix}\right)}{wz-xy};$ (6.16)
$\delta^{2}\mathbf{M}^{-1}\simeq$ $\;\frac{\left(\begin{matrix}z^{4},x^{2}z^{2}\\ y^{2}z^{2},x^{2}y^{2}\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}y^{2}z^{2},w^{2}z^{2}\\ y^{4},w^{2}y^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{4}}+$ (6.17)
$\;\frac{\left(\begin{matrix}x^{2}z^{2},x^{4}\\ w^{2}z^{2},w^{2}x^{2}\end{matrix}\right)(\delta y)^{2}+\left(\begin{matrix}x^{2}y^{2},w^{2}x^{2}\\ w^{2}y^{2},w^{4}\end{matrix}\right)(\delta z)^{2}}{(wz-xy)^{4}};$
$\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}\simeq$ $\;\frac{\left(\begin{matrix}2z^{3},-2xz^{2}\\ -2yz^{2},2xyz\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}2y^{2}z,-2wyz\\ -2y^{3},2wy^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{3}}+$ (6.18)
$\frac{\left(\begin{matrix}2x^{2}z,-2x^{3}\\ -2wxz,2wx^{2}\end{matrix}\right)(\delta y)^{2}+\left(\begin{matrix}2wxy,-2w^{2}x\\ -2w^{2}y,2w^{3}\end{matrix}\right)(\delta z)^{2}}{(wz-xy)^{3}};$
Variance arithmetic is path-independent, so it computes inverse matrix directly from the definition of Formula (6.13). For example, for the inverse matrix $\mathbf{M}^{-1}$ of size $2$ in Formula (6.16), Formula (6.17) and (6.18) present first-order approximations for variance $\delta^{2}\mathbf{M}^{-1}$, and bias $\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}$, respectively. The resulting uncertainties of determinant $\delta^{2}|\mathbf{M}|$ and inversion $\delta^{2}\mathbf{M}^{-1}$, as well as the resulting bias $\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}$, are not linear to each input uncertainty $\delta x_{i,j}$. Inverse variance $\delta^{2}\mathbf{M}^{-1}$ contains sum of all input variance $(\delta x_{i,j})^{2}$ and their higher-order permutation products, such that the uncertainty response ratio for matrix inversion should roughly equal the matrix size. Thus, an uncertainty response ratio for matrix inversion is inherently large, independent of computational path. A seemly small uncertainty response ratio using Gaussian elimination [13][37] is probably a path-dependent artifact; while the conventionally “bad” method of applying Formula (6.13) for matrix inversion [13][37] is actually the correct one, except missing resulting uncertainty when using floating-point arithmetic. All conventional path-dependent results are questionable. Only statistical Taylor expansion presents a complete picture of resulting uncertainty for an analytic expression.
It is doubtful if Formula (6.15) still holds for uncertainty in statistical Taylor expansion, because it seems that uncertainty response ratio can only increase in matrix inversion according to Formula (6.17). On the other hand, because the bias in Formula (6.18) can be either positive or negative, it is expected that Formula (6.15) still holds for value.
7 Moving-Window Linear Regression
7.1 Moving-Window Linear Regression Algorithm
Figure 29: Result of fitting $\alpha+\beta\;Y$ to a time-series input $Y$ within a moving window of size 22+12*2+1. The x-axis indicates the time index. The y-axis on the left corresponds to the value of $Y$, $\alpha$, and $\beta$, while the y-axis on the right corresponds to the uncertainty of $\alpha$ and $\beta$. The uncertainty for $Y$ is fixed at $0.2$. In the legend, Unadjusted refers to results obtained by directly applying Formula (7.5) and (7.6) using variance arithmetic, whereas Adjusted refers to using Formula (7.5) and (7.6) for $\alpha$ and $\beta$ values but Formula (7.7) and (7.8) for their variances.
Figure 30: Error deviations of the $\alpha+\beta\;Y$ fit vs time index. The x-axis represents the time index. The y-axis on the left corresponds to the error deviation. For reference, the input time-series signal $Y$ is also plotted, with its values indicated on the y-axis on the right.
Formula (7.1) and (7.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [13].
$\alpha$ $=\frac{\sum_{j}Y_{j}}{\sum_{j}1};$ (7.1)
$\beta$ $=\frac{\sum_{j}X_{j}Y_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}Y_{j}}{\sum_{j}X_{j}X_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}X_{j}};$ (7.2)
In many applications, data set $Y_{j}$} represents an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is referred to as a time-series input, where $j$ corresponds to $X_{j}$. A moving window algorithm [13] is applied within a small window centered on each $j$. For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying window’s half width. This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (7.1) and (7.2) into Formula (7.3) and (7.4), respectively [1]:
$\alpha_{j}$ $=\alpha\;2H=\sum_{X=-H+1}^{H}Y_{j-H+X};$ (7.3)
$\beta_{j}$ $=\beta\;\frac{H(H+1)(2H+1)}{3}=\sum_{X=-H}^{H}XY_{j-H+X};$ (7.4)
The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (7.3) and (7.4) to be reformulated into the progressive moving-window calculation given by Formula (7.5) and (7.6), respectively [1]
$\beta_{j}$ $=\beta_{j-1}-\alpha_{j-1}+H\left(Y_{j-2H-1}+Y_{j}\right);$ (7.5)
$\alpha_{j}$ $=\alpha_{j-1}-Y_{j-2H-1}+Y_{j};$ (7.6)
7.2 Variance Adjustment
When the time series contains uncertainty, directly applying Formula (7.5) and (7.6) will lead to a loss of precision, as both formulas reuse each input multiple times, accumulating the variance of that input at every usage. To avoid this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (7.6) and (7.5), respectively, while the variance should instead be computed using Formula (7.7) and (7.8), respectively. Notably, Formula (7.8) is no longer progressive.
$\delta^{2}\alpha_{j}$ $=\sum_{X=-H+1}^{H}(\delta Y_{j-H+X})^{2}=\delta^{2}\alpha_{j-1}+(\delta Y_{j})^{2}-(\delta Y_{j-2H})^{2};$ (7.7)
$\delta^{2}\beta_{j}$ $=\sum_{X=-H+1}^{H}X^{2}(\delta Y_{j-H+X})^{2};$ (7.8)
Figure 29 shows that the input signal $Y_{j}$ consists of the following components:
  1. 1.
    An increasing slope for $j=0\dots 9$.
  2. 2.
    A decreasing slope for $j=1\dots 39$.
  3. 3.
    A sudden jump of magnitude +10+10 at $j=40$
  4. 4.
    A decreasing slope for $j=41\dots 49$.
For each increment of $j$, the increasing and the decreasing rates are +1+1 and 1-1, respectively.
The specified input uncertainty is consistently $0.2$. Normal noises with a deviation of $0.2$ are added to the slopes, except for the segment $j=10\dots 19$ where Normal noises with a deviation of $2$ are introduced, representing a $10$-fold actual value over the specified uncertainty.
Figure 29 also presents the result of moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$. The fitted values of $\alpha$ and $\beta$ follow the expected behave with a characteristic delay of $H$ in $j$. When (7.3) and (7.4) are applied to compute the uncertainties for $\alpha$ and $\beta$, they both increase exponentially with the time index $j$. In contrast, when Formula (7.3) and (7.4) are used exclusively for value calculation, while Formula (7.7) and (7.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$. Both are less than the input uncertainty $\delta Y$, due to the averaging effect of the moving window.
7.3 Unspecified Input Error
To determine the error deviation of $\alpha$ and $\beta$, the fitting procedure is applied to multiple time-series data sets, each generated with independent noise realizations. Figure 30 illustrates the resulting error deviation as a function of the time index $j$, which remains close to $1$ except in the range $j=10\dots 19$ where the actual noise is $10$-times greater than the specified value. This observation indicates that a larger than $1$ error deviation may signal the presence of unspecified additional input errors beyond rounding errors, such as the numerical errors in mathematical library functions.
8 FFT (Fast Fourier Transformation)
8.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the DFT (discrete Fourier transformation) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is given by Formula (8.1) and (8.2) [13], respectively. In these expressions, $k$ denotes the time index while $n$ represents the frequency index. The frequency index and time index are not necessarily related to time unit and frequency unit, respectively. The naming is just a convenient way to distinguish the two opposite domains in DFT: the waveform domain $h[k]$ versus the frequency domain $H[n]$.
$H[n]=\sum_{k=0}^{N-1}h[k]\;e^{\frac{i2\pi}{N}kn};$ (8.1)
$h[k]=\frac{1}{N}\sum_{n=0}^{N-1}H[n]\;e^{-\frac{i2\pi}{N}nk};$ (8.2)
8.2 FFT (Fast Fourier Transformation)
When $N=2^{L}$, where $L$ is a natural number, the generalized Danielson-Lanczos lemma [13] can be applied to DFT to produce FFT [13].
  • For each output, each input is used only once, so there is no dependency problem when decomposing FFT into arithmetic operations involving Formula (2.11), (2.12), (2.13), and (2.14).
  • When $L$ is large, the large volume of input and output data enables high-quality statistical analysis.
  • The computational complexity is proportional to $L$, because increasing $L$ by 1 adds one more step involving a sum of multiplications.
  • Each step in the forward transformation increases the variance by a factor of $2$, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty mean decreases with $L$ as $\sqrt{1/2}^{L}$. The uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
  • Forward and reverse transformations are identical except for a sign difference, so they are essentially the same algorithm, and any observed difference arises purely from the input data.
In normal usage, forward and reverse FFT transforms differ in their data prospective of time domain versus frequency domain:
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum where most values are 0, causing its uncertainties to grow faster.
  • The reverse transformation spreads the precise frequency-domain spectrum (with most values being 0) back into a time-domain sine or cosine signals, causing its uncertainties to grow more slowly.
8.3 Modeling Errors of DFT and FFT
Although mathematically self-consistent, by implying a periodic boundary condition in the time domain [1], DFT and its efficient implementation FFT are only approximate for the mathematically defined continuous Fourier transformation [1]. They show no modeling error only when the input signal frequency $f$ satisfies $f=j\frac{2\pi}{N}$, and exhibit different degrees of modeling errors otherwise, with the modeling errors peaking at $f=(j+\frac{1}{2})\frac{2\pi}{N}$. Because of these modeling errors, using DFT and FFT as the digital implementation of continuous Fourier transformation is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [13].
To avoid the modeling errors of FFT, only Formula (8.1) and (8.2) are used in this paper.
8.4 Testing Signals
The following signals are used for testing:
  • Sin: $h[k]=\sin(2\pi kf/N),f=1,2,...N/2-1$.
  • Cos: $h[k]=\cos(2\pi kf/N),f=1,2,...N/2-1$.
  • Linear: $h[k]=k$, whose DFT is given by Formula (8.3).
    $y\equiv i2\pi\frac{n}{N}:\;\;\;G(y)=\sum_{k=0}^{N-1}e^{yk}=\frac{e^{Ny}-1}{e^{y}-1};$
    $H[n]$ $=\frac{dG}{dy}=\begin{cases}n=0:\;\;\;\frac{N(N-1)}{2}\\ n\neq 0:\;\;\;-\frac{N}{2}(1+i\frac{\cos(n\frac{\pi}{N})}{\sin(n\frac{\pi}{N})})\end{cases};$ (8.3)
Empirically, except when using trigonometric library directly for FFT transformations of clean Sin and Cos signals:
  • Results from Sin and Cos signals are statistically indistinguishable.
  • Results from Sin signals at different frequencies are also statistically indistinguishable.
Therefore, the results for Sin and Cos signals at all frequencies are pooled together for statistical analysis, under the combined category Sin/Cos signals.
8.5 Trigonometric Library Errors
Figure 31: Difference between library and Quart $\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend). The uncertainties of the Quart $\sin(x)$ is $\sin(x)$ ULP, which shows a periodicity of $\pi$.
Figure 32: Difference between the library and the Quart $\cos(x)/\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend).
Formula (8.1) and (8.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, in which $L$ is the FFT order. To minimize numerical errors in computing $\sin(x)$, indexed sine can replace standard library sine functions:
  1. 1.
    Instead of a floating-point value $x$ as input for $\sin(x)$, an integer index $j$ defines the input as $\sin(\pi j/2^{L})$, thereby eliminating the floating-point rounding error of $x$.
  2. 2.
    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
  3. 3.
    The values of $\sin(\pi j/2^{L})$ are extended from $j\in[0,2^{L-1}]$ to $j\in[0,2^{L+1}]$ by exploiting the symmetry of $\sin(\pi j/2^{L})$.
  4. 4.
    The values of $\sin(\pi j/2^{L})$ are extended to all the integer value of $j$ by leveraging the periodicity of $\sin(2\pi j/2^{L})$.
Such constructed indexed $\sin(x)$ is named as the Quart indexed sine function. In contrast, the direct application of library $\sin(x)$ is named as the Library sine function.
Because the Quart sine function obeys the symmetry and periodicity of sine function strictly, it is superior to the library function.
  • Figure 31 shows that the value difference between Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases roughly linearly with $|x|$.
  • Figure 32 shows the value difference between Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but they are $10^{2}$ times larger than those of $\sin(x)$. Therefore, the linear spectrum in Formula (8.3) may contain huge numerical errors using library sine functions.
For both sine functions, the uncertainty of each $\sin(x)$ is assumed to be its ULP. Because Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in library $\sin(x)$, its true numerical errors is probably less than ULP, therefore its uncertainty should be sightly overestimated. In contrast, Figure 31 shows that library sine function has underestimated uncertainties.
8.6 Using Quart Sine for Sin/Cos Signals
Figure 33: FFT spectrum of $\sin(2\pi 3j/2^{6})$ using Quart sine function after the forward transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 34: FFT waveform of $\sin(2\pi 3j/2^{6})$ using the Quart sine function after the reverse transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 35: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal at frequency 33 as a function of FFT order (x axis) and transformations (legend) using Quart sine function.
Figure 36: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using Quart sine function. The FFT order is $18$.
Using Quart sine function, for a sine wave with $3$ as frequency, Figure 33 displays the spectrum for the forward transformation, while Figure 34 presents the waveform for the reverse transformation.
  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation is $0.31$. In the reverse transformation, the error deviation is $0.43$. Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
  • The uncertainty mean of the forward transformation is $7$ times larger than that of the reverse transformation.
Figure 35 illustrates that the forward transformation differs from the reverse transformation significantly.
  1. 1.
    When FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that in the reverse transformation. At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times more than that of the reverse transformation.
  2. 2.
    At FFT order $18$, Figure 36 shows that the normalized error distribution for the reverse transformation is wider than that for the forward transformation, but the two distributions do not differ by order-of-magnitude. Therefore, the value error grows at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
  3. 3.
    In forward transformation, error deviations reach their stable values quickly when FFT order $L\geq 4$. In contrast, in the reverse transformation, error deviations stabilize slowly with increasing FFT order until $L\geq 15$. The slow growth of normalized errors in the reverse transformation is attributed to its input data which contains all precise zeros except at the two frequency indexes.
Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases.
8.7 Using Library Sine for Sin/Cos Signals
Figure 37: FFT spectrum of $\sin(2\pi 3j/2^{6})$ using Library sine function after the forward transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 38: FFT waveform of $\sin(2\pi 3j/2^{6})$ using Library sine function after the reverse transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 39: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin signal at frequency 33 as a function of FFT order (x axis) and transformations (legend) using Library sine function.
Figure 40: Histograms of normalized errors of Sin signal at frequency 33 for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
Using library sine function, for a sine wave with $3$ as frequency, Figure 37 displays the spectrum for the forward transformation, while Figure 38 presents the waveform for the reverse transformation:
  • The uncertainties are identical to the corresponding values using Quart sine function, because the input uncertainties are identical for these two sine functions. This identical relationship presents in all FFT orders when comparing the corresponding uncertainty means in Figure 35 and 39.
  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in Library sine function are more than those using Quart sine function in Figure 31.
When compared with Figure 36, Figure 40 also confirms that at FFT order $18$, normalized errors have much wider distributions using Library sine function than those using Quart sine function.
Due to numerical errors in Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 39.
8.8 Numerical Error Resonance Using Library Sine
Figure 41: Error deviation (z-axis) of FFT forward transformation of $sin(2\pi fj/2^{L})$ as a function of frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
Figure 42: Error deviation (z-axis) of FFT reverse transformation of $sin(2\pi fj/2^{L})$ as a function of frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
Value errors in the reverse transformation show a clear trend of increasing with time index, as in Figure 41. These large value errors looks systematic and do not appear as typical noise visually; or rather, they look more like a resonant pattern. Such increases present in other frequencies and FFT order, as well as in computational result using mathematical libraries such as SciPy. A FFT spectrum of the forward transformation shows a sudden increase of intensity at frequency index which is also consistent with the result of SciPy, as in Figure 37. Figure 41 and 42 show that error deviations of such value errors increase with sine or cosine frequencies, the resonant beats between the signal and the numerical library errors become stronger. In contrast, such increase does not appear at all when using Quart sine function, such as in Figure 33 and 34. Figure 31 shows that the numerical errors using Library sine increase with a periodicity of $\pi$, which may resonate with a signal which has a periodicity of an integer fold of $\pi$, to produce the resonant pattern in Figure 37 and 38. To avoid the numerical error resonance in Figure 41 and 42, $10^{-14}$ input noise is required to add to the input sine or cosine signals.
8.9 Using Quart Sine for Linear Signal
Figure 43: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal as a function of FFT order (x axis) and transformations (legend) using Quart sine function.
Figure 44: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) using Quart sine function. The FFT order is $18$.
Figure 43 shows that using Quart sine function can track the result uncertainties of Linear signals with proper coverage. Because the input to reverse transform no longer contains mostly precise zeros, the output uncertainty increases far more rapidly with FFT order than its counterpart in Figure 35.
Figure 44 shows that for each transformation, the normalized error distribution in reverse transformation is similar to its counterpart of the Sin/Cos signals using Quart sine function in Figure 36, while the distribution in the forward transformation is similar to the Linear signal using Library sine function in Figure 46. The reason for this disparity in similarity of the normalized error distributions is not clear.
8.10 Using Library Sine for Linear Signal
Figure 45: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal as a function of FFT order (x axis) and transformations (legend) using Library sine function.
Figure 46: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
As shown as the difference between Figure 31 and 32, linear signals using Library sine function may introduce $10^{3}$ times larger numerical errors into the results through Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these extra errors or not.
Figure 45 shows that proper coverage can no longer be achieved because value error outpaces uncertainty as FFT order increases. Figure 46 shows that the normalized error distribution for the reverse transformation is no longer bound effectively within three deviations. Although the normalized error distribution for the forward transformation using Library sine function looks similar to its counterpart of using Quart sine function in Figure 46 looks similar to that using Quart sine function in Figure 44, detailed analysis shows that the normalized error distribution using Library sine function contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal, respectively. The presence of these extreme values consistent with those delta-like numerical errors from Figure 32.
The observed difficulty of variance arithmetic in tracking Linear signals using Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified errors. A potential solution is to develop a new sin library implemented within variance arithmetic framework, ensuring that all numerical calculation errors are explicitly accounted for, as what Quart sine function has achieved in FFT transformations.
8.11 Ideal Coverage
Figure 47: Histograms of normalized errors of Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
Figure 48: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals as a function of FFT order (x axis) and transformations (legend) using Library sine function.
Figure 49: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for forward transformations of Linear signals using Library sine function.
Figure 50: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Library sine function.
Figure 51: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for forward transformations of Linear signals using Quart sine function.
Figure 52: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Quart sine function.
Adding sufficient noise to the input can overpower unspecified input errors, thereby achieving ideal coverage; for example, applying an input noise of $10^{-3}$ to a Linear signal when using library sine function.
  • Figure 47 presents the corresponding histogram. As expected, the normalized errors for the forward and reverse transformations follow Normal distributed, while those for the roundtrip transformations are Delta distributed at 0zero, indicating perfect recovery of the input uncertainties.
  • Figure 48 illustrates the corresponding error deviations and uncertainty means:
    • As expected, the result uncertainty means for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
    • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $\sqrt{1/2}^{L}$.
    • As expected, the result uncertainty means for the roundtrip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
    • As expected, the result error deviations for the forward and reverse transformations are constant at $1$, whereas those for the roundtrip transformation decay exponentially to 0 with increasing FFT order.
Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, which is expected because FFT is a linear algorithm [13].
The range of ideal coverage depends on how accurately the input uncertainty reflects the actual input noise. For Linear signals using Library sine function, Figure 49 and 50 present the error deviation as a function of the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation is $1$. Outside this region, proper coverage cannot be achievable. Because uncertainties grow more slowly in reverse transformation than in forward transformation, the reverse transformations exhibit a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that yields ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually, beyond FFT order $25$ for the reverse transformation, ideal coverage may not be achievable. Although FFT is regarded as one of the most robust numerical algorithms very insensitive to input errors, it can still fail due to numerical errors in library sine function. Such deterioration of the calculation is not easily detectable using conventional floating-point arithmetic.
In contrast, for Linear signals using Quart sine functions, as shown in Figure 51 and 52 for the forward and reverse transformations, respectively, the ideal coverage region is significantly larger, and proper coverage is also achieved in other regions. Forward transformations show a larger ideal coverage region than that of reverse transformations.
As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders. The difference is that using Quart sine functions, within the proper coverage region, error deviations differ from $1$ only marginally.
8.12 Prec Sine Functions
Figure 53: Value Errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) as a function of $x$ (x-axis) and $L$ using either Prec or Quart sine function (legend).
Figure 54: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals as a function of FFT order (x axis) and transformations (legend) using Prec sine function.
Figure 55: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals using Prec sine functions.
Figure 56: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Prec sine function.
Quart sine function over-estimates the uncertainty, while Library sine function underestimates the uncertainty. Is it possible to have a better sine function? If Quart sine function is constructed using the 128-bit floating-point library of gcc, then converted to 64-bit values, using the difference between a 128-bit and the corresponding 64-bit value as the uncertainty for the 64-bit value, the resulting indexed sine function is named as the Prec sine function.
Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare Prec and Quart sine functions, Figure 53 illustrates that Prec sine function produces slightly more value errors, while having half uncertainties. Figure 54 and 55 both show that Prec sine function underestimates uncertainties slight, with stable error deviation around $1.8$ for both forward and reverse transformations without added noise. Compared with 52, Figure 56 illustrates that the ideal coverage using Prec sine function is identical to using Quart sine function. The observation that 128-bit sine function does not improve over 64-bit sine function in any noticeable way challenges the legitimacy of using the difference as the uncertainty, the correctness of the 128-bit gcc numerical library, or the algorithm in converting a 128-bit floating-point value to 64-bit.
Both Quart and Prec sine function are not ideal. An ideal sine function should have error deviations closer to $1$ without added noise.
8.13 Summary
The library sine functions implemented with conventional floating-point arithmetic have been shown to contain numerical errors equivalent to $10^{-3}$ of input precision in worst cases for FFT transformations. These library $\sin(x)$ errors increase periodically with $x$, which may resonate with periodic input signals. The computational result using floating-point arithmetic is very sensitive to input data, amount of calculation, and library function implementation. This means that a small-scale test cannot properly qualify the result of a large-scale calculation. The impact of numerical errors within mathematical libraries has not received sufficient attention. The numerical errors using floating-point arithmetic are very difficult to detect and quantify.
Variance arithmetic offers a robust alternative, as its computed values closely match those from conventional floating-point arithmetic while its associated uncertainties trace all input errors. Moreover, its resulting error deviations enable objective classification of calculation quality as ideal, proper, or suspicious.
9 Regressive Generation of Sin and Cos
Figure 57: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ as a function of $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
Figure 58: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ as a function of regression order (x-axis) when compared with Quart sine function (legend)
Formula (9.2) and Formula (9.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (9.1). Formula (9.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, so that value errors will not accumulate when the regression order increases.
$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (9.1)
$\sin(\frac{\alpha+\beta}{2})=\sqrt{\frac{1-\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1-\cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)}{2}};$ (9.2)
$\cos(\frac{\alpha+\beta}{2})=\sqrt{\frac{1+\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1+\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)}{2}};$ (9.3)
$\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (9.4)
Formula (9.2) is not suitable to compute $\sin(x)$ when $x\rightarrow 0$ because it fails similarly as catastrophic cancellation, obtaining excessive uncertainty. In Figure 57, Quart sine has uncertainties proportional to $\sin(\pi j/2^{L})$, while regression sine has its uncertainties increase when $x\rightarrow 0$.
Figure 58 tests Quart and Regression sin/cos for $\sin(x)^{2}+\cos(x)^{2}-1=0$. In Figure 58, the result uncertainties are nearly a constant independent of the regression order $L$ when $L\geq 4$ for both Regression sine and Quart sine. It also shows that the value errors for both are quite comparable, resulting in the error deviation of the Regression sin/cos closer to $1$.
10 Conclusion and Discussion
10.1 Summary of Statistical Taylor Expansion
The starting point of statistical Taylor expansion is the assumption that all input variables have uncorrelated uncertainties. This is considered a very reasonable statistical requirement on the input data [1]. Once this assumption is met, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors. It can track the variable dependence in the intermediate steps using standard statistical methods. While it eliminates the dependency problem, it requires a more rigorous process to determine resulting mean and deviation for an analytic expression. In addition, it can reject calculations on unified statistical and mathematical grounds. It also incorporates sample count into statistical analyses.
10.2 Summary of Variance Arithmetic
Variance arithmetic simplifies statistical Taylor expansion by assuming that all inputs are Gaussian distributed and with enough sample counts for ideal variance. Ill-formed problems can invalidate resulting variance in several ways: non-convergent, unstable, negative, infinite, and unreliable.
The presence of ideal coverage is a necessary condition for a numerical algorithm in variance arithmetic to be considered correct. Ideal coverage also defines the algorithm’s optimal applicable range for the algorithm. Under ideal coverage, the calculated uncertainty equals the deviation of the value errors, and the normalized errors follow either a Normal or Delta distribution depending on the context.
Variance arithmetic can provide proper coverage for floating-point rounding errors.
The applicability of variance arithmetic has been demonstrated in diverse scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
10.3 Improvements Needed
This paper presents variance arithmetic which is still in its early developmental stage. From a theoretical perspective, several important questions remain to be addressed.
Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty. Without this, the value errors in the library functions can produce unpredictable and potentially large effects on numerical results.
bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized.
Determining the uncertainty upper bounds for each analytic expression analytically is important. Under current numerical framework, the measured upper bounds in Figure 7 for $(x+\delta x)^{c}$, and Figure 9 for $\sin(x+\delta x)$ may change with implementation such as between 32-bit and 64-bit floating-point computations.
Cases where $\delta^{2}f<\widehat{\delta^{2}f}$ due to low sample count require further theoretical study and experimental verification.
The performance of variance arithmetic has to be improved. The fundamental formulas for statistical Taylor expansion, Formula (2.6), (2.3), (2.9), and (2.3) contains large number of independent summations, making them excellent candidates for parallel processing, such as direct hardware implementation. The procedural nature of these formulas allows statistical Taylor expansion to be implemented directly in hardware.
When an analytic expression undergoes statistical Taylor expansion, the result expression may be very complex, such as in matrix inversion. However, modern symbolic computation tools such as SymPy can greatly facilitate these calculations. This suggests that it may be time to shift from purely numerical programming toward analytic programming for problems with inherently analytic formulations.
As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty. This capability can help engineers pinpoint the primary drivers of measurement inaccuracy and thus guide targeted improvements in data acquisition and processing strategies.
A key open question is whether variance arithmetic can be adapted to provide ideal coverage for floating-point rounding errors. Once the mechanism for rounding error is thoroughly understood, variance arithmetic with ideal coverage to floating-point rounding errors is quite valuable because many theoretical calculations lack explicit input uncertainty.
Because traditional numerical approaches are based on floating-point arithmetic, they need to be reexamined or even reinvented in variance arithmetic. For example, most conventional numerical algorithms aim to select optimal computational paths, whereas variance arithmetic fundamentally rejects conceptually all path-dependent calculations. Reconciling these two paradigms may present a significant challenge.
Establishing theoretical foundation to apply statistical Taylor expansion in the absence of analytic solution, or with only limited low-order numerical derivatives, such as in solving differential equations, remains an important research direction.
10.4 Acknowledgments
As an independent researcher without institutional affiliation, the author expresses deep gratitude for the encouragement and valuable discussions provided by Dr. Zhong Zhong (Brookhaven National Laboratory), Prof Weigang Qiu (Hunter College). Special thanks are extended to the organizers of AMCS 2005, particularly Prof. Hamid R. Arabnia (University of Georgia), and organizers of NKS Mathematica Forum 2007. Prof Dongfeng Wu (Louisville University) is acknowledged for insightful guidance on statistical topics. The author is also indebted to the editors and reviewers of Reliable Computing for their substantial assistance in shaping and accepting a previous version of this work, with special recognition to Managing Editor Prof. Rolph Baker Kearfott.
11 Statements and Declarations
Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.
12 Data Availability Statement
The datasets used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic. The execution assistance and explanation of the above code are available from the author upon reasonable request.

### v12

@@
 Abstract
-As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions, to compute means and standard deviations of the results.
-Statistical Taylor expansion traces the dependency of the input uncertainties in the intermediate steps, so that the variables in the intermediate analytic expressions can no longer be regarded as independent of each other, and the result of the analytic expression is path independent.
-Thus, it differs fundamentally from the conventional common approaches in applied mathematics which optimize execution path for each calculation.
-In fact, statistical Taylor expansion may standardize numerical calculations for analytic expressions.
-Its statistical nature allows religious testing of its result when the sample size is large enough.
-This paper also introduces an implementation of statistical Taylor expansion called variance arithmetic and presents corresponding test results in a very wide range of mathematical applications.
-Another important conclusion of this paper is that the numerical errors in the library function can have significant effects on the result.
-For example, the periodic numerical errors in the trigonometric library functions can resonate with periodic signals, producing large numerical errors in the results.
+As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions to compute the means and standard deviations of the results.
+Statistical Taylor expansion tracks the propagation of the input uncertainties through intermediate steps, causing the variables in intermediate analytic expressions to become interdependent, while the final analytic result becomes path independent.
+This fundamentally distinguishes it from common approaches in applied mathematics that optimize computational path for each calculation.
+In essence, statistical Taylor expansion may standardize numerical computations for analytic expressions.
+Its statistical nature enables reliable testing of results when the sample size is sufficiently large, as well as includes sample count in result qualification.
+This study also introduces an implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
+Another important conclusion of this study is that numerical errors in library functions can significantly affect results.
+For instance, periodic numerical errors in trigonometric library functions may resonate with periodic signals, producing large numerical errors in the outcomes.
 Keywords: computer arithmetic, error analysis, interval arithmetic, uncertainty, numerical algorithms.
@@
 Except for the most basic counting, scientific and engineering measurements never yield completely precise results [2][3].
-In such measurements, the uncertainty of a quantity $x$ is typically expressed by either the sample deviation $\delta x$ or the uncertainty range $\Delta x$ [2][3].
+In such measurements, the uncertainty of a quantity $x$ is typically expressed either by the sample deviation $\delta x$ or by the uncertainty range $\Delta x$ [2][3].
   • If $\delta x=0$ or $\Delta x=0$, $x$ is a precise value.
   • Otherwise, $x$ is an imprecise value.
-$P(x)\equiv\delta x/|x|$ is defined as the statistical precision (hereafter referred to simply as precision) of the measurement, where $x$ denotes the value, and $\delta x$ represents the uncertainty deviation.
+$P(x)\equiv\delta x/|x|$ is defined as the statistical precision (hereafter referred to as precision) of the measurement, where $x$ denotes the value, and $\delta x$ represents the uncertainty deviation.
 A larger precision indicates a coarser measurement whereas a smaller precision indicates a finer measurement.
 The precision of measured values can range from order-of-magnitude estimates to $10^{-2}$ to $10^{-4}$ in common measurements, and to $10^{-15}$ in state-of-the-art determinations of fundamental physical constants [4].
-The focus of this paper is on determining the result value and uncertainty of a general analytic expression with imprecise input values.
+This study focuses on determining the result value and uncertainty of a general analytic expression with imprecise input values.
 1.2 Extension to Existing Statistics
-Rather than focusing on the result uncertainty distribution when applying a function to a random variable with known uncertainty distribution [5], statistical Taylor expansion provides the result mean and variance for general analytic expressions without the need for the result uncertainty distribution.
-  • Previous work has addressed the effect of input uncertainties to output values for special cases [6].
-Statistical Taylor expansion generalizes this as the uncertainty bias as shown in Formula (2.6), and (2.9) in this paper.
-  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [5][6], whereas statistical Taylor expansion extends this to higher-order interactions as expressed in Formula (2.3) in this paper.
+Rather than focusing on the uncertainty distribution of the result when applying a function to a random variable with a known uncertainty distribution [5], statistical Taylor expansion provides the mean and variance of the result for general analytic expressions.
+  • Previous studies have examined the effect of input uncertainties on output values for specific cases [6].
+Statistical Taylor expansion generalizes these effects as uncertainty bias, as shown in Formula (2.9) and (2.12) in this paper.
+  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [5][6], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper.
+As shown in Formula (7.9), (7.17), and (7.18), such extension is important even for linear algebra.
 1.3 Problem of the Related Numerical Arithmetics
@@
 1.3.1 Conventional Floating-Point Arithmetic
-Conventional floating-point arithmetic [7][8][9] assumes that every bit of its significand is valid at all times, and continuously either generates artificial information or erase useful information to maintain this assumption during its normalization process [10].
+Conventional floating-point arithmetic [7][8][9] treats every bit of its significand to be valid at all times, and continuously either generates artificial information or erases useful information to maintain this assumption during the normalization process [10].
 For example, the following calculation is performed exactly in integer format:
@@
   1. 1.
-    The normalization of the input values may pad zero bits artificially in the least significant positions until every bit in the 53-bit significand are used, for example, converting $64919224$ to $64919224.000000000$.
+    The normalization of input values may artificially pad zero bits in the least significant positions until every bit in the 53-bit significand is used, for example, converting $64919224$ to $64919224.000000000$.
   2. 2.
-    The addition or multiplication results may exceed the bit range of the 53-bit significand, which is round off from the least significant bits, to fit into the 53-bit significand.
-This process may generate rounding errors, for example, from $13316075197586561$ to $13316075197586560$, to generate a rounding error of $1$ at the least significant bit.
+    The results of addition or multiplication may exceed the bit range of the 53-bit significand, requiring rounding off from the least significant bits to fit within the 53-bit significand.
+This process can generate rounding errors, for example, converting $13316075197586561$ to $13316075197586560$, which produces a rounding error of $1$ in the least significant bit.
   3. 3.
-    The subtraction or division results can cancel values from high significant bits, exposing the rounding errors in the low significant bits, for example, 13316075197586562.13316075197586560.=2.13316075197586562.-13316075197586560.=2..
+    The results of subtraction or division can cancel values from higher significant bits, exposing rounding errors in the lower significant bits.
+For example, 13316075197586562.13316075197586560.=213316075197586562.-13316075197586560.=2.
   4. 4.
-    The normalization of the subtraction or division results can amplify the rounding error toward the most significant bit by padding zeros, producing catastrophic cancellation [11][12] in this case.
+    The normalization of subtraction or division results can amplify rounding errors toward the most significant bit by padding zeros, such as converting $2.$ to $2.0000000000000000$.
+Formula (1.2) demonstrates a classic case of catastrophic cancellation [11][12].
 Because rounding errors from lower digits can propagate to higher digits, the $10^{-7}$ significance of the 32-bit IEEE floating-point format [7][8][9] is usually insufficient for calculations involving input data with a precision of $10^{-2}$ to $10^{-4}$.
-For more complex calculations, even the $10^{-16}$ significance of the 64-bit IEEE floating-point format [7][8][9] may not enough for inputs with $10^{-2}$ to $10^{-4}$ precision.
-This is a fundamental controversy of the conventional floating-point arithmetic.
-Because rounding-error is path-dependent, a major effort of conventional numerical methods is to identify optimal computation strategies to minimize rounding errors, such as in Gaussian elimination [12][13][14], even though all other paths of root finding methods in linear algebra are equivalent mathematically [37].
-Self-censoring rules have been developed to limit such rounding error propagation [13][14], such as avoiding the subtraction of results from large multiplications, as in Formula (1.2).
+For more complex calculations, even the $10^{-16}$ significance of the 64-bit IEEE floating-point format [7][8][9] may be sufficient for inputs with $10^{-2}$ to $10^{-4}$ precision.
+This represents a fundamental controversy of conventional floating-point arithmetic.
+Because rounding error is path dependent, a major objective of conventional numerical methods is to identify optimal computation strategies that minimize rounding errors, such as Gaussian elimination [12][13][14], even though all alternative paths of root finding methods in linear algebra are mathematically equivalent [37].
+Self-censoring rules have been developed to limit such rounding error propagation [13][14], for example, by avoiding the subtraction of results from large multiplications, as shown in Formula (1.2).
 However, these rules are neither enforceable nor easily adoptable, and they are even more difficult to quantify.
-To date, research on rounding error propagation has focused mainly either on linear calculations [11][12][15], or special cases [13][16][17], whereas in practice, rounding errors often appear as the pervasive and mysterious numerical instability [18].
-The forward rounding error study [12] compares 1) the result with rounding error and 2) the ideal result without rounding error, such as comparing the result using 32-bit IEEE floating-point arithmetic with the corresponding result using 64-bit IEEE floating-point arithmetic [19].
-The most recent study of this type presents an extremely optimistic view of numerical library errors, reporting them as a fraction of the least significant bit of the floating-point significands [19].
-However, such optimism contradicts statistical tests on numerical library functions as presented in this paper.
-The backward rounding error study [11][12][15] only estimates the result uncertainty due to rounding errors, thereby overlooking the bias caused by rounding errors on the result value.
-This analysis is typically restricted to very small uncertainty uncertainties because it relies on perturbation theory, and it tailored to each specific algorithm [11][12][15].
-In contrast, variance arithmetic traces the rounding error directly as part of the uncertainty.
-Statistical Taylor expansion applies generically to any analytic function, for both result mean and deviation, and accommodates input uncertainties of any magnitude.
-By demonstrating that the analytic result should be path-independent, statistical Taylor expansion fundamentally challenges the conventional methodology for seeking optimal execution strategy for a given analytic expression.
+To date, research on rounding error propagation has focused primarily on linear calculations [11][12][15], or on special cases [13][16][17], whereas in practice, rounding errors often manifest as pervasive and mysterious numerical instabilities [18].
+The forward rounding error study [12] compares (1) the result containing rounding error and (2) the ideal result without rounding error, for example, by comparing a result obtained using 32-bit IEEE floating-point arithmetic with the corresponding result computed using 64-bit IEEE floating-point arithmetic [19].
+The most recent study of this type presents an extremely optimistic view of numerical library errors, reporting them as fractions of the least significant bit of the floating-point significand [19].
+However, such optimism contradicts the statistical tests on numerical library functions presented in this paper.
+The backward rounding error study [11][12][15] estimates only the result uncertainty caused by rounding errors, thereby overlooking the bias that rounding errors introduce into the result value.
+This analysis is typically limited to very small uncertainties because it relies on perturbation theory and it is tailored to each specific algorithm [11][12][15].
+In contrast, variance arithmetic traces rounding error directly as part of uncertainty.
+Statistical Taylor expansion applies generally to any analytic function, providing both result mean and deviation, and accommodates input uncertainties of any magnitude.
+By demonstrating that the analytic result should be path independent, statistical Taylor expansion fundamentally challenges the conventional methodology of seeking an optimal execution strategy for a given analytic expression.
 1.3.2 Interval Arithmetic
-Interval arithmetic [14][20][21][22][23][24] is currently a standard method to track calculation uncertainty.
-Its goal is to ensure that a value x remains absolutely bound within its bounding range throughout the computation.
-However, the bounding range in interval arithmetic is not compatible with the approach commonly employed in scientific and engineering measurements, which instead characterizes uncertainty using the statistical mean and deviation [2][3]  [^2] There is one attempt to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics [25].
+Interval arithmetic [14][20][21][22][23][24] is currently a standard method for tracking computational uncertainty.
+Its objective is to ensure that a value remains strictly bounded within its bounding range throughout the computation.
+However, the bounding range in interval arithmetic is not compatible with the approach commonly used in scientific and engineering measurements, which instead characterizes uncertainty in terms of the statistical mean and deviation [2][3]  [^2] There is one attempt to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics [25].
 Because this attempt seems to rely heavily on 1) specific properties of the uncertainty distribution within the interval and/or 2) specific properties of the functions upon which the interval arithmetic is used, this attempt does not seem to be generic.
@@
 Interval arithmetic represents only the worst-case scenario of uncertainty propagation.
-For example, in addition, it assumes that the two input variables are perfectly positively correlated [26], thereby yielding the widest possible bounding range.
-In contrast, if the variables were perfectly negatively correlated, the bounding range after addition would reduce [27]  [^3] Such case is called the best case in random interval arithmetic.
+For example, in addition, it assumes that the two input variables are perfectly positively correlated [26], thereby producing the widest possible bounding range.
+In contrast, if the variables were perfectly negatively correlated, the bounding range after addition would decrease [27]  [^3] Such case is called the best case in random interval arithmetic.
 The vast overestimation of bounding ranges in these two worst cases prompts the development of affine arithmetic [26][28], which traces error sources using a first-order model.
@@
 .
-This worst-case assumption can lead to order-of-magnitude overestimates [1].
+This worst-case assumption can lead to order-of-magnitude over-estimations [1].
 The results of interval arithmetic can depend strongly on the specific algebraic form of an analytic function $f(x)$, a phenomenon known as the dependency problem.
-This issue is amplified in interval arithmetic [23] but also presents in conventional floating-point arithmetic [13].
+This issue is amplified in interval arithmetic [23] but also exists in conventional floating-point arithmetic [13].
 Furthermore, interval arithmetic lacks a mechanism to reject invalid calculations, even though every mathematical operation has a valid input range.
 For example, it produces branched results for $1/(x\pm\Delta x)$ or $\sqrt{x\pm\Delta x}$ when $0\in[x-\Delta x,x+\Delta x]$, whereas a context-sensitive uncertainty bearing arithmetic should reject such calculation naturally.
-In contrast, variance arithmetic specifies a value with mean and deviation.
-It has no dependency problem.
-Its statistical context rejects certain input intervals mathematically, such as inversion and square root when the statistical bounding range contains zero.
+In contrast, variance arithmetic specifies each value by its mean and deviation.
+It does not suffer from the dependency problem.
+Its statistical framework naturally rejects certain input intervals on mathematical grounds, such as inversion and square root operations when the statistical bounding range includes zero.
 1.3.3 Statistical Propagation of Uncertainty
-Statistical propagation of uncertainty treats the uncertainty of a value as the random part of each imprecise value.
-When the statistical correlation between any two input variables is known, statistical propagation can be used to determine result mean and variance [29][30].
-In contrast, statistical Taylor expansion treats the uncertainty of a value as the limitation of obtaining the precise value.
+Statistical propagation of uncertainty treats each imprecise value as a random variable and seeks correlations among such random variables to calculate resulting mean and deviation [29].
+In contrast, statistical Taylor expansion interprets the uncertainty of an imprecise value as a limitation in obtaining its precise value.
 This assumption aligns with most error analyses in the literature [2][3].
-Its statistical foundation is the uncorrelated uncertainty assumption [1]: Each input variable is measured independently with fine enough precision, so that their uncertainties are independent of each other, even though the input variables could be highly correlated.
-This statistical foundation can be turned into a quantitative test on the input variables on the precision vs the correlation [1].
-Statistical propagation of uncertainty seems to have applied a wrong statistical context for uncertainty.
-For example, a time series is a random variable while each value in the time series is just an imprecise value.
-If two such imprecise values have correlated uncertainty, they have systematic errors [2].
-On the other hand, the uncorrelated uncertainty assumption can hold for each value in the time series, while the time series can be self-correlated [5] such as being periodic.
+Its statistical foundation is the uncorrelated uncertainty assumption [1]: any two imprecise values do not correlate in their uncertainties.
+Statistical propagation of uncertainty appears to have applied an incorrect statistical context to uncertainty.
+For example, a time series is a random variable whereas each imprecise value within the time series is merely an imprecise measurement.
+The uncorrelated uncertainty assumption can hold for each imprecise value in a time series, even though the time series can highly correlate to another time series at their signal level, or even correlated to itself as a periodic time series.
+If the uncertainties of some imprecise values are correlated, these imprecise values contain systematic errors among them [2].
+Systematic errors should be treated as unwanted signals but not as normal uncertainties [2].
+In short, uncertainty should not be treated as signal.
 1.3.4 Significance Arithmetic
-Significance arithmetic [31] seeks to track the number of reliable bits in an imprecise value through calculation.
-In two early implementations [32][33], significance arithmetic was based on simple operational rules applied to reliable bit counts, rather than on formal statistical approaches.
-In these methods, the reliable bit count is treated as an integer, even though in practice it can take a fractional value [34].
-This constraint can cause an artificial step-wise reduction in significance.
+Significance arithmetic [31] seeks to track the number of reliable bits in an imprecise value throughout a calculation.
+In its early implementations [32][33], significance arithmetic was based on simple operational rules applied to reliable bit counts rather than on formal statistical methods.
+In these approaches, the reliable bit count is treated as an integer, even though in practice it can take a fractional value [34].
+This constraint can produce an artificial step-wise reduction in significance.
 The implementation of significance arithmetic in Mathematica [34] employs a linear error model consistent with the first-order approximation of interval arithmetic [14][22][23].
 One limitation of significance arithmetic is its inability to specify uncertainty accurately [1].
-For example, if the least significant bit of significand is used to represent uncertainty, the result precision can be very coarse as in $1\pm 10^{-3}=1024\times 2^{-10}$ [1].
-Introducing a limited number of bits calculated inside uncertainty does not fully resolve this issue.
-Therefore, various attempts for floating-point arithmetic without normalization [33] are not widely adopted, and the conventional floating-point arithmetic dominates the numerical world.
-For this reason, variance arithmetic has abandoned the significance arithmetic principle of its predecessor [1].
+For example, if the least significant bit of the significand is used to represent uncertainty, the result precision can be very coarse, as in $1\pm 10^{-3}=1024\times 2^{-10}$ [1].
+Introducing a limited number of bits to represent uncertainty does not fully resolve this issue.
+Consequently, various attempts to develop floating-point arithmetic without normalization [33] have not been widely adopted, and conventional floating-point arithmetic continues to dominate the numerical world.
+For this reason, variance arithmetic has abandoned the significance arithmetic principles of its predecessor [1].
 1.3.5 Stochastic Arithmetic
-Stochastic arithmetic [35][36] randomizes the least significant bits (LSB) of each input floating-point value, repeats the same calculation multiple times, and then uses statistical analysis to identify invariant digits among the calculation results as significant digits.
-However, this approach can require excessive computation as the number of repetitions needed for each input is algorithm-independent, particularly when the algorithm contains branches.
+Stochastic arithmetic [35][36] randomizes the least significant bits of each input floating-point value, repeats the same calculation multiple times, and then applies statistical analysis to identify the invariant digits among the results as significant digits.
+However, this approach can be computationally expensive as the number of repetitions required for each input depends on the algorithm and can increase substantially when the algorithm contains branching operations.
 In contrast, statistical Taylor expansion provides a direct characterization of the result’s mean and deviation without sampling.
@@
 This paper presents the theory of statistical Taylor expansion, its implementation as variance arithmetic, and the corresponding validation tests.
-Section 1 compares statistical Taylor expansion and variance arithmetic with other established uncertainty-bearing arithmetic.
+Section 1 compares statistical Taylor expansion and variance arithmetic with other established uncertainty bearing arithmetic.
 Section 2 develops the theoretical foundation of statistical Taylor expansion.
-Section 3 describes the variance arithmetic as an implementation of the statistical Taylor expansion.
-Section 4 discusses the standards and methodologies used to validate variance arithmetic.
-Section 5 evaluates variance arithmetic for common mathematical library functions.
+Section 3 describes variance arithmetic as an implementation of statistical Taylor expansion.
+Section 4 outlines the standards and methodologies used to validate variance arithmetic.
+Section 5 illustrates variance arithmetic in computing polynomial.
+Section 6 evaluates variance arithmetic on common mathematical library functions.
 Section 7 applies variance arithmetic to adjugate matrix and matrix inversion.
-Section 7 applies variance arithmetic to process time-series data.
+Section 8 demonstrates its application to time-series data.
 Section 9 examines the impact of numerical library errors and shows that these errors can be significant.
@@
 2.1 The Uncorrelated Uncertainty Assumption
-The uncorrelated uncertainty assumption [1] states that the uncertainty of any input is statistically uncorrelated with the uncertainty of any other input.
-This condition is satisfied when uncorrelated noise is the primary contributor to the uncertainty of the signals.
-It is consistent with standard methods for processing experimental data [2][3].
-The uncorrelated uncertainty assumption permits the signals themselves to be highly correlated.
-Suppose two signals have a correlation coefficient $\gamma$, and measured precision $P_{1}$ and $P_{2}$, respectively.
-Let $P$ be the coarser of $P_{1}$ and $P_{2}$.
-At the level of $P$, the correlation is reduced to $\gamma_{P}$ according to Formula (2.1) [1].
-The value of $\gamma_{P}$ decreases rapidly as $P$ becomes finer, so that when $P$ is sufficiently fine, the correlation $\gamma_{P}$ of the uncertainties of two signals are effectively zero [1].
-$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}};$ (2.1)
-The uncorrelated uncertainty assumption ensures no systematic error [2].
+Figure 1: Effect of noise on bit values of a measured value.
+The triangular wave signal and the added white noise are shown at top using the thin black line and the grey area, respectively.
+The values are measured by a theoretical 4-bit Digital-to-Analog Converter in ideal condition, assuming LSB is the 0th bit.
+The measured 3rd and 2nd bits without the added noise are shown using thin black lines, while the mean values of the measured 3rd and 2nd bits with the added noise are shown using thin grey lines.
+This figure is a reproduction of Figure 1 in [1]
+Figure 2: $\gamma$ vs $P$ for different $\gamma_{P}$ in Formula (2.4).
+If $\gamma_{P}$ is the maximal allowed correlation on the uncertainty level for uncorrelated uncertainty assumption to hold, then $\gamma$ is the maximal allowed correlation for two input to hold.
+For a given $\gamma_{P}$, $\gamma$ increases with finer $P$.
+This figure is a reproduction of Figure 2 in [1].
+The uncorrelated uncertainty assumption [1] states that the uncertainties of any two input imprecise values are uncorrelated.
+This assumption is satisfied when there is no systematic errors in the inputs [2] and is consistent with standard methods for processing experimental data [2][3].
+The uncorrelated uncertainty assumption permits the signals themselves to be highly correlated in the following statistical test [1].
+Let $X$, $Y$, and $Z$ denote three mutually independent random variables [5] with variance $V(X)$, $V(Y)$ and $V(Z)$, respectively.
+Let $\alpha$ denote a constant.
+Let $Cov()$ denote the covariance function.
+Let $\gamma$ denote the correlation between $(X+Y)$ and $(\alpha X+Z)$.
+And let:
+$\eta_{1}^{2}\equiv\frac{V(Y)}{V(X)};\;\;\;\eta_{2}^{2}\equiv\frac{V(Z)}{V(\alpha X)}=\frac{V(Z)}{\alpha^{2}V(X)};$ (2.1)
+$\gamma=\frac{Cov(X+Y,\alpha X+Z)}{\sqrt{V(X+Y)}\sqrt{V(\alpha X+Z)}}=\frac{\alpha/|\alpha|}{\sqrt{1+\eta_{1}^{2}}\sqrt{1+\eta_{2}^{2}}}\equiv\frac{\alpha/|\alpha|}{1+\eta^{2}};$ (2.2)
+Formula (2.2) gives the correlation $\gamma$ between two random variables, each of which contains a completely uncorrelated part and a completely correlated part $X$, with $\eta$ being the average ratio between these two parts.
+Formula (2.2) can also be interpreted reversely: if two random variables are correlated by $\gamma$, each of them can be viewed hypothetically as containing a completely uncorrelated part and a completely correlated part, with $\eta$ being the average ratio between these two parts  [^4] The widely used common factor analysis is also based on this principle..
+The correlated parts common to different measurements are regarded as signals, which can either be desired or unwanted.
+One special application of Formula (2.2) is the correlation between a measured signal and its true signal, in which noise is the uncorrelated part between the two.
+Figure 1 shows the effect of noise on the most significant two bits of a 4-bit measured signal when $\eta=1/4$.
+Its top chart shows a triangular waveform between 0 and 16 as a black line, and a white noise between -2 and +2, using the grey area.
+The measured signal is the sum of the triangle waveform and the noise.
+The middle chart of Figure 1 shows the values of the 3rd digit of the true signal as a black line, and the mean values of the 3rd bit of the measurement as a grey line.
+The 3rd bit is affected by the noise during its transition between 0 and 1.
+For example, when the signal is slightly below 8, only a small positive noise can turn the 3rd digit from 0 to 1.
+The bottom chart of Figure 1 shows the values of the 2nd digit of the signal and the measurement as a black line and a grey line, respectively.
+Figure 1 clearly shows that the correlation between the measurement and the true signal is less at the 2nd digit than at the 3rd digit.
+Quantitatively:
+  • The overall measurement is 99.2% correlated to the signal with $\eta=1/8$;
+  • The 3rd digit of the measurement is 97.0% correlated to the signal with $\eta=1/4$;
+  • The 2nd digit of the measurement is 89.4% correlated to the signal with $\eta=1/2$;
+  • The 1st digit of the measurement is 70.7% correlated to the signal with $\eta=1$;
+  • The 0th digit of the measurement is 44.7% correlated to the signal with $\eta=2$.
+The above conclusion agrees with the common experiences that, below the noise level of measured signals, noises rather than true signals dominate each digit.
+Similarly, while the correlated portion between two values has the same value at each bit of the two values, the ratio of the uncorrelated portion to the correlated portion increases by 2-fold for each bit down from MSB of the two values.
+Quantitatively, let $P$ denote the precision of an imprecise value, and let $\eta_{P}$ denote the ratio of the uncorrelated portion to the correlated portion at level of uncertainty; then $\eta_{P}$ increases with decreased $P$ according to Formula (2.3).
+According to Formula (2.2), if two significant values are overall correlated with $\gamma$, at the level of uncertainty the correlation between the two values decreases to $\gamma_{P}$ according to Formula (2.4).
+$\eta_{P}=\frac{\eta}{P},\;\;\;P<1;$ (2.3)
+$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}},\;\;\;P<1;$ (2.4)
+Figure 2 plots the relation of $\gamma$ vs.
+$P$ for each given $\gamma_{P}$ in Formula (2.4).
+When $\gamma_{P}$ is less than a predefined maximal threshold (e.g., 2%, 5% or 10%), the two values can be deemed uncorrelated to each other at the level of uncertainty.
+For each independence standard $\gamma_{P}$, there is a maximal allowed correlation between two values below which the uncorrelated uncertainty assumption of statistical Taylor expansion holds.
+The maximal allowed correlation is a function of the larger precision of the two values according to Formula (2.4).
+Figure 2 shows that for two precisely measured values, their correlation $\gamma$ is allowed to be quite high.
 2.2 Distributional Zero and Distributional Pole
@@
 […]
-[…]
-$\rho(\tilde{z})d\tilde{z}$= $\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y}$; (2.2)
+$\rho(\tilde{x},\mu,\sigma)d\tilde{x}$ $=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y}$; (2.5)
 $y=x^{c}:\;\;\;\rho(\tilde{y},\mu_{y},\sigma_{y})=c\tilde{y}^{\frac{1}{c}-1}\frac{1}{\sigma}N(\frac{\tilde{y}^{\frac{1}{c}}-\mu}{\sigma});$ (2.6)
 […]
-  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, sine $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [5].
-A linear transformation generates neither a distributional zero nor a distributional pole, in accordance with to Formula (2.2).
-  • Figure 1 illustrates the probability density function for $(x\pm 1)^{2}$ according to Formula (2.3), which exhibits a distributional pole at $x=0$.
+  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, since $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [5].
+A linear transformation generates neither a distributional zero nor a distributional pole, according to Formula (2.5).
+  • Figure 3 shows the probability density function for $(x\pm 1)^{2}$ according to Formula (2.6), which exhibits a distributional pole at $x=0$.
 The distribution $(0\pm 1)^{2}$ corresponds to the $\chi^{2}$ distribution [2].
@@
 At the distributional zero, the probability density function is zero.
-In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ more when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero.
-This resemblance allows for a generic characterization of the output.
+In both Figure 3 and 4, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, thereby allowing for a generic characterization of the output.
 2.3 Statistical Taylor Expansion
 Formula (2.5) provides the uncertainty distribution of an analytic function.
-However, in most scientific and engineering calculations, the primary interest is not the full result distribution, but rather just few summary statistics of the result, such as the mean and deviation [2][3].
+However, in most scientific and engineering calculations, the primary interest lies not in the full result distribution but in few summary statistics of the result, such as the mean and deviation [2][3].
 These simplified statistics can be obtained through a statistical Taylor expansion.
-Let $\overline{x}$ and $\delta x$ denote the sample mean and deviation of a random variable.
-Let $\mu$ and $\sigma$ denote the distribution mean and deviation of the same variable.
-An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.5).
-Formula (2.4) defines the bound moment $\zeta(n,\kappa)$, where $0<\varrho,\kappa$ specify the bounding ranges.
-Using Formula (2.4), Formula (2.6) and Formula (2.3) yields the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively.
-The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the result value.
+[…]
 $\zeta(n)\equiv$ $\int_{\mu-\varrho\sigma}^{\mu+\kappa\sigma}\tilde{x}^{n}\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
@@
 $=$ $\sum_{n=1}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.10)
-Formula (2.4) defines a bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ for the underlying distribution, where $\mu$ and $\sigma$ are the distribution mean and distribution deviation, respectively.
-As discussed later in this paper, $\kappa$ statistically determines $\varrho$, while $\kappa$ itself is determined by both the sample size and the underlying distribution of the input.
-Certain $f(x)$ such as $\log(x)$ and $x^{c}$ also require an upper bound on $\kappa$ to ensure the convergence of Formula (2.6) and (2.3).
-The probability $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
-Under the uncorrelated uncertainty assumption, Formula (2.9) and (2.3) compute the mean and variance of the Taylor expansion in Formula (2.8), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ are the variance moments for $x$ and $y$, respectively:
+An confidence interval [5] $\tilde{x}\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ can describe a sampling from an underlying distribution, where $0<\varrho,\kappa$ specify the bounding ranges.
+Formula (2.7) defines the bound moment $\zeta(n)$.
+As discussed later, $\kappa$ determines $\varrho$, while $\kappa$ itself is determined by both the sample size $N$ and the underlying distribution of the input.
+When $N\rightarrow\infty$, $\varrho,\kappa\rightarrow\infty$, so that by definition $\zeta(0)=1$, $\zeta(1)=0$, and $\zeta(2)=1$.
+When $N$ is bounded, $\varrho$ and $\kappa$ are also bounded, so that $\zeta(0)<1$, and $\zeta(2)<1$.
+The probability of $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
+An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.8).
+Using Formula (2.7), Formula (2.9) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively.
+The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
 $f(x+\tilde{x},y+\tilde{y})=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\frac{f^{(m,n)}_{(x,y)}}{m!n!}\tilde{x}^{m}\tilde{y}^{n};$ (2.11)
@@
 $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\big(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\big);$ (2.13)
+Under the uncorrelated uncertainty assumption, Formula (2.12) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.11), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively.
 Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
 2.4 Bounding Symmetry
-A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbound distribution.
+A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbounded distribution.
 To achieve mean preserving, $\kappa$ determines $\varrho$, so that $\zeta(n)$ becomes $\zeta(n,\kappa)$.
-Under a mean-preserving bounding, Formula (2.11) and (2.12) provide the result for $x\pm y$, while Formula (2.13) and (2.14) give the result for $xy$:
+Under a mean preserving bounding, Formula (2.14) and (2.15) provide the result for $x\pm y$, while Formula (2.16) and (2.17) give the result for $xy$:
 $\overline{x\pm y}$ $=\zeta(0,\kappa_{x})x\pm\zeta(0,\kappa_{x})y;$ (2.14)
@@
 $\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\sigma_{y})^{2}+\zeta(2,\kappa_{x})(\sigma_{x})^{2}\;\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.17)
-When $4\leq\kappa$, $\zeta(0,\kappa)\simeq 1$ and $\zeta(2,\kappa)\simeq 1$, making Formula (2.11) and (2.12) the convolution results for $x\pm y$ [5], and Formula (2.13) and (2.14) the corresponding results of the product distribution for $xy$ [5].
+When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.14) and (2.15) the convolution results for $x\pm y$ [5], and Formula (2.16) and (2.17) the corresponding results of the product distribution for $xy$ [5].
 For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
+Both Gaussian and Uniform distributions are symmetric.
 2.5 Bounding Asymptote
@@
 For uniform distribution, the bounding range is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. Formula (2.18) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
-$\eta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.15)
+$\zeta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.18)
 2.5.2 Gaussian Input Uncertainty
@@
 This convergence occurs rapidly [1].
-In a digital computer, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sum of expansion terms [8][9].
+In digital computation, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sums of expansion terms [8][9].
 Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [2][3][5].
-For Gaussian input uncertainty, $\zeta(2n+1)=0$.
-If $\kappa=\infty$, then $\zeta(2n)=(2n-1)!!$, which causes Formula (2.29) for $\log(x\pm\delta x)$ and Formula (2.35) for $(x\pm\delta x)^{c}$ to diverge.
-As discussed, $\kappa$ is determine by the sample count $N$.
-When $\kappa$ has an upper bound, Formula (2.4) reduces to Formula (2.16):
+Formula (2.7) reduces to Formula (2.19) and (2.20):
 $\zeta(2n,\kappa)=$ $\;(2n-1)!!\left(\xi(\frac{\kappa}{\sqrt{2}})-2N(\kappa)\sum_{j=0}^{n-1}\frac{\kappa^{2j+1}}{(2j+1)!!}\right);$ (2.19)
@@
   • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$ according to Formula (2.21).
-When $\kappa=5$, and $n<5$, the relative error satisfies $\left|\eta(2n)/(2n-1)!!-1\right|<10^{-3}$.
+When $\kappa=5$, and $n<5$, the relative error $\left|\zeta(2n)/(2n-1)!!-1\right|$ is less than $10^{-3}$.
   • For large $2n$, Formula (2.20) reduces to Formula (2.22), showing that $\zeta(2n)$ increases more slowly than $\kappa^{2n}$ as $2n$ grows.
@@
 $e^{-\varrho}\varrho^{2}$ $=e^{-\kappa}\kappa^{2};$ (2.24)
-$\zeta(n)$ $=(n+1)!\left(e^{-\varrho}\sum_{j=0}^{n+1}\frac{\varrho^{j}}{j!}-e^{-\kappa}\sum_{j=0}^{n+1}\frac{\kappa^{j}}{j!}\right);$ (2.22)
-$\lim_{n\rightarrow+\infty}\zeta(n)=\frac{\kappa^{n+2}}{n+2}e^{-\kappa};$ (2.23)
-Formula (2.20) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<1<\kappa$.
-Formula (2.21) gives its mean-preserving equation.
-Formula (2.22) expresses its bound moment.
+$\lim_{n\rightarrow+\infty}\zeta(n,\kappa)$ $=\frac{\frac{\kappa}{\sqrt{2}}^{n}}{n+2}e^{-\kappa};$ (2.25)
+Formula (2.23) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<2<\kappa$.
+Formula (2.24) gives its mean preserving equation.
 And Formula (2.25) describes its asymptotic behavior.
@@
 $\frac{\delta^{2}e^{x}}{(e^{x})^{2}}\simeq$ $\;\sigma^{2}+\frac{3}{2}\sigma^{4}+\frac{7}{6}\sigma^{6}+\frac{5}{8}\sigma^{8}+o((\delta x)^{10});$ (2.38)
+$\delta^{2}\log(x)\simeq$ $\;P(x)^{2}+P(x)^{4}\frac{9}{8}+P(x)^{6}\frac{119}{24}+P(x)^{8}\frac{991}{32}+o(P(x)^{10});$ (2.39)
 $\delta^{2}\sin(x)\simeq$ $\;\sigma^{2}\cos(x)^{2}-(\delta x)^{4}(\cos(x)^{2}\frac{3}{2}-\frac{1}{2})$
 $+\sigma^{6}(\cos(x)^{2}\frac{7}{6}-\frac{1}{2})-\sigma^{8}(\cos(x)^{2}\frac{5}{8}-\frac{7}{24})+o((\delta x)^{10});$ (2.40)
-$\delta^{2}\log(x)\simeq$ $\;P(x)^{2}+P(x)^{4}\frac{9}{8}+P(x)^{6}\frac{119}{24}+P(x)^{8}\frac{991}{32}+o(P(x)^{10});$ (2.38)
 $\frac{\delta^{2}x^{c}}{(x^{c})^{2}}\simeq$ $\;c^{2}P(x)^{2}+\frac{3}{2}c^{2}(c-1)(c-\frac{5}{3})P(x)^{4}$
@@
 Formula (2.28) for $e^{x\pm\delta x}$ and Formula (2.34) for $\sin(x\pm\delta x)$ both converge unconditionally.
-However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can take negative values for large $\delta x$, which imposes a constraint upper bound on the input $\delta x$.
-Formula (2.29) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) when $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$ for a Gaussian distribution, or $P(x)<1/\sqrt{3}$ for a uniform distribution.
+However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can become negative for large $\delta x$, which imposes an upper-bound constraint on the input $\delta x$.
+Formula (2.31) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) as $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$.
 $\delta^{2}\log(x\pm\delta x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}$
@@
 Formula (2.37) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.46) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ .
-This expression converges when $P(x)\lesssim 1/\kappa$ although the precisse upper bound for $P(x)$ varies with $c$.
+This expression converges when $P(x)\lesssim 1/\kappa$ although the precis upper bound for $P(x)$ varies with $c$.
 $\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.46)
-2.9 Bounding Range
-Figure 3: The result variance ratio $\delta^{2}f/\widehat{\delta^{2}f}$ (left y-axis) and the bounding leakage (right y-axis) for varying bounding range $\kappa$ (x-axis) for the selected function $f(x)$ (legend), where $\widehat{\delta^{2}f}$ denotes the corresponding variance for $\kappa=5$.
-Figure 4: The ratio of adjusted $\widehat{\delta^{2}f}$ to the original $\widehat{\delta^{2}f}$ (as shown by the y-axis) for varying bounding range $\kappa$ (as shown by the x-axis) for the selected $f(x)$ (as shown by the legend), where $\widehat{\delta^{2}f}$ is approximated by normalizing the input variance with $\zeta(2,\kappa)$ such that $\delta^{2}x=(\delta x)^{2}$.
-The bounding range $\kappa$ influences the results of statistical Taylor expansion unless it is sufficiently large.
-Figure 3 shows that $\epsilon(\kappa)$ decreases as $\kappa$ increases, while $\delta^{2}f$ for the selected functions rises with $\kappa$ until reaching stable value when $\kappa\geq 4$.
-When $\kappa\geq 5$, the results show negligible difference for varying $\kappa$.
-In practice, based on the 5-$\sigma$ rule [3] for reliably distinguishing two random variables, let $\widehat{\delta^{2}f}$ denote the ideal variance obtained once the result variance stabilizes for $\kappa=5$.
-Figure 3 indicates that $\delta^{2}f\leq\widehat{\delta^{2}f}$ due to $\epsilon(\kappa)$, and when $\epsilon(\kappa)\simeq 0$, $\delta^{2}f\simeq\widehat{\delta^{2}f}$.
-The variance ratio is defined as $\delta^{2}f/\widehat{\delta^{2}f}$.
-Figure 3 also demonstrates that variance ratio remains approximately a constant for a $\kappa$ value.
-To approach $\widehat{\delta^{2}f}$ when $\kappa<5$, instead of $(\delta x)^{2}$, the adjusted input uncertainty variance $(\delta x)^{2}/\zeta(2,\kappa)$ is used in Formula (2.3) to obtain the adjusted variance whose $(\delta x)^{2}$-term has the value as if $\zeta(2,\kappa)=1$, which is same as that of the ideal variance $\widehat{\delta^{2}f}$.
-Figure 4 shows that the adjusted variances are close to the corresponding ideal variance, however as a first order approximation, their ratios depend on the specific form of $f(x)$.
-Figure 4 also indicates a universal worst $\kappa\simeq 1.75$, below which the tail effect of Gaussian distributions vanishes.
-$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.45)
-$\zeta(2,\kappa)\equiv$ $\frac{\delta^{2}f}{\widehat{\delta^{2}f}}$; (2.46)
-$x\pm y:$ $\;\;\;\zeta(2,\kappa)=\frac{\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.47)
-Formula (2.45) is a special case for Formula (2.3), which can be reinterpreted as obtaining the bounding range $\kappa$ from a variance ratio.
-This relation between variance ratio and bounding range can be generalized as in Formula (2.46).
-For example, Formula (2.47) shows that the result $\kappa$ for $x\pm y$ is a variance-weighted average of the input $\kappa_{x}$ and $\kappa_{y}$ via $\zeta(2,\kappa)$.
 2.9 Statistical Bounding
 Figure 5: Measured bounding leakage $\epsilon(\kappa_{s},N)$ (y-axis) for varying measuring bounding range $\kappa_{s}$ (x-axis) and sample count NN (legend).
-Figure 6: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range for the normal distribution.
+Figure 6: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range $\kappa_{s}$ for the normal distribution.
 When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [5].
@@
 Because $\epsilon(\kappa)\neq\epsilon(\kappa,N)$ for finite $N$, let $\epsilon(\kappa)=\epsilon(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bonding range, and $\kappa(\kappa_{s},N)$ is the measured bounding range.
+When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to bounding leakage $\epsilon(N)$.
+Figure 6 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically.
+The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.7).
 $\epsilon(\kappa)=$  $\;1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.47)
 $\epsilon(\kappa_{s},N)=$  $\;1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.48)
-[…]
-When the underlying distribution is uniform, $\kappa_{s}=\sqrt{3}$ and $\epsilon=0$.
-Figure 6 shows that $0<\epsilon(N)\sim N^{-0.564}$, resulting in a measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.4).
-In Figure 6, when the underlying distribution is Gaussian, the bounding range $\kappa(5,N)$ start to reach it stable value at sample count $N=10^{3}$.
-In contrast, when the underlying distribution is uniform, such sample count $N$ is $10^{2}$.
-This difference of the sample count upper bounds necessary for computing adjusted variances is caused by the underlying distributions.
+When the underlying distribution is Normal, Formula (2.47) and (2.48) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Gaussian error function [5].
+Figure 5 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$.
+It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$.
+Figure 6 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$).
+Figure 6 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$.
+Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 5 should be used as $\kappa$ in Formula (2.7).
+2.10 Ideal Statistics
+Figure 7: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Gaussian.
+Figure 8: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Uniform.
+$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.49)
+$\alpha(x\pm y)=$ $\frac{\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.50)
+When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}\rightarrow\sigma$ in Formula (2.49).
+Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
+  • When the underlying uncertainty distribution is Gaussian, Figure 7 shows that the resulting variance $\delta^{2}f$ for a selected functions rise with NN until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 20$.
+When $N=20$ and $\kappa_{s}=5$, $\kappa=3.8$ according to Figure 5.
+Therefore it is sufficient to compute the stable variances $\widehat{\delta^{2}}f$ with $\zeta(n,5)$.
+  • When the underlying uncertainty distribution is Uniform, Figure 8 shows the same trend, however the stable variances are reached only when $N>10^{5}$.
+The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.18).
+Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input.
+Ideal variances are consistent with traditional variances [5].
+In both Figure 7 and 8, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
+The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$.
+All functions approach their corresponding ideal variances with exactly the same trend.
+These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$.
+Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution.
+When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.50) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
+An ideal ratio also applies to the corresponding uncertainty bias.
+The trends of uncertainty biases $\overline{f}-f$ versus sample count NN are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$.
+This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
 2.11 Dependency Tracing
+$\delta^{2}(f+g)=$ $\;\delta^{2}f+\delta^{2}g+2(\overline{fg}-\overline{f}\overline{g});$ (2.51)
+$\delta^{2}(fg)=$ $\;\overline{f^{2}g^{2}}-\overline{fg}^{2};$ (2.52)
+$\delta^{2}f(g)=$ $\;\overline{f(g)^{2}}-\overline{f(g)}^{2};$ (2.53)
+$\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.54)
+$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;\sigma^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+\sigma^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.55)
+$+2\sigma^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
 When all inputs satisfy the uncorrelated uncertainty assumption, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
@@
   • Formula (2.54) gives the variance of the linear transformation of a function, which can be applied to Formula (2.51) and (2.52) for more general dependency tracing.
-$\delta^{2}(f+g)=$ $\;\delta^{2}f+\delta^{2}g+2(\overline{fg}-\overline{f}\overline{g});$ (2.50)
-$\delta^{2}(fg)=$ $\;\overline{f^{2}g^{2}}-\overline{fg}^{2};$ (2.51)
-$\delta^{2}f(g)=$ $\;\overline{f(g)^{2}}-\overline{f(g)}^{2};$ (2.52)
-$\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.53)
-$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;\sigma^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+\sigma^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.54)
-$+2\sigma^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
-Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance fit statistics rigorously.
-However, dependency tracing comes with a price: variance calculations are generally more complex than value calculations, exhibiting narrower convergence range for input variables.
-Dependency tracing also implies that the results of statistical Taylor expansion should remain path independent.
+Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously.
+However, dependency tracing comes at a cost: variance calculations are generally more complex than value calculations and exhibits a narrower convergence range for input variables.
+Dependency tracing also implies that the results of statistical Taylor expansion must remain path independent.
 2.12 Traditional Execution and Dependency Problem
 Dependency tracing requires an analytic form of the function to apply statistical Taylor expansion for the result mean and variance, as in Formula (2.9), (2.3), (2.12), and (2.3).
-This requirement often conflicts with conventional numerical methods for on analytic functions:
-  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships between original input variables.
-  • Similarly, conditional executions are commonly employed to optimize performance and minimize rounding errors, such as using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [37].
-For dependency tracing, such conditional executions should be replaced by direct matrix inversion as described in Section 6.
-  • Furthermore, traditional approaches frequently apply approximations to the result values during executions.
-Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than the value in Taylor expansion.
+This requirement often conflicts with conventional numerical methods for analytic functions:
+  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships among the original input variables.
+  • Similarly, conditional executions are often employed to optimize performance and minimize rounding errors, for example, using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [37].
+For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 7.
+  • Furthermore, traditional approaches frequently apply approximations to result values during execution.
+Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than value in statistical Taylor expansion.
 Consequently, approximation strategies should prioritize variances than values.
-Section 6 illustrates this with a first-order approximation for computing a matrix determinant.
-  • Traditionally, results from mathematical library functions are accepted without question, with accuracy assumed down to the last bit.
-As demonstrated in Section 8, statistical Taylor expansion enables the detection of numerical errors within these functions and necessitates recalculating them with uncertainty explicitly included in the output.
-  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, and division  [^5] Sometimes square root is also regarded as an arithmetic operation..
+Section 7 illustrates this principle through a first-order approximation used in computing a matrix determinant.
+  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last bit.
+As demonstrated in Section 9, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty explicitly incorporated into the output.
+  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls.
 However, this decomposition introduces dependency problems in floating-point arithmetic, interval arithmetic, and statistical Taylor expansion.
@@
 The path dependence of $f(y)|_{y=g(x)}$ and $\delta^{2}f(y)|_{y=g(x)}$ are evident in cases such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
-Dependency tracing thus removes nearly all flexibility in traditional numerical executions, eliminating associated dependency problems.
-Consequently, all conventional numerical algorithms require reevaluation or redesign to align with the principles of statistical Taylor expansion.
+Dependency tracing therefore removes nearly all flexibility from traditional numerical executions, effectively eliminating the associated dependency problems.
+Consequently, all conventional numerical algorithms must be reevaluated or redesigned to align with the principles of statistical Taylor expansion.
 3 Variance Arithmetic
-This paper confines the calculating of variances to ideal variances $\widehat{\delta^{2}f}$ when $\kappa=5$, and assumes that that the input distribution is Gaussian, although variance arithmetic also implements statistical Taylor expansion for other cases.
-Furthermore, for discussion simplicity, Taylor coefficients in Formula (2.5) and (2.8) are all assumed to be precise.
+Variance arithmetic implements statistical Taylor expansion.
+Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms.
+Consequently, the following numerical rules are introduced:
+  • finite: The resulting value and variance must remain finite.
+  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion exhibiting an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$.
+Unless all the remaining terms in the expansion are known to be precisely zeros, each expansion is executed to the full 448 terms for the monotonicity check.
+  • positive: At every order, the expansion variance must be positive.
+  • stable: To avoid truncation error [13], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value.
+This rule ensures sufficiently fast convergence in the context of monotonic convergence.
+  • reliable: At every order, the uncertainty of the variance must be less than $1/5$ times of the value of the variance.
+For simplicity of discussion, This paper confines the calculation of variances to ideal variances $\widehat{\delta^{2}}f$ and assumes that the input distribution is Gaussian with $\kappa_{s}=5$.
+Furthermore,the Taylor coefficients in Formula (2.8) and (2.11) are assumed to be precise.
 3.1 Numerical Representation
-Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. Other conventional numerical numbers need to convert to this format.
-If the least $20$ significant bits of its significand are all 0 in a standard floating-point number, the value is considered precise, representing a 2’s fractional with a possibility no less than $1-2^{-20}=1-2.384\;10^{-7}$.
-Otherwise, a standard floating-point value is considered imprecise with the uncertainty given as $1/\sqrt{3}$-fold of the ULP of the value, where ULP stands for Unit in the Last Place for a conventional floating-point number [9].
-This follows from the fact that the pure rounding error of round-to-nearest is uniformly distributed within half bit of the significand of the floating-point value [1].
+Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. All other conventional numerical numbers must be converted to this format.
+If the least $20$ significant bits of the significand in a standard floating-point number are all 0, the value is considered precise, representing a 2’s fractional with a probability no less than $1-2^{-20}=1-2.384\;10^{-7}$.
+Otherwise, a standard floating-point value is considered imprecise with its uncertainty defined as $1/\sqrt{3}$ times of the ULP of the value, where ULP refers to the Unit in the Last Place in conventional floating-point representation [9].
+This follows from the fact that the pure rounding error in round-to-nearest mode is uniformly distributed within half bit of the significand of a floating-point value [1].
 If an integer number is within the range $2^{53}+1,+2^{53}-1]$ of the significand of a 64-bit standard floating-point number, its uncertainty is zero.
-Otherwise, it is first converted to a conventional floating-point value before being converted to an imprecise value.
-3.2 Calculation Rules
-Variance arithmetic implements statistical Taylor expansion with several practical assumptions.
-Due to finite precision and range of conventional floating-point representation, $\zeta(n)$ can be computed only for $n<448$ before becoming infinite under the IEEE 64-bit floating-point format, Consequently, the following numerical rules are introduced:
-  • finite: The result must remain finite.
-  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion to an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$.
-Unless all the remaining terms in the expansion are known to be precise zeros, an expansion is always executed to the full 448 terms for the monotonic check.
-  • positive: At any order, the expansion variance must be positive.
-  • stable: For sufficiently fast convergence, the value of the last expansion term must be less than the $\epsilon(5)=5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value.
-  • reliable: At any order, the value of the variance must be less than $1/\kappa$-fold of the uncertainty of the variance.
-The validity of these numerical rules of variance arithmetic will be demonstrated in subsequent sections.
-3.2.1 Finite
+Otherwise, it is first converted to a conventional floating-point value before being transformed into an imprecise value.
+Variance arithmetic uses floating-point arithmetic for computation.
+3.2 Finite
 For $(1\pm\delta x)^{-2}$, the Taylor coefficient increases with the expansion order $n$ as $(-1)^{n}(n+1)$. When $\delta x=0.5$, this growth causes the result variance to diverge to infinity.
-3.2.2 Monotonic
+3.3 Monotonic
 Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty.
@@
 Beyond this upper bound, the expansion is no longer monotonic.
-Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.29) mathematically, with $\zeta(2n)$ as the connection of these two perspectives.
+Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.31) mathematically, with $\zeta(2n)$ providing the connection between these two perspectives.
 For $e^{x\pm\delta x}$ the convergence holds for $\delta x\lesssim 19.864$ regardless of $x$, while the result $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$.
@@
 From Formula (2.46), and except when $c$ is a natural number, Formula (2.37) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$.
-This trend is approximately confirmed in Figure 7, and beyond the upper bound $P(x)$, the expansion is no longer monotonic.
+This trend is approximately confirmed in Figure 9, and the expansion is no longer monotonic beyond the upper bound $P(x)$.
 Qualitatively, $\delta^{2}1/x$ converges more slowly than $\delta^{2}\sqrt{x}$, consistent with Formula (2.44) and (2.43).
-When the input uncertainty is Uniformly distributed, Figure 8 illustrates that the input uncertainty upper bound is close to $1/\sqrt{3}$ according to Formula (2.44).
-The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such the expansion in Formula (2.44) contains $652$ terms instead of $448$ terms.
-The upper bounds of input uncertainty in Figure 7 and 8 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 7 are $3$ order-of-magnitude smaller than those in Figure 8 when the exponent is less than 0.
-This suggests that convergence is not the only consideration in statistical Taylor expansion because uncertainty bias or uncertainty can also invalidate a result in practice.
-3.2.3 Positive
+When the input uncertainty is Uniformly distributed, Figure 10 illustrates that the input uncertainty upper bound is close to $1/\kappa=/\sqrt{3}$ according to Formula (2.46).
+The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms.
+The upper bounds of input uncertainty in Figure 9 and 10 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 9 are $3$ order-of-magnitude smaller than those in Figure 10 when the exponent $c$ is less than 0.
+Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
+3.4 Positive
 Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty.
@@
 The x-axis is expressed in units of $\pi$.
-In some cases, the variance expansion may produce negative results, as in Formula (2.32) for $sin(x\pm\delta x)$.
-Figure 9 demonstrates that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$, which is less than $\pi$ of $\sin(x)$.
-Beyond this upper bound, the expansion is no longer positive in Figure 9.
+In some cases, the variance expansion may yield negative results, as in Formula (2.34) for $sin(x\pm\delta x)$.
+Figure 11 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$.
+Beyond this upper bound, the expansion is no longer positive.
 In Figure 12, when the bounded momentum $\zeta(2n)$ for Uniform distribution is used instead of that for Normal distribution, the input uncertainty upper bound is $4$ times larger than that in Figure 11 but is still periodic.
 The resulting uncertainty increases not much and still remains less than $1$.
-3.2.4 Stable
+3.5 Stable
 The unstable condition happens rarely.
-3.2.5 Reliable
+3.6 Reliable
 The condition of not being reliable seldom happens.
-3.3 Continuity
-Figure 11: Histograms of normalized errors for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
-In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space.
-For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number.
-The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remains continuous around $c=n$.
-Table 1 shows that the result of $(1\pm\delta x)^{n\pm d},0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former had upper bound for $\delta x$, while the latter does not.
-A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity.
-The presence of such poles does not disrupt continuity in of the result mean, variance, or histogram.
-Figure 12 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
-  • When the second derivative is zero, the resulting distribution is symmetric two-sided and Delta-like, such as when $n=3,x=0$.
-  • When the second derivative is positive, the resulting distribution is right-sided Delta-like, such as the distribution when $n=2,x=0$, or when $n=2,x=\pm 0.2$, or when $n=3,x=0.2$.
-  • When the second derivative is negative, the resulted distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
-  • In each case, the transition from $x=0$ to $x=0.2$ is continuous.
-A statistical bounding range inf variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative.
-An illustrative example is $\sin(x)$ in Figure 9.
-A statistical bounding range in variance arithmetic cannot include any distributional zero because the result will diverge, such as at $x=0$ for both $(x\pm\delta x)^{c},c<1$ and $\log(x\pm\delta x)$.
-$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
-Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
-Value $1\mp 2.155\;10^{-8}$ $1\pm 2.073\;10^{-8}$ $041\pm 6.065\;10^{-8} 1.1$$122\pm 1.328\;10^{-7}$
-Uncertainty $0+2.127\;10^{-7}$ $0.201-1.358\;10^{-6}$ $0.407\pm 2.201\;10^{-7}$ $0.654\pm 2.784\;10^{-7}$
-Table 1: The result value and uncertainty of $(1\pm\delta x)^{n\pm d}$ vs $(1\pm\delta x)^{n}$, in which $n$ is a natural number, $0<d\ll 1$, and $\delta x$ is the upper bound for $(1\pm\delta x)^{n\pm d}$.
-The value and the uncertainty are expressed as the difference with the corresponding value and uncertainty of those of $(1\pm\delta x)^{n}$.
-3.4 Tracking Rounding Error
-Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x axis).
-The y-axis to the left shows both the value and the uncertainty of the residual errors.
-The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
-$\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}=\sum_{j=0}^{N}\tilde{x}^{j}\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (3.1)
-Variance arithmetic can track rounding errors effectively without the need for additional rules.
-Figure 12 illustrates the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (3.1), $\frac{1}{1-x}$ is computed using Formula (2.33), and $x$ is initiated as a floating-point value.
-Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (3.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of $\frac{1}{1-x}$.
-Figure 12 shows:
-  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, meaning that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$.
-A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$.
-The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
-  • When $x\not\in[-0.74,+0.75]$, for example when x=0.75,0.76x=-0.75,0.76, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$.
-The residual error magnitude increases as $|x|\rightarrow 1$, reaching about $50$ when $x=0.98$.
 3.7 Comparison
-Two imprecise values can be compared statistically for their difference.
-When the value difference is zero, the two imprecise values are equal.
-In statistics, such two imprecise values have $50\%$ possibility to be less than or greater to each other but no chance to be equal to each other [5].
-In variance arithmetic, however, they are treated as neither less nor greater than each other, thus they are equal.
-Otherwise, the standard z-statistic method [5] is applied to determine whether they are equal, less than, or more than each other.
-For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$, and the probability for them to be not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative density function for Normal distribution [5].
-If the threshold probability for inequality is $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$.
+Two imprecise values can be compared statistically based on their difference.
+When the value difference is zero, the two imprecise values are considered equal.
+In statistics, such two values have a $50\%$ possibility of being either less than or greater to each other but zero probability of being exactly equal [5].
+In variance arithmetic, however, they are treated as neither less nor greater than each other and therefore are considered equal.
+Otherwise, the standard z-statistic method [5] is applied to determine whether two imprecise values are statistically equal, less than, or greater than each other.
+For example, the difference between $1.002\pm 0.001$ and $1.000\pm 0.002$ is $0.002\pm 0.00224$, yielding $z=0.002/0.00224$.
+The probability that they are not equal is $\xi(|z|/\sqrt{2})=62.8\%$, in which $\xi(z)$ is the cumulative distribution function for Normal distribution [5].
+If the threshold probability for inequality is set at $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$.
 Alternatively, an equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for an equal probability threshold of $50\%$.
-4 Verification of Statistical Taylor Expansion
+Because the result of comparison depends on threshold probability which is application specific, comparison is not part of variance arithmetic.
+4 Verification of Variance Arithmetic
 4.1 Verification Methods and Standards
-Statistical Taylor expansion is verified through the validation of variance arithmetic.
-Analytic functions or algorithms with precisely known results are used to compare the outputs of variance arithmetic using the following statistical properties:
+Analytic functions or algorithms with precisely known results are used to evaluate the outputs of variance arithmetic based on the following statistical properties:
   • Value error: the difference between the numerical result and the corresponding known precise analytic result.
@@
   • Calculation response: the relationship between the amount of calculation and output uncertainties.
-One objective of uncertainty-based calculation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
+One objective of uncertainty-based computation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
   1. 1.
@@
     The error distribution should follow Normal distribution when an imprecise value is expected, or Delta distribution when a precise value is expected.
-  3. 3.
+If the precise result is unknown, the resulting normalized error distribution can be used to assess whether ideal coverage is achieved.
+However, if the input uncertainty is known only to order of magnitude, proper coverage is achieved when the error deviations fall within the range $[0.1,10]$.
+When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained.
+The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainties.
+Achieving ideal coverage serves as a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly within the given context.
+The input noise range that yields ideal coverage defines the ideal application range for input uncertainties.
+Besides coverage, other validation criteria are also used:
   • The uncertainty response should match the expected functional form.
-For a linear function, the output uncertainties should increase linearly with input uncertainties, with the ratio of the output to input uncertainty means defined as uncertainty response ratio.
-  4. 4.
-    The calculation response should follow the expected trend; for instance, increasing the amount of calculations should result in larger uncertainties.
-If the precise result is unknown, as outlined in Section 6, the result error distribution can be used to determine whether ideal coverage is achieved.
-If however, the input uncertainty is only accurate to the order of magnitude, the proper coverage is obtained when the error deviations fall within the range of $[0.1,10]$.
-When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained.
-The minimal noise deviation required provides a good estimation of the magnitude of unspecified input uncertainties.
-The presence of ideal coverage is a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly in the given context.
-The input noise range yielding ideal coverage defines the ideal application range for input uncertainties.
-Gaussian noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
+For a linear function, the output uncertainties should increase linearly with the input uncertainties, with the ratio of output to input uncertainty means defined as uncertainty response ratio.
+  • The calculation response should follow the expected trend; for instance, increasing the number of calculations should result in larger output uncertainties.
+Noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
 4.2 Types of Uncertainties
 There are five primary sources of result uncertainty in a calculation [2][13][1]:
-  • Input uncertainties: The examples in this paper demonstrate that when input uncertainty precision is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
-  • Rounding errors: Empirical results indicate that variance arithmetic can provide proper coverage for rounding errors.
+  • Input uncertainties: The examples presented in this paper demonstrate that when the precision of input uncertainties is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
+  • Rounding errors: Empirical results indicate that variance arithmetic provides proper coverage for rounding errors.
   • Truncation errors: Variance arithmetic avoids truncation errors with its stable rule.
-  • External errors: External errors are the value errors which are not specified in input uncertainties, such as numerical errors in library functions.
-Section 8 examines the effects of numerical errors in $\sin$ and $\tan$, showing that when these external errors are sufficiently large enough, neither ideal coverage nor proper coverage can be achieved.
-This finding indicates that library functions must be recalculated to include the corresponding uncertainty of each value.
-  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to obtain a solution.
-For example, the predecessor of statistical Taylor expansion [1] demonstrated that the discrete Fourier transformation is only an approximation for the mathematically defined continuous Fourier transformation, containing modeling errors.
-Conceptually, modeling errors originate in mathematics and are therefore outside the domain of statistical Taylor expansion.
+However, using Formula (5.1) for polynomial expansion can result in truncation errors when the choice of the expansion order is insufficient, as illustrated in Figure 14 .
+  • External errors: External errors are value errors not specified in the input uncertainties, such as numerical errors in library functions.
+Section 9 examines the effects of numerical errors of library sine function, showing that when these external errors are sufficiently large, neither ideal coverage nor proper coverage can be achieved.
+This finding indicates that library functions must be recalculated to explicitly include the corresponding uncertainty for each computed value.
+  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to make a solution tractable.
+For example, Section 9 demonstrates that the discrete Fourier transform (DFT) is only an approximation of the mathematically defined continuous Fourier transform (FT), and therefore contains modeling errors.
+Conceptually, modeling errors originate in mathematics within and are thus outside the domain of statistical Taylor expansion.
 4.3 Types of Calculations to Verify
-Algorithms of distinct nature each representative of its category are required to test the broad applicability of variance arithmetic [1].
+Algorithms of distinct natures, with each representative of its respective category, are required to test the broad applicability of variance arithmetic [1].
 An algorithm can be categorized by comparing the amount of its input and output data as [1]:
@@
 An application algorithm computes numerical values from an analytic formula. Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems.
-A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information contained remaining largely unchanged.
-For reversible transformations, a unique requirement is to recover the original input uncertainties after a round-trip transformation: performing a forward transformation followed by its reverse transformation.
-DFT (discrete Fourier transformation) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back to the input using essentially the same process.
-A test of variance arithmetic using FFT (fast Fourier transformation, which is an implementation of DFT) algorithms is presented in Section 8.
+A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged.
+For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a round-trip transformation that is performing a forward transformation followed by its reverse transformation.
+The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process.
+A test of variance arithmetic using the fast Fourier transform (FFT, which is an implementation of DFT) algorithms is presented in Section 9.
 A generation algorithm produces substantially more output data than input data.
 Such algorithms encode mathematical knowledge into data.
-Some generation algorithms are theoretical calculations which involve no imprecise input, so all resulting uncertainty arises from rounding errors.
-Section 9 presents a generation algorithm, which generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
-A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or statistical characterization of a data set.
+Certain generation algorithms are purely theoretical calculations that involve no imprecise input, so all resulting uncertainty arises solely from rounding errors.
+Section 10 presents a generation algorithm that generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
+A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or in the statistical characterization of a data set.
 In this process, certain information is lost while other information is extracted.
-Conventional wisdom is that a reduction algorithm generally benefits from a larger input data set [5].
-A test of statistical characterization through sampling is described in Section 5.
+As a statistical approach, variance arithmetic inherently carries out statistical reduction.
+For example, the result of averaging $N$ imprecise values with each uncertainty close to $\delta x$ is $\overline{x}\pm\frac{\delta x}{\sqrt{N}}$, reflecting central limit theorem [5] naturally.
+5 Polynomial
+Formula (5.1) presents polynomial Taylor expansion:
+$\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}$ $=\sum_{j=0}^{N}\tilde{x}^{j}P_{j},\;\;\;P_{j}\equiv\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (5.1)
+5.1 Tracking Rounding Error
+Figure 13: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x-axis).
+The y-axis to the left shows both the value and the uncertainty of the residual errors.
+The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
+Variance arithmetic can track rounding errors effectively without the need for additional rules.
+Figure 14 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (5.1), $\frac{1}{1-x}$ is computed using Formula (2.35), and $x$ is initiated as a floating-point value.
+Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (5.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of the statistical Taylor expansion of $\frac{1}{1-x}$.
+Figure 14 shows:
+  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, which indicates that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$.
+A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$.
+The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
+  • When $x\not\in[-0.74,+0.75]$, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$.
+The residual error magnitude increases as $|x|\rightarrow 1$, reaching approximately $50$ when $x=0.98$.
+5.2 Continuity
+Figure 14: Histograms of normalized errors for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
+$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
+Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
+Value $1\mp 2.155\;10^{-8}$ $1\pm 2.073\;10^{-8}$ $041\pm 6.065\;10^{-8} 1.1$$122\pm 1.328\;10^{-7}$
+Uncertainty $0+2.127\;10^{-7}$ $0.201-1.358\;10^{-6}$ $0.407\pm 2.201\;10^{-7}$ $0.654\pm 2.784\;10^{-7}$
+Table 1: The result value and uncertainty of $(1\pm\delta x)^{n\pm d}$ vs $(1\pm\delta x)^{n}$, in which $n$ is a natural number, $0<d\ll 1$, and $\delta x$ is the upper bound for $(1\pm\delta x)^{n\pm d}$.
+The value and the uncertainty are expressed as the difference with the corresponding value and uncertainty of those of $(1\pm\delta x)^{n}$.
+In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space.
+For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number.
+The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$.
+Table 2 shows that the result of $(1\pm\delta x)^{n\pm d}$ where $0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former has an upper bound for $\delta x$, while the latter does not.
+A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity.
+The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram.
+Figure 14 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
+  • When the second derivative is zero, the resulting distribution is symmetric two-sided and Delta-like, such as when $n=3,x=0$.
+  • When the second derivative is positive, the resulting distribution is right-sided Delta-like, such as the distribution when $n=2,x=0$, or when $n=2,x=\pm 0.2$, or when $n=3,x=0.2$.
+  • When the second derivative is negative, the resulted distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
+In each case, the transition from $x=0$ to $x=0.2$ is continuous.
+A statistical bounding range in variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative.
+An illustrative example is $\sin(x)$ shown in Figure 11.
+A statistical bounding range in variance arithmetic cannot include any distributional zero because the result will diverge, such as at $x=0$ for both $(x\pm\delta x)^{c},c<1$ and $\log(x\pm\delta x)$.
 6 Mathematical Library Functions
@@
 Formula (2.28), (2.31), (2.34), and (2.37) are evaluated using the corresponding mathematical library functions exp, log, sin, and pow, respectively.
-At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as follows:
+At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as:
   1. 1.
@@
   2. 2.
-    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between using $\tilde{x}$ and using $x$ as the input.
+    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between the outputs obtained using $\tilde{x}$ and $x$ as inputs.
   3. 3.
@@
 6.1 Exponential
-Figure 13 demonstrates that the calculated uncertainties using Formula (2.26) align closely with the measured value deviations for $e^{x+\delta x}$.
+Figure 15 shows that the calculated uncertainties obtained using Formula (2.28) align closely with the measured value deviations for $e^{x+\delta x}$.
 Consequently, all error deviations remain very close to $1$, even though both the uncertainty and the value deviations increase exponentially with $x$ and $\delta x$.
@@
 Because $\log(x)$ has a distributional zero at $x=0$ all $log(x\pm\delta x)$ values are rejected when $P(x)>1/5$.
-Figure 14 demonstrates that the uncertainties calculated using Formula (2.29) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
+Figure 16 shows that the uncertainties calculated using Formula (2.31) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
 6.3 Power
-Figure 15 demonstrates that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.35) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11.
-Figure 16 reveals that the error deviations of $\sin(x+\delta x)$ align close to 11 for all the input exponent $c$ and the input uncertainty $\delta x>10^{15}$.
+Figure 17 shows that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.37) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11.
+Figure 18 reveals that the error deviations of $\sin(x+\delta x)$ remain close to 11 for all input exponents $c$ and input uncertainties $\delta x>10^{15}$.
 6.4 Sine
-Figure 17 demonstrate that the uncertainties calculated using Formula (2.32) correspond closely to the measured value deviations for $\sin(x+\delta x)$.
-It also reveals that $\delta^{2}\sin(x)$ has the same periodicity as $\sin(x)$:
+Figure 19 shows that the uncertainties calculated using Formula (2.34) correspond closely to the measured value deviations for $\sin(x+\delta x)$.
+It also reveals that $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$:
   • When $x=0$, $\sin(x)\simeq x$, so that $\delta^{2}\sin(x)\simeq(\delta x)^{2}$.
@@
 The combined numerical error of the library functions $e^{x}$ and $\log(x)$ is evaluated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$.
-Using either variance arithmetic or conventional floating-point library functions produce identical value errors.
-Figure 19 demonstrates the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$.
-Figure 19 also shows that $e^{\log(x)}$ has much less error than $\log(e^{x})$.
+Using either variance arithmetic or conventional floating-point library functions produces identical value errors.
+Figure 21 shows the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$.
+Figure 21 also shows that $e^{\log(x)}$ exhibits much smaller error than $\log(e^{x})$.
 For $\log(e^{x})-x$, the error deviation is $0.409$ when $|x|\leq 1$ or approaches zero otherwise.
-The reason for surprising small value errors for $e^{\log(x)}-x$ is not clear at the moment.
+The reason for the unexpectedly small value errors for $e^{\log(x)}-x$ is not yet clear.
 The numerical error of the library function $x^{p}$ is computed as $(x^{p})^{1/p}-x$.
@@
 6.6 Summary
-Formula (2.26), (2.29), (2.35), and (2.32) provide effective estimates for their respective uncertainties for library functions.
-With added noise larger than $10^{-15}$, ideal coverage is achieved unless near a distributional pole where the error deviation approaches 0.
-In other cases, proper coverage is attainable.
+Formula (2.28), (2.31), (2.37), and (2.34) provide effective estimates of the corresponding library functions.
+When added noise exceeds $10^{-15}$ precision, ideal coverage is achieved except near a distributional pole where the error deviation approaches zero.
+In all other cases, proper coverage is attainable.
 7 Matrix Calculations
@@
 The difference between $\widetilde{\mathbf{M}}^{A}$ and $\mathbf{M}^{A}$ defines the Adjugate Test.
-Figure 21 shows that the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$.
+As seen in Figure 23, the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$.
 Figure 24 shows that the ideal coverage is achieved for the input precision except at matrix size $8$ and input precision $10^{-17}$.
@@
 Figure 28: Histograms of normalized errors of the adjugate matrix as a function of matrix size with $10^{-11}$ input noise (legend).
-The significand of the conventional floating-point representation [9] has $53$-bit resolution.
-Figure 25 shows that the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$.
+The significance of the conventional floating-point representation [9] has a $53$-bit resolution.
+As shown in Figure 27, the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$.
 When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs, which results in non-Delta distribution in Figure 27.
@@
 7.5 Matrix Inversion
-Figure 28: The linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis).
+Figure 30: Linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis).
 The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
@@
 $\beta$ $=\frac{\sum_{j}X_{j}Y_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}Y_{j}}{\sum_{j}X_{j}X_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}X_{j}};$ (8.2)
-In many applications, data set $Y_{j}$} represents an input data stream where $j$ represents the time index or sequence index.
-$Y_{j}$} is referred to as a time-series input, where $j$ corresponds to $X_{j}$.
+In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index.
+$Y_{j}$} is thus referred to as a time-series input, with $j$ corresponding to $X_{j}$.
 A moving window algorithm [13] is applied within a small window centered on each $j$.
-For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying window’s half width.
-This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (7.1) and (7.2) into Formula (7.3) and (7.4), respectively [1]:
+For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window.
+This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (8.1) and (8.2) into Formula (8.3) and (8.4), respectively [1].
 $\alpha_{j}$ $=\alpha\;2H=\sum_{X=-H+1}^{H}Y_{j-H+X};$ (8.3)
 $\beta_{j}$ $=\beta\;\frac{H(H+1)(2H+1)}{3}=\sum_{X=-H}^{H}XY_{j-H+X};$ (8.4)
-The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (7.3) and (7.4) to be reformulated into the progressive moving-window calculation given by Formula (7.5) and (7.6), respectively [1]
+According to Figure LABEL:fig:_Moving_Linear_Fit_Coeff, in which $H$ takes an example value of 2, the values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [1].
 $\beta_{j}$ $=\beta_{j-1}-\alpha_{j-1}+H\left(Y_{j-2H-1}+Y_{j}\right);$ (8.5)
@@
 8.2 Variance Adjustment
-When the time series contains uncertainty, directly applying Formula (7.5) and (7.6) will lead to a loss of precision, as both formulas reuse each input multiple times, accumulating the variance of that input at every usage.
-To avoid this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (7.6) and (7.5), respectively, while the variance should instead be computed using Formula (7.7) and (7.8), respectively.
-Notably, Formula (7.8) is no longer progressive.
-$\delta^{2}\alpha_{j}$ $=\sum_{X=-H+1}^{H}(\delta Y_{j-H+X})^{2}=\delta^{2}\alpha_{j-1}+(\delta Y_{j})^{2}-(\delta Y_{j-2H})^{2};$ (7.7)
-$\delta^{2}\beta_{j}$ $=\sum_{X=-H+1}^{H}X^{2}(\delta Y_{j-H+X})^{2};$ (7.8)
+$\delta^{2}\alpha_{j}$ $=\sum_{X=-H+1}^{H}(\delta Y_{j-H+X})^{2}=\delta^{2}\alpha_{j-1}-(\delta Y_{j-2H})^{2}+(\delta Y_{j})^{2};$ (8.7)
+$\delta^{2}\beta_{j}$ $=\sum_{X=-H}^{H}X^{2}(\delta Y_{j-H+X})^{2};$ (8.8)
+When the time series contains uncertainty, directly applying Formula (8.5) and (8.6) results in a loss of precision since both formulas reuse each input multiple times, thereby accumulating the variance of that input with every reuse.
+To prevent this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (8.6) and (8.5), respectively, while the variances should instead be computed using Formula (8.7) and (8.8), respectively.
+Formula (8.8) is not progressive because the progressive form of $\delta^{2}\beta_{j}$ is more expensive in computation than Formula (8.8).
 Figure 31 shows that the input signal $Y_{j}$ consists of the following components:
@@
 For each increment of $j$, the increasing and the decreasing rates are +1+1 and 1-1, respectively.
-The specified input uncertainty is consistently $0.2$.
-Normal noises with a deviation of $0.2$ are added to the slopes, except for the segment $j=10\dots 19$ where Normal noises with a deviation of $2$ are introduced, representing a $10$-fold actual value over the specified uncertainty.
-Figure 29 also presents the result of moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$.
-The fitted values of $\alpha$ and $\beta$ follow the expected behave with a characteristic delay of $H$ in $j$.
-When (7.3) and (7.4) are applied to compute the uncertainties for $\alpha$ and $\beta$, they both increase exponentially with the time index $j$.
+The specified input uncertainty is fixed at $0.2$.
+Normal noise with a deviation of $0.2$ is added to the slopes, except for the segment $j=10\dots 19$ where Normal noise with a deviation of $2$ is introduced, representing actual uncertainty $10$ times larger than the specified uncertainty.
+Figure 31 also presents the results of the moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$.
+The fitted values of $\alpha$ and $\beta$ follow the expected behave, exhibiting a characteristic delay of $H$ in $j$.
+When (8.3) and (8.4) are applied to compute the uncertainties of $\alpha$ and $\beta$, both uncertainties increase exponentially with the time index $j$.
 In contrast, when Formula (8.3) and (8.4) are used exclusively for value calculation, while Formula (8.7) and (8.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$.
@@
 8.3 Unspecified Input Error
-To determine the error deviation of $\alpha$ and $\beta$, the fitting procedure is applied to multiple time-series data sets, each generated with independent noise realizations.
-Figure 30 illustrates the resulting error deviation as a function of the time index $j$, which remains close to $1$ except in the range $j=10\dots 19$ where the actual noise is $10$-times greater than the specified value.
-This observation indicates that a larger than $1$ error deviation may signal the presence of unspecified additional input errors beyond rounding errors, such as the numerical errors in mathematical library functions.
+To determine the error deviations of $\alpha$ and $\beta$, the fitting procedure is applied to multiple time-series data sets, each generated with independent noise realizations.
+Figure 32 illustrates the resulting error deviation as a function of the time index $j$, which remains close to $1$ except within the range $j=10\dots 19$ where the actual noise is ten times greater than the specified value.
+This observation suggests that an error deviation exceeding 1 may indicate the presence of unspecified additional input errors beyond rounding errors, such as numerical errors in mathematical library functions.
 9 FFT (Fast Fourier Transformation)
 9.1 DFT (Discrete Fourier Transformation)
-For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the DFT (discrete Fourier transformation) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is given by Formula (8.1) and (8.2) [13], respectively.
+For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (9.1) and (9.2), respectively [13].
 In these expressions, $k$ denotes the time index while $n$ represents the frequency index.
-The frequency index and time index are not necessarily related to time unit and frequency unit, respectively.
-The naming is just a convenient way to distinguish the two opposite domains in DFT: the waveform domain $h[k]$ versus the frequency domain $H[n]$.
+The frequency index and time index are not necessarily associated with physical time unit or frequency unit, respectively; rather, the naming convention provides a convenient way to distinguish between the two opposite domains in DFT: the waveform domain $h[k]$ and the frequency domain $H[n]$.
 $H[n]=\sum_{k=0}^{N-1}h[k]\;e^{\frac{i2\pi}{N}kn};$ (9.1)
 $h[k]=\frac{1}{N}\sum_{n=0}^{N-1}H[n]\;e^{-\frac{i2\pi}{N}nk};$ (9.2)
+9.2 Modeling Errors of DFT
+Figure 33: The DFT spectrum $H[n]$ of signal $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, as intensity (y-axis) and phase (embedded y-axis) versus frequency index $n\in[0,18]$ (x-axis and embedded x-axis) for different signal frequency $f$ (legend).
+This result agrees with both theoretical formula [1] and numerical computation from any mathematical libraries such as SciPy.
+Although mathematically self-consistent, DFT implies a periodic boundary condition in the time domain [1].
+Consequently, it is only an approximation for the mathematically defined continuous Fourier transform (FT) [1].
+For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [13].
+Figure 33 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency.
+If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise.
+Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [13].
+To avoid the modeling errors inherent in DFT, only Formula (9.1) and (9.2) are used in this paper.
 9.3 FFT (Fast Fourier Transformation)
 When $N=2^{L}$, where $L$ is a natural number, the generalized Danielson-Lanczos lemma [13] can be applied to DFT to produce FFT [13].
-  • For each output, each input is used only once, so there is no dependency problem when decomposing FFT into arithmetic operations involving Formula (2.11), (2.12), (2.13), and (2.14).
-  • When $L$ is large, the large volume of input and output data enables high-quality statistical analysis.
-  • The computational complexity is proportional to $L$, because increasing $L$ by 1 adds one more step involving a sum of multiplications.
-  • Each step in the forward transformation increases the variance by a factor of $2$, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$.
+  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations as Formula (2.14), (2.15), (2.16), and (2.17).
+  • When $L$ is large, the substantial volume of input and output data enables high-quality statistical analysis.
+  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of multiplications.
+  • Each step in the forward transformation doubles the variance, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$.
 Because the reverse transformation divides the result by $2^{L}$, its uncertainty mean decreases with $L$ as $\sqrt{1/2}^{L}$.
-The uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
-  • Forward and reverse transformations are identical except for a sign difference, so they are essentially the same algorithm, and any observed difference arises purely from the input data.
+Consequently, the uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
+  • The forward and reverse transformations are identical except for a sign difference, meaning that they are essentially the same algorithm, and any observed difference arises solely from the input data.
 In normal usage, forward and reverse FFT transforms differ in their data prospective of time domain versus frequency domain:
-  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum where most values are 0, causing its uncertainties to grow faster.
-  • The reverse transformation spreads the precise frequency-domain spectrum (with most values being 0) back into a time-domain sine or cosine signals, causing its uncertainties to grow more slowly.
-8.3 Modeling Errors of DFT and FFT
-Although mathematically self-consistent, by implying a periodic boundary condition in the time domain [1], DFT and its efficient implementation FFT are only approximate for the mathematically defined continuous Fourier transformation [1].
-They show no modeling error only when the input signal frequency $f$ satisfies $f=j\frac{2\pi}{N}$, and exhibit different degrees of modeling errors otherwise, with the modeling errors peaking at $f=(j+\frac{1}{2})\frac{2\pi}{N}$.
-Because of these modeling errors, using DFT and FFT as the digital implementation of continuous Fourier transformation is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [13].
-To avoid the modeling errors of FFT, only Formula (8.1) and (8.2) are used in this paper.
+  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zeros, causing its uncertainties to grow more rapidly.
+  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zeros) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
 9.4 Testing Signals
@@
 Empirically, except when using trigonometric library directly for FFT transformations of clean Sin and Cos signals:
-  • Results from Sin and Cos signals are statistically indistinguishable.
-  • Results from Sin signals at different frequencies are also statistically indistinguishable.
-Therefore, the results for Sin and Cos signals at all frequencies are pooled together for statistical analysis, under the combined category Sin/Cos signals.
+  • The results obtained from Sin and Cos signals are statistically indistinguishable.
+  • Similarly, the results from Sin and Cos signals at different frequencies does not show significant differences statistically except in the situation of numerical error resonance.
+Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals.
 9.5 Trigonometric Library Errors
@@
 Figure 35: Difference between the library and the Quart $\cos(x)/\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend).
-Formula (8.1) and (8.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, in which $L$ is the FFT order.
-To minimize numerical errors in computing $\sin(x)$, indexed sine can replace standard library sine functions:
+Formula (9.1) and (9.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, where $L$ is the FFT order.
+To minimize numerical errors in computing $\sin(x)$, the following indexed sine can be used in place of standard library sine functions:
   1. 1.
@@
     The values of $\sin(\pi j/2^{L})$ are extended to all the integer value of $j$ by leveraging the periodicity of $\sin(2\pi j/2^{L})$.
-Such constructed indexed $\sin(x)$ is named as the Quart indexed sine function. In contrast, the direct application of library $\sin(x)$ is named as the Library sine function.
-Because the Quart sine function obeys the symmetry and periodicity of sine function strictly, it is superior to the library function.
-  • Figure 31 shows that the value difference between Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases roughly linearly with $|x|$.
-  • Figure 32 shows the value difference between Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but they are $10^{2}$ times larger than those of $\sin(x)$.
-Therefore, the linear spectrum in Formula (8.3) may contain huge numerical errors using library sine functions.
-For both sine functions, the uncertainty of each $\sin(x)$ is assumed to be its ULP.
-Because Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in library $\sin(x)$, its true numerical errors is probably less than ULP, therefore its uncertainty should be sightly overestimated.
-In contrast, Figure 31 shows that library sine function has underestimated uncertainties.
+The constructed indexed $\sin(x)$ is referred to as the Quart indexed sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
+Because the the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides numerical accuracy compared to Library function.
+  • Figure 34 shows that the value difference between the Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
+  • Figure 35 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but are $10^{2}$ times larger than those observed for $\sin(x)$.
+Therefore, the linear spectrum in Formula (9.3) may contain significant numerical errors when computed using library sine functions.
+For both sine functions, the uncertainty of each $\sin(x)$ is assumed to equal its ULP, which is displayed in Figure 34.
+Because the Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in calls to the library $\sin(x)$, its true numerical errors are likely smaller than ULP, therefore its uncertainty is sightly overestimated.
+In contrast, Figure 34 indicates that the Library sine function exhibits underestimated uncertainties.
 9.6 Using Quart Sine for Sin/Cos Signals
-Figure 33: FFT spectrum of $\sin(2\pi 3j/2^{6})$ using Quart sine function after the forward transformation computed by variance arithmetic.
-The legend distinguishes between the uncertainty and the value errors.
+Figure 36: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation.
+The legend distinguishes between the uncertainty and the value error.
 The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
-Figure 34: FFT waveform of $\sin(2\pi 3j/2^{6})$ using the Quart sine function after the reverse transformation computed by variance arithmetic.
-The legend distinguishes between the uncertainty and the value errors.
+Figure 37: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation.
+The legend distinguishes between the uncertainty and the value error.
 The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
-Figure 35: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal at frequency 33 as a function of FFT order (x axis) and transformations (legend) using Quart sine function.
-Figure 36: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using Quart sine function. The FFT order is $18$.
-Using Quart sine function, for a sine wave with $3$ as frequency, Figure 33 displays the spectrum for the forward transformation, while Figure 34 presents the waveform for the reverse transformation.
-  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation is $0.31$.
+Figure 38: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
+Figure 39: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using the Quart sine function. The FFT order is $18$.
+Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 36 displays the spectrum for the forward transformation, while Figure 37 presents the waveform for the reverse transformation.
+  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.31$.
 In the reverse transformation, the error deviation is $0.43$.
@@
   • The uncertainty mean of the forward transformation is $7$ times larger than that of the reverse transformation.
+Even with this difference, the uncertainty tracks the corresponding value error effectively.
 Figure 38 illustrates that the forward transformation differs from the reverse transformation significantly.
   1. 1.
-    When FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that in the reverse transformation.
-At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times more than that of the reverse transformation.
+    As the FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that of the reverse transformation.
+At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times greater than that of the reverse transformation.
   2. 2.
-    At FFT order $18$, Figure 36 shows that the normalized error distribution for the reverse transformation is wider than that for the forward transformation, but the two distributions do not differ by order-of-magnitude.
-Therefore, the value error grows at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
+    At FFT order $18$, Figure 39 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, although the two distributions do not differ by an order-of-magnitude.
+Therefore, the value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
   3. 3.
-    In forward transformation, error deviations reach their stable values quickly when FFT order $L\geq 4$.
-In contrast, in the reverse transformation, error deviations stabilize slowly with increasing FFT order until $L\geq 15$.
-The slow growth of normalized errors in the reverse transformation is attributed to its input data which contains all precise zeros except at the two frequency indexes.
-Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases.
+    In forward transformation, error deviations reach their stable values rapidly once the FFT order $L\geq 4$.
+In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching stability only when $L\geq 15$.
+This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
+Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
 9.7 Using Library Sine for Sin/Cos Signals
-Figure 37: FFT spectrum of $\sin(2\pi 3j/2^{6})$ using Library sine function after the forward transformation computed by variance arithmetic.
-The legend distinguishes between the uncertainty and the value errors.
+Figure 40: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
+The legend distinguishes between the uncertainty and the value error.
 The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
-Figure 38: FFT waveform of $\sin(2\pi 3j/2^{6})$ using Library sine function after the reverse transformation computed by variance arithmetic.
-The legend distinguishes between the uncertainty and the value errors.
+Figure 41: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
+The legend distinguishes between the uncertainty and the value error.
 The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
-Figure 39: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin signal at frequency 33 as a function of FFT order (x axis) and transformations (legend) using Library sine function.
-Figure 40: Histograms of normalized errors of Sin signal at frequency 33 for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
-Using library sine function, for a sine wave with $3$ as frequency, Figure 37 displays the spectrum for the forward transformation, while Figure 38 presents the waveform for the reverse transformation:
-  • The uncertainties are identical to the corresponding values using Quart sine function, because the input uncertainties are identical for these two sine functions.
-This identical relationship presents in all FFT orders when comparing the corresponding uncertainty means in Figure 35 and 39.
-  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in Library sine function are more than those using Quart sine function in Figure 31.
-When compared with Figure 36, Figure 40 also confirms that at FFT order $18$, normalized errors have much wider distributions using Library sine function than those using Quart sine function.
-Due to numerical errors in Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 39.
+Figure 42: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
+Figure 43: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
+Using the Library sine function, for a sine wave with a frequency of $3$, Figure 40 displays the spectrum for the forward transformation, while Figure 41 presents the waveform for the reverse transformation:
+  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions.
+This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and 42.
+  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 34.
+  • When compared with Figure 39, Figure 43 also confirms that at FFT order $18$, the normalized errors distribution is significantly broader when using the Quart sine function.
+  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 42 that that in Figure 38.
+Using the Library sine function, variance arithmetic also achieves proper coverage for sine and cosine signals.
 9.8 Numerical Error Resonance Using Library Sine
-Figure 41: Error deviation (z-axis) of FFT forward transformation of $sin(2\pi fj/2^{L})$ as a function of frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
-Figure 42: Error deviation (z-axis) of FFT reverse transformation of $sin(2\pi fj/2^{L})$ as a function of frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
-Value errors in the reverse transformation show a clear trend of increasing with time index, as in Figure 41.
-These large value errors looks systematic and do not appear as typical noise visually; or rather, they look more like a resonant pattern.
-Such increases present in other frequencies and FFT order, as well as in computational result using mathematical libraries such as SciPy.
-A FFT spectrum of the forward transformation shows a sudden increase of intensity at frequency index which is also consistent with the result of SciPy, as in Figure 37.
-Figure 41 and 42 show that error deviations of such value errors increase with sine or cosine frequencies, the resonant beats between the signal and the numerical library errors become stronger.
-In contrast, such increase does not appear at all when using Quart sine function, such as in Figure 33 and 34.
-Figure 31 shows that the numerical errors using Library sine increase with a periodicity of $\pi$, which may resonate with a signal which has a periodicity of an integer fold of $\pi$, to produce the resonant pattern in Figure 37 and 38.
-To avoid the numerical error resonance in Figure 41 and 42, $10^{-14}$ input noise is required to add to the input sine or cosine signals.
+Figure 44: Error deviation (z-axis) of FFT forward transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
+Figure 45: Error deviation (z-axis) of FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
+In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 44.
+These large value errors appear systematic rather than random and visually resemble a resonant pattern.
+Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy.
+In the forward transformation, although there is no visible resonant pattern in FFT spectrum, the intensity of value errors increase with FFT order.
+Figure 44 and 45 demonstrate that the error deviations increase with sine or cosine frequency.
+In contrast, this increase is completely absent when using the Quart sine function, as seen in Figure 36 and 37.
+Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 40 and 41.
+At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
+To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
 9.9 Using Quart Sine for Linear Signal
-Figure 43: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal as a function of FFT order (x axis) and transformations (legend) using Quart sine function.
-Figure 44: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) using Quart sine function. The FFT order is $18$.
-Figure 43 shows that using Quart sine function can track the result uncertainties of Linear signals with proper coverage.
-Because the input to reverse transform no longer contains mostly precise zeros, the output uncertainty increases far more rapidly with FFT order than its counterpart in Figure 35.
-Figure 44 shows that for each transformation, the normalized error distribution in reverse transformation is similar to its counterpart of the Sin/Cos signals using Quart sine function in Figure 36, while the distribution in the forward transformation is similar to the Linear signal using Library sine function in Figure 46.
-The reason for this disparity in similarity of the normalized error distributions is not clear.
+Figure 46: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Quart sine function.
+Figure 47: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
+Figure 46 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage.
+Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases much more rapidly with FFT order than its counterpart in Figure 38.
+This increase enables the proper coverage in Figure 46, resulting in error deviations similar to those Figure 38.
+The normalized error histogram of the reverse transformation in Figure 49 is narrower than that of Figure 43, confirming Figure 46.
+Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
 9.10 Using Library Sine for Linear Signal
-Figure 45: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal as a function of FFT order (x axis) and transformations (legend) using Library sine function.
-Figure 46: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
-As shown as the difference between Figure 31 and 32, linear signals using Library sine function may introduce $10^{3}$ times larger numerical errors into the results through Library $\cos(x)/sin(x)$.
-The key question is whether variance arithmetic can effectively track these extra errors or not.
-Figure 45 shows that proper coverage can no longer be achieved because value error outpaces uncertainty as FFT order increases.
-Figure 46 shows that the normalized error distribution for the reverse transformation is no longer bound effectively within three deviations.
-Although the normalized error distribution for the forward transformation using Library sine function looks similar to its counterpart of using Quart sine function in Figure 46 looks similar to that using Quart sine function in Figure 44, detailed analysis shows that the normalized error distribution using Library sine function contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal, respectively.
-The presence of these extreme values consistent with those delta-like numerical errors from Figure 32.
-The observed difficulty of variance arithmetic in tracking Linear signals using Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified errors.
-A potential solution is to develop a new sin library implemented within variance arithmetic framework, ensuring that all numerical calculation errors are explicitly accounted for, as what Quart sine function has achieved in FFT transformations.
+Figure 48: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
+Figure 49: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
+As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$.
+The key question is whether variance arithmetic can effectively track these additional errors.
+Figure 48 shows that proper coverage can no longer be achieved, as the value error increases faster than the uncertainty with rising FFT order.
+Figure 49 demonstrates that the normalized error distribution for the reverse transformation is no longer effectively bounded within three standard deviations.
+Although the normalized error distribution for the forward transformation using the Library sine function appears visually similar to that obtained using the Quart sine function in Figure 49, a detailed analysis reveals that the normalized error distribution contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal of the normalized errors, respectively.
+The presence of these extreme values is consistent with those delta-like numerical errors observed in Figure 35.
+The observed difficulty of variance arithmetic in tracking Linear signals computed using the Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified numerical errors.
+A potential solution is to develop a new sine function library implemented directly within the variance arithmetic framework, ensuring that all numerical calculations are performed with explicit uncertainty tracking.
 9.11 Ideal Coverage
-Figure 47: Histograms of normalized errors of Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
-Figure 48: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals as a function of FFT order (x axis) and transformations (legend) using Library sine function.
-Figure 49: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for forward transformations of Linear signals using Library sine function.
-Figure 50: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Library sine function.
-Figure 51: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for forward transformations of Linear signals using Quart sine function.
-Figure 52: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Quart sine function.
-Adding sufficient noise to the input can overpower unspecified input errors, thereby achieving ideal coverage; for example, applying an input noise of $10^{-3}$ to a Linear signal when using library sine function.
-  • Figure 47 presents the corresponding histogram.
-As expected, the normalized errors for the forward and reverse transformations follow Normal distributed, while those for the roundtrip transformations are Delta distributed at 0zero, indicating perfect recovery of the input uncertainties.
-  • Figure 48 illustrates the corresponding error deviations and uncertainty means:
+Figure 50: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
+Figure 51: Histograms of normalized errors for Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) computed using the Library sine function.
+The FFT order is $18$.
+Figure 52: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Library sine function.
+Figure 53: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Library sine function.
+Figure 54: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Quart sine function.
+Figure 55: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Quart sine function.
+Adding sufficient noise to the input can suppress unspecified input errors and thereby achieve ideal coverage.
+For example, applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage.
+Figure 50 illustrates the error deviations and uncertainty means:
   • As expected, the result uncertainty means for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
@@
   • As expected, the result uncertainty means for the roundtrip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
-    • As expected, the result error deviations for the forward and reverse transformations are constant at $1$, whereas those for the roundtrip transformation decay exponentially to 0 with increasing FFT order.
-Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, which is expected because FFT is a linear algorithm [13].
-The range of ideal coverage depends on how accurately the input uncertainty reflects the actual input noise.
-For Linear signals using Library sine function, Figure 49 and 50 present the error deviation as a function of the added noise and FFT order for the forward and reverse transformations, respectively.
-Ideal coverage corresponds to the region where the error deviation is $1$.
-Outside this region, proper coverage cannot be achievable.
-Because uncertainties grow more slowly in reverse transformation than in forward transformation, the reverse transformations exhibit a smaller ideal coverage region.
-Furthermore, as numerical errors increase with computational load, the range of input noise that yields ideal coverage decreases with increasing FFT order.
-At sufficiently high FFT orders, visually, beyond FFT order $25$ for the reverse transformation, ideal coverage may not be achievable.
-Although FFT is regarded as one of the most robust numerical algorithms very insensitive to input errors, it can still fail due to numerical errors in library sine function.
-Such deterioration of the calculation is not easily detectable using conventional floating-point arithmetic.
-In contrast, for Linear signals using Quart sine functions, as shown in Figure 51 and 52 for the forward and reverse transformations, respectively, the ideal coverage region is significantly larger, and proper coverage is also achieved in other regions.
-Forward transformations show a larger ideal coverage region than that of reverse transformations.
-As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders.
-The difference is that using Quart sine functions, within the proper coverage region, error deviations differ from $1$ only marginally.
-8.12 Prec Sine Functions
-Figure 53: Value Errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) as a function of $x$ (x-axis) and $L$ using either Prec or Quart sine function (legend).
-Figure 54: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals as a function of FFT order (x axis) and transformations (legend) using Prec sine function.
-Figure 55: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals using Prec sine functions.
-Figure 56: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Prec sine function.
-Quart sine function over-estimates the uncertainty, while Library sine function underestimates the uncertainty.
-Is it possible to have a better sine function?
-If Quart sine function is constructed using the 128-bit floating-point library of gcc, then converted to 64-bit values, using the difference between a 128-bit and the corresponding 64-bit value as the uncertainty for the 64-bit value, the resulting indexed sine function is named as the Prec sine function.
-Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare Prec and Quart sine functions, Figure 53 illustrates that Prec sine function produces slightly more value errors, while having half uncertainties.
-Figure 54 and 55 both show that Prec sine function underestimates uncertainties slight, with stable error deviation around $1.8$ for both forward and reverse transformations without added noise.
-Compared with 52, Figure 56 illustrates that the ideal coverage using Prec sine function is identical to using Quart sine function.
-The observation that 128-bit sine function does not improve over 64-bit sine function in any noticeable way challenges the legitimacy of using the difference as the uncertainty, the correctness of the 128-bit gcc numerical library, or the algorithm in converting a 128-bit floating-point value to 64-bit.
-Both Quart and Prec sine function are not ideal. An ideal sine function should have error deviations closer to $1$ without added noise.
+  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$, whereas those for the roundtrip transformation decay exponentially toward 0 as the FFT order increases.
+Figure 51 presents the corresponding histogram.
+As expected, the normalized errors for the forward and reverse transformations follow Normal distributions, while those for the roundtrip transformation are Delta distributed around zero, indicating perfect recovery of the input uncertainties.
+According to Figure 51, adding either Gaussian or white noise to input results in the same Gaussian distribution of the normalized errors.
+Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, as expected since FFT is a linear algorithm [13].
+The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise.
+For Linear signals computed using the Library sine function, Figure 52 and 53 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
+Ideal coverage corresponds to the region where the error deviation equals $1$.
+Outside this region, proper coverage cannot be achieved.
+Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region.
+Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order.
+At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable.
+Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function.
+Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
+In contrast, for Linear signals computed using the Quart sine functions, Figure 54 and 55 show the results for the forward and reverse transformations, respectively.
+The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions.
+The forward transformations exhibits a broader ideal coverage region than the reverse transformations.
+As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or the Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders.
+The key difference is that when using the Quart sine functions, the error deviations within the proper coverage region differ from $1$ only marginally.
+9.12 Prec Sine Function
+Figure 56: Value errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) versus $x$ (x-axis) and $L$ computed using either the Prec or the Quart sine function (legend).
+Figure 57: Error deviations (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT orders (x-axis) and transformation types (legend) computed using the the Prec sine function.
+Figure 58: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals computed using the the Prec sine function.
+Figure 59: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for reverse transformations of Linear signals computed using the the Prec sine function.
+Signal vs Sine Transformation Quart Prec Library
+Sin/Cos Forward $10^{-15}$ $10^{-15}$ $10^{-14}$
+Sin/Cos Reverse $10^{-12}$ $10^{-12}$ $10^{-11}$
+Linear Forward $10^{-9}$ $10^{-9}$ $10^{-6}$
+Linear Reverse $10^{-7}$ $10^{-7}$ $10^{-3}$
+Table 2: The minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different sine functions.
+The Quart sine function tends to overestimate the uncertainty, while the Library sine function underestimates it.
+This raises the question of whether a sine function with a more accurate uncertainty can be developed.
+If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation.
+The resulting indexed sine function is referred to as the Prec sine function.
+Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare the Prec and the Quart sine functions, Figure 56 shows that the the Prec sine function produces slightly larger value errors, while having half uncertainties.
+Figure 57 and 58 demonstrate that the the Prec sine function slightly underestimates uncertainties, maintaining a stable error deviation around $1.8$ for both forward and reverse transformations without added noise.
+Compared with Figure 55, Figure 59 shows that the ideal coverage achieved using the the Prec sine function is identical to that obtained with the Quart sine function.
+The observation that the 128-bit sine function provides no noticeable improvement over the 64-bit sine function raises concerns about the statistical legitimacy of directly using the difference as the uncertainty, the accuracy of the gcc 128-bit numerical library, or the algorithm used to convert a 128-bit floating-point values to 64-bit.
+Table 2 compares using different sine functions for Sin/Cos or Linear signals.
+Both Quart and the the Prec sine function are not ideal.
+An ideal sine function should have error deviations closer to $1$ without added noise.
 9.13 Summary
-The library sine functions implemented with conventional floating-point arithmetic have been shown to contain numerical errors equivalent to $10^{-3}$ of input precision in worst cases for FFT transformations.
-These library $\sin(x)$ errors increase periodically with $x$, which may resonate with periodic input signals.
-The computational result using floating-point arithmetic is very sensitive to input data, amount of calculation, and library function implementation.
-This means that a small-scale test cannot properly qualify the result of a large-scale calculation.
+The Library sine functions implemented using conventional floating-point arithmetic have been shown to contain numerical errors equivalent up to $10^{-3}$ of the input precision in the worst cases of FFT transformations.
+These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals.
+Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the specific implementation of the mathematical library.
+This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations.
 The impact of numerical errors within mathematical libraries has not received sufficient attention.
-The numerical errors using floating-point arithmetic are very difficult to detect and quantify.
-Variance arithmetic offers a robust alternative, as its computed values closely match those from conventional floating-point arithmetic while its associated uncertainties trace all input errors.
-Moreover, its resulting error deviations enable objective classification of calculation quality as ideal, proper, or suspicious.
+Moreover, such errors in floating-point arithmetic are inherently difficult to detect and quantify.
+Variance arithmetic offers a robust alternative, as its computed values closely align with those obtained from conventional floating-point arithmetic while its associated uncertainties systematically trace all input errors.
+Furthermore, its resulting error deviations provide an objective basis for classifying of calculation quality as ideal, proper, or suspicious.
 10 Regressive Generation of Sin and Cos
-Figure 57: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ as a function of $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
-Figure 58: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ as a function of regression order (x-axis) when compared with Quart sine function (legend)
+Figure 60: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ versus $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
+Figure 61: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ versus regression order (x-axis) when compared with the Quart sine function (legend)
 Formula (10.2) and Formula (10.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1).
@@
 $\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (10.4)
-Formula (9.2) is not suitable to compute $\sin(x)$ when $x\rightarrow 0$ because it fails similarly as catastrophic cancellation, obtaining excessive uncertainty.
-In Figure 57, Quart sine has uncertainties proportional to $\sin(\pi j/2^{L})$, while regression sine has its uncertainties increase when $x\rightarrow 0$.
-Figure 58 tests Quart and Regression sin/cos for $\sin(x)^{2}+\cos(x)^{2}-1=0$.
-In Figure 58, the result uncertainties are nearly a constant independent of the regression order $L$ when $L\geq 4$ for both Regression sine and Quart sine.
-It also shows that the value errors for both are quite comparable, resulting in the error deviation of the Regression sin/cos closer to $1$.
+Formula (10.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation, resulting in excessive uncertainty.
+In Figure 60, the Quart sine function exhibits uncertainties proportional to $\sin(\frac{\pi}{2^{L}})$, whereas the regression sine function shows increasing uncertainties as $x\rightarrow 0$.
+Unlike catastrophic cancellation in floating-point arithmetic, variance arithmetic uses very coarse precision to signal that the regression algorithm is unfit to compute $\sin(x\rightarrow 0)$.
+Figure 61 evaluates the Quart and Regression sine functions using $\sin(x)^{2}+\cos(x)^{2}-1=0$.
+The figure shows that the resulting uncertainties remain nearly constant and independent of the regression order $L$ when $L\geq 4$ for both the Regression sine and the Quart sine functions.
+It also demonstrates that the value errors for both are comparable.
+The Regression sine function has error deviations closer to $1$.
 11 Conclusion and Discussion
 11.1 Summary of Statistical Taylor Expansion
-The starting point of statistical Taylor expansion is the assumption that all input variables have uncorrelated uncertainties.
-This is considered a very reasonable statistical requirement on the input data [1].
-Once this assumption is met, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors.
-It can track the variable dependence in the intermediate steps using standard statistical methods.
-While it eliminates the dependency problem, it requires a more rigorous process to determine resulting mean and deviation for an analytic expression.
-In addition, it can reject calculations on unified statistical and mathematical grounds.
-It also incorporates sample count into statistical analyses.
+The starting point of statistical Taylor expansion is the assumption that all input variables obeys uncorrelated uncertainties.
+This is generally regarded as a reasonable and practical statistical requirement for input data [1].
+Once this assumption is satisfied, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors and traces track the variable dependencies in intermediate steps using standard statistical methods.
+While this approach eliminates the dependency problem, it also requires a more rigorous process to determine the resulting mean and deviation for any analytic expression.
+Furthermore, it can reject invalid calculations on unified statistical and mathematical grounds and explicitly incorporates sample counts and the corresponding underlying uncertainty distributions into the statistical analysis.
 11.2 Summary of Variance Arithmetic
-Variance arithmetic simplifies statistical Taylor expansion by assuming that all inputs are Gaussian distributed and with enough sample counts for ideal variance.
-Ill-formed problems can invalidate resulting variance in several ways: non-convergent, unstable, negative, infinite, and unreliable.
+Variance arithmetic simplifies statistical Taylor expansion by assuming that the sample counts for all inputs are sufficiently large to achieve ideal variance.
+Ill-formed problems can invalidate the resulting variance in several ways by producing non-convergent, unstable, negative, infinite, or unreliable results.
 The presence of ideal coverage is a necessary condition for a numerical algorithm in variance arithmetic to be considered correct.
-Ideal coverage also defines the algorithm’s optimal applicable range for the algorithm.
+Ideal coverage also defines the algorithm’s optimal applicable range.
 Under ideal coverage, the calculated uncertainty equals the deviation of the value errors, and the normalized errors follow either a Normal or Delta distribution depending on the context.
-Variance arithmetic can provide proper coverage for floating-point rounding errors.
-The applicability of variance arithmetic has been demonstrated in diverse scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
+Variance arithmetic provides proper coverage for floating-point rounding errors.
+The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
 The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
 11.3 Improvements Needed
-This paper presents variance arithmetic which is still in its early developmental stage. From a theoretical perspective, several important questions remain to be addressed.
+This paper presents variance arithmetic which is still in its early stage of development. From a theoretical standpoint, several important questions still need to be addressed.
 Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty.
-Without this, the value errors in the library functions can produce unpredictable and potentially large effects on numerical results.
+Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results.
+How to compute and validate these new uncertainty-bearing library functions will need innovation.
+For example, the uncertainty of a 64-bit value can not directly determined by its difference with the corresponding 128-bit value,
 Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized.
-Determining the uncertainty upper bounds for each analytic expression analytically is important.
-Under current numerical framework, the measured upper bounds in Figure 7 for $(x+\delta x)^{c}$, and Figure 9 for $\sin(x+\delta x)$ may change with implementation such as between 32-bit and 64-bit floating-point computations.
 Cases where $\delta^{2}f<\widehat{\delta^{2}}f$ due to low sample count require further theoretical study and experimental verification.
-The performance of variance arithmetic has to be improved.
-The fundamental formulas for statistical Taylor expansion, Formula (2.6), (2.3), (2.9), and (2.3) contains large number of independent summations, making them excellent candidates for parallel processing, such as direct hardware implementation.
-The procedural nature of these formulas allows statistical Taylor expansion to be implemented directly in hardware.
-When an analytic expression undergoes statistical Taylor expansion, the result expression may be very complex, such as in matrix inversion.
+Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step.
+These results can validate variance arithmetic, and provide performance enhancement to variance arithmetic when the convergence of an analytic expression has to be deduced numerically from the full expansion each time.
+The performance of variance arithmetic must be improved for broader practical adoption.
+The fundamental formulas of statistical Taylor expansion, Formula (2.9), (2.3), (2.12), and (2.3) contain a large number of independent summations, making them excellent candidates for parallel processing.
+Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
+When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion.
 However, modern symbolic computation tools such as SymPy can greatly facilitate these calculations.
-This suggests that it may be time to shift from purely numerical programming toward analytic programming for problems with inherently analytic formulations.
+This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
 As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty.
-This capability can help engineers pinpoint the primary drivers of measurement inaccuracy and thus guide targeted improvements in data acquisition and processing strategies.
-A key open question is whether variance arithmetic can be adapted to provide ideal coverage for floating-point rounding errors.
-Once the mechanism for rounding error is thoroughly understood, variance arithmetic with ideal coverage to floating-point rounding errors is quite valuable because many theoretical calculations lack explicit input uncertainty.
-Because traditional numerical approaches are based on floating-point arithmetic, they need to be reexamined or even reinvented in variance arithmetic.
-For example, most conventional numerical algorithms aim to select optimal computational paths, whereas variance arithmetic fundamentally rejects conceptually all path-dependent calculations.
-Reconciling these two paradigms may present a significant challenge.
-Establishing theoretical foundation to apply statistical Taylor expansion in the absence of analytic solution, or with only limited low-order numerical derivatives, such as in solving differential equations, remains an important research direction.
+This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies.
+A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors.
+Once the mechanism of rounding error is thoroughly understood, developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
+Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic.
+For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas variance arithmetic conceptually rejects all path-dependent calculations.
+Reconciling these two paradigms may present a significant and ongoing challenge.
+Establishing theoretical foundation for applying statistical Taylor expansion in the absence of a closed-form analytic solution, or when only limited low-order numerical derivatives are available, as in solving differential equations, remains an important direction for future research.
 11.4 Acknowledgments
-As an independent researcher without institutional affiliation, the author expresses deep gratitude for the encouragement and valuable discussions provided by Dr.
-Zhong Zhong (Brookhaven National Laboratory), Prof Weigang Qiu (Hunter College).
+As an independent researcher without institutional affiliation, the author expresses sincere gratitude to Dr.
+Zhong Zhong (Brookhaven National Laboratory) and Prof Weigang Qiu (Hunter College) for their encouragement and valuable discussions.
 Special thanks are extended to the organizers of AMCS 2005, particularly Prof.
 Hamid R.
-Arabnia (University of Georgia), and organizers of NKS Mathematica Forum 2007.
-Prof Dongfeng Wu (Louisville University) is acknowledged for insightful guidance on statistical topics.
-The author is also indebted to the editors and reviewers of Reliable Computing for their substantial assistance in shaping and accepting a previous version of this work, with special recognition to Managing Editor Prof.
+Arabnia (University of Georgia), and to the organizers of the NKS Mathematica Forum 2007.
+The author also gratefully acknowledges Prof Dongfeng Wu (Louisville University) for her insightful guidance on statistical topics.
+Finally, heartfelt appreciation is extended to the editors and reviewers of Reliable Computing for their substantial assistance in shaping and accepting an earlier version of this work, with special recognition to Managing Editor Prof.
 Rolph Baker Kearfott.

### v13

@@
 $\epsilon(\kappa_{s},N)=$  $\;1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.48)
-When the underlying distribution is Normal, Formula (2.47) and (2.48) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Gaussian error function [5].
+When the underlying distribution is Normal, Formula (2.47) and (2.48) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
 Figure 5 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$.
@@
 Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 5 should be used as $\kappa$ in Formula (2.7).
+$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges for the uniform and Gaussian distributions respectively.
 2.10 Ideal Statistics
@@
 Therefore it is sufficient to compute the stable variances $\widehat{\delta^{2}}f$ with $\zeta(n,5)$.
-  • When the underlying uncertainty distribution is Uniform, Figure 8 shows the same trend, however the stable variances are reached only when $N>10^{5}$.
+  • When the underlying uncertainty distribution is Uniform, Figure 8 shows the same trend, however the stable variances are reached only when $N>10^{4}$.
 The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.18).
@@
 Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution.
+Define the ideal ratio to be zero when the sample count $N$ is $1$.
 When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.50) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
 An ideal ratio also applies to the corresponding uncertainty bias.
-The trends of uncertainty biases $\overline{f}-f$ versus sample count NN are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$.
+The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$.
 This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
+Statistical Taylor expansion outputs ideal uncertainty, ideal uncertainty bias, and ideal ratio for the chosen ideal bounding ranges, in which the ideal ratio quantifies the reliability of the ideal uncertainty and ideal uncertainty bias.
 2.11 Dependency Tracing
@@
   • positive: At every order, the expansion variance must be positive.
-  • stable: To avoid truncation error [13], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value.
+  • stable: To avoid truncation error [12], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value, in which $5.73\;10^{-7}$ is the bounding leakage for Gaussian distribution with bounding range $\kappa=5$.
 This rule ensures sufficiently fast convergence in the context of monotonic convergence.
@@
   • Reduction.
-An application algorithm computes numerical values from an analytic formula. Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems.
+An application algorithm computes numerical values from an analytic formula.
+Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems.
+For example, for the catastrophic cancellation example of Formula (1.2), Formula (4.3) shows that variance arithmetic bound the rounding error of $1$ with uncertainty $1$:
+$64919224$ $\times 205117922-159018721\times 83739041=$
+$13316075197586562.-13316075197586560.\pm 1=2.\pm 1;$ (4.1)
 A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged.
-For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a round-trip transformation that is performing a forward transformation followed by its reverse transformation.
+For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a roundtrip transformation that is performing a forward transformation followed by its reverse transformation.
 The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process.
@@
 Let vector $p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [14].
-Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$37].
+Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$14].
 Formula (7.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [14].
@@
 $\beta_{j}$ $=\beta\;\frac{H(H+1)(2H+1)}{3}=\sum_{X=-H}^{H}XY_{j-H+X};$ (8.4)
-According to Figure LABEL:fig:_Moving_Linear_Fit_Coeff, in which $H$ takes an example value of 2, the values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [1].
+The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [29].
 $\beta_{j}$ $=\beta_{j-1}-\alpha_{j-1}+H\left(Y_{j-2H-1}+Y_{j}\right);$ (8.5)
@@
 9.6 Using Quart Sine for Sin/Cos Signals
-Figure 36: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation.
+Figure 36: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation.
 The legend distinguishes between the uncertainty and the value error.
 The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
-Figure 37: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation.
+Figure 37: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation.
 The legend distinguishes between the uncertainty and the value error.
@@
 Figure 38: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
+The variance arithmetic is implemented using python.
 Figure 39: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using the Quart sine function. The FFT order is $18$.
-Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 36 displays the spectrum for the forward transformation, while Figure 37 presents the waveform for the reverse transformation.
-  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.31$.
-In the reverse transformation, the error deviation is $0.43$.
+Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 36 displays the value error spectrum for the forward transformation, while Figure 37 presents the value error waveform for the reverse transformation.
+  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.60$.
+In the reverse transformation, the error deviation is $0.55$.
 Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
@@
 Even with this difference, the uncertainty tracks the corresponding value error effectively.
+  • The value error waveform for the roundtrip transformation is qualitatively similar to that in Figure 37 with a similar error deviation of $0.55$.
 Figure 38 illustrates that the forward transformation differs from the reverse transformation significantly.
-  1. 1.
-    As the FFT order increases, the uncertainty mean of the forward transformation grows exponentially faster than that of the reverse transformation.
-At FFT order $18$, the uncertainty mean of the forward transformation is $10^{3}$ times greater than that of the reverse transformation.
-  2. 2.
-    At FFT order $18$, Figure 39 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, although the two distributions do not differ by an order-of-magnitude.
-Therefore, the value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
-  3. 3.
-    In forward transformation, error deviations reach their stable values rapidly once the FFT order $L\geq 4$.
-In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching stability only when $L\geq 15$.
+  • In the forward transformation, error deviations reach their stable values around $0.65$ rapidly once the FFT order $L\geq 6$.
+  • In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching their stable values around $0.90$ only when $L\geq 12$.
 This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
-Despite these differences, error deviations are between $0.2$ and $0.5$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
+  • At FFT order $18$, Figure 39 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, confirming the larger error deviation.
+The value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
+Despite these differences, error deviations are $0.65$ and $0.90$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
+9.7 Sensitivity to Quart sine functions
+Figure 40: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
+The variance arithmetic is implemented using C++.
+Figure 41: Histograms of normalized errors between the Quart sine functions of C++, Java, and Python in the unit of least significant bits.
+The reverse transformation of Sin/Cos signals is very sensitive to the underlying Quart sine function.
+If variance arithmetic is implemented using C++ instead of Python, Figure 40 shows that the error deviations of the reverse transform quickly reaches the stable values around $0.89$, while the forward transformation remains similar to its counterpart in Figure 38.
+Figure 40 also shows identical uncertainty means to their counterparts in Figure 38.
+Figure 41 shows that the difference of the Quart sine functions between Python and C++ is only about $1\%$ LSBs.
+Figure 41 also reveals that the difference of the Quart sine functions between Java and C++ is only about $0.1\%$ LSBs.
+Indeed, the results between Java and C++ are much more similar.
+Because the Python implementation gives smaller error deviations, its Quart sine function has less numerical errors on average.
+For discussion simplicity, the rest results are only based on the python implementation.
 9.8 Using Library Sine for Sin/Cos Signals
-Figure 40: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
+Figure 42: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
 The legend distinguishes between the uncertainty and the value error.
 The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
-Figure 41: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
+Figure 43: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
 The legend distinguishes between the uncertainty and the value error.
@@
 Figure 45: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
-Using the Library sine function, for a sine wave with a frequency of $3$, Figure 40 displays the spectrum for the forward transformation, while Figure 41 presents the waveform for the reverse transformation:
+Using the Library sine function, for a sine wave with a frequency of $3$, Figure 42 displays the value error spectrum for the forward transformation, while Figure 43 presents the vale error waveform for the reverse transformation:
   • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions.
-This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and 42.
+This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and $4$.
   • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 34.
   • When compared with Figure 39, Figure 45 also confirms that at FFT order $18$, the normalized errors distribution is significantly broader when using the Quart sine function.
-  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 42 that that in Figure 38.
-Using the Library sine function, variance arithmetic also achieves proper coverage for sine and cosine signals.
+  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 44 that these in Figure 38.
+  • The stable error deviations using the Library sine function are much larger than their counterparts using Quart library functions: $10.1>0.65$ for the forward transformation, $6.3>0.90$ for the reverse transformation, and $1.47>0.94$ for the roundtrip transformation.
+Using the Library sine function, variance arithmetic barely achieves proper coverage for sine and cosine signals.
 9.9 Numerical Error Resonance Using Library Sine
+Figure 46: FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
+The legend distinguishes between the uncertainty and the value error.
+The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
+Figure 47: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation.
+The legend distinguishes between the uncertainty and the value error.
+The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
 Figure 48: Error deviation (z-axis) of FFT forward transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
@@
 Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy.
-In the forward transformation, although there is no visible resonant pattern in FFT spectrum, the intensity of value errors increase with FFT order.
-Figure 44 and 45 demonstrate that the error deviations increase with sine or cosine frequency.
-In contrast, this increase is completely absent when using the Quart sine function, as seen in Figure 36 and 37.
-Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 40 and 41.
+For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 43.
+In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order.
+In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 47.
+Also, it is completely absent when using the Quart sine function, as seen in Figure 36 and 37.
+Figure 48 and Figure 49 demonstrate that the error deviations increase with sine or cosine frequency.
+Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 42, 43, and 46.
 At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
@@
 Figure 50 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage.
-Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases much more rapidly with FFT order than its counterpart in Figure 38.
-This increase enables the proper coverage in Figure 46, resulting in error deviations similar to those Figure 38.
-The normalized error histogram of the reverse transformation in Figure 49 is narrower than that of Figure 43, confirming Figure 46.
+Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 38.
+Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized value errors in Figure 51 versus the corresponding histograms in Figure 39.
 Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
@@
 As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$.
-The key question is whether variance arithmetic can effectively track these additional errors.
+The key question is whether variance arithmetic can effectively track these additional unspecified errors.
 Figure 52 shows that proper coverage can no longer be achieved, as the value error increases faster than the uncertainty with rising FFT order.
@@
 Figure 59: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Quart sine function.
-Adding sufficient noise to the input can suppress unspecified input errors and thereby achieve ideal coverage.
-For example, applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage.
-Figure 50 illustrates the error deviations and uncertainty means:
+Code Signal Sine function Forward Reverse Roundtrip
+Python Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
+Java, C++ Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
+Python Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-12}$
+Java, C++ Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-13}$
+Python Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
+Java, C++ Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
+Python Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
+Java, C++ Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
+Table 2: The measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
+Adding noise to the input can suppress unspecified input errors and improve the portion of specified versus unspecified input uncertainties.
+Table 2 shows the measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
+Figure 54 illustrates the error deviations and uncertainty means when applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage for all the transformations:
   • As expected, the result uncertainty means for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
-  • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $\sqrt{1/2}^{L}$.
+  • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
   • As expected, the result uncertainty means for the roundtrip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
-  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$, whereas those for the roundtrip transformation decay exponentially toward 0 as the FFT order increases.
+  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$.
 Figure 55 presents the corresponding histogram.
@@
 The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise.
-For Linear signals computed using the Library sine function, Figure 52 and 53 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
+For Linear signals computed using the Library sine function, Figure 56 and 57 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
 Ideal coverage corresponds to the region where the error deviation equals $1$.
-Outside this region, proper coverage cannot be achieved.
+Outside this region, proper coverage cannot be achieved for the reverse transformation.
 Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region.
@@
 Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
-In contrast, for Linear signals computed using the Quart sine functions, Figure 54 and 55 show the results for the forward and reverse transformations, respectively.
+In contrast, for Linear signals computed using the Quart sine functions, Figure 58 and 59 show the results for the forward and reverse transformations, respectively.
 The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions.
-The forward transformations exhibits a broader ideal coverage region than the reverse transformations.
+The forward transformation exhibits a broader ideal coverage region than the reverse transformations.
+While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region.
 As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or the Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders.
 The key difference is that when using the Quart sine functions, the error deviations within the proper coverage region differ from $1$ only marginally.
+Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$.
+For example, Figure 58 and 59 are almost identical using either C++, or Java, or Python.
+9.13 Input Uncertainty Recovery
+Figure 60: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
+Figure 61: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
+With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 38, 40, 44, and 50, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 52.
+With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
+With input uncertainty specified, roundtrip transformation can always recover input uncertainty numerically, and the quality of the recover increases with the quality of input uncertainty specification.
+For example, when using the Library sine function:
+  • Figure 54 shows that the error deviations for the roundtrip transformation is less than $1$, suggesting that each input uncertainty is recovered numerically.
+These error deviations increases with FFT orders, because the error deviations without added noises increases with the FFT orders according to Figure 52.
+  • As a comparison, for Sin/Cos signal, because Figure 42 shows little increase with the FFT orders of the error deviations without added noises, the improvement of the numerical recovering in Figure 61 is almost independent of FFT orders.
+In both cases, error deviations decreases with the strength of the added noise.
 9.14 Prec Sine Function
@@
 Figure 65: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for reverse transformations of Linear signals computed using the the Prec sine function.
-Signal vs Sine Transformation Quart Prec Library
-Sin/Cos Forward $10^{-15}$ $10^{-15}$ $10^{-14}$
-Sin/Cos Reverse $10^{-12}$ $10^{-12}$ $10^{-11}$
-Linear Forward $10^{-9}$ $10^{-9}$ $10^{-6}$
-Linear Reverse $10^{-7}$ $10^{-7}$ $10^{-3}$
-Table 2: The minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different sine functions.
-The Quart sine function tends to overestimate the uncertainty, while the Library sine function underestimates it.
-This raises the question of whether a sine function with a more accurate uncertainty can be developed.
+The Quart sine function slightly overestimates the uncertainty, while the Library sine function underestimates it significantly.
+This raises the question of whether a sine function with a more accurate uncertainty quantification can be developed.
 If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation.
 The resulting indexed sine function is referred to as the Prec sine function.
-Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare the Prec and the Quart sine functions, Figure 56 shows that the the Prec sine function produces slightly larger value errors, while having half uncertainties.
-Figure 57 and 58 demonstrate that the the Prec sine function slightly underestimates uncertainties, maintaining a stable error deviation around $1.8$ for both forward and reverse transformations without added noise.
-Compared with Figure 55, Figure 59 shows that the ideal coverage achieved using the the Prec sine function is identical to that obtained with the Quart sine function.
-The observation that the 128-bit sine function provides no noticeable improvement over the 64-bit sine function raises concerns about the statistical legitimacy of directly using the difference as the uncertainty, the accuracy of the gcc 128-bit numerical library, or the algorithm used to convert a 128-bit floating-point values to 64-bit.
-Table 2 compares using different sine functions for Sin/Cos or Linear signals.
-Both Quart and the the Prec sine function are not ideal.
-An ideal sine function should have error deviations closer to $1$ without added noise.
+Figure 62 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting.
+This difference indicates that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic, and the Prec sine function is expected to under-perform the Quart sine function due to under-estimation of the value errors by the uncertainties.
+For Sin/Cos signals, Figure 63 is very similar Figure 38, provided that the error deviations using Prec sine functions are about $1.7$ times larger than their counterparts using the Quart sine functions.
+For Linear signals, when comparing Figure 64 and 51 for the normalized value error distributions:
+  • For the forward transformations, using either Prec or Quart sine function has very similar distributions.
+  • For the reverse transformations, using Prec sine function results in wider distributions that using Quart sine function.
+Although the ideal coverage regions for the reverse transformations remains identical in Figure 65 when compared with Figure 59, using Prec sine function results in larger error deviations in the proper coverage regions.
+Thus, using Prec sine function is slightly inferior than using Quart sine function.
+In another word, casting values from higher floating-point representation to a lower floating-point arithmetic is not necessarily superior.
+9.15 Calculation Response
+Figure 66: The FFT execution time (y-axis) versus the FFT order (x-axis) for the implementation using either C++, Java, or Python (legend), with each calculation consisted of $36$ roundtrip FFT transformations.
+The C++ executable is complied with -O3 -march=native optimization.
+Without such optimization, the C++ -O0 implementation is about a order of magnitude slower than the Java implementation.
+Figure 66 plots the execution time for the roundtrip FFT calculations using either C++, Java, or Python.
+As expected, in Figure 66, the execution time increases exponentially with the FFT order $L$ as $2^{L}$ for all the three implementations when the FFT order $L>10$, validating the calculation response.
+At first, it is surprising that in Figure 66 the Java implementation is 3 times faster than the C++ implementation which has been fully optimized for speed.
+Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer.
+Perhaps it is necessary to rethink multi-threading as a way to manually increase the time performance, because the dynamic just-in-time optimization of Java is superior to the static optimization of C++ for repeated and heavy computations with preload data such as FFT.
 9.16 Summary
@@
 These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals.
-Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the specific implementation of the mathematical library.
-This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations.
+Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the numerical errors in library functions.
+This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations, but needs rigorous analysis for the ideal or proper coverage regions in each parameter space.
 The impact of numerical errors within mathematical libraries has not received sufficient attention.
@@
 For example, the uncertainty of a 64-bit value can not directly determined by its difference with the corresponding 128-bit value,
-Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions. Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized.
+Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions.
+Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized.
+Perhaps the asymptotic behavior implies a generic new requirements on the probability distributions.
+The choice of ideal bounding range needs further study.
+For example, the choice of $\kappa_{s}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only.
+Such choice certainly can not be applied to other distributions.
 Cases where $\delta^{2}f<\widehat{\delta^{2}}f$ due to low sample count require further theoretical study and experimental verification.
+The dependence of the convergence of $\log(x\pm\delta x)$ and $(x\pm\delta x)^{c}$ on the chose of ideal bounding range $\kappa$ needs further discussion.
 Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step.
-These results can validate variance arithmetic, and provide performance enhancement to variance arithmetic when the convergence of an analytic expression has to be deduced numerically from the full expansion each time.
 The performance of variance arithmetic must be improved for broader practical adoption.
@@
 This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies.
+For example, Formula (2.50) can guide how to improve the ideal ratio of $x\pm y$.
 A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors.
-Once the mechanism of rounding error is thoroughly understood, developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
+Except for integer multiplication, variance arithmetic does not track rounding errors during computation, because tracking rounding errors for every operation is too expensive to achieve in software.
+The sensitivity to the library functions increases the difficulty to solve these types of problems.
+Developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
 Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic.
@@
 Establishing theoretical foundation for applying statistical Taylor expansion in the absence of a closed-form analytic solution, or when only limited low-order numerical derivatives are available, as in solving differential equations, remains an important direction for future research.
+12 Statements and Declarations
 12.1 Acknowledgments
@@
 Rolph Baker Kearfott.
-12 Statements and Declarations
-Competing Interests: The authors have no competing interests to declare that are relevant to the content of this article.
-13 Data Availability Statement
-The datasets used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
+12.2 Data Availability Statement
+The data set used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
 The execution assistance and explanation of the above code are available from the author upon reasonable request.
+12.3 Competing Interests
+The author has no competing interests to declare that are relevant to the content of this article.
+12.4 Founding
+No funding was received from any organization or agency in support of this research.

### v14

@@
 Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 5 should be used as $\kappa$ in Formula (2.7).
-$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges for the uniform and Gaussian distributions respectively.
+$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges $\hat{\kappa}$ for the uniform and Gaussian distributions respectively.
 2.10 Ideal Statistics
@@
 Therefore, the linear spectrum in Formula (9.3) may contain significant numerical errors when computed using library sine functions.
-For both sine functions, the uncertainty of each $\sin(x)$ is assumed to equal its ULP, which is displayed in Figure 34.
-Because the Quart sine function has minimal Taylor expansion error due to its minimal range of $x$ in calls to the library $\sin(x)$, its true numerical errors are likely smaller than ULP, therefore its uncertainty is sightly overestimated.
-In contrast, Figure 34 indicates that the Library sine function exhibits underestimated uncertainties.
+9.6 Refine Quart Sine Function
+Figure 36: Value errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) versus $x$ (x-axis) and $L$ computed using either the Prec or the Quart sine function (legend).
+Figure 37: The value error histogram for the Quart sine function when it is calculated using 128-bit floating-point arithmetic and cast to 64-bit floating-point representation.
+If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the resulting indexed sine function is referred to as the Prec sine function.
+  • Figure 36 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting.
+As suggested in Figure 36, the Prec sine function has about $5\%$ more non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors, and $14\%$ non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors exceeding the corresponding uncertainties.
+These difference indicate that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic.
+  • Figure 37 demonstrates that the value errors of the Prec sine function after normalized by the corresponding value ULP is uniform distributed with in the half bit of ULP.
+Because variance arithmetic assumes that the value error is uniformly distributed within ULP of each value, the Quart sine function overestimates value errors by one times.
 9.7 Using Quart Sine for Sin/Cos Signals
@@
   • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions.
-This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and $4$.
+This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 40 and 46.
   • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 34.
@@
 In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 49.
-Also, it is completely absent when using the Quart sine function, as seen in Figure 36 and 37.
+Also, it is completely absent when using the Quart sine function, as seen in Figure 38 and 39.
 Figure 50 and Figure 51 demonstrate that the error deviations increase with sine or cosine frequency.
-Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 42, 43, and 46.
+Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 44, 45, and 48.
 At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
@@
 Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 40.
-Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized value errors in Figure 51 versus the corresponding histograms in Figure 39.
+Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized errors in Figure 53 versus the corresponding histograms in Figure 41.
 Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
@@
 The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise.
-For Linear signals computed using the Library sine function, Figure 56 and 57 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
+For Linear signals computed using the Library sine function, Figure 58 and 59 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
 Ideal coverage corresponds to the region where the error deviation equals $1$.
@@
 Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
-In contrast, for Linear signals computed using the Quart sine functions, Figure 58 and 59 show the results for the forward and reverse transformations, respectively.
+In contrast, for Linear signals computed using the Quart sine functions, Figure 60 and 61 show the results for the forward and reverse transformations, respectively.
 The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions.
@@
 While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region.
-As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or the Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders.
-The key difference is that when using the Quart sine functions, the error deviations within the proper coverage region differ from $1$ only marginally.
+Figure 61 shows that the rounding errors balances the over-estimation of value errors in the Quart sine function in Figure 37, so that the error deviations within the proper coverage region differ from $1$ only marginally.
 Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$.
-For example, Figure 58 and 59 are almost identical using either C++, or Java, or Python.
+For example, Figure 60 and 61 are almost identical using either C++, or Java, or Python.
 9.14 Input Uncertainty Recovery
@@
 Figure 63: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
-With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 38, 40, 44, and 50, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 52.
+With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 40, 42, 46, and 52, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 54.
 With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
@@
 In both cases, error deviations decreases with the strength of the added noise.
-9.14 Prec Sine Function
-Figure 62: Value errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) versus $x$ (x-axis) and $L$ computed using either the Prec or the Quart sine function (legend).
-Figure 63: Error deviations (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT orders (x-axis) and transformation types (legend) computed using the the Prec sine function.
-Figure 64: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals computed using the the Prec sine function.
-Figure 65: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for reverse transformations of Linear signals computed using the the Prec sine function.
-The Quart sine function slightly overestimates the uncertainty, while the Library sine function underestimates it significantly.
-This raises the question of whether a sine function with a more accurate uncertainty quantification can be developed.
-If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation.
-The resulting indexed sine function is referred to as the Prec sine function.
-Figure 62 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting.
-This difference indicates that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic, and the Prec sine function is expected to under-perform the Quart sine function due to under-estimation of the value errors by the uncertainties.
-For Sin/Cos signals, Figure 63 is very similar Figure 38, provided that the error deviations using Prec sine functions are about $1.7$ times larger than their counterparts using the Quart sine functions.
-For Linear signals, when comparing Figure 64 and 51 for the normalized value error distributions:
-  • For the forward transformations, using either Prec or Quart sine function has very similar distributions.
-  • For the reverse transformations, using Prec sine function results in wider distributions that using Quart sine function.
-Although the ideal coverage regions for the reverse transformations remains identical in Figure 65 when compared with Figure 59, using Prec sine function results in larger error deviations in the proper coverage regions.
-Thus, using Prec sine function is slightly inferior than using Quart sine function.
-In another word, casting values from higher floating-point representation to a lower floating-point arithmetic is not necessarily superior.
 9.15 Calculation Response
@@
 As expected, in Figure 64, the execution time increases exponentially with the FFT order $L$ as $2^{L}$ for all the three implementations when the FFT order $L>10$, validating the calculation response.
-At first, it is surprising that in Figure 66 the Java implementation is 3 times faster than the C++ implementation which has been fully optimized for speed.
+At first, it is surprising that in Figure 64 the Java implementation is 2 times faster than the C++ implementation which has been fully optimized for speed.
 Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer.
@@
 Perhaps the asymptotic behavior implies a generic new requirements on the probability distributions.
-The choice of ideal bounding range needs further study.
-For example, the choice of $\kappa_{s}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only.
+The choice of ideal bounding range $\hat{\kappa}$ needs further study.
+For example, the choice of $\hat{\kappa}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only.
 Such choice certainly can not be applied to other distributions.
+Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding.
 Cases where $\delta^{2}f<\widehat{\delta^{2}}f$ due to low sample count require further theoretical study and experimental verification.

### v15

@@
 […]
-$\zeta(n)\equiv$ $\int_{\mu-\varrho\sigma}^{\mu+\kappa\sigma}\tilde{x}^{n}\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
+$\zeta(n)\equiv$ $\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
 […]
@@
 Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
-2.4 Bounding Symmetry
+2.4 Bound Moment Symmetry
 A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbounded distribution.
@@
 For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
-Both Gaussian and Uniform distributions are symmetric.
-2.5 Bounding Asymptote
-Empirically, Formula (2.18), (2.22), and (2.25) demonstrate that as $2n\rightarrow+\infty$, $\zeta(2n)\rightarrow\kappa^{2n}/(2n)$.
-2.5.1 Uniform Input Uncertainty
-For uniform distribution, the bounding range is $\kappa=\sqrt{3}$, and the bounding leakage $\epsilon=0$. Formula (2.18) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
-$\zeta(2n)=\int_{-\sqrt{3}}^{+\sqrt{3}}\frac{1}{2\sqrt{3}}\tilde{z}^{2n}d\tilde{z}=\frac{(\sqrt{3})^{2n}}{2n+1};$ (2.18)
-2.5.2 Gaussian Input Uncertainty
-The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [4].
-This convergence occurs rapidly [29].
-In digital computation, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sums of expansion terms [7][8].
-Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [1][2][4].
-Formula (2.7) reduces to Formula (2.19) and (2.20):
-$\zeta(2n,\kappa)=$ $\;(2n-1)!!\left(\xi(\frac{\kappa}{\sqrt{2}})-2N(\kappa)\sum_{j=0}^{n-1}\frac{\kappa^{2j+1}}{(2j+1)!!}\right);$ (2.19)
-$=$  $\;2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.20)
-$=$ $\;(2n-1)\zeta(2n-2,\kappa)-2N(\kappa)\kappa^{2n-1};$ (2.21)
-$\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)\simeq 2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.22)
-  • For small $2n$, $\zeta(2n)$ can be approximated by $\zeta(2n)=(2n-1)!!$ according to Formula (2.21).
-When $\kappa=5$, and $n<5$, the relative error $\left|\zeta(2n)/(2n-1)!!-1\right|$ is less than $10^{-3}$.
-  • For large $2n$, Formula (2.20) reduces to Formula (2.22), showing that $\zeta(2n)$ increases more slowly than $\kappa^{2n}$ as $2n$ grows.
-2.5.3 An Input Uncertainty with Limited Range
-$\rho(\tilde{x},\mu,\sigma)$ $=\frac{\tilde{x}}{\lambda^{2}}e^{-\frac{\tilde{x}}{\lambda}};\;\;\;\mu=2\lambda;\;\;\;\sigma=\sqrt{2}\lambda;$ (2.23)
-$e^{-\varrho}\varrho^{2}$ $=e^{-\kappa}\kappa^{2};$ (2.24)
-$\lim_{n\rightarrow+\infty}\zeta(n,\kappa)$ $=\frac{\frac{\kappa}{\sqrt{2}}^{n}}{n+2}e^{-\kappa};$ (2.25)
-Formula (2.23) shows a probability density function with $\tilde{x}\in[0,+\infty)$, whose bounding range $\varrho\lambda,\kappa\lambda]$ satisfying $0<\varrho<2<\kappa$.
-Formula (2.24) gives its mean preserving equation.
-And Formula (2.25) describes its asymptotic behavior.
 2.5 One-Dimensional Examples
@@
 $\frac{\delta^{2}1/x}{(1/x)^{2}}\simeq$ $\;P(x)^{2}+8P(x)^{4}+69P(x)^{6}+o(P(x)^{8});$ (2.36)
-2.8 Convergence
+2.7 Asymptotic Behavior for Bound Moment
+Formula (2.38) is a special case of Formula (2.37) for symmetric probability distribution.
+Empirically, Formula (2.38) is confirmed by Formula (2.40), (2.42), and (2.43), and Formula (2.37) is confirmed by Formula (2.45) and (2.51).
+$\kappa\ll n$ $:\;\;\;\zeta(n,\kappa)\simeq\rho(\kappa)\frac{\kappa^{n}}{n};$ (2.37)
+[…]
+2.7.1 Gaussian Input Uncertainty
+Figure 5: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus Formula (2.41) (left y-axis), and their ratio (right y-axis).
+The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [4].
+This convergence occurs rapidly [29].
+In digital computation, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sums of expansion terms [7][8].
+Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [1][2][4].
+With Formula (2.39) [Double_Factorial], for Gaussian distribution, Formula (2.40) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, while Formula (2.41) demonstrates its asymptotic behavior.
+Figure 5 shows the transition to Formula (2.41) with increasing order $2n$, in which the ratio is the ratio of Formula (2.41) when $\kappa=5$ to $\zeta(2n,5)$.
+$\sum_{j=0}^{+\infty}\frac{\kappa^{2j+1}}{(2j+1)!!}=\sqrt{\frac{\pi}{2}}\xi(\frac{\kappa}{\sqrt{2}})e^{\frac{1}{2}\kappa^{2}};$ (2.39)
+$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}N(\tilde{x})d\tilde{x}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.40)
+$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
+2.7.2 Uniform Input Uncertainty
+As shown in Figure 38, rounding errors can be uniformly distributed [29].
+For uniform distribution, the bounding range $\kappa$ is less than the maximal range $\sqrt{3}$. Formula (2.42) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
+$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2\sqrt{3}}d\tilde{x}=2\frac{1}{2\sqrt{3}}\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.42)
+2.7.3 Laplace Input Uncertainty
+For Laplace distribution [4], Formula (2.43) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, and Formula (2.44) demonstrates its asymptotic behavior.
+$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2}e^{-|\tilde{x}|}d\tilde{x}=e^{-\kappa}\kappa^{2n}\sum_{\kappa=1}^{\infty}\kappa^{j}\frac{(2n)!}{(2n+j)!};$ (2.43)
+$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2\frac{1}{2}e^{-\kappa}\frac{\kappa^{2n+1}}{2n+1};$ (2.44)
+2.7.4 Exponential Input Uncertainty
+An exponential distribution [4] is no longer symmetric.
+Formula (2.45) provides $\zeta(2n,\kappa)$, Formula (2.46) shows the equation for the mean preserving condition, and Formula (2.47) demonstrates its asymptotic behavior.
+$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}(\tilde{x}-1)^{n}e^{-\tilde{x}}d\tilde{x}$ (2.45)
+$=e^{-\kappa}(\kappa-1)^{n}\sum_{j=1}^{\infty}(\kappa-1)^{j}\frac{n!}{(n+j)!}-e^{-\varrho}(\varrho-1)^{n}\sum_{j=1}^{\infty}(\varrho-1)^{j}\frac{n!}{(n+j)!};$
+$\varrho e^{-\varrho}=\kappa e^{-\kappa}\Longrightarrow 0<\varrho<1<\kappa;$ (2.46)
+$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=e^{-\kappa}\frac{\kappa^{n+1}}{n+1};$ (2.47)
+2.7.5 Simplified Gamma Input Uncertainty
+The gamma distribution [4] $\rho(\tilde{x})=x^{\alpha-1}e^{-x/\lambda}\lambda^{\alpha}/\Gamma(\alpha),\tilde{x},\alpha,\lambda>0$ is simplified as Formula (2.48), whose mean and deviation are $m+1$ and $\sqrt{m+1}$, respectively.
+Formula (2.49) provides $\zeta(2n,\kappa)$, Formula (2.50) shows the equation for the mean preserving condition, and Formula (2.51) demonstrates its asymptotic behavior.
+[…]
+$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}\left(\frac{\tilde{x}-(m+1)}{\sqrt{m+1}}\right)^{n}\rho(\tilde{x})d\tilde{x}$ (2.49)
+$e^{-\varrho}\varrho^{m+1}=e^{-\kappa}\kappa^{m+1};\Longrightarrow 0<\varrho<m+1<\kappa;$ (2.50)
+$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=\frac{1}{m!}e^{-\kappa}\frac{\kappa^{m+n+1}}{m+n+1};$ (2.51)
+2.8 Convergence of Variance
+When $n\rightarrow\infty$, $\zeta(j,\kappa)\zeta(n-j,\kappa)$ will be ignored in Formula (2.3) because $\zeta(j,\kappa)\zeta(n-j,\kappa)\ll\zeta(n,\kappa)$.
+The asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
 Formula (2.20) for $e^{x\pm\delta x}$ and Formula (2.26) for $\sin(x\pm\delta x)$ both converge unconditionally.
@@
 Ideal variances are consistent with traditional variances [4].
-In both Figure 7 and 8, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
+In both Figure 8 and 9, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
 The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$.
@@
 An ideal ratio also applies to the corresponding uncertainty bias.
-The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 7 and 8, except that $\overline{x}-x=0$.
+The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 8 and 9, except that $\overline{x}-x=0$.
 This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
@@
 The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms.
-The upper bounds of input uncertainty in Figure 9 and 10 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 9 are $3$ order-of-magnitude smaller than those in Figure 10 when the exponent $c$ is less than 0.
+The upper bounds of input uncertainty in Figure 10 and 11 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 10 are $3$ order-of-magnitude smaller than those in Figure 11 when the exponent $c$ is less than 0.
 Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
@@
   • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions.
-This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 40 and 46.
+This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 41 and 47.
   • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 35.
@@
 In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 50.
-Also, it is completely absent when using the Quart sine function, as seen in Figure 38 and 39.
+Also, it is completely absent when using the Quart sine function, as seen in Figure 39 and 40.
 Figure 51 and Figure 52 demonstrate that the error deviations increase with sine or cosine frequency.
-Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 44, 45, and 48.
+Figure 35 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 45, 46, and 49.
 At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
@@
 Figure 56: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
-As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$.
+As shown by the difference between Figure 35 and 36, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$.
 The key question is whether variance arithmetic can effectively track these additional unspecified errors.
@@
 The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise.
-For Linear signals computed using the Library sine function, Figure 58 and 59 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
+For Linear signals computed using the Library sine function, Figure 59 and 60 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
 Ideal coverage corresponds to the region where the error deviation equals $1$.
@@
 Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
-In contrast, for Linear signals computed using the Quart sine functions, Figure 60 and 61 show the results for the forward and reverse transformations, respectively.
+In contrast, for Linear signals computed using the Quart sine functions, Figure 61 and 62 show the results for the forward and reverse transformations, respectively.
 The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions.
@@
 Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$.
-For example, Figure 60 and 61 are almost identical using either C++, or Java, or Python.
+For example, Figure 61 and 62 are almost identical using either C++, or Java, or Python.
 9.14 Input Uncertainty Recovery
@@
 Figure 64: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
-With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 40, 42, 46, and 52, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 54.
+With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 41, 43, 47, and 53, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 55.
 With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
@@
 Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions.
-Its asymptotic behavior $2n\rightarrow+\infty:\zeta(2n)\rightarrow\kappa^{2n}/(2n)$ needs to be generalized.
-Perhaps the asymptotic behavior implies a generic new requirements on the probability distributions.
+The necessary and sufficient conditions for Formula (2.37) need to be found, to extend the applicable types of distributions for the statistical Taylor expansion from the most common ones as enumerated in Section 2.
+The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.
 The choice of ideal bounding range $\hat{\kappa}$ needs further study.

### v16

@@
 2.7 Asymptotic Behavior for Bound Moment
-Formula (2.38) is a special case of Formula (2.37) for symmetric probability distribution.
-Empirically, Formula (2.38) is confirmed by Formula (2.40), (2.42), and (2.43), and Formula (2.37) is confirmed by Formula (2.45) and (2.51).
-$\kappa\ll n$ $:\;\;\;\zeta(n,\kappa)\simeq\rho(\kappa)\frac{\kappa^{n}}{n};$ (2.37)
+When $\rho(\tilde{z})>0$ and $\frac{d\rho(\tilde{z})}{d\tilde{z}}$ is bounded in $\tilde{z}\in[0,\kappa]$, Formula (2.38) and (2.39) leads to Formula (2.40).
+Formula (2.37) and (2.40) conclude Formula (2.41) and (2.42).
+$\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$ $=\rho(\kappa)\frac{\kappa^{n+1}}{n+1}-\frac{1}{n+1}\int_{0}^{\kappa}\tilde{z}^{n+1}\frac{d\rho(\tilde{z})}{d\tilde{z}}d\tilde{z}$; (2.37)
 […]
+[…]
+$\lim_{n\rightarrow+\infty}$ $\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}=\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.40)
+$|\varrho|=\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
+$|\varrho|<\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(n,\kappa)=\rho(\varrho)\frac{\varrho^{n+1}}{n+1}+\rho(\kappa)\frac{\kappa^{n+1}}{n+1}\simeq\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.42)
 2.7.1 Gaussian Input Uncertainty
@@
 $\sum_{j=0}^{+\infty}\frac{\kappa^{2j+1}}{(2j+1)!!}=\sqrt{\frac{\pi}{2}}\xi(\frac{\kappa}{\sqrt{2}})e^{\frac{1}{2}\kappa^{2}};$ (2.43)
-$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}N(\tilde{x})d\tilde{x}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.40)
-$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
+$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}N(\tilde{z})d\tilde{z}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.44)
+$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.45)
 2.7.2 Uniform Input Uncertainty
@@
 For uniform distribution, the bounding range $\kappa$ is less than the maximal range $\sqrt{3}$. Formula (2.46) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
-$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2\sqrt{3}}d\tilde{x}=2\frac{1}{2\sqrt{3}}\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.42)
-2.7.3 Laplace Input Uncertainty
-For Laplace distribution [4], Formula (2.43) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, and Formula (2.44) demonstrates its asymptotic behavior.
-$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{x}^{2n}\frac{1}{2}e^{-|\tilde{x}|}d\tilde{x}=e^{-\kappa}\kappa^{2n}\sum_{\kappa=1}^{\infty}\kappa^{j}\frac{(2n)!}{(2n+j)!};$ (2.43)
-$1<\kappa^{2}\ll 2n:\;\;\;\zeta(2n,\kappa)=2\frac{1}{2}e^{-\kappa}\frac{\kappa^{2n+1}}{2n+1};$ (2.44)
-2.7.4 Exponential Input Uncertainty
-An exponential distribution [4] is no longer symmetric.
-Formula (2.45) provides $\zeta(2n,\kappa)$, Formula (2.46) shows the equation for the mean preserving condition, and Formula (2.47) demonstrates its asymptotic behavior.
-$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}(\tilde{x}-1)^{n}e^{-\tilde{x}}d\tilde{x}$ (2.45)
-$=e^{-\kappa}(\kappa-1)^{n}\sum_{j=1}^{\infty}(\kappa-1)^{j}\frac{n!}{(n+j)!}-e^{-\varrho}(\varrho-1)^{n}\sum_{j=1}^{\infty}(\varrho-1)^{j}\frac{n!}{(n+j)!};$
-$\varrho e^{-\varrho}=\kappa e^{-\kappa}\Longrightarrow 0<\varrho<1<\kappa;$ (2.46)
-$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=e^{-\kappa}\frac{\kappa^{n+1}}{n+1};$ (2.47)
-2.7.5 Simplified Gamma Input Uncertainty
-The gamma distribution [4] $\rho(\tilde{x})=x^{\alpha-1}e^{-x/\lambda}\lambda^{\alpha}/\Gamma(\alpha),\tilde{x},\alpha,\lambda>0$ is simplified as Formula (2.48), whose mean and deviation are $m+1$ and $\sqrt{m+1}$, respectively.
-Formula (2.49) provides $\zeta(2n,\kappa)$, Formula (2.50) shows the equation for the mean preserving condition, and Formula (2.51) demonstrates its asymptotic behavior.
-[…]
-$\zeta(n,\kappa)$ $=\int_{\varrho}^{\kappa}\left(\frac{\tilde{x}-(m+1)}{\sqrt{m+1}}\right)^{n}\rho(\tilde{x})d\tilde{x}$ (2.49)
-$e^{-\varrho}\varrho^{m+1}=e^{-\kappa}\kappa^{m+1};\Longrightarrow 0<\varrho<m+1<\kappa;$ (2.50)
-$1<\kappa\ll n:\;\;\;\zeta(n,\kappa)=\frac{1}{m!}e^{-\kappa}\frac{\kappa^{m+n+1}}{m+n+1};$ (2.51)
+$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.46)
 2.8 Convergence of Variance
-When $n\rightarrow\infty$, $\zeta(j,\kappa)\zeta(n-j,\kappa)$ will be ignored in Formula (2.3) because $\zeta(j,\kappa)\zeta(n-j,\kappa)\ll\zeta(n,\kappa)$.
-The asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
+$\lim_{n\rightarrow+\infty}$ $\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.47)
+$\delta^{2}\log(x\pm\delta x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}$
+$\simeq$  $\;2\nu(\kappa)\log(2)\sum_{n=1}^{+\infty}\frac{(P(x)\kappa)^{2n}}{(2n)^{2}},\begin{cases}\text{Gaussian}:\nu(\kappa)=N(\kappa)\kappa\\ \text{Uniform}:\nu(\kappa)=1,\;\;\;\kappa=\sqrt{3}\end{cases};$ (2.48)
+$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.49)
+For input distribution without bounding, Formula (2.3) may not converge, e.g., when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$.
+Thus, it is necessary that the input uncertainty is bounded.
+Formula (2.47) shows that the asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
 Formula (2.20) for $e^{x\pm\delta x}$ and Formula (2.26) for $\sin(x\pm\delta x)$ both converge unconditionally.
@@
 Formula (2.23) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) as $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$.
-$\delta^{2}\log(x\pm\delta x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}$
-$\simeq$  $\;2\nu(\kappa)\log(2)\sum_{n=1}^{+\infty}\frac{(P(x)\kappa)^{2n}}{(2n)^{2}},\begin{cases}\text{Gaussian}:\nu(\kappa)=N(\kappa)\kappa\\ \text{Uniform}:\nu(\kappa)=1,\;\;\;\kappa=\sqrt{3}\end{cases};$ (2.52)
 Formula (2.29) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.49) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ .
 This expression converges when $P(x)\lesssim 1/\kappa$ although the precis upper bound for $P(x)$ varies with $c$.
-$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.53)
 2.9 Statistical Bounding
@@
 Because $\epsilon(\kappa)\neq\epsilon(\kappa,N)$ for finite $N$, let $\epsilon(\kappa)=\epsilon(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bonding range, and $\kappa(\kappa_{s},N)$ is the measured bounding range.
+$\epsilon(\kappa)=$  $\;1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.50)
+$\epsilon(\kappa_{s},N)=$  $\;1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.51)
+When the underlying distribution is Normal, Formula (2.50) and (2.51) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
+Figure 6 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction.
+It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
+Figure 7 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$).
+Figure 7 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$.
+Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 6 should be used as $\kappa$ in Formula (2.7).
 When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to bounding leakage $\epsilon(N)$.
@@
 The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.7).
-$\epsilon(\kappa)=$  $\;1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.54)
-$\epsilon(\kappa_{s},N)=$  $\;1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.55)
-When the underlying distribution is Normal, Formula (2.54) and (2.55) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
-Figure 6 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$.
-It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$.
-Figure 7 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$).
-Figure 7 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$.
-Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 6 should be used as $\kappa$ in Formula (2.7).
-$\kappa_{s}=\sqrt{3}$ and $\kappa_{s}=5$ are defined as the ideal bounding ranges $\hat{\kappa}$ for the uniform and Gaussian distributions respectively.
+[…]
 2.10 Ideal Statistics
@@
 Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input.
-Ideal variances are consistent with traditional variances [4].
 In both Figure 8 and 9, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
@@
 This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
-Statistical Taylor expansion outputs ideal uncertainty, ideal uncertainty bias, and ideal ratio for the chosen ideal bounding ranges, in which the ideal ratio quantifies the reliability of the ideal uncertainty and ideal uncertainty bias.
+In statistical Taylor expansion, every calculation should be carried out twice:
+  1. 1.
+[…]
+  2. 2.
+    With the real $\zeta(n,\kappa)$ of each input given the sample count and input uncertainty distribution, to obtain the ideal ratio to quantify the reliability of the ideal variance and ideal uncertainty bias as a value in the range of $[0,1)$.
 2.11 Dependency Tracing
@@
 $\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.57)
-$\delta^{2}\left(\frac{f^{(m)}_{x}\tilde{x}^{m}}{m!}+\frac{f^{(n)}_{x}\tilde{x}^{n}}{n!}\right)=$ $\;\sigma^{2m}(\frac{f^{(m)}_{x}}{m!})^{2}\eta(2m)+\sigma^{2n}(\frac{f^{(n)}_{x}}{n!})^{2}\eta(2n)$ (2.62)
-$+2\sigma^{m+n}\left(\frac{f^{(m)}_{x}f^{(n)}_{x}}{m!\;n!}\eta(m+n)-\frac{f^{(m)}_{x}}{m!}\eta(m)\;\frac{f^{(n)}_{x}}{n!}\eta(n)\right);$
 When all inputs satisfy the uncorrelated uncertainty assumption, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
   • Formula (2.54) expresses $\delta^{2}(f+g)$, whose dependency tracing is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15).
-Formula (2.62) shows that Formula (2.3) applies Formula (2.58) between any two terms in the Taylor expansion in Formula (2.8).
+Formula (LABEL:eqn:_Taylor_inner_dependency) shows that Formula (2.3) applies Formula (2.54) between any two terms in the Taylor expansion in Formula (2.8).
   • Formula (2.55) expresses $\delta^{2}(fg)$, illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.17).
@@
 Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy.
-For example, Figure LABEL:fig:_FFT_Cost_Clean_6_6_Wave_Lib shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46.
+For example, Figure 49 shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46.
 In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order.
@@
 Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions.
-The necessary and sufficient conditions for Formula (2.37) need to be found, to extend the applicable types of distributions for the statistical Taylor expansion from the most common ones as enumerated in Section 2.
 The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.
@@
 Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding.
+The condition for Formula (2.47) seems to be as generic as the definition in Formula (2.7) which needs further proof.
 Cases where $\delta^{2}f<\widehat{\delta^{2}}f$ due to low sample count require further theoretical study and experimental verification.

### v17

@@
 Abstract
-As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions to compute the means and standard deviations of the results.
-Statistical Taylor expansion tracks the propagation of the input uncertainties through intermediate steps, causing the variables in intermediate analytic expressions to become interdependent, while the final analytic result becomes path independent.
-This fundamentally distinguishes it from common approaches in applied mathematics that optimize computational path for each calculation.
-In essence, statistical Taylor expansion may standardize numerical computations for analytic expressions.
-Its statistical nature enables reliable testing of results when the sample size is sufficiently large, as well as includes sample count in result qualification.
-This study also introduces an implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
+As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions and sample counts to compute the mean, the standard deviation, and the reliable factor of each result.
+It tracks the propagation of the input uncertainties through intermediate steps, so that the final analytic result becomes path independent.
+Therefore, it differs fundamentally from common approaches in applied mathematics that optimize computational path for each calculation.
+Statistical Taylor expansion may standardize numerical computations for analytic expressions.
+This study also introduces the implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
 Another important conclusion of this study is that numerical errors in library functions can significantly affect results.
-For instance, periodic numerical errors in trigonometric library functions may resonate with periodic signals, producing large numerical errors in the outcomes.
+It is desirable that each value from library functions be accomplished by an uncertainty deviation.
 Keywords: computer arithmetic, error analysis, interval arithmetic, uncertainty, numerical algorithms.
@@
 1 Introduction
-1.1 Measurement Uncertainty
-Except for the most basic counting, scientific and engineering measurements never yield completely precise results [1][2].
-In such measurements, the uncertainty of a quantity $x$ is typically expressed either by the sample deviation $\delta x$ or by the uncertainty range $\Delta x$ [1][2].
-  • If $\delta x=0$ or $\Delta x=0$, $x$ is a precise value.
-  • Otherwise, $x$ is an imprecise value.
-$P(x)\equiv\delta x/|x|$ is defined as the statistical precision (hereafter referred to as precision) of the measurement, where $x$ denotes the value, and $\delta x$ represents the uncertainty deviation.
-A larger precision indicates a coarser measurement whereas a smaller precision indicates a finer measurement.
-The precision of measured values can range from order-of-magnitude estimates to $10^{-2}$ to $10^{-4}$ in common measurements, and to $10^{-15}$ in state-of-the-art determinations of fundamental physical constants [3].
-This study focuses on determining the result value and uncertainty of a general analytic expression with imprecise input values.
-1.2 Extension to Existing Statistics
-Rather than focusing on the uncertainty distribution of the result when applying a function to a random variable with a known uncertainty distribution [4], statistical Taylor expansion provides the mean and variance of the result for general analytic expressions.
+Let $x$ and $\delta x$ be a value and its uncertainty deviation, respectively.
+As an input, $x$ and $\delta x$ are the mean and standard deviation of a measurement [1][2].
+If $\delta x=0$, $x$ is a precise value, otherwise, the pair specifies an imprecise value $x\pm\delta x$.
+Let $P(x)\equiv\delta x/|x|$ be the statistical precision (hereafter referred to as precision) of $x\pm\delta x$.
+A smaller $P(x)$ indicate a higher measurement quality of $x\pm\delta x$, and when $P(x)\ll 1$, it is unclear whether $\delta x$ is too poor to measure $x$, or it is a measurement of $x=0$.
+Statistical Taylor expansion determines the result $f\pm\delta f$ of a general analytic expression $f(x,\dots)$, and the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ based on the inputs $x\pm\delta x,\dots$ and their corresponding sample counts.
   • Previous studies have examined the effect of input uncertainties on output values for specific cases [3].
 Statistical Taylor expansion generalizes these effects as uncertainty bias, as shown in Formula (2.5) and (2.8) in this paper.
-  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper.
-As shown in Formula (7.9), (7.17), and (7.18), such extension is important even for linear algebra.
-1.3 Problem of the Related Numerical Arithmetics
-Variance arithmetic implements statistical Taylor expansion and surpasses all existing related numerical arithmetic methods.
-1.3.1 Conventional Floating-Point Arithmetic
-Conventional floating-point arithmetic [6][7][8] treats every bit of its significand to be valid at all times, and continuously either generates artificial information or erases useful information to maintain this assumption during the normalization process [9].
-For example, the following calculation is performed exactly in integer format:
-$64919224\times 205117922-159018721\times 83739041=\\ 13316075197586562-13316075197586561=1;$ (1.1)
-If Formula (1.1) is carried out using conventional 64-bit floating-point arithmetic:
-$64919224\times 205117922-159018721\times 83739041=\\ 64919224.000000000\times 205117922.000000000-159018721.000000000\times 83739041.000000000=\\ 13316075197586562.-13316075197586560.=2.=2.0000000000000000;$ (1.2)
+  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper.
+  • Calculating the reliability of $f\pm\delta f$ seems completely new.
+Statistical Taylor expansion is superior to all existing numerical arithmetic.
+Conventional floating-point arithmetic [6][7][8] can only calculate the result $f$ which contains unknown amount of rounding error [9][10][11].
+As a consequence, a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision.
+While statistical Taylor expansion can track rounding errors.
+The bounding range in interval arithmetic [12][13][14][15][16] is inconsistent with the statistical nature of an $x\pm\delta x$ pair and it tends to over-estimate the result uncertainty due to its worst-case assumption [17].
+On the other hand, statistical Taylor expansion is precise statistically.
+Both conventional floating-point arithmetic and interval arithmetic depend strongly on the specific algebraic form of an analytic function, a phenomenon known as the dependency problem [10][11][12][14], which makes the numerical calculations of analytic problems sometimes more like an art than science.
+In contrast, statistical Taylor expansion is path independent.
+To be rigorous in mathematics and statistics, statistical Taylor expansion abandons the significance arithmetic nature of its processor [17].
+As a statistical sampling process, stochastic arithmetic [18][19] is computationally expensive, while statistical Taylor expansion provides a direct characterization without sampling.
+In this paper:
   1. 1.
-    The normalization of input values may artificially pad zero bits in the least significant positions until every bit in the 53-bit significand is used, for example, converting $64919224$ to $64919224.000000000$.
+    Section 1 compares briefly statistical Taylor expansion with all other numerical arithmetic.
   2. 2.
-    The results of addition or multiplication may exceed the bit range of the 53-bit significand, requiring rounding off from the least significant bits to fit within the 53-bit significand.
-This process can generate rounding errors, for example, converting $13316075197586561$ to $13316075197586560$, which produces a rounding error of $1$ in the least significant bit.
+    Section 2 develops the theoretical foundation of statistical Taylor expansion.
   3. 3.
-    The results of subtraction or division can cancel values from higher significant bits, exposing rounding errors in the lower significant bits.
-For example, 13316075197586562.13316075197586560.=213316075197586562.-13316075197586560.=2.
+    Section 3 describes variance arithmetic as a numerical implementation of statistical Taylor expansion.
   4. 4.
-    The normalization of subtraction or division results can amplify rounding errors toward the most significant bit by padding zeros, such as converting $2.$ to $2.0000000000000000$.
-Formula (1.2) demonstrates a classic case of catastrophic cancellation [10][11].
-Because rounding errors from lower digits can propagate to higher digits, the $10^{-7}$ significance of the 32-bit IEEE floating-point format [6][7][8] is usually insufficient for calculations involving input data with a precision of $10^{-2}$ to $10^{-4}$.
-For more complex calculations, even the $10^{-16}$ significance of the 64-bit IEEE floating-point format [6][7][8] may be sufficient for inputs with $10^{-2}$ to $10^{-4}$ precision.
-This represents a fundamental controversy of conventional floating-point arithmetic.
-Because rounding error is path dependent, a major objective of conventional numerical methods is to identify optimal computation strategies that minimize rounding errors, such as Gaussian elimination [11][12][13], even though all alternative paths of root finding methods in linear algebra are mathematically equivalent [14].
-Self-censoring rules have been developed to limit such rounding error propagation [12][13], for example, by avoiding the subtraction of results from large multiplications, as shown in Formula (1.2).
-However, these rules are neither enforceable nor easily adoptable, and they are even more difficult to quantify.
-To date, research on rounding error propagation has focused primarily on linear calculations [10][11][15], or on special cases [12][16][17], whereas in practice, rounding errors often manifest as pervasive and mysterious numerical instabilities [18].
-The forward rounding error study [11] compares (1) the result containing rounding error and (2) the ideal result without rounding error, for example, by comparing a result obtained using 32-bit IEEE floating-point arithmetic with the corresponding result computed using 64-bit IEEE floating-point arithmetic [19].
-The most recent study of this type presents an extremely optimistic view of numerical library errors, reporting them as fractions of the least significant bit of the floating-point significand [19].
-However, such optimism contradicts the statistical tests on numerical library functions presented in this paper.
-The backward rounding error study [10][11][15] estimates only the result uncertainty caused by rounding errors, thereby overlooking the bias that rounding errors introduce into the result value.
-This analysis is typically limited to very small uncertainties because it relies on perturbation theory and it is tailored to each specific algorithm [10][11][15].
-In contrast, variance arithmetic traces rounding error directly as part of uncertainty.
-Statistical Taylor expansion applies generally to any analytic function, providing both result mean and deviation, and accommodates input uncertainties of any magnitude.
-By demonstrating that the analytic result should be path independent, statistical Taylor expansion fundamentally challenges the conventional methodology of seeking an optimal execution strategy for a given analytic expression.
-1.3.2 Interval Arithmetic
-Interval arithmetic [13][20][21][22][23][24] is currently a standard method for tracking computational uncertainty.
-Its objective is to ensure that a value remains strictly bounded within its bounding range throughout the computation.
-However, the bounding range in interval arithmetic is not compatible with the approach commonly used in scientific and engineering measurements, which instead characterizes uncertainty in terms of the statistical mean and deviation [1][2]  [^2] There is one attempt to connect intervals in interval arithmetic to confidence interval or the equivalent so-called p-box in statistics [25].
-Because this attempt seems to rely heavily on 1) specific properties of the uncertainty distribution within the interval and/or 2) specific properties of the functions upon which the interval arithmetic is used, this attempt does not seem to be generic.
-If probability model is introduced to interval arithmetic to allow tiny bounding leakage, the bounding range is much less than the corresponding pure bounding range [15].
-Anyway, these attempts seem to be outside the main course of interval arithmetic.
-.
-Interval arithmetic represents only the worst-case scenario of uncertainty propagation.
-For example, in addition, it assumes that the two input variables are perfectly positively correlated [26], thereby producing the widest possible bounding range.
-In contrast, if the variables were perfectly negatively correlated, the bounding range after addition would decrease [27]  [^3] Such case is called the best case in random interval arithmetic.
-The vast overestimation of bounding ranges in these two worst cases prompts the development of affine arithmetic [26][28], which traces error sources using a first-order model.
-Being expensive in execution and depending on approximate modeling even for such basic operations as multiplication and division, affine arithmetic has not been widely used.
-In another approach, random interval arithmetic [27] randomly chooses between the best-case and the worst-case intervals, so that it can no longer guarantee bounding without leakage.
-Anyway, these attempts seem to be outside the main course of interval arithmetic.
-.
-This worst-case assumption can lead to order-of-magnitude over-estimations [29].
-The results of interval arithmetic can depend strongly on the specific algebraic form of an analytic function $f(x)$, a phenomenon known as the dependency problem.
-This issue is amplified in interval arithmetic [23] but also exists in conventional floating-point arithmetic [12].
-Furthermore, interval arithmetic lacks a mechanism to reject invalid calculations, even though every mathematical operation has a valid input range.
-For example, it produces branched results for $1/(x\pm\Delta x)$ or $\sqrt{x\pm\Delta x}$ when $0\in[x-\Delta x,x+\Delta x]$, whereas a context-sensitive uncertainty bearing arithmetic should reject such calculation naturally.
-In contrast, variance arithmetic specifies each value by its mean and deviation.
-It does not suffer from the dependency problem.
-Its statistical framework naturally rejects certain input intervals on mathematical grounds, such as inversion and square root operations when the statistical bounding range includes zero.
-1.3.3 Statistical Propagation of Uncertainty
-Statistical propagation of uncertainty treats each imprecise value as a random variable and seeks correlations among such random variables to calculate resulting mean and deviation [30].
-In contrast, statistical Taylor expansion interprets the uncertainty of an imprecise value as a limitation in obtaining its precise value.
-This assumption aligns with most error analyses in the literature [1][2].
-Its statistical foundation is the uncorrelated uncertainty assumption [29]: any two imprecise values do not correlate in their uncertainties.
-Statistical propagation of uncertainty appears to have applied an incorrect statistical context to uncertainty.
-For example, a time series is a random variable whereas each imprecise value within the time series is merely an imprecise measurement.
-The uncorrelated uncertainty assumption can hold for each imprecise value in a time series, even though the time series can highly correlate to another time series at their signal level, or even correlated to itself as a periodic time series.
-If the uncertainties of some imprecise values are correlated, these imprecise values contain systematic errors among them [1].
-Systematic errors should be treated as unwanted signals but not as normal uncertainties [1].
-In short, uncertainty should not be treated as signal.
-1.3.4 Significance Arithmetic
-Significance arithmetic [31] seeks to track the number of reliable bits in an imprecise value throughout a calculation.
-In its early implementations [32][33], significance arithmetic was based on simple operational rules applied to reliable bit counts rather than on formal statistical methods.
-In these approaches, the reliable bit count is treated as an integer, even though in practice it can take a fractional value [34].
-This constraint can produce an artificial step-wise reduction in significance.
-The implementation of significance arithmetic in Mathematica [34] employs a linear error model consistent with the first-order approximation of interval arithmetic [13][22][23].
-One limitation of significance arithmetic is its inability to specify uncertainty accurately [29].
-For example, if the least significant bit of the significand is used to represent uncertainty, the result precision can be very coarse, as in $1\pm 10^{-3}=1024\times 2^{-10}$ [29].
-Introducing a limited number of bits to represent uncertainty does not fully resolve this issue.
-Consequently, various attempts to develop floating-point arithmetic without normalization [33] have not been widely adopted, and conventional floating-point arithmetic continues to dominate the numerical world.
-For this reason, variance arithmetic has abandoned the significance arithmetic principles of its predecessor [29].
-1.3.5 Stochastic Arithmetic
-Stochastic arithmetic [35][36] randomizes the least significant bits of each input floating-point value, repeats the same calculation multiple times, and then applies statistical analysis to identify the invariant digits among the results as significant digits.
-However, this approach can be computationally expensive as the number of repetitions required for each input depends on the algorithm and can increase substantially when the algorithm contains branching operations.
-In contrast, statistical Taylor expansion provides a direct characterization of the result’s mean and deviation without sampling.
-1.4 An Overview of This Paper
-This paper presents the theory of statistical Taylor expansion, its implementation as variance arithmetic, and the corresponding validation tests.
-Section 1 compares statistical Taylor expansion and variance arithmetic with other established uncertainty bearing arithmetic.
-Section 2 develops the theoretical foundation of statistical Taylor expansion.
-Section 3 describes variance arithmetic as an implementation of statistical Taylor expansion.
-Section 4 outlines the standards and methodologies used to validate variance arithmetic.
-Section 5 illustrates variance arithmetic in computing polynomial.
-Section 6 evaluates variance arithmetic on common mathematical library functions.
-Section 7 applies variance arithmetic to adjugate matrix and matrix inversion.
-Section 8 demonstrates its application to time-series data.
+    Section 4 lays out standard to validate variance arithmetic.
+  5. 5.
+    Section 5 illustrates variance arithmetic in computing polynomial, demonstrating its capability in tracing floating-point rounding errors, and its continuity in parameter space.
+  6. 6.
+    Section 6 applies variance arithmetic to adjugate matrix and matrix inversion, showing that the uncertainty calculation for a linear calculation is no longer linear.
+It also distinguishes between a distribution test and a value test.
+  7. 7.
+    Section 7 evaluates variance arithmetic on common mathematical library functions. It also shows that the error deviation should be close to $1$ except near a distributional pole.
+  8. 8.
+    Section 8 reveals the accumulation of numerical errors by reusing an input multiple time in a moving-window progressive algorithm.
+  9. 9.
+    Section 9 showcases variance arithmetic in catching catastrophic cancellation in a regression algorithm.
+  10. 10.
     Section 10 examines the impact of numerical library errors and shows that these errors can be significant.
-Section 10 applies variance arithmetic to regression analysis.
-Section 11 concludes with a summary and discussion.
+  11. 11.
+    Section 11 concludes with a summary and a discussion.
 2 Statistical Taylor Expansion
-2.1 The Uncorrelated Uncertainty Assumption
-Figure 1: Effect of noise on bit values of a measured value.
-The triangular wave signal and the added white noise are shown at top using the thin black line and the grey area, respectively.
-The values are measured by a theoretical 4-bit Digital-to-Analog Converter in ideal condition, assuming LSB is the 0th bit.
-The measured 3rd and 2nd bits without the added noise are shown using thin black lines, while the mean values of the measured 3rd and 2nd bits with the added noise are shown using thin grey lines.
-This figure is a reproduction of Figure 1 in [29]
-Figure 2: $\gamma$ vs $P$ for different $\gamma_{P}$ in Formula (2.4).
-If $\gamma_{P}$ is the maximal allowed correlation on the uncertainty level for uncorrelated uncertainty assumption to hold, then $\gamma$ is the maximal allowed correlation for two input to hold.
-For a given $\gamma_{P}$, $\gamma$ increases with finer $P$.
-This figure is a reproduction of Figure 2 in [29].
-The uncorrelated uncertainty assumption [29] states that the uncertainties of any two input imprecise values are uncorrelated.
-This assumption is satisfied when there is no systematic errors in the inputs [1] and is consistent with standard methods for processing experimental data [1][2].
-The uncorrelated uncertainty assumption permits the signals themselves to be highly correlated in the following statistical test [29].
-Let $X$, $Y$, and $Z$ denote three mutually independent random variables [4] with variance $V(X)$, $V(Y)$ and $V(Z)$, respectively.
-Let $\alpha$ denote a constant.
-Let $Cov()$ denote the covariance function.
-Let $\gamma$ denote the correlation between $(X+Y)$ and $(\alpha X+Z)$.
-And let:
-$\eta_{1}^{2}\equiv\frac{V(Y)}{V(X)};\;\;\;\eta_{2}^{2}\equiv\frac{V(Z)}{V(\alpha X)}=\frac{V(Z)}{\alpha^{2}V(X)};$ (2.1)
-$\gamma=\frac{Cov(X+Y,\alpha X+Z)}{\sqrt{V(X+Y)}\sqrt{V(\alpha X+Z)}}=\frac{\alpha/|\alpha|}{\sqrt{1+\eta_{1}^{2}}\sqrt{1+\eta_{2}^{2}}}\equiv\frac{\alpha/|\alpha|}{1+\eta^{2}};$ (2.2)
-Formula (2.2) gives the correlation $\gamma$ between two random variables, each of which contains a completely uncorrelated part and a completely correlated part $X$, with $\eta$ being the average ratio between these two parts.
-Formula (2.2) can also be interpreted reversely: if two random variables are correlated by $\gamma$, each of them can be viewed hypothetically as containing a completely uncorrelated part and a completely correlated part, with $\eta$ being the average ratio between these two parts  [^4] The widely used common factor analysis is also based on this principle..
-The correlated parts common to different measurements are regarded as signals, which can either be desired or unwanted.
-One special application of Formula (2.2) is the correlation between a measured signal and its true signal, in which noise is the uncorrelated part between the two.
-Figure 1 shows the effect of noise on the most significant two bits of a 4-bit measured signal when $\eta=1/4$.
-Its top chart shows a triangular waveform between 0 and 16 as a black line, and a white noise between -2 and +2, using the grey area.
-The measured signal is the sum of the triangle waveform and the noise.
-The middle chart of Figure 1 shows the values of the 3rd digit of the true signal as a black line, and the mean values of the 3rd bit of the measurement as a grey line.
-The 3rd bit is affected by the noise during its transition between 0 and 1.
-For example, when the signal is slightly below 8, only a small positive noise can turn the 3rd digit from 0 to 1.
-The bottom chart of Figure 1 shows the values of the 2nd digit of the signal and the measurement as a black line and a grey line, respectively.
-Figure 1 clearly shows that the correlation between the measurement and the true signal is less at the 2nd digit than at the 3rd digit.
-Quantitatively:
-  • The overall measurement is 99.2% correlated to the signal with $\eta=1/8$;
-  • The 3rd digit of the measurement is 97.0% correlated to the signal with $\eta=1/4$;
-  • The 2nd digit of the measurement is 89.4% correlated to the signal with $\eta=1/2$;
-  • The 1st digit of the measurement is 70.7% correlated to the signal with $\eta=1$;
-  • The 0th digit of the measurement is 44.7% correlated to the signal with $\eta=2$.
-The above conclusion agrees with the common experiences that, below the noise level of measured signals, noises rather than true signals dominate each digit.
-Similarly, while the correlated portion between two values has the same value at each bit of the two values, the ratio of the uncorrelated portion to the correlated portion increases by 2-fold for each bit down from MSB of the two values.
-Quantitatively, let $P$ denote the precision of an imprecise value, and let $\eta_{P}$ denote the ratio of the uncorrelated portion to the correlated portion at level of uncertainty; then $\eta_{P}$ increases with decreased $P$ according to Formula (2.3).
-According to Formula (2.2), if two significant values are overall correlated with $\gamma$, at the level of uncertainty the correlation between the two values decreases to $\gamma_{P}$ according to Formula (2.4).
-$\eta_{P}=\frac{\eta}{P},\;\;\;P<1;$ (2.3)
-$\frac{1}{\gamma_{P}}-1=\left(\frac{1}{\gamma}-1\right)\frac{1}{P^{2}},\;\;\;P<1;$ (2.4)
-Figure 2 plots the relation of $\gamma$ vs.
-$P$ for each given $\gamma_{P}$ in Formula (2.4).
-When $\gamma_{P}$ is less than a predefined maximal threshold (e.g., 2%, 5% or 10%), the two values can be deemed uncorrelated to each other at the level of uncertainty.
-For each independence standard $\gamma_{P}$, there is a maximal allowed correlation between two values below which the uncorrelated uncertainty assumption of statistical Taylor expansion holds.
-The maximal allowed correlation is a function of the larger precision of the two values according to Formula (2.4).
-Figure 2 shows that for two precisely measured values, their correlation $\gamma$ is allowed to be quite high.
+2.1 The Uncorrelated Uncertainty Condition
+When there is no systematic error [1][2] among the inputs, the uncertainties of the inputs are uncorrelated to each other, even though the inputs may have significant correlations among them.
+This condition can be characterized quantitatively in statistics as the uncorrelated uncertainty condition [17].
 2.2 Distributional Zero and Distributional Pole
@@
 $\rho(\tilde{x},\mu,\sigma)d\tilde{x}$ $=\rho(f^{-1}(\tilde{y}),\mu,\sigma)\frac{d\tilde{x}}{d\tilde{y}}d\tilde{y}=\rho(\tilde{y},\mu_{y},\sigma_{y})d\tilde{y}$; (2.1)
-$y=x^{c}:\;\;\;\rho(\tilde{y},\mu_{y},\sigma_{y})=c\tilde{y}^{\frac{1}{c}-1}\frac{1}{\sigma}N(\frac{\tilde{y}^{\frac{1}{c}}-\mu}{\sigma});$ (2.6)
 […]
-  • If $\tilde{y}=\alpha+\beta\tilde{x}$, the resulting distribution is identical to the original distribution, since $\rho(\tilde{y},\mu_{y},\sigma_{y})=\rho(\tilde{y},\alpha+\beta\mu,\beta\sigma)$ [4].
-A linear transformation generates neither a distributional zero nor a distributional pole, according to Formula (2.5).
-  • Figure 3 shows the probability density function for $(x\pm 1)^{2}$ according to Formula (2.6), which exhibits a distributional pole at $x=0$.
-The distribution $(0\pm 1)^{2}$ corresponds to the $\chi^{2}$ distribution [1].
-At the distributional pole, the probability density function resembles a Delta distribution.
-  • Figure 4 illustrates the probability density function for $\sqrt{x\pm 1}$ according to Formula (2.6), which has a distributional zero at $x=0$.
-At the distributional zero, the probability density function is zero.
-In both Figure 3 and 4, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, thereby allowing for a generic characterization of the output.
+  • A distributional pole occurs when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$.
+For example, $(\mu\pm 1)^{2}$ has a distributional pole at $\tilde{x}=0$, as shown in Figure 1.
+  • A distributional zero occurs when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$.
+For example $\sqrt{\mu\pm 1}$ has a distributional zero at $\tilde{x}=0$, as shown in Figure 2.
+In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, such as $(5\pm 1)^{2}$ in Figure 1 and $\sqrt{5\pm 1}$ in Figure 2, thereby allowing for characterization of the output $f(x)$ using mean $\overline{f(x)}$ and deviation $\delta f(x)$ only.
 2.3 Statistical Taylor Expansion
-Formula (2.5) provides the uncertainty distribution of an analytic function.
-However, in most scientific and engineering calculations, the primary interest lies not in the full result distribution but in few summary statistics of the result, such as the mean and deviation [1][2].
-These simplified statistics can be obtained through a statistical Taylor expansion.
 […]
-$\zeta(n)\equiv$ $\int_{-\varrho}^{+\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.7)
+$\zeta(n)$ $\equiv\int_{\varrho}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.2)
+$\zeta(1)$ $=0;$ (2.3)
+Let $\tilde{z}\in[\varrho,\kappa]$ where $\varrho,\kappa$ specify the bounding ranges.
+Formula (2.2) defines the corresponding bound moment $\zeta(n)$, which further satisfies the mean-reverting condition of Formula (2.3) so that $\kappa$ determines $\varrho$.
+For any symmetric $\rho(-\tilde{z})=\rho(\tilde{z}):\varrho=-\kappa,\zeta(2n+1)=0$.
+The probability of $\tilde{z}\not\in[\varrho,\kappa]$ is defined as the bounding leakage $\epsilon(\kappa)=1-\zeta(0,\kappa)$.
 […]
-$\overline{f(x)}=$ $\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=\sum_{n=0}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.9)
+$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=f(x)+\sum_{n=1}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.5)
 $\delta^{2}f(x)$ $=\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$
 $=\sum_{n=1}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.6)
-An confidence interval [4] $\tilde{x}\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ can describe a sampling from an underlying distribution, where $0<\varrho,\kappa$ specify the bounding ranges.
-Formula (2.7) defines the bound moment $\zeta(n)$.
-As discussed later, $\kappa$ determines $\varrho$, while $\kappa$ itself is determined by both the sample size $N$ and the underlying distribution of the input.
-When $N\rightarrow\infty$, $\varrho,\kappa\rightarrow\infty$, so that by definition $\zeta(0)=1$, $\zeta(1)=0$, and $\zeta(2)=1$.
-When $N$ is bounded, $\varrho$ and $\kappa$ are also bounded, so that $\zeta(0)<1$, and $\zeta(2)<1$.
-The probability of $\tilde{x}\not\in[\mu-\varrho\sigma,\mu+\kappa\sigma]$ is defined as the bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
 An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.4).
-Using Formula (2.7), Formula (2.9) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively.
+Formula (2.5) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively.
 The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
-$f(x+\tilde{x},y+\tilde{y})=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}\frac{f^{(m,n)}_{(x,y)}}{m!n!}\tilde{x}^{m}\tilde{y}^{n};$ (2.11)
-$\overline{f(x,y)}=$ $\int\int f(x+\tilde{x},y+\tilde{y})\rho(\tilde{x},\mu_{x},\sigma_{x})\rho(\tilde{y},\mu_{y},\sigma_{y})\;d\tilde{x}d\tilde{y}$ (2.12)
-$=$ $\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$
+[…]
+$\overline{f(x,y)}$ $=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$ (2.8)
 $\delta^{2}f(x,y)$ $=\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
 $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\big(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\big);$ (2.9)
-Under the uncorrelated uncertainty assumption, Formula (2.12) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.11), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively.
+Under the uncorrelated uncertainty condition, Formula (2.8) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.7), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively.
 Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
-2.4 Bound Moment Symmetry
-A bounding range $\mu-\varrho\sigma,\mu+\kappa\sigma]$ is mean-preserving if $\zeta(1)=0$, meaning that it has the same mean as the unbounded distribution.
-To achieve mean preserving, $\kappa$ determines $\varrho$, so that $\zeta(n)$ becomes $\zeta(n,\kappa)$.
-Under a mean preserving bounding, Formula (2.14) and (2.15) provide the result for $x\pm y$, while Formula (2.16) and (2.17) give the result for $xy$:
+With the mean reverting condition of Formula (2.3):
 $\overline{x\pm y}$ $=\zeta(0,\kappa_{x})x\pm\zeta(0,\kappa_{x})y;$ (2.10)
@@
 $\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\sigma_{y})^{2}+\zeta(2,\kappa_{x})(\sigma_{x})^{2}\;\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.13)
-When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.14) and (2.15) the convolution results for $x\pm y$ [4], and Formula (2.16) and (2.17) the corresponding results of the product distribution for $xy$ [4].
-For any input distribution $\rho(\tilde{x},\mu,\sigma)$ that is symmetric about its mean $\mu$, any bounding range $\mu-\kappa\sigma,\mu+\kappa\sigma]$ satisfies $\zeta(2n+1)=0$, which further simplifies the statistical Taylor expansion.
+When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formula (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
 2.4 One-Dimensional Examples
 Formula (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
-$e^{x+\tilde{x}}=e^{x}\sum_{n=0}^{\infty}\frac{\tilde{x}^{n}}{n!};$ (2.18)
-$\frac{\overline{e^{x}}}{e^{x}}$ $=\sum_{n=0}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.19)
+$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.14)
 $\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.15)
 Formula (2.16) and (2.17) give the mean and variance for $\log(x)$, respectively:
-$\log(x+\tilde{x})-\log(x)=\log(1+\frac{\tilde{x}}{x})=\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j}\frac{\tilde{x}^{j}}{x^{j}};$ (2.21)
 $\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n)}{n};$ (2.16)
@@
 Formula (2.21) and (2.22) give the mean and variance for $x^{c}$, respectively:
-$(x+\tilde{x})^{c}=x^{c}(1+\frac{\tilde{x}}{x})^{c}=x^{c}+x^{c}\sum_{n=1}^{\infty}\frac{\tilde{x}^{n}}{x^{n}}\begin{pmatrix}c\\ n\end{pmatrix};\;\;\;\begin{pmatrix}c\\ n\end{pmatrix}\equiv\frac{\prod_{j=0}^{n-1}(c-j)}{n!};$ (2.27)
-$\frac{\overline{x^{c}}}{x^{c}}$ $=1+1\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.28)
+$\frac{\overline{x^{c}}}{x^{c}}$ $=1+\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.21)
 $\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.22)
-The result variance in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
-2.6 Low-Order Approximation
-When $n<5\leq\kappa$, $\eta(n)\simeq n!!$. Under these conditions, Formula (2.20), (2.23), (2.26), and (2.29) can be simplified as Formula (2.30), (2.31), (2.6), and (2.6), respectively.
-$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}\simeq$ $\;\sigma^{2}+\frac{3}{2}\sigma^{4}+\frac{7}{6}\sigma^{6}+\frac{5}{8}\sigma^{8}+o((\delta x)^{10});$ (2.30)
-$\delta^{2}\log(x)\simeq$ $\;P(x)^{2}+P(x)^{4}\frac{9}{8}+P(x)^{6}\frac{119}{24}+P(x)^{8}\frac{991}{32}+o(P(x)^{10});$ (2.31)
-$\delta^{2}\sin(x)\simeq$ $\;\sigma^{2}\cos(x)^{2}-(\delta x)^{4}(\cos(x)^{2}\frac{3}{2}-\frac{1}{2})$
-$+\sigma^{6}(\cos(x)^{2}\frac{7}{6}-\frac{1}{2})-\sigma^{8}(\cos(x)^{2}\frac{5}{8}-\frac{7}{24})+o((\delta x)^{10});$ (2.32)
-$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}\simeq$ $\;c^{2}P(x)^{2}+\frac{3}{2}c^{2}(c-1)(c-\frac{5}{3})P(x)^{4}$
-$+\frac{7}{6}c^{2}(c-1)(c-2)^{2}(c-\frac{16}{7})P(x)^{6}+o(P(x)^{8});$ (2.33)
-Formula (2.34), (2.35), and (2.36) are special cases of Formula (2.6).
-$\delta^{2}x^{2}\simeq$  $\;4x^{2}(\delta x)^{2}+2(\delta x)^{4};$ (2.34)
-$\frac{\delta^{2}\sqrt{x}}{(\sqrt{x})^{2}}\simeq$ $\;\frac{1}{4}P(x)^{2}+\frac{7}{32}P(x)^{4}+\frac{75}{128}P(x)^{6}+o(P(x)^{8});$ (2.35)
-$\frac{\delta^{2}1/x}{(1/x)^{2}}\simeq$ $\;P(x)^{2}+8P(x)^{4}+69P(x)^{6}+o(P(x)^{8});$ (2.36)
-2.7 Asymptotic Behavior for Bound Moment
-When $\rho(\tilde{z})>0$ and $\frac{d\rho(\tilde{z})}{d\tilde{z}}$ is bounded in $\tilde{z}\in[0,\kappa]$, Formula (2.38) and (2.39) leads to Formula (2.40).
-Formula (2.37) and (2.40) conclude Formula (2.41) and (2.42).
-$\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$ $=\rho(\kappa)\frac{\kappa^{n+1}}{n+1}-\frac{1}{n+1}\int_{0}^{\kappa}\tilde{z}^{n+1}\frac{d\rho(\tilde{z})}{d\tilde{z}}d\tilde{z}$; (2.37)
+The input and output in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
+2.5 Convergence of Variance
+Figure 3: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus $2N(\kappa)\frac{\kappa^{2n+1}}{2n+1}$ (left y-axis), and their ratio (right y-axis).
+$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.23)
+$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.24)
+$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.25)
 […]
 […]
-$\lim_{n\rightarrow+\infty}$ $\int_{0}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}=\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.40)
-$|\varrho|=\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.41)
-$|\varrho|<\kappa$ $:\;\;\;\lim_{n\rightarrow+\infty}\zeta(n,\kappa)=\rho(\varrho)\frac{\varrho^{n+1}}{n+1}+\rho(\kappa)\frac{\kappa^{n+1}}{n+1}\simeq\rho(\kappa)\frac{\kappa^{n+1}}{n+1};$ (2.42)
-2.7.1 Gaussian Input Uncertainty
-Figure 5: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus Formula (2.45) (left y-axis), and their ratio (right y-axis).
-The central limit theorem states that the sum of many independent and identically distributed random variables converges toward a Gaussian distribution [4].
-This convergence occurs rapidly [29].
-In digital computation, multiplication is implemented as a sequence of shifts and additions, division as a sequence of shifts and subtractions, and general functions are calculated as sums of expansion terms [7][8].
-Consequently, uncertainty without explicit bounds is generally assumed to follow a Gaussian distribution [1][2][4].
-With Formula (2.43) [Double_Factorial], for Gaussian distribution, Formula (2.44) provides $\zeta(2n,\kappa)$ while $\zeta(2n+1,\kappa)=0$, while Formula (2.45) demonstrates its asymptotic behavior.
-Figure 5 shows the transition to Formula (2.45) with increasing order $2n$, in which the ratio is the ratio of Formula (2.45) when $\kappa=5$ to $\zeta(2n,5)$.
-$\sum_{j=0}^{+\infty}\frac{\kappa^{2j+1}}{(2j+1)!!}=\sqrt{\frac{\pi}{2}}\xi(\frac{\kappa}{\sqrt{2}})e^{\frac{1}{2}\kappa^{2}};$ (2.43)
-$\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}N(\tilde{z})d\tilde{z}=2N(\kappa)\kappa^{2n}\sum_{j=1}^{\infty}\kappa^{2j-1}\frac{(2n-1)!!}{(2n-1+2j)!!}$ (2.44)
-$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2N(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.45)
-2.7.2 Uniform Input Uncertainty
-As shown in Figure 38, rounding errors can be uniformly distributed [29].
-For uniform distribution, the bounding range $\kappa$ is less than the maximal range $\sqrt{3}$. Formula (2.46) provides $\zeta(2n)$ while $\zeta(2n+1)=0$.
-$\zeta(2n)=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.46)
-2.8 Convergence of Variance
-$\lim_{n\rightarrow+\infty}$ $\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.47)
-$\delta^{2}\log(x\pm\delta x)\simeq$ $\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\frac{1}{j}\frac{1}{2n-j}=\sum_{n=1}^{+\infty}P(x)^{2n}\zeta(2n)\frac{1}{n}\sum_{j=1}^{2n-1}\frac{1}{j}$
-$\simeq$  $\;2\nu(\kappa)\log(2)\sum_{n=1}^{+\infty}\frac{(P(x)\kappa)^{2n}}{(2n)^{2}},\begin{cases}\text{Gaussian}:\nu(\kappa)=N(\kappa)\kappa\\ \text{Uniform}:\nu(\kappa)=1,\;\;\;\kappa=\sqrt{3}\end{cases};$ (2.48)
-$\frac{\delta^{2}(x\pm\delta x)^{c}}{(x^{c})^{2}}\simeq$ $\sum_{n=1}^{\infty}P(x)^{2n}\zeta(2n)\sum_{j=1}^{2n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ 2n-j\end{pmatrix}\simeq\nu(\kappa)\sum_{n=1}^{+\infty}(P(x)\kappa)^{2n}\frac{\begin{pmatrix}2c\\ 2n\end{pmatrix}}{2n};$ (2.49)
-For input distribution without bounding, Formula (2.3) may not converge, e.g., when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$.
-Thus, it is necessary that the input uncertainty is bounded.
-Formula (2.47) shows that the asymptotic behavior of $\zeta(n,\kappa)$ when $n\rightarrow\infty$ determines if Formula (2.3) will converge.
+[…]
   • Formula (2.15) for $e^{x\pm\delta x}$ and Formula (2.20) for $\sin(x\pm\delta x)$ both converge unconditionally.
-However, as shown later in this pap, $\delta^{2}\sin(x\pm\delta x)$ can become negative for large $\delta x$, which imposes an upper-bound constraint on the input $\delta x$.
-Formula (2.23) for $\log(x\pm\delta x)$ can be approximated by Formula (2.8) as $n\rightarrow\infty$, which converges when $P(x)<1/\kappa$.
-Formula (2.29) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.49) after applying Vandermonde’s identity $\sum_{k=0}^{r}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}n\\ r-k\end{pmatrix}=\begin{pmatrix}m+n\\ r\end{pmatrix}$ .
-This expression converges when $P(x)\lesssim 1/\kappa$ although the precis upper bound for $P(x)$ varies with $c$.
+[…]
+  • Formula (2.22) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.27). which converges when $P(x)\lesssim 1/\kappa$ although the precise upper bound for $P(x)$ varies with $c$.
 2.6 Statistical Bounding
@@
 It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
-Figure 7 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases (e.g., $\kappa(2,100)\simeq 2$ vs $\kappa(5,1000)\simeq 5$), but converges to a larger stable bounding leakage (e.g., $\epsilon(2)=4.55\;10^{-2}$ vs $\epsilon(5)=5.73\;10^{-7}$).
-Figure 7 also indicates that when $N\geq 30$, the difference between $\epsilon(4,N)$ and $\epsilon(5,N)$ is less than $10^{-3}$, suggesting that $\kappa(\kappa_{s},N)$ becomes stable when $\kappa_{s}\geq 4$.
-Moreover, according to the 5-$\sigma$ rule, $\kappa(5,N)$ in Figure 6 should be used as $\kappa$ in Formula (2.7).
-When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to bounding leakage $\epsilon(N)$.
+Figure 5 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases, but converges to a larger stable bounding leakage $\epsilon(\kappa_{s})$.
+When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to the bounding leakage $\epsilon(N)$.
 Figure 5 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically.
-The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$, which should be used as $\kappa$ in Formula (2.7).
-[…]
+The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$.
 2.7 Ideal Statistics
@@
 $\alpha(x\pm y)=$ $\frac{\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.31)
-When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}\rightarrow\sigma$ in Formula (2.52).
+When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}$ in Formula (2.30).
 Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
-  • When the underlying uncertainty distribution is Gaussian, Figure 8 shows that the resulting variance $\delta^{2}f$ for a selected functions rise with NN until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 50$.
-When $N=20$ and $\kappa_{s}=5$, $\kappa=3.8$ according to Figure 6.
-Therefore it is sufficient to compute the stable variances $\widehat{\delta^{2}}f$ with $\zeta(n,5)$.
-  • When the underlying uncertainty distribution is Uniform, Figure 9 shows the same trend, however the stable variances are reached only when $N>10^{4}$.
-The stable variance $\widehat{\delta^{2}}f$ is computed using Formula (2.46).
-Define the stable variance when the sample count $N$ is sufficient large as the ideal variance $\widehat{\delta^{2}}f$, which assumes that the sample count $N$ is effectively infinite for each input.
-In both Figure 8 and 9, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
-The stable resulting variances $\delta^{2}f$ is reached only when $\epsilon(s,N)<10^{-3}$.
-All functions approach their corresponding ideal variances with exactly the same trend.
-These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer confidential interval range $\kappa$ from bounding leakage $\epsilon(N)$.
-Define $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f\in[0,1]$ as the ideal ratio, which quantifies the reliability for $\widehat{\delta^{2}}f$ due to insufficient sample count $N$ for the given underlying distribution.
-Define the ideal ratio to be zero when the sample count $N$ is $1$.
-When there are multiple inputs each with its own ideal ratio $\alpha=\zeta(2,\kappa)$, Formula (2.53) demonstrates an example on how to calculate the resulting ideal ratio $\alpha$ for $x\pm y$.
-An ideal ratio also applies to the corresponding uncertainty bias.
-The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 8 and 9, except that $\overline{x}-x=0$.
-This similarity is expected because the first-order approximation in both Formula (2.9) and Formula (2.3) contains $\sigma^{2}\zeta(2)$.
-In statistical Taylor expansion, every calculation should be carried out twice:
-  1. 1.
-[…]
-  2. 2.
-    With the real $\zeta(n,\kappa)$ of each input given the sample count and input uncertainty distribution, to obtain the ideal ratio to quantify the reliability of the ideal variance and ideal uncertainty bias as a value in the range of $[0,1)$.
+  • When the underlying uncertainty distribution is Gaussian and $\kappa_{s}=5$, Figure 6 shows that the resulting variances $\delta^{2}f$ for a selected functions rise with $N$ until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 50$.
+  • When the underlying uncertainty distribution is Uniform, $\kappa_{s}=\sqrt{3}$, Figure 7 shows the same trend, however the stable variances $\widehat{\delta^{2}}f$ are reached only when $N\geq 10^{4}$.
+The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 6 and 7 respectively, except that $\overline{x}-x=0$.
+This similarity is expected because the first-order approximation in both Formula (2.5) and Formula (2.3) contains $\sigma^{2}\zeta(2,\kappa)$.
+In both Figure 6 and 7, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
+The stable resulting variances $\widehat{\delta^{2}}f$ is reached only when $\epsilon(s,N)<10^{-3}$.
+All functions approach their corresponding stable variances with exactly the same trend in each figure.
+These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer the confidential interval range $\kappa$ from bounding leakage $\epsilon(\kappa_{s})=1-\zeta(0,\kappa_{s})$ in Figure 4, or more precisely $1-\zeta(2,\kappa_{s})$ for $\delta^{2}f$.
+Define the stable variance when the sample count $N$ is sufficiently large as the ideal variance $\widehat{\delta^{2}}f$.
+The choice of $\kappa_{s}$ to calculate $\widehat{\delta^{2}}f$ is defined as the ideal bounding range $\hat{\kappa}$.
+For Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$.
+When the input distribution is unbounded, the choice of $\hat{\kappa}$ is a trade-off: a larger $\hat{\kappa}$ means a smaller bounding leakage for the result but a narrower convergence range for the input.
+For Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$.
+When the sample count $N$ is insufficient, $\widehat{\delta^{2}}f$ is calculated using $\hat{\kappa}$, to assume that $N$ is sufficient.
+$\delta^{2}f$ is calculated with the actual $N$.
+Define the ideal ratio as $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f$, which is 0 when $N=1$ and $1$ when $N$ is sufficiently large.
+The ideal ratio $\alpha$ quantifies the reliability of $\widehat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.31) as an example for $\alpha(x\pm y)$ .
 2.8 Dependency Tracing
 $\delta^{2}(f+g)=$ $\;\delta^{2}f+\delta^{2}g+2(\overline{fg}-\overline{f}\overline{g});$ (2.32)
-$\delta^{2}(fg)=$ $\;\overline{f^{2}g^{2}}-\overline{fg}^{2};$ (2.55)
-$\delta^{2}f(g)=$ $\;\overline{f(g)^{2}}-\overline{f(g)}^{2};$ (2.56)
 $\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.33)
 When all inputs satisfy the uncorrelated uncertainty assumption, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
-  • Formula (2.54) expresses $\delta^{2}(f+g)$, whose dependency tracing is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.15).
-Formula (LABEL:eqn:_Taylor_inner_dependency) shows that Formula (2.3) applies Formula (2.54) between any two terms in the Taylor expansion in Formula (2.8).
-  • Formula (2.55) expresses $\delta^{2}(fg)$, illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.17).
-  • Formula (2.56) shows $\delta^{2}f(g(x))$, whose dependency tracing is demonstrated by $\delta^{2}(f^{-1}(f))=(\delta x)^{2}$.
-  • Formula (2.57) gives the variance of the linear transformation of a function, which can be applied to Formula (2.54) and (2.55) for more general dependency tracing.
-Variance arithmetic employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously.
+  • Formula (2.32) expresses $\delta^{2}(f+g)$, The dependency tracing of is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
+  • The dependence tracing of $\delta^{2}(fg)$ is illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.13).
+  • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$.
+For a reversible transformation such as matrix reversion or FFT, after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
+  • Formula (2.33) gives the variance of the linear transformation of a function, which can be applied to other formulas for more general dependency tracing.
+Statistical Taylor expansion employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously.
+Dependency tracing also implies that the results of statistical Taylor expansion must remain path independent.
 However, dependency tracing comes at a cost: variance calculations are generally more complex than value calculations and exhibits a narrower convergence range for input variables.
-Dependency tracing also implies that the results of statistical Taylor expansion must remain path independent.
 2.9 Traditional Execution and Dependency Problem
@@
 Consequently, approximation strategies should prioritize variances than values.
-Section 7 illustrates this principle through a first-order approximation used in computing a matrix determinant.
-  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last bit.
-As demonstrated in Section 9, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty explicitly incorporated into the output.
+  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last digit.
+As demonstrated in Section 10, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty deviations explicitly incorporated into the output.
   • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls.
-However, this decomposition introduces dependency problems in floating-point arithmetic, interval arithmetic, and statistical Taylor expansion.
+However, this decomposition introduces dependency problems.
 For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
   • Similarly, large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$.
-This approach fails in statistical Taylor expansion because dependency tracing within $g(x)$ affects $f(g(x))$.
-In this context, $\overline{f(g(x))}\neq\overline{f(y)|_{y=g(x)}}$ and $\delta^{2}f(g(x))\neq\delta^{2}f(y)|_{y=g(x)}$.
-The path dependence of $f(y)|_{y=g(x)}$ and $\delta^{2}f(y)|_{y=g(x)}$ are evident in cases such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
+This approach also introduces the dependency problem by ignoring dependency tracing within $g(x)$ affecting $f(g(x))$, such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
 Dependency tracing therefore removes nearly all flexibility from traditional numerical executions, effectively eliminating the associated dependency problems.
@@
 Variance arithmetic implements statistical Taylor expansion.
+It represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers and uses floating-point arithmetic for computation.
 Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms.
@@
   • finite: The resulting value and variance must remain finite.
-  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion exhibiting an absolute increase is no more than $2^{-20}=9.53\;10^{-7}$.
-Unless all the remaining terms in the expansion are known to be precisely zeros, each expansion is executed to the full 448 terms for the monotonicity check.
-  • positive: At every order, the expansion variance must be positive.
-  • stable: To avoid truncation error [12], the value of the last expansion term must be less than $5.73\;10^{-7}$ times of both the result uncertainty and the result absolute value, in which $5.73\;10^{-7}$ is the bounding leakage for Gaussian distribution with bounding range $\kappa=5$.
+  • monotonic: As a necessary condition for convergence, the last $20$ terms of the expansion must decrease monotonically in absolute value, ensuring that the probability of the expansion exhibiting an absolute increase is no more than $2^{-20}\simeq 9.53\;10^{-7}$.
+  • stable: To avoid truncation error [5], the absolute value of the last expansion term must be less than $\epsilon$ times of both the result deviation and the result absolute value, in which $\epsilon\simeq 5.73\;10^{-7}$ is the bounding leakage for Gaussian distribution with $\hat{\kappa}=5$.
 This rule ensures sufficiently fast convergence in the context of monotonic convergence.
-  • reliable: At every order, the uncertainty of the variance must be less than $1/5$ times of the value of the variance.
-For simplicity of discussion, This paper confines the calculation of variances to ideal variances $\widehat{\delta^{2}}f$ and assumes that the input distribution is Gaussian with $\kappa_{s}=5$.
-Furthermore,the Taylor coefficients in Formula (2.8) and (2.11) are assumed to be precise.
-3.1 Numerical Representation
-Variance arithmetic represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers. All other conventional numerical numbers must be converted to this format.
-If the least $20$ significant bits of the significand in a standard floating-point number are all 0, the value is considered precise, representing a 2’s fractional with a probability no less than $1-2^{-20}=1-2.384\;10^{-7}$.
-Otherwise, a standard floating-point value is considered imprecise with its uncertainty defined as $1/\sqrt{3}$ times of the ULP of the value, where ULP refers to the Unit in the Last Place in conventional floating-point representation [8].
-This follows from the fact that the pure rounding error in round-to-nearest mode is uniformly distributed within half bit of the significand of a floating-point value [29].
-If an integer number is within the range $2^{53}+1,+2^{53}-1]$ of the significand of a 64-bit standard floating-point number, its uncertainty is zero.
-Otherwise, it is first converted to a conventional floating-point value before being transformed into an imprecise value.
-Variance arithmetic uses floating-point arithmetic for computation.
-3.2 Finite
-For $(1\pm\delta x)^{-2}$, the Taylor coefficient increases with the expansion order $n$ as $(-1)^{n}(n+1)$. When $\delta x=0.5$, this growth causes the result variance to diverge to infinity.
+  • positive: At every expansion order, the expansion variance must be positive.
+  • reliable: At every order, the deviation of the variance must be less than $1/5$ times the value of the variance.
+For simplicity of discussion, the Taylor coefficients in Formula (2.4) and (2.7) are assumed to be precise.
 3.1 Monotonic
@@
 When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
-From Formula (2.8), the convergence condition applied to $\log(x\pm\delta x)$ is $P(x)\leq 1/\kappa=1/5$, which is numerically confirmed as $P(x)\lesssim 0.20086$.
-Beyond this upper bound, the expansion is no longer monotonic.
-Variance arithmetic rejects the distributional zero of $\log(x)$ in the range of $x-\delta x,x+\delta x]$ statistically due to the divergence of Formula (2.23) mathematically, with $\zeta(2n)$ providing the connection between these two perspectives.
-For $e^{x\pm\delta x}$ the convergence holds for $\delta x\lesssim 19.864$ regardless of $x$, while the result $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$.
-These limits follow directly from the relationship $\delta x\rightarrow P(e^{x})$ and $P(x)\rightarrow\delta\log(x)$, as indicated in Formula (2.20) and (2.23).
-From Formula (2.49), and except when $c$ is a natural number, Formula (2.29) for $(x\pm\delta x)^{c}$ converges near $P(x)\simeq 1/\kappa=1/5$, with the upper bound $P(x)$ increasing with $c$.
-This trend is approximately confirmed in Figure 10, and the expansion is no longer monotonic beyond the upper bound $P(x)$.
-Qualitatively, $\delta^{2}1/x$ converges more slowly than $\delta^{2}\sqrt{x}$, consistent with Formula (2.36) and (2.35).
-When the input uncertainty is Uniformly distributed, Figure 11 illustrates that the input uncertainty upper bound is close to $1/\kappa=/\sqrt{3}$ according to Formula (2.49).
-The bounded momentum $\zeta(2n)$ for Uniform distribution is much less than that for Normal distribution, such that the expansion contains $652$ terms instead of $448$ terms.
-The upper bounds of input uncertainty in Figure 10 and 11 looks identical, however, the resulting uncertainty biases and uncertainties in Figure 10 are $3$ order-of-magnitude smaller than those in Figure 11 when the exponent $c$ is less than 0.
-Even a result converges, it can still be abandoned due to too large uncertainty bias or uncertainty.
+When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
+Beyond an upper bound $\delta x$, the expansion is no longer monotonic for $e^{x\pm\delta x}$, $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c}$.
+For Gaussian input uncertainty with $\hat{\kappa}=5$:
+  • For $e^{x\pm\delta x}$, $\delta x\lesssim 19.864$ and $P(e^{x\pm\delta x})\lesssim 1681.767$ regardless of $x$.
+These limits follow directly from the relationship $\delta x\rightarrow P(e^{x})$, as indicated in Formula (2.15).
+  • For $\log(x\pm\delta x)$, $P(x)\lesssim 0.20086$ and $\delta\log(x\pm\delta x)\lesssim 0.213$ regardless of $x$.
+These limits follow directly from the relationship $P(x)\rightarrow\delta\log(x)$, as indicated in Formula (2.17).
+  • For $(x\pm\delta x)^{c}$, except when $c$ is a natural number, the upper bound $P(x)$ is close to $1/5$ but increasing with $c$. This trend is displayed in Figure 8.
+Similar trends holds when the input uncertainty is uniform but with different values.
+For example, after the respective upper bound $P(x)$ for $(x\pm\delta x)^{c}$ are normalized by the corresponding $\hat{\kappa}$, they show almost the same trends of increasing with $c$.
+Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$.
 3.2 Positive
@@
 The x-axis is expressed in units of $\pi$.
-In some cases, the variance expansion may yield negative results, as in Formula (2.26) for $sin(x\pm\delta x)$.
-Figure 12 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$.
-Beyond this upper bound, the expansion is no longer positive.
-In Figure 13, when the bounded momentum $\zeta(2n)$ for Uniform distribution is used instead of that for Normal distribution, the input uncertainty upper bound is $4$ times larger than that in Figure 12 but is still periodic.
-The resulting uncertainty increases not much and still remains less than $1$.
-3.5 Stable
-The unstable condition happens rarely.
-3.6 Reliable
-The condition of not being reliable seldom happens.
+Besides convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$.
+Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty.
+Beyond this upper bound, the result deviation is no longer positive at some expansion order.
+Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which Figure 10 satisfies.
+Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
+3.3 Floating-Point Rounding Errors
+Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$.
+Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP of $x$, where ULP refers to the Unit in the Last Place in conventional floating-point representation [8], because rounding errors are shown to be uniformly distributed within ULP [17].
 3.4 Comparison
@@
 4 Verification of Variance Arithmetic
-4.1 Verification Methods and Standards
 Analytic functions or algorithms with precisely known results are used to evaluate the outputs of variance arithmetic based on the following statistical properties:
   • Value error: the difference between the numerical result and the corresponding known precise analytic result.
-  • Value deviation: the standard deviation of the value errors.
-  • Normalized error: the ratio of a value error to the corresponding uncertainty.
+  • Normalized error: the ratio of a value error to the corresponding result deviation from statistical Taylor expansion.
   • Error deviation: the standard deviation of normalized errors.
   • Error distribution: the histogram of the normalized errors.
-  • Uncertainty mean: the sample mean of the result uncertainties.
-  • Uncertainty response: the relationship between input uncertainties and output uncertainties.
-  • Calculation response: the relationship between the amount of calculation and output uncertainties.
-One objective of uncertainty-based computation is to account precisely for all input errors from every source, thereby achieving ideal coverage:
-  1. 1.
-    The error deviation is exactly $1$.
-  2. 2.
-    The error distribution should follow Normal distribution when an imprecise value is expected, or Delta distribution when a precise value is expected.
-If the precise result is unknown, the resulting normalized error distribution can be used to assess whether ideal coverage is achieved.
+Once input error from every source is accounted precisely, ideal coverage is achieved in either contexts:
+  • Distribution Test: When comparing the calculated mean and deviation with the result data set, the error deviation is exactly $1$ and the error distribution is Normal.
+Empirically, error distributions are close to Normal when error deviation are close to 1, regardless of input uncertainty distribution, and such convergence occurs rapidly [17].
+This is due to central limit theorem [4].
+  • Value Test: When comparing the calculated value and the result value in a corresponding data set, the error deviation is much less than $1$ and the error distribution is Delta.
+For example, a round-trip test is a value test.
 However, if the input uncertainty is known only to order of magnitude, proper coverage is achieved when the error deviations fall within the range $[0.1,10]$.
-When an input contains unspecified errors, such as numerical errors in library functions, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained.
-The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainties.
-Achieving ideal coverage serves as a necessary verification step to ensure that Formula (2.3) or (2.3) have been applied correctly within the given context.
-The input noise range that yields ideal coverage defines the ideal application range for input uncertainties.
-Besides coverage, other validation criteria are also used:
-  • The uncertainty response should match the expected functional form.
-For a linear function, the output uncertainties should increase linearly with the input uncertainties, with the ratio of output to input uncertainty means defined as uncertainty response ratio.
-  • The calculation response should follow the expected trend; for instance, increasing the number of calculations should result in larger output uncertainties.
-Noises of varying deviations can also be added to the input to evaluate a function’s uncertainty response.
-4.2 Types of Uncertainties
-There are five primary sources of result uncertainty in a calculation [1][12][29]:
-  • Input uncertainties: The examples presented in this paper demonstrate that when the precision of input uncertainties is $10^{-15}$ or larger, variance arithmetic can achieve ideal coverage for input uncertainties.
-  • Rounding errors: Empirical results indicate that variance arithmetic provides proper coverage for rounding errors.
-  • Truncation errors: Variance arithmetic avoids truncation errors with its stable rule.
-However, using Formula (5.1) for polynomial expansion can result in truncation errors when the choice of the expansion order is insufficient, as illustrated in Figure 14 .
-  • External errors: External errors are value errors not specified in the input uncertainties, such as numerical errors in library functions.
-Section 9 examines the effects of numerical errors of library sine function, showing that when these external errors are sufficiently large, neither ideal coverage nor proper coverage can be achieved.
-This finding indicates that library functions must be recalculated to explicitly include the corresponding uncertainty for each computed value.
-  • Modeling errors: Modeling errors arise when an approximate analytic solution is used, or when a real-world problem is simplified to make a solution tractable.
-For example, Section 9 demonstrates that the discrete Fourier transform (DFT) is only an approximation of the mathematically defined continuous Fourier transform (FT), and therefore contains modeling errors.
-Conceptually, modeling errors originate in mathematics within and are thus outside the domain of statistical Taylor expansion.
-4.3 Types of Calculations to Verify
-Algorithms of distinct natures, with each representative of its respective category, are required to test the broad applicability of variance arithmetic [29].
-An algorithm can be categorized by comparing the amount of its input and output data as [29]:
-  • Application,
-  • Transformation,
-  • Generation,
-  • Reduction.
-An application algorithm computes numerical values from an analytic formula.
-Through statistical Taylor expansion, variance arithmetic applies directly to analytic problems.
-For example, for the catastrophic cancellation example of Formula (1.2), Formula (4.3) shows that variance arithmetic bound the rounding error of $1$ with uncertainty $1$:
-$64919224$ $\times 205117922-159018721\times 83739041=$
-$13316075197586562.-13316075197586560.\pm 1=2.\pm 1;$ (4.1)
-A transformation algorithm produces output data of approximately the same quantity as its input, with the overall information content remaining largely unchanged.
-For reversible transformations, a unique requirement is to recover every original input for both value and uncertainty after a roundtrip transformation that is performing a forward transformation followed by its reverse transformation.
-The discrete Fourier transform (DFT) is a typical reversible transformation algorithm: it has the same amount of input and output data, and its output can be transformed back into the input using essentially the same process.
-A test of variance arithmetic using the fast Fourier transform (FFT, which is an implementation of DFT) algorithms is presented in Section 9.
-A generation algorithm produces substantially more output data than input data.
-Such algorithms encode mathematical knowledge into data.
-Certain generation algorithms are purely theoretical calculations that involve no imprecise input, so all resulting uncertainty arises solely from rounding errors.
-Section 10 presents a generation algorithm that generates a sine function table using trigonometric relations and two precise inputs: $sin(0)=0$ and $sin(\pi/2)=1$.
-A reduction algorithm yields significantly fewer output data than input data, as in numerical integration, or in the statistical characterization of a data set.
-In this process, certain information is lost while other information is extracted.
-As a statistical approach, variance arithmetic inherently carries out statistical reduction.
-For example, the result of averaging $N$ imprecise values with each uncertainty close to $\delta x$ is $\overline{x}\pm\frac{\delta x}{\sqrt{N}}$, reflecting central limit theorem [4] naturally.
+When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained.
+The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations.
+Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context.
+The input noise range that yields ideal coverage defines the ideal application range for the analytic function.
 5 Polynomial
@@
 $\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}$ $=\sum_{j=0}^{N}\tilde{x}^{j}P_{j},\;\;\;P_{j}\equiv\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (5.1)
-5.1 Tracking Rounding Error
+Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can only reach half of the maximal expansion order of Formula (2.3), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian.
+5.1 Residual Error
 Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x-axis).
@@
 The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
-Variance arithmetic can track rounding errors effectively without the need for additional rules.
-Figure 14 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$, where the polynomial $\sum_{j=0}^{224}x^{j}$ is computed using Formula (5.1), $\frac{1}{1-x}$ is computed using Formula (2.27), and $x$ is initiated as a floating-point value.
-Because $\eta(2n)$ is limited to $2n\leq 448$, Formula (5.1) for polynomial evaluation is restricted to $N\leq 224$, so that $\sum_{j=0}^{224}x^{j}$ has lower expansion order than that of the statistical Taylor expansion of $\frac{1}{1-x}$.
-Figure 14 shows:
-  • When $x\in[-0.73,0.75]$, the required expansion order is no more than $224$, which indicates that the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$.
-A detailed analysis indicates that the maximal residual error is $4$ times the ULP of $\frac{1}{1-x}$.
-The calculated uncertainty bounds the residual error effectively for all $x\in[-0.73,0.75]$.
-  • When $x\not\in[-0.74,+0.75]$, the required expansion order exceeds $224$, so that the residual error arises from the insufficient expansion order of $\sum_{j=0}^{224}x^{j}$.
-The residual error magnitude increases as $|x|\rightarrow 1$, reaching approximately $50$ when $x=0.98$.
+Figure 12 shows the residual error of $\sum_{j=0}^{224}x^{j}-1(1-x)$.
+It also displays the required expansion orders for 1/(1x)1/(1-x), which are all less than $224$.
+Therefore, the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$.
+A detailed analysis indicates that the maximal residual error is $4$ times the ULP of 1/(1x)1/(1-x).
+In all cases, the calculated uncertainty bounds the residual error effectively for all $x$, with an error deviation of $2.60$ when the expansion order is less than $224$.
+Variance arithmetic can provide proper coverage for rounding errors.
 5.2 Continuity
+In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space.
+For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number $n$.
+The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$.
+5.3 Distributional Hole
 Figure 13: Histograms of normalized errors for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
-$n\pm d$ $0\pm 10^{-6}$ $1\pm 10^{-6}$ $2\pm 10^{-6}$ $3\pm 10^{-6}$
-Upper Bound $\delta x$ $0.2006$ $0.2014$ $0.2018$ $0.2020$
-Value $1\mp 2.155\;10^{-8}$ $1\pm 2.073\;10^{-8}$ $041\pm 6.065\;10^{-8} 1.1$$122\pm 1.328\;10^{-7}$
-Uncertainty $0+2.127\;10^{-7}$ $0.201-1.358\;10^{-6}$ $0.407\pm 2.201\;10^{-7}$ $0.654\pm 2.784\;10^{-7}$
-Table 1: The result value and uncertainty of $(1\pm\delta x)^{n\pm d}$ vs $(1\pm\delta x)^{n}$, in which $n$ is a natural number, $0<d\ll 1$, and $\delta x$ is the upper bound for $(1\pm\delta x)^{n\pm d}$.
-The value and the uncertainty are expressed as the difference with the corresponding value and uncertainty of those of $(1\pm\delta x)^{n}$.
-In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space.
-For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number.
-The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$.
-Table 1 shows that the result of $(1\pm\delta x)^{n\pm d}$ where $0<d\ll 1$ is very close to that of $(1\pm\delta x)^{n}$, even though the former has an upper bound for $\delta x$, while the latter does not.
-A statistical bounding range in variance arithmetic can include a distributional pole if the analytic function is defined in its vicinity.
+Unlike a distributional zero, a statistical bounding range in variance arithmetic can include a distributional pole, such as around $(0\pm\delta x)^{c},c>1$.
 The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram.
-Figure 15 illustrates the histograms of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
+Figure 13 illustrates the error distributions of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
   • When the second derivative is zero, the resulting distribution is symmetric two-sided and Delta-like, such as when $n=3,x=0$.
@@
   • When the second derivative is negative, the resulted distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
-In each case, the transition from $x=0$ to $x=0.2$ is continuous.
-A statistical bounding range in variance arithmetic cannot encompass more than one distributional pole, as this condition causes the corresponding statistical Taylor expansion becomes negative.
-An illustrative example is $\sin(x)$ shown in Figure 12.
-A statistical bounding range in variance arithmetic cannot include any distributional zero because the result will diverge, such as at $x=0$ for both $(x\pm\delta x)^{c},c<1$ and $\log(x\pm\delta x)$.
+In each case, the transition from $x=0$ to $x=0.2$ is continuous for the resulting mean and deviation.
+6 Matrix Calculations
+Figure 14: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
+Figure 15: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (6.3).
+Figure 16: Linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis).
+The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
+$\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (6.1)
+$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (6.2)
+Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20].
+Formula (6.1) and (6.2) gives the statistical Taylor expansion of the determinant of a square matrix $\mathbf{M}$ of size $n$, when the uncertainties of matrix elements are all independent of each other.
+In Formula (6.2), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices.
+The square matrix whose element is $(-1)^{i+j}|m_{j,i}|$ is defined as the adjugate matrix [20] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$, in which $m_{j,i}$ is the element of $\mathbf{M}$ at row index $j$ and column index $i$.
+Each $m_{j,i}$ is an integer between $2^{8},+2^{8}]$ so that $\mathbf{M}^{A}$ can be calculated precisely using integer arithmetic.
+By adding Gaussian noise of deviation $\delta x$ to each $m_{j,i}$, $\widetilde{\mathbf{M}}$ is created.
+The value error of $\mathbf{M}^{A}$ is the difference between $\widetilde{\mathbf{M}^{A}}$ and $\mathbf{M}^{A}$, while the result deviation is calculated by Formula (6.2), so that $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ is a distribution test.
+Figure 14 shows that the error deviations of $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ are very close to $1$ except when the matrix size is $8$ and the input uncertainty is $\delta x=10^{-17}$.
+This abnormality is caused by floating-point rounding errors only at this point, when the element values exceed the significand range of 64-bit floating-point representation, to result in a error deviation of $3.2$.
+Figure 14 demonstrates a typical distribution test of error deviation versus input noise deviation $\delta x$ and another specific dimension, with ideal coverage for all region except when $\delta x\rightarrow 0$.
+$\mathbf{M}\times\mathbf{M}^{A}$ $=\mathbf{M}^{A}\times\mathbf{M}=|\mathbf{M}|\mathbf{I};$ (6.3)
+$\mathbf{M}^{-1}$ $\equiv\;\mathbf{M}^{A}/|\mathbf{M}|;$ (6.4)
+$\delta^{2}\left(\begin{matrix}w,x\\ y,z\end{matrix}\right)^{-1}\simeq$ $\;\frac{\left(\begin{matrix}z^{4},x^{2}z^{2}\\ y^{2}z^{2},x^{2}y^{2}\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}y^{2}z^{2},w^{2}z^{2}\\ y^{4},w^{2}y^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{4}}+$ (6.5)
+$\;\frac{\left(\begin{matrix}x^{2}z^{2},x^{4}\\ w^{2}z^{2},w^{2}x^{2}\end{matrix}\right)(\delta y)^{2}+\left(\begin{matrix}x^{2}y^{2},w^{2}x^{2}\\ w^{2}y^{2},w^{4}\end{matrix}\right)(\delta z)^{2}}{(wz-xy)^{4}};$
+Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [20].
+Formula (6.3) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ which leads to the definition of inverse matrix $\mathbf{M}^{-1}$ in Formula (6.4) [20].
+$\tilde{\mathbf{M}}\times\tilde{\mathbf{M}}^{A}-|\tilde{\mathbf{M}}|\mathbf{I}$ tests each element value so that it is a value test.
+Figure 15 shows the error deviation of a typical value test: the error deviation decreases linearly with increasing input uncertainty deviation $\delta x$, because the value should match unless at $\delta x\rightarrow 0$ when floating-point rounding errors dominate.
+The matrix size is the specific dimension in Figure 15.
+Because an element of the original matrix $\mathbf{M}$ appears multiple times in Formula (6.4), its result variance is very complicated using Formula (2.3).
+For example, Formula (6.5) shows the simplest case for Formula (6.4): the first-order approximation of a 2x2 matrix.
+Contrary to traditional approach, statistical Taylor expansion uses Formula (6.4) for matrix inversion instead of Gaussian elimination [5], because logically, the result should be symmetric for all matrix elements, as what Formula (6.5) has demonstrated.
+In Formula (6.4), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$.
+Figure 16 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices.
+As a reference, Figure 16 presents the Hilbert matrix [20] for each matrix size and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number.
+Thus, determinant precision can replace matrix condition number.
 7 Mathematical Library Functions
-Figure 16: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $e^{x\pm\delta x}$, for different $x$ (x-axis), and different $\delta x$ (legend).
-Figure 17: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $\log(x\pm\delta x)$, for different $x$ (x-axis), and different $\delta x$ (legend).
-Figure 18: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $(1\pm\delta x)^{c}$, for different $c$ (x-axis), and different $\delta x$ (legend).
-Figure 19: Error deviation for $(1\pm\delta x)^{c}$ as a function of $c$ and $\delta x$.
-The x-axis represents $c$ value between 2-2 and +3+3.
-The y-axis represents $\delta x$ value between $10^{-16}$ and $1$.
-The z-axis shows the corresponding error deviation.
-Figure 20: Calculated uncertainties versus measured value deviations (left y-axis), along with the measured error deviations (right y-axis) for $\sin(x\pm\delta x)$, for different $x$ (x-axis), and different $\delta x$ (legend).
+Basic Function $e^{x\pm\delta x}$ $\log(x\pm\delta x)$ $(1\pm\delta x)^{c}$ $\sin(x\pm\delta x)$
+Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
+Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.978\pm 0.150$
+Table 1: The result error deviation of selected basic functions using variance arithmetic when $\delta x>10^{-15}$.
+The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverge regions are excluded.
+The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
 Figure 17: Error deviation for $\sin(x\pm\delta x)$ as a function of $x$ and $\delta x$.
@@
 Figure 19: Normalized errors of $(x^{p})^{\frac{1}{p}}-x$ as functions of $x$ and $p$.
-Formula (2.20), (2.23), (2.26), and (2.29) are evaluated using the corresponding mathematical library functions exp, log, sin, and pow, respectively.
-At each point $x$ with an input uncertainty $\delta x$, the result uncertainty is calculated by variance arithmetic. The corresponding error deviation is determined by sampling as:
-  1. 1.
-    Generate $10000$ samples from a Gaussian noise distribution, each with $\delta x$ as the distributional deviation, and construct $\tilde{x}$ by adding the sampled noise to $x$.
-  2. 2.
-    For each $\tilde{x}$, use the corresponding library function to compute the value error as the difference between the outputs obtained using $\tilde{x}$ and $x$ as inputs.
-  3. 3.
-    Divide the value error by the result uncertainty to obtain the normalized error.
-  4. 4.
-    Calculate the standard deviation of the normalized errors to determine the error deviation.
-6.1 Exponential
-Figure 16 shows that the calculated uncertainties obtained using Formula (2.20) align closely with the measured value deviations for $e^{x+\delta x}$.
-Consequently, all error deviations remain very close to $1$, even though both the uncertainty and the value deviations increase exponentially with $x$ and $\delta x$.
-6.2 Logarithm
-Because $\log(x)$ has a distributional zero at $x=0$ all $log(x\pm\delta x)$ values are rejected when $P(x)>1/5$.
-Figure 17 shows that the uncertainties calculated using Formula (2.23) align closely with the measured value deviations for $\log(x+\delta x)$, and the resulting error deviations remain very close to $1$ except when $P(x)>0.20086$.
-6.3 Power
-Figure 18 shows that the uncertainties calculated for $(1\pm\delta x)^{c}$ using Formula (2.29) fit closely with the measured value deviations for $(1+\delta x)^{c}$, with the error deviations remaining near 11.
-Figure 19 reveals that the error deviations of $\sin(x+\delta x)$ remain close to 11 for all input exponents $c$ and input uncertainties $\delta x>10^{15}$.
-6.4 Sine
-Figure 20 shows that the uncertainties calculated using Formula (2.26) correspond closely to the measured value deviations for $\sin(x+\delta x)$.
-It also reveals that $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$:
-  • When $x=0$, $\sin(x)\simeq x$, so that $\delta^{2}\sin(x)\simeq(\delta x)^{2}$.
-  • When $x=\pi/2$, $\sin(x)\simeq 1$, so that $\delta^{2}\sin(x)\simeq 0$.
-Since $\sin(x)$ has distributional poles at $x=\pm\pi/2$, Figure 21 shows that the error deviation for $\sin(x+\delta x)$ equals $1$ except when $x=\pm\pi/2$ and $\delta x<10^{-8}$, matching the expected Delta distribution these poles.
-Elsewhere, the error deviations remain close to $1$.
-6.5 Numerical Errors for Library Functions
-The combined numerical error of the library functions $e^{x}$ and $\log(x)$ is evaluated as either $\log(e^{x})-x$ or $e^{\log(x)}-x$.
-Using either variance arithmetic or conventional floating-point library functions produces identical value errors.
-Figure 22 shows the corresponding result uncertainties, which reach a minimum when the input is a 2’s fraction such as $1$ or $1/2$.
-Figure 22 also shows that $e^{\log(x)}$ exhibits much smaller error than $\log(e^{x})$.
-For $\log(e^{x})-x$, the error deviation is $0.409$ when $|x|\leq 1$ or approaches zero otherwise.
-The reason for the unexpectedly small value errors for $e^{\log(x)}-x$ is not yet clear.
-The numerical error of the library function $x^{p}$ is computed as $(x^{p})^{1/p}-x$.
-Figure 23 shows that the normalized errors do not depend on either $x$ or $p$, resulting in an error deviation $0.548$.
-The numerical errors of the library functions $\sin(x)$, $\cos(x)$, and $\tan(x)$ will be examined in greater detail in Section 9.
-6.6 Summary
-Formula (2.20), (2.23), (2.29), and (2.26) provide effective estimates of the corresponding library functions.
-When added noise exceeds $10^{-15}$ precision, ideal coverage is achieved except near a distributional pole where the error deviation approaches zero.
-In all other cases, proper coverage is attainable.
-7 Matrix Calculations
-7.1 Matrix Determinant
-Let vector $p_{1},p_{2}\dots p_{n}]_{n}$ denote a permutation of the vector $(1,2\dots n)$ [14].
-Let $\mathdollar[p_{1},p_{2}\dots p_{n}]_{n}$ denote the permutation sign of $p_{1},p_{2}\dots p_{n}]_{n} [$14].
-Formula (7.1) defines the determinant of a $n$-by-$n$ square matrix $\mathbf{M}$ with the element $x_{i,j},i,j=1\dots n$ [14].
-The sub matrix $\mathbf{M_{i,j}}$ at index $(i,j)$ is formed by deleting the row $i$ and column $j$ of $M$, whose determinant is given by Formula (7.2) [14].
-For discussion simplicity, sub determinant $|\mathbf{M}|_{i,j}$ in Formula (7.2) contains the permutation sign, which is different from the determinant of the sub matrix $|\mathbf{M}_{i,j}|$ [14] that treats the sub matrix $\mathbf{M_{i,j}}$ as an independent matrix.
-Formula (7.3) holds for the arbitrary row index $i$ or the arbitrary column index $j$ [14].
-$|\mathbf{M}|$ $\equiv\sum_{[p_{1}\dots p_{n}]_{n}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}x_{k,p_{k}};$ (7.1)
-$|\mathbf{M}|_{i,j}$ $\equiv\sum_{[p_{1}\dots p_{n}]_{n}}^{p_{i}=j}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}^{k\neq i}x_{k,p_{k}};$ (7.2)
-$|\mathbf{M}|$ $=\sum_{j=1\dots n}|\mathbf{M}_{i,j}|x_{i,j}=\sum_{i=1\dots n}|\mathbf{M}|_{i,j}x_{i,j};$ (7.3)
-Let $i_{1},i_{2},\dots>$} denote an ordered permutation of a subset from $1\dots n$, and $i_{1},i_{2},\dots]$} an unordered permutation [14].
-Apply Formula (7.3) progressively to $M_{i,j}$, to expand Formula (7.2) as (7.4), and Formula (7.1) as (7.5).
-The (nm)(n-m)-by-(nm)(n-m) sub matrix in $|\mathbf{M}_{<i_{1}\dots i_{m}>_{n}[j_{1}\dots j_{m}]_{n}}|$ is obtained by deleting the rows in $\{i_{1}\dots i_{m}\}$ and the columns in $\{j_{1}\dots j_{m}\}$.
-This leads to Formula (7.6).
-$|\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\equiv\sum_{[p_{1}\dots p_{n}]_{n}}^{k\in\{i_{1}\dots i_{m}\}:p_{k}=j_{k}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}^{k\not\in\{i_{1}\dots i_{m}\}}x_{k,p_{k}};$ (7.4)
-$|\mathbf{M}|$ $=\sum_{{[j_{1}\dots j_{m}]}_{n}}|\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{k=1}^{m}x_{i_{k},j_{k}};$ (7.5)
-$\left||\mathbf{M}|_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\right|=\left||\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<j_{1}\dots j_{m}>_{n}|}\right|;$ (7.6)
-[…]
-[…]
-$=\sum_{m=0\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$
-$\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (7.8)
-$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<j_{1}\dots j_{m}>_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (7.9)
-Formula (7.8) and (7.9) assume that the uncertainties of matrix elements are all independent of each other.
-This assumption maximized the result uncertainties.
-For discussion simplicity, other uncertainty assumptions are ignored in this paper.
-7.2 Adjugate Matrix
-Figure 24: Uncertainty means (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
-Figure 25: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
-Figure 26: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (7.10).
-Figure 27: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (7.11).
-The square matrix whose element is $a_{i,j}=(-1)^{i+j}|\mathbf{M}_{j,i}|$ is defined as the adjugate matrix [14] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$.
-Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [14].
-Formula (7.10) and (7.11) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ [14].
-$\mathbf{M}\times\mathbf{M}^{A}$ $=\mathbf{M}^{A}\times\mathbf{M}=|\mathbf{M}|\mathbf{I};$ (7.10)
-$|\mathbf{M}|\;\mathbf{M}^{A}$ $=|\mathbf{M}^{A}|\;\mathbf{M};$ (7.11)
-To test Formula (7.9):
-  1. 1.
-    A matrix $\mathbf{M}$ is constructed using random integers uniformly distributed in the range of $2^{8},+2^{8}]$, which has a distribution deviation of $2^{8}/\sqrt{3}$.
-The integer arithmetic ensures that $|\mathbf{M}|$, $\mathbf{M}^{A}$, and $|\mathbf{M}^{A}|$ are precise.
-  2. 2.
-[…]
-The difference between $\widetilde{\mathbf{M}}^{A}$ and $\mathbf{M}^{A}$ defines the Adjugate Test.
-As seen in Figure 24, the uncertainty means increase exponentially with both the input noises and the matrix size; however, such linear increase is segmented into two areas at input precision $10^{-10}$.
-Figure 25 shows that the ideal coverage is achieved for the input precision except at matrix size $8$ and input precision $10^{-17}$.
-The existence of ideal coverage validates Formula (7.9).
-Formula (7.10) defines the Forward Test.
-Figure 26 shows that the difference of the two sides of Formula (7.10) is precise zero, whether it is $\widetilde{\mathbf{M}}\times\widetilde{\mathbf{M}}^{A}-|\widetilde{\mathbf{M}}|\mathbf{I}$, or $\widetilde{\mathbf{M}}^{A}\times\widetilde{\mathbf{M}}-|\widetilde{\mathbf{M}}|\mathbf{I}$.
-The difference in Formula (7.10) is closer to Delta distribution for larger input precision, because without input noise, rounding errors dominate value errors.
-The validation of Formula (7.9) leads naturally to the validation of Formula (7.10).
-Formula (7.11) defines the Roundtrip Test.
-Similarly, Figure 27 validates Formula (7.11).
-Because roundtrip test is no longer linear, error deviation in Figure 27 no longer increases linearly with input precision.
-7.3 Floating Point Rounding Errors
-Figure 28: Histograms of normalized errors of the adjugate matrix as a function of matrix size without input noise (legend).
-Figure 29: Histograms of normalized errors of the adjugate matrix as a function of matrix size with $10^{-11}$ input noise (legend).
-The significance of the conventional floating-point representation [8] has a $53$-bit resolution.
-As shown in Figure 28, the histogram of the normalized errors is Delta distributed for the matrix size less than $8$, because the adjugate matrix calculation involves about $8\times 6=48$ significand bits for a matrix size $7$.
-When the matrix size is $8$, $8\times 7=56$ significand bits are needed so rounding occurs, which results in non-Delta distribution in Figure 28.
-The rounding error is also the reason why only at matrix size 8 and input noise precision $10^{-10}$, error deviation is no longer $1$ in Figure 25.
-With $10^{-11}$ noises added to the input, Figure 29, the distribution becomes Gaussian with a hint of Delta distribution.
-Such Delta-like distribution persists until the input noise precision reaches $10^{-10}$, which is also the transition of the two trends in Figure 24.
-Figure 29 shows the distribution when the input noise precision is $10^{-11}$, where the distinction due to matrix size vanishes.
-7.4 First Order Approximation
-[…]
-[…]
-$|\widetilde{\mathbf{M}}|\simeq\sum_{[p_{1}\dots p_{n}]_{n}}\mathdollar[p_{1}\dots p_{n}]_{n}\prod_{k=1\dots n}(x_{k,p_{k}}+\tilde{x}_{k,p_{k}});\;\;\;\Rightarrow\;\;\;\delta^{2}|\mathbf{M}|\simeq\sum_{i}^{n}\sum_{j}^{n}M_{i,j}(\delta x_{i,j})^{2};$ (7.12)
-Figure 30 contains the result of applying Formula (7.12). It is very similar to Figure 25, validating Formula (7.12).
-7.5 Matrix Inversion
-Figure 31: Linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis).
-The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
-$\mathbf{M}^{-1}\equiv$ $\;\mathbf{M}^{A}/|\mathbf{M};$ (7.13)
-$\mathbf{M}^{-1}\times\mathbf{M}=\mathbf{M}\times\mathbf{M}^{-1}=\mathbf{I};$ (7.14)
-$(\mathbf{M}^{-1})^{-1}=\mathbf{M};$ (7.15)
-An inverse matrix is defined by Formula (7.13), which satisfies Formula (7.14) and (7.15) [14].
-However, this definition is seldom used conventionally to calculate inverse matrix due to large uncertainty response ratio which appears as small input uncertainties causing large output uncertainties [12][14].
-Traditionally, matrix condition number [14] is a proxy for uncertainty response ratio of matrix inversion.
-In Formula (7.13), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$.
-Figure 31 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices.
-As a reference, Figure 31 presents the Hilbert matrix [14] for each matrix size, and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number.
-Thus, determinant precision can replace condition number to estimate uncertainty response ratio of matrix inversion.
-$\mathbf{M}=$ $\;\left(\begin{matrix}w,x\\ y,z\end{matrix}\right);\;\;\;\mathbf{M}^{-1}=\frac{\left(\begin{matrix}z,-x\\ -y,w\end{matrix}\right)}{wz-xy};$ (7.16)
-$\delta^{2}\mathbf{M}^{-1}\simeq$ $\;\frac{\left(\begin{matrix}z^{4},x^{2}z^{2}\\ y^{2}z^{2},x^{2}y^{2}\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}y^{2}z^{2},w^{2}z^{2}\\ y^{4},w^{2}y^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{4}}+$ (7.17)
-$\;\frac{\left(\begin{matrix}x^{2}z^{2},x^{4}\\ w^{2}z^{2},w^{2}x^{2}\end{matrix}\right)(\delta y)^{2}+\left(\begin{matrix}x^{2}y^{2},w^{2}x^{2}\\ w^{2}y^{2},w^{4}\end{matrix}\right)(\delta z)^{2}}{(wz-xy)^{4}};$
-$\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}\simeq$ $\;\frac{\left(\begin{matrix}2z^{3},-2xz^{2}\\ -2yz^{2},2xyz\end{matrix}\right)(\delta w)^{2}+\left(\begin{matrix}2y^{2}z,-2wyz\\ -2y^{3},2wy^{2}\end{matrix}\right)(\delta x)^{2}}{(wz-xy)^{3}}+$ (7.18)
-$\frac{\left(\begin{matrix}2x^{2}z,-2x^{3}\\ -2wxz,2wx^{2}\end{matrix}\right)(\delta y)^{2}+\left(\begin{matrix}2wxy,-2w^{2}x\\ -2w^{2}y,2w^{3}\end{matrix}\right)(\delta z)^{2}}{(wz-xy)^{3}};$
-Variance arithmetic is path-independent, so it computes inverse matrix directly from the definition of Formula (7.13).
-For example, for the inverse matrix $\mathbf{M}^{-1}$ of size $2$ in Formula (7.16), Formula (7.17) and (7.18) present first-order approximations for variance $\delta^{2}\mathbf{M}^{-1}$, and bias $\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}$, respectively.
-The resulting uncertainties of determinant $\delta^{2}|\mathbf{M}|$ and inversion $\delta^{2}\mathbf{M}^{-1}$, as well as the resulting bias $\overline{\mathbf{M}^{-1}}-\mathbf{M}^{-1}$, are not linear to each input uncertainty $\delta x_{i,j}$.
-Inverse variance $\delta^{2}\mathbf{M}^{-1}$ contains sum of all input variance $(\delta x_{i,j})^{2}$ and their higher-order permutation products, such that the uncertainty response ratio for matrix inversion should roughly equal the matrix size.
-Thus, an uncertainty response ratio for matrix inversion is inherently large, independent of computational path.
-A seemly small uncertainty response ratio using Gaussian elimination [12][14] is probably a path-dependent artifact; while the conventionally “bad” method of applying Formula (7.13) for matrix inversion [12][14] is actually the correct one, except missing resulting uncertainty when using floating-point arithmetic.
-All conventional path-dependent results are questionable.
-Only statistical Taylor expansion presents a complete picture of resulting uncertainty for an analytic expression.
-It is doubtful if Formula (7.15) still holds for uncertainty in statistical Taylor expansion, because it seems that uncertainty response ratio can only increase in matrix inversion according to Formula (7.17).
-On the other hand, because the bias in Formula (7.18) can be either positive or negative, it is expected that Formula (7.15) still holds for value.
+Table 1 shows that Formula (2.15), (2.17), (2.20), and (2.22) provide almost perfect deviation for the result values which are calculated using corresponding math library functions by sampling from $x\pm\delta x$ in which $\delta x$ is Gaussian input noise deviation.
+Table 1 shows the results from Python, which differs slightly from those from either C++ or Java.
+The result error deviation is similar to Figure 14 but with the specific dimension as $x$ in $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$, or $c$ in $(1+\delta x)^{c}$.
+The coverage is only proper when $\delta x<10^{-15}$.
+Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$.
+Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation.
+As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$.
+The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ will be examined in greater detail in Section 10.
+To test $f^{-1}(f(x))-x=0$ when $\delta x=0$ for the library functions:
+  • Figure 18 shows that the value errors in $e^{\log(x)}-x$ are much less than those in $\log(e^{x})-x$. For $\log(e^{x})-x$, the error deviation is $0.41$ when $|x|\leq 1$, or 0 otherwise.
+  • Figure 19 shows that the error deviation for $(x^{p})^{1/p}-x$ is $0.56$, independent on neither $x$ nor $p$.
+The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has almost no value error, are not clear.
 8 Moving-Window Linear Regression
@@
 This observation suggests that an error deviation exceeding 1 may indicate the presence of unspecified additional input errors beyond rounding errors, such as numerical errors in mathematical library functions.
+9 Regressive Generation of Sin and Cos
+Figure 22: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ versus $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
+Formula (9.2) and Formula (9.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (9.1).
+Formula (9.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, so that value errors will not accumulate when the regression order increases.
+$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (9.1)
+$\sin(\frac{\alpha+\beta}{2})=\sqrt{\frac{1-\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1-\cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)}{2}};$ (9.2)
+$\cos(\frac{\alpha+\beta}{2})=\sqrt{\frac{1+\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1+\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)}{2}};$ (9.3)
+$\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (9.4)
+Formula (9.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation, resulting in excessive uncertainty.
+In Figure 22, the Quart sine function exhibits uncertainties proportional to $\sin(\frac{\pi}{2^{L}})$, whereas the regression sine function shows increasing uncertainties as $x\rightarrow 0$.
+Unlike catastrophic cancellation in floating-point arithmetic, variance arithmetic uses very coarse precision to signal that the regression algorithm is unfit to compute $\sin(x\rightarrow 0)$.
 10 FFT (Fast Fourier Transformation)
 10.1 DFT (Discrete Fourier Transformation)
-For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (9.1) and (9.2), respectively [12].
-In these expressions, $k$ denotes the time index while $n$ represents the frequency index.
-The frequency index and time index are not necessarily associated with physical time unit or frequency unit, respectively; rather, the naming convention provides a convenient way to distinguish between the two opposite domains in DFT: the waveform domain $h[k]$ and the frequency domain $H[n]$.
 $H[n]$ $=\sum_{k=0}^{N-1}h[k]\;e^{\frac{i2\pi}{N}kn};$ (10.1)
 $h[k]$ $=\frac{1}{N}\sum_{n=0}^{N-1}H[n]\;e^{-\frac{i2\pi}{N}nk};$ (10.2)
-9.2 Modeling Errors of DFT
+For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (10.1) and (10.2), respectively [5].
+As a convention, $k$ denotes time index for $h[k]$ as a waveform, while $n$ represents frequency index for $H[n]$ as a spectrum.
 Figure 23: The DFT spectrum $H[n]$ of signal $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, as intensity (y-axis) and phase (embedded y-axis) versus frequency index $n\in[0,18]$ (x-axis and embedded x-axis) for different signal frequency $f$ (legend).
 This result agrees with both theoretical formula [17] and numerical computation from any mathematical libraries such as SciPy.
-Although mathematically self-consistent, DFT implies a periodic boundary condition in the time domain [29].
-Consequently, it is only an approximation for the mathematically defined continuous Fourier transform (FT) [29].
 For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [5].
@@
 If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise.
-Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [12].
+Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics
 To avoid the modeling errors inherent in DFT, only Formula (10.1) and (10.2) are used in this paper.
 10.2 FFT (Fast Fourier Transformation)
-When $N=2^{L}$, where $L$ is a natural number, the generalized Danielson-Lanczos lemma [12] can be applied to DFT to produce FFT [12].
-  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations as Formula (2.14), (2.15), (2.16), and (2.17).
+When $N=2^{L}$, where $L$ is a natural number named FFT order, the generalized Danielson-Lanczos lemma can be applied to DFT to produce FFT [5].
+  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations such as Formula (2.10), (2.11), (2.12), and (2.13).
   • When $L$ is large, the substantial volume of input and output data enables high-quality statistical analysis.
-  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of multiplications.
-  • Each step in the forward transformation doubles the variance, so the uncertainty mean increases with the FFT order $L$ as $\sqrt{2}^{L}$.
-Because the reverse transformation divides the result by $2^{L}$, its uncertainty mean decreases with $L$ as $\sqrt{1/2}^{L}$.
-Consequently, the uncertainty mean for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
+  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of two multiplications.
+  • Each step in the forward transformation doubles the variance, so the uncertainty deviation increases with the FFT order $L$ as $\sqrt{2}^{L}$.
+Because the reverse transformation divides the result by $2^{L}$, its uncertainty deviation decreases with $L$ as $\sqrt{1/2}^{L}$.
+Consequently, the uncertainty deviation for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
   • The forward and reverse transformations are identical except for a sign difference, meaning that they are essentially the same algorithm, and any observed difference arises solely from the input data.
 In normal usage, forward and reverse FFT transforms differ in their data prospective of time domain versus frequency domain:
-  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zeros, causing its uncertainties to grow more rapidly.
-  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zeros) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
+  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly by cancellation.
+  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly by spreading.
+The question is whether variance arithmetic can work properly in these two contrary cases.
 10.3 Testing Signals
@@
     $H[n]$ $=\frac{dG}{dy}=\begin{cases}n=0:\;\;\;\frac{N(N-1)}{2}\\ n\neq 0:\;\;\;-\frac{N}{2}(1+i\frac{\cos(n\frac{\pi}{N})}{\sin(n\frac{\pi}{N})})\end{cases};$ (10.3)
-Empirically, except when using trigonometric library directly for FFT transformations of clean Sin and Cos signals:
-  • The results obtained from Sin and Cos signals are statistically indistinguishable.
-  • Similarly, the results from Sin and Cos signals at different frequencies does not show significant differences statistically except in the situation of numerical error resonance.
-Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals.
 10.4 Trigonometric Library Errors
@@
   2. 2.
-    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
+    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
   3. 3.
@@
     The values of $\sin(\pi j/2^{L})$ are extended to all the integer value of $j$ by leveraging the periodicity of $\sin(2\pi j/2^{L})$.
-The constructed indexed $\sin(x)$ is referred to as the Quart indexed sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
-Because the the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides numerical accuracy compared to Library function.
-  • Figure 35 shows that the value difference between the Quart and Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
-  • Figure 36 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but are $10^{2}$ times larger than those observed for $\sin(x)$.
-Therefore, the linear spectrum in Formula (9.3) may contain significant numerical errors when computed using library sine functions.
-9.6 Refine Quart Sine Function
-Figure 37: Value errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) versus $x$ (x-axis) and $L$ computed using either the Prec or the Quart sine function (legend).
-Figure 38: The value error histogram for the Quart sine function when it is calculated using 128-bit floating-point arithmetic and cast to 64-bit floating-point representation.
-If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the resulting indexed sine function is referred to as the Prec sine function.
-  • Figure 37 compares the value errors and the uncertainties of $\sin(x)^{2}+\cos(x)^{2}-1$ between the Prec sine function and the Quart sine function, and demonstrates that the Prec sine functions has smaller uncertainties numerically than the Quart sine functions, but larger value errors statically, presumably due to rounding errors introduced by the casting.
-As suggested in Figure 37, the Prec sine function has about $5\%$ more non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors, and $14\%$ non-zero $\sin(x)^{2}+\cos(x)^{2}-1$ value errors exceeding the corresponding uncertainties.
-These difference indicate that casting a 128-bit floating-point value to a 64-bit floating-point value is less accurate in the 64-bit floating-point arithmetic.
-  • Figure 38 demonstrates that the value errors of the Prec sine function after normalized by the corresponding value ULP is uniform distributed with in the half bit of ULP.
-Because variance arithmetic assumes that the value error is uniformly distributed within ULP of each value, the Quart sine function overestimates value errors by one times.
+The constructed indexed $\sin(x)$ is referred to as the Quart sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
+Because the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides superior numerical accuracy compared to the Library sine function.
+  • Figure 24 shows that the absolute value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
+  • Figure 25 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$ but are $10^{2}$ times larger than those observed for $\sin(x)$.
+Therefore, the linear spectrum in Formula (10.3) contains significantly larger numerical errors when computed using library sine functions.
 10.5 Using Quart Sine for Sin/Cos Signals
-Figure 39: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the Quart sine function after the forward transformation.
-The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
-Figure 40: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using the the Quart sine function after the reverse transformation.
-The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
-Figure 41: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
-The variance arithmetic is implemented using python.
 Figure 26: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using the Quart sine function. The FFT order is $18$.
-Using the Quart sine function, for a sine wave with a frequency of $3$, Figure 39 displays the value error spectrum for the forward transformation, while Figure 40 presents the value error waveform for the reverse transformation.
-  • In the forward transformation, the value errors are slightly smaller than the corresponding result uncertainties, with an error deviation of $0.60$.
-In the reverse transformation, the error deviation is $0.55$.
-Both error deviations are less than $1$, confirming a slight overestimation of the sine uncertainty by ULP.
-  • The uncertainty mean of the forward transformation is $7$ times larger than that of the reverse transformation.
-Even with this difference, the uncertainty tracks the corresponding value error effectively.
-  • The value error waveform for the roundtrip transformation is qualitatively similar to that in Figure 40 with a similar error deviation of $0.55$.
-Figure 41 illustrates that the forward transformation differs from the reverse transformation significantly.
-  • In the forward transformation, error deviations reach their stable values around $0.65$ rapidly once the FFT order $L\geq 6$.
-  • In contrast, in the reverse transformation, error deviations stabilize more slowly with increasing FFT order, approaching their stable values around $0.90$ only when $L\geq 12$.
-This slow convergence of normalized errors in the reverse transformation is attributed to its input data which consists entirely of precise zeros except at the frequency indexes.
-  • At FFT order $18$, Figure 42 shows that the normalized error distribution for the reverse transformation is wider than that of the forward transformation, confirming the larger error deviation.
-The value error increases at least $10^{2}$ times faster in the forward transformation than in the reverse transformation.
-Despite these differences, error deviations are $0.65$ and $0.90$, suggesting that in variance arithmetic, uncertainties track value errors effectively in all cases, achieving proper coverage.
-9.8 Sensitivity to Quart Sine Functions
-Figure 43: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) using the Quart sine function.
-The variance arithmetic is implemented using C++.
-Figure 44: Histograms of normalized errors between the Quart sine functions of C++, Java, and Python in the unit of least significant bits.
-The reverse transformation of Sin/Cos signals is very sensitive to the underlying Quart sine function.
-If variance arithmetic is implemented using C++ instead of Python, Figure 43 shows that the error deviations of the reverse transform quickly reaches the stable values around $0.89$, while the forward transformation remains similar to its counterpart in Figure 41.
-Figure 43 also shows identical uncertainty means to their counterparts in Figure 41.
-Figure 44 shows that the difference of the Quart sine functions between Python and C++ is only about $1\%$ LSBs.
-Figure 44 also reveals that the difference of the Quart sine functions between Java and C++ is only about $0.1\%$ LSBs.
-Indeed, the results between Java and C++ are much more similar.
-Because the Python implementation gives smaller error deviations, its Quart sine function has less numerical errors on average.
-For discussion simplicity, the rest results are only based on the python implementation.
+With FFT order as the specific dimension, the error deviations for forward and reverse transformations resemble Figure 14, while those for round-trip transformation is almost identical to Figure 15, independent of Sin or Cos signals, or frequency of the signals.
+Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals.
+When the FFT order $L$ is less than 8, the error deviations oscillate around $1$ due to insufficient sample count of $2^{L}$.
+Even data for forward and reverse transformations are drastically different, variance arithmetic works effectively in all cases.
+When $L=18$ and $\delta x=0$, Figure 26 shows that the error distributions of Sin/Cos signal resemble Normal distribution, with an additional Delta-like distribution at $\tilde{z}=0$.
+The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in sine function.
 10.6 Using Library Sine for Sin/Cos Signals
-Figure 45: FFT value error specturm of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
-The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
+Figure 27: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
+Figure 28: FFT value error spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
+The legend distinguishes between uncertainty and value error.
+The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
 Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
-The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
-Figure 47: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
-Figure 48: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
-Using the Library sine function, for a sine wave with a frequency of $3$, Figure 45 displays the value error spectrum for the forward transformation, while Figure 46 presents the vale error waveform for the reverse transformation:
-  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions.
-This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 41 and 47.
-  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 35.
-  • When compared with Figure 42, Figure 48 also confirms that at FFT order $18$, the normalized errors distribution is significantly broader when using the Quart sine function.
-  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 47 that these in Figure 41.
-  • The stable error deviations using the Library sine function are much larger than their counterparts using Quart library functions: $10.1>0.65$ for the forward transformation, $6.3>0.90$ for the reverse transformation, and $1.47>0.94$ for the roundtrip transformation.
-Using the Library sine function, variance arithmetic barely achieves proper coverage for sine and cosine signals.
-9.10 Numerical Error Resonance Using Library Sine
-Figure 49: FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
-The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
+The legend distinguishes between uncertainty and value error.
+The x-axis represents the time index, while the y-axis represents both uncertainty and value error.
 Figure 30: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation.
 The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
-Figure 51: Error deviation (z-axis) of FFT forward transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
+The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
 Figure 31: Error deviation (z-axis) of FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
-In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 51.
+With FFT order as the specific dimension, the error deviations for the forward and the reverse transformations resemble Figure 14, while those for the round-trip transformation is almost identical to Figure 15.
+In addition, the error deviations are larger than $1$ when $\delta x<10^{-15}$ for the forward transformation and $\delta x<10^{-14}$ for the reverse transformation.
+When $\delta x=0$, the error deviations for the reverse transformation increase with FFT order, to $6.2$ at FFT order $18$.
+As shown in Figure 24, the Library sine function contains more numerical errors, so the error distribution for the reverse transformation using Library sine function in Figure 27 is more structured and broader than that in Figure 26 using Quart sine function, while the error distribution for the forward transformations is more similar.
+This difference is consistent with a larger error deviation $6.2>1$ for the reverse transformation, but a comparable error deviation $1.1>1$ for the forward transformation when $\delta x=0$.
+Using the Library sine function, for a sine wave with a frequency of $3$, Figure 28, 29, and 30 presents the value errors for the forward, the reverse, and the round-trip transformations, respectively.
+In the reverse transformation, value errors exhibit a clear trend of increasing with the time index.
 These large value errors appear systematic rather than random and visually resemble a resonant pattern.
-Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy.
-For example, Figure 49 shows that the FFT value error waveform of $\cos(6\frac{2\pi}{2^{6}}j)$ has much stronger resonant pattern than that of $\sin(3\frac{2\pi}{2^{6}}j)$ in Figure 46.
-In the forward transformation, although there is no visible resonant pattern in FFT value error specturm, the intensity of value errors increase with FFT order.
-In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 50.
-Also, it is completely absent when using the Quart sine function, as seen in Figure 39 and 40.
-Figure 51 and Figure 52 demonstrate that the error deviations increase with sine or cosine frequency.
-Figure 35 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 45, 46, and 49.
+Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy.
+In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 30, as well as when using the Quart sine function.
+Figure 31 demonstrates that the error deviations increase with sine or cosine frequency.
+Figure 24 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern in Figure 29.
 At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
-To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
+To suppress these numerical error resonances, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals.
+Such resonance of numerical errors can easily and mistakenly be taken as signals.
 10.7 Using Quart Sine for Linear Signal
-Figure 53: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Quart sine function.
+Figure 32: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Quart sine function.
+Figure 33: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Quart sine function.
 Figure 34: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
-Figure 53 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage.
-Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases more rapidly with FFT order than its counterpart in Figure 41.
-Such increase results in larger error deviations: At FFT order $18$, $0.89>0.65$ for the forward transformation, $1.08>0.90$ for the reverse transformation, $1.67>0.90$ for the roundtrip transformation, and wider distributions of the normalized errors in Figure 54 versus the corresponding histograms in Figure 42.
-Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
+Figure 32 and 33 show the error deviation for the forward and the reverse transformations, respectively.
+The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation.
+In other areas, both transformations achieve proper coverage with error values around $1$.
+When $L=18$ and $\delta x=0$, the error distribution of the reverse transformation in Figure 34 is narrower than that in Figure 27. The corresponding error deviations are $1.5<6.2$, respectively.
 10.8 Using Library Sine for Linear Signal
-Figure 55: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
 Figure 35: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
-As shown by the difference between Figure 35 and 36, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$.
-The key question is whether variance arithmetic can effectively track these additional unspecified errors.
-Figure 55 shows that proper coverage can no longer be achieved, as the value error increases faster than the uncertainty with rising FFT order.
-Figure 56 demonstrates that the normalized error distribution for the reverse transformation is no longer effectively bounded within three standard deviations.
-Although the normalized error distribution for the forward transformation using the Library sine function appears visually similar to that obtained using the Quart sine function in Figure 56, a detailed analysis reveals that the normalized error distribution contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal of the normalized errors, respectively.
-The presence of these extreme values is consistent with those delta-like numerical errors observed in Figure 36.
-The observed difficulty of variance arithmetic in tracking Linear signals computed using the Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified numerical errors.
-A potential solution is to develop a new sine function library implemented directly within the variance arithmetic framework, ensuring that all numerical calculations are performed with explicit uncertainty tracking.
-9.13 Ideal Coverage
-Figure 57: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
-Figure 58: Histograms of normalized errors for Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) computed using the Library sine function.
-The FFT order is $18$.
 Figure 36: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Library sine function.
 Figure 37: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Library sine function.
-Figure 61: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Quart sine function.
-Figure 62: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Quart sine function.
-Code Signal Sine function Forward Reverse Roundtrip
-Python Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
-Java, C++ Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
-Python Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-12}$
-Java, C++ Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-13}$
-Python Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
-Java, C++ Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
-Python Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
-Java, C++ Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
+Figure 38: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
+Figure 35 shows that the reverse error distribution when $\delta x=0$ is no longer bound.
+The difference between Figure 35 and Figure 34 is consistent with the large numerical errors as demonstrated in Figure 25.
+Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine.
+Figure 36 and 37 shows a much smaller ideal coverage areas than those in Figure 32 and 33.
+Because result deviations grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region.
+Outside the ideal coverage region, proper coverage cannot be achieved for the reverse transformation.
+Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order.
+At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable.
+Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function.
+Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
+Figure 38 shows that, even when variance arithmetic can no longer effectively track the value errors for the forward and the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as the plateau region at high $L$ and low $\delta x$ in Figure 38.
+Such error cancellation is due to dependence tracing in statistical Taylor expansion.
+10.9 Ideal Coverage
+Figure 39: Result error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
+Adding noise to the input can suppress unspecified input errors.
+After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, while the error distribution for the round-trip transformation becomes Delta.
+Figure 39 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
+  • As expected, the resulting uncertainty deviations for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
+  • As expected, the resulting uncertainty deviations for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
+  • As expected, the resulting uncertainty deviations for the round-trip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
+  • As expected, the resulting error deviations for the forward and the reverse transformations remain constant at $1$.
+  • As expected, the resulting error deviations for the round-trip transformations are far less than $1$ but increase with FFT order $L$ exponentially due to increasing calculation.
+Signal Sine function Forward Reverse Roundtrip
+Sin/Cos Quart $10^{-16}$ $10^{-12}$ $10^{-14}$
+Sin/Cos Library $10^{-16}$ $10^{-11}$ $10^{-12}$
+Linear Quart $10^{-11}$ $10^{-7}$ $10^{-8}$
+Linear Library $10^{-11}$ $10^{-3}$ $10^{-8}$
 Table 2: The measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
-Adding noise to the input can suppress unspecified input errors and improve the portion of specified versus unspecified input uncertainties.
-Table 2 shows the measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
-Figure 57 illustrates the error deviations and uncertainty means when applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage for all the transformations:
-  • As expected, the result uncertainty means for the forward transformations increase with FFT order $L$ as $\sqrt{2}^{L}$.
-  • As expected, the result uncertainty means for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
-  • As expected, the result uncertainty means for the roundtrip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
-  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$.
-Figure 58 presents the corresponding histogram.
-As expected, the normalized errors for the forward and reverse transformations follow Normal distributions, while those for the roundtrip transformation are Delta distributed around zero, indicating perfect recovery of the input uncertainties.
-According to Figure 58, adding either Gaussian or white noise to input results in the same Gaussian distribution of the normalized errors.
-Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, as expected since FFT is a linear algorithm [12].
-The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise.
-For Linear signals computed using the Library sine function, Figure 59 and 60 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively.
-Ideal coverage corresponds to the region where the error deviation equals $1$.
-Outside this region, proper coverage cannot be achieved for the reverse transformation.
-Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region.
-Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order.
-At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable.
-Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function.
-Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
-In contrast, for Linear signals computed using the Quart sine functions, Figure 61 and 62 show the results for the forward and reverse transformations, respectively.
-The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions.
-The forward transformation exhibits a broader ideal coverage region than the reverse transformations.
-While the forward transformation shows consistent over-estimation of the value errors in the proper coverage region, the reverse transformation demonstrates both over-estimation and under-estimation of the value errors in the proper coverage region.
-Figure 62 shows that the rounding errors balances the over-estimation of value errors in the Quart sine function in Figure 38, so that the error deviations within the proper coverage region differ from $1$ only marginally.
-Adding noises has another obvious benefit: Different implementations of the variance arithmetic have almost identical result when the input noise is larger than $10^{-16}$.
-For example, Figure 61 and 62 are almost identical using either C++, or Java, or Python.
-9.14 Input Uncertainty Recovery
-Figure 63: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
-Figure 64: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
-With input uncertainty unspecified, the value error waveform for the roundtrip transformation is qualitatively similar to that of the the reverse transformation, with error deviations close to $1$, such as in Figure 41, 43, 47, and 53, even when proper coverage can no longer achieve for both forward and reverse transformations in Figure 55.
-With input uncertainty unspecified, roundtrip transformation can always recover input uncertainty statistically.
-With input uncertainty specified, roundtrip transformation can always recover input uncertainty numerically, and the quality of the recover increases with the quality of input uncertainty specification.
-For example, when using the Library sine function:
-  • Figure 57 shows that the error deviations for the roundtrip transformation is less than $1$, suggesting that each input uncertainty is recovered numerically.
-These error deviations increases with FFT orders, because the error deviations without added noises increases with the FFT orders according to Figure 55.
-  • As a comparison, for Sin/Cos signal, because Figure 45 shows little increase with the FFT orders of the error deviations without added noises, the improvement of the numerical recovering in Figure 64 is almost independent of FFT orders.
-In both cases, error deviations decreases with the strength of the added noise.
-9.15 Calculation Response
-Figure 65: The FFT execution time (y-axis) versus the FFT order (x-axis) for the implementation using either C++, Java, or Python (legend), with each calculation consisted of $36$ roundtrip FFT transformations.
-The C++ executable is complied with -O3 -march=native optimization.
-Without such optimization, the C++ -O0 implementation is about a order of magnitude slower than the Java implementation.
-Figure 65 plots the execution time for the roundtrip FFT calculations using either C++, Java, or Python.
-As expected, in Figure 65, the execution time increases exponentially with the FFT order $L$ as $2^{L}$ for all the three implementations when the FFT order $L>10$, validating the calculation response.
-At first, it is surprising that in Figure 65 the Java implementation is 2 times faster than the C++ implementation which has been fully optimized for speed.
-Because FFT contains a large amount of independent and logically identical branches, the dynamic just-in-time Java compiler [37] can optimize according to data in real-time, compile the bytecodes into native machine codes, and parallelize the execution to the multiple cores and GPUs on a modern computer.
-Perhaps it is necessary to rethink multi-threading as a way to manually increase the time performance, because the dynamic just-in-time optimization of Java is superior to the static optimization of C++ for repeated and heavy computations with preload data such as FFT.
+Table 2 shows the minimal required noise to achieve ideal coverage for FFT transformations at FFT order $L=18$ for different signals and sine functions, which is consistent with the corresponding error distributions in Figure 26, 27, 34, and 35.
+This shows that an error distribution can qualify its input uncertainty quantification.
+Without knowing precise result, similar histogram can be constructed from the calculated mean $\overline{f}$ and deviation $\delta f$, and the result data set for $f$.
+It is worth investigating if such empirical histogram has similar power to reveal tracking of value errors by variance arithmetic.
+11 Conclusion and Discussion
 11.1 Summary
-The Library sine functions implemented using conventional floating-point arithmetic have been shown to contain numerical errors equivalent up to $10^{-3}$ of the input precision in the worst cases of FFT transformations.
-These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals.
-Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the numerical errors in library functions.
-This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations, but needs rigorous analysis for the ideal or proper coverage regions in each parameter space.
-The impact of numerical errors within mathematical libraries has not received sufficient attention.
-Moreover, such errors in floating-point arithmetic are inherently difficult to detect and quantify.
-Variance arithmetic offers a robust alternative, as its computed values closely align with those obtained from conventional floating-point arithmetic while its associated uncertainties systematically trace all input errors.
-Furthermore, its resulting error deviations provide an objective basis for classifying of calculation quality as ideal, proper, or suspicious.
-10 Regressive Generation of Sin and Cos
-Figure 66: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ versus $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
-Figure 67: Uncertainties, value errors (left y-axis), and error deviation (right y-axis) for $\sin(x)^{2}+\cos(x)^{2}-1,x\in[0,\pi/4]$ versus regression order (x-axis) when compared with the Quart sine function (legend)
-Formula (10.2) and Formula (10.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1).
-Formula (10.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, so that value errors will not accumulate when the regression order increases.
-$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (10.1)
-$\sin(\frac{\alpha+\beta}{2})=\sqrt{\frac{1-\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1-\cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)}{2}};$ (10.2)
-$\cos(\frac{\alpha+\beta}{2})=\sqrt{\frac{1+\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1+\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)}{2}};$ (10.3)
-$\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (10.4)
-Formula (10.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation, resulting in excessive uncertainty.
-In Figure 66, the Quart sine function exhibits uncertainties proportional to $\sin(\frac{\pi}{2^{L}})$, whereas the regression sine function shows increasing uncertainties as $x\rightarrow 0$.
-Unlike catastrophic cancellation in floating-point arithmetic, variance arithmetic uses very coarse precision to signal that the regression algorithm is unfit to compute $\sin(x\rightarrow 0)$.
-Figure 67 evaluates the Quart and Regression sine functions using $\sin(x)^{2}+\cos(x)^{2}-1=0$.
-The figure shows that the resulting uncertainties remain nearly constant and independent of the regression order $L$ when $L\geq 4$ for both the Regression sine and the Quart sine functions.
-It also demonstrates that the value errors for both are comparable.
-The Regression sine function has error deviations closer to $1$.
-11 Conclusion and Discussion
-11.1 Summary of Statistical Taylor Expansion
-The starting point of statistical Taylor expansion is the assumption that all input variables obeys uncorrelated uncertainties.
-This is generally regarded as a reasonable and practical statistical requirement for input data [29].
-Once this assumption is satisfied, statistical Taylor expansion quantifies uncertainty as the deviation of the value errors and traces track the variable dependencies in intermediate steps using standard statistical methods.
-While this approach eliminates the dependency problem, it also requires a more rigorous process to determine the resulting mean and deviation for any analytic expression.
-Furthermore, it can reject invalid calculations on unified statistical and mathematical grounds and explicitly incorporates sample counts and the corresponding underlying uncertainty distributions into the statistical analysis.
-11.2 Summary of Variance Arithmetic
-Variance arithmetic simplifies statistical Taylor expansion by assuming that the sample counts for all inputs are sufficiently large to achieve ideal variance.
-Ill-formed problems can invalidate the resulting variance in several ways by producing non-convergent, unstable, negative, infinite, or unreliable results.
-The presence of ideal coverage is a necessary condition for a numerical algorithm in variance arithmetic to be considered correct.
-Ideal coverage also defines the algorithm’s optimal applicable range.
-Under ideal coverage, the calculated uncertainty equals the deviation of the value errors, and the normalized errors follow either a Normal or Delta distribution depending on the context.
-Variance arithmetic provides proper coverage for floating-point rounding errors.
-The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios, including analytic calculations, progressive computation, regressive generalization, polynomial expansion, statistical sampling, and transformations.
-The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
-11.3 Improvements Needed
-This paper presents variance arithmetic which is still in its early stage of development. From a theoretical standpoint, several important questions still need to be addressed.
+When uncorrelated uncertainty condition is met, statistical Taylor expansion outputs the mean, deviation, and reliability of an analytic expression.
+It tracks the variable dependencies in intermediate steps, rejects invalid calculations, and explicitly incorporates the sample counts and the input uncertainty distributions into the statistical analysis.
+While statistical Taylor expansion eliminates the dependency problem, it also removes most execution flexibility.
+The presence of ideal coverage is a necessary condition for a numerical algorithm using statistical Taylor expansion to be considered correct.
+The ideal coverage defines the optimal applicable range for an algorithm.
+Variance arithmetic simplifies statistical Taylor expansion by introducing numerical rules to elimination invalid results, such as divergent, negative variance, unstable, infinite, or unreliable results.
+It provides proper coverage for floating-point rounding errors.
+The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios.
 Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty.
 Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results.
-How to compute and validate these new uncertainty-bearing library functions will need innovation.
-For example, the uncertainty of a 64-bit value can not directly determined by its difference with the corresponding 128-bit value,
+The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
+A more detailed description of this paper is presented at https://arxiv.org/abs/2410.01223.
+11.2 Improvements Needed
+This paper presents statistical Taylor expansion and variance arithmetic, which is still in early stage of development, with several important questions remaining.
 Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions.
-The statistical bounding to find the bounding range $\kappa$ for the bounded moment needs to be extended to discrete distributions.
-The choice of ideal bounding range $\hat{\kappa}$ needs further study.
-For example, the choice of $\hat{\kappa}=5$ for Normal distribution is based on the $5-\sigma$ rule in practice only.
-Such choice certainly can not be applied to other distributions.
-Also, the choice of ideal bounding range $\hat{\kappa}$ determines the convergence of both $\log(x)$ and $x^{c}$ as $P(x)\lesssim 1/\hat{\kappa}$, which needs better mathematical understanding.
-The condition for Formula (2.47) seems to be as generic as the definition in Formula (2.7) which needs further proof.
-Cases where $\delta^{2}f<\widehat{\delta^{2}}f$ due to low sample count require further theoretical study and experimental verification.
-The dependence of the convergence of $\log(x\pm\delta x)$ and $(x\pm\delta x)^{c}$ on the chose of ideal bounding range $\kappa$ needs further discussion.
-Determining the uncertainty upper bounds for each analytic expression through analytic methods is an important next step.
+The statistical bounding to find the bounding range $\kappa$ for a sample count needs to be extended to discrete distributions.
+The choice of ideal bounding range $\hat{\kappa}$ needs to extend to other distributions.
+The statistical meaning for ideal ratio $\alpha$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\alpha$ the possibility for the result to be bound by $\hat{\kappa}$?
 The performance of variance arithmetic must be improved for broader practical adoption.
-The fundamental formulas of statistical Taylor expansion, Formula (2.9), (2.3), (2.12), and (2.3) contain a large number of independent summations, making them excellent candidates for parallel processing.
+The fundamental formulas of statistical Taylor expansion, Formula (2.5), (2.3), (2.8), and (2.3), contain a large number of independent summations, making them excellent candidates for parallel processing.
 Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
+A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors, because many theoretical calculations lack explicit input uncertainties.
+Variance arithmetic does not adjust uncertainty variance when floating-point rounding error occurs during calculation.
+This leads to error deviations larger than 1.
+Detecting floating-point rounding error also needs hardware implementation for efficiency.
+In variance arithmetic, deviations are comparable with values, but variances are used in calculation.
+This effectively limits the range of deviations to the square root of that of values.
+If the sign bit of the floating type can be re-purposed as an exponent bit in a new unsigned floating-point representation, the range of the deviation will be identical to that of values.
 When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion.
-However, modern symbolic computation tools such as SymPy can greatly facilitate these calculations.
+However, modern symbolic computation tools such as SymPy and Mathematica can greatly facilitate these calculations.
 This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
@@
 For example, Formula (2.31) can guide how to improve the ideal ratio of $x\pm y$.
-A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors.
-Except for integer multiplication, variance arithmetic does not track rounding errors during computation, because tracking rounding errors for every operation is too expensive to achieve in software.
-The sensitivity to the library functions increases the difficulty to solve these types of problems.
-Developing variance arithmetic with ideal coverage for such errors would be quite valuable because many theoretical calculations lack explicit input uncertainties.
 Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic.
-For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas variance arithmetic conceptually rejects all path-dependent calculations.
+For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas statistical Taylor expansion conceptually rejects all path-dependent calculations.
 Reconciling these two paradigms may present a significant and ongoing challenge.
 Establishing theoretical foundation for applying statistical Taylor expansion in the absence of a closed-form analytic solution, or when only limited low-order numerical derivatives are available, as in solving differential equations, remains an important direction for future research.
+11.3 Possible Connections to Quantum Physics
+For input distribution without bounding, Formula (2.3) may not converge, for example, when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$.
+Formula (2.3) converges in all cases only when the input uncertainty is bounded due to sampling, which resembles the need for re-normalization in quantum field theory [21].
+The quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: what it is depends on how it is measured.
+For example, because mathematical divergence seems to correspond to distributional zero, it is even tempting to postulate if a localized experiment changes $\hat{\kappa}$ to produce distributional zero and the quantum result as “waveform collapse” [21].
+Both resemblances are worth further investigation.
+It is also worthwhile investigating if uncertainty bias produces Lamb shift [21].
 12 Statements and Declarations
@@
 The data set used in this study are all generated in the open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
-The execution assistance and explanation of the above code are available from the author upon reasonable request.
+The execution assistance and explanation of the above code are available from the author upon request.
 12.3 Competing Interests

### v18

@@
 Abstract
-As a rigorous statistical approach, statistical Taylor expansion extends the conventional Taylor expansion by replacing precise input variables with random variables of known distributions and sample counts to compute the mean, the standard deviation, and the reliable factor of each result.
-It tracks the propagation of the input uncertainties through intermediate steps, so that the final analytic result becomes path independent.
-Therefore, it differs fundamentally from common approaches in applied mathematics that optimize computational path for each calculation.
-Statistical Taylor expansion may standardize numerical computations for analytic expressions.
-This study also introduces the implementation of statistical Taylor expansion termed variance arithmetic and presents corresponding test results across a wide range of mathematical applications.
-Another important conclusion of this study is that numerical errors in library functions can significantly affect results.
-It is desirable that each value from library functions be accomplished by an uncertainty deviation.
+As a rigorous extension of conventional Taylor expansion, statistical Taylor expansion replaces precise input variables with random input variables of known distributions and sample counts, to compute the mean, standard deviation, and reliability of each result.
+This approach tracks the propagation of input uncertainties through intermediate steps, making the final result path independent.
+Therefore, it differs fundamentally from conventional approaches that are usually path dependent, such that it may standardize numerical computations for analytic expressions.
+This study further presents an implementation of statistical Taylor expansion termed variance arithmetic, and demonstrates its performance across diverse mathematical applications.
+The finding of this study also reveals the potential and substantial impact of numerical errors in library functions.
+The possible link between statistical Taylor expansion and quantum physics is discussed as well.
 Keywords: computer arithmetic, error analysis, interval arithmetic, uncertainty, numerical algorithms.
@@
 1 Introduction
-Let $x$ and $\delta x$ be a value and its uncertainty deviation, respectively.
-As an input, $x$ and $\delta x$ are the mean and standard deviation of a measurement [1][2].
-If $\delta x=0$, $x$ is a precise value, otherwise, the pair specifies an imprecise value $x\pm\delta x$.
+Let $x$ and $\delta x$ denote a value and its uncertainty deviation, respectively.
+As input, $x$ and $\delta x$ are usually the mean and standard deviation of a measurement, respectively [1][2].
+If $\delta x=0$, then $x$ is a precise value; otherwise, the pair specifies an imprecise value $x\pm\delta x$.
 Let $P(x)\equiv\delta x/|x|$ be the statistical precision (hereafter referred to as precision) of $x\pm\delta x$.
-A smaller $P(x)$ indicate a higher measurement quality of $x\pm\delta x$, and when $P(x)\ll 1$, it is unclear whether $\delta x$ is too poor to measure $x$, or it is a measurement of $x=0$.
-Statistical Taylor expansion determines the result $f\pm\delta f$ of a general analytic expression $f(x,\dots)$, and the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ based on the inputs $x\pm\delta x,\dots$ and their corresponding sample counts.
+A smaller $P(x)$ indicates a higher measurement quality of $x\pm\delta x$.
+Statistical Taylor expansion determines the result $f\pm\delta f$ and reliability of a general analytic expression $f(x,\dots)$ on the basis of inputs $x\pm\delta x,\dots$ and their corresponding distributions and sample counts.
   • Previous studies have examined the effect of input uncertainties on output values for specific cases [3].
-Statistical Taylor expansion generalizes these effects as uncertainty bias, as shown in Formula (2.5) and (2.8) in this paper.
-  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this paper.
-  • Calculating the reliability of $f\pm\delta f$ seems completely new.
-Statistical Taylor expansion is superior to all existing numerical arithmetic.
-Conventional floating-point arithmetic [6][7][8] can only calculate the result $f$ which contains unknown amount of rounding error [9][10][11].
-As a consequence, a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision.
-While statistical Taylor expansion can track rounding errors.
-The bounding range in interval arithmetic [12][13][14][15][16] is inconsistent with the statistical nature of an $x\pm\delta x$ pair and it tends to over-estimate the result uncertainty due to its worst-case assumption [17].
-On the other hand, statistical Taylor expansion is precise statistically.
-Both conventional floating-point arithmetic and interval arithmetic depend strongly on the specific algebraic form of an analytic function, a phenomenon known as the dependency problem [10][11][12][14], which makes the numerical calculations of analytic problems sometimes more like an art than science.
+Statistical Taylor expansion generalizes these effects as uncertainty bias, as shown in Formulas (2.5) and (2.8) of this study.
+  • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this study.
+  • Calculating the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ seems completely new.
+As a new approach, statistical Taylor expansion is superior to all existing numerical arithmetic.
+Conventional floating-point arithmetic [6][7][8] computes only the result value $f$.
+Because $f$ contains unknown amount of rounding error [9][10][11], a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision.
+However, statistical Taylor expansion can account for rounding errors as part of the result deviation $\delta f$.
+The bounding range in interval arithmetic [12][13][14][15][16] is inconsistent with the statistical nature of an $x\pm\delta x$ pair and it tends to over-estimate the result uncertainty because of its worst-case assumption [17].
+Conversely, statistical Taylor expansion is precise statistically.
+Both conventional floating-point arithmetic and interval arithmetic depend strongly on the specific algebraic form of an analytic function, a phenomenon known as the dependency problem [10][11][12][14], which can make conventional numerical computation more an art than a science.
 In contrast, statistical Taylor expansion is path independent.
-To be rigorous in mathematics and statistics, statistical Taylor expansion abandons the significance arithmetic nature of its processor [17].
-As a statistical sampling process, stochastic arithmetic [18][19] is computationally expensive, while statistical Taylor expansion provides a direct characterization without sampling.
-In this paper:
-  1. 1.
-    Section 1 compares briefly statistical Taylor expansion with all other numerical arithmetic.
-  2. 2.
+To ensure mathematical and statistical rigor, statistical Taylor expansion abandons the significance arithmetic nature of its processor [17].
+As a statistical sampling process, stochastic arithmetic [18][19] is computationally expensive, whereas statistical Taylor expansion provides a direct characterization without sampling.
+The remainder of this paper is organized as follows:
   • Section 2 develops the theoretical foundation of statistical Taylor expansion.
-  3. 3.
   • Section 3 describes variance arithmetic as a numerical implementation of statistical Taylor expansion.
-  4. 4.
-    Section 4 lays out standard to validate variance arithmetic.
-  5. 5.
-    Section 5 illustrates variance arithmetic in computing polynomial, demonstrating its capability in tracing floating-point rounding errors, and its continuity in parameter space.
-  6. 6.
-    Section 6 applies variance arithmetic to adjugate matrix and matrix inversion, showing that the uncertainty calculation for a linear calculation is no longer linear.
-It also distinguishes between a distribution test and a value test.
-  7. 7.
-    Section 7 evaluates variance arithmetic on common mathematical library functions. It also shows that the error deviation should be close to $1$ except near a distributional pole.
-  8. 8.
+  • Section 4 presents standards for validating variance arithmetic.
+  • Section 5 illustrates variance arithmetic in polynomial computation, demonstrating its ability to trace floating-point rounding errors and its continuity in parameter space.
+  • Section 6 describes the applications of variance arithmetic to matrix inversion, distinguishing between distribution tests and value tests.
+  • Section 7 discusses the evaluation of variance arithmetic on common mathematical library functions, showing the effect of distribution pole.
   • Section 8 reveals the accumulation of numerical errors by reusing an input multiple time in a moving-window progressive algorithm.
-  9. 9.
+  • Section 9 examines the impact of numerical library errors and shows that these errors can be significant.
   • Section 10 showcases variance arithmetic in catching catastrophic cancellation in a regression algorithm.
-  10. 10.
-    Section 10 examines the impact of numerical library errors and shows that these errors can be significant.
-  11. 11.
-    Section 11 concludes with a summary and a discussion.
+  • Section 11 concludes with a summary and a discussion of the findings.
 2 Statistical Taylor Expansion
-2.1 The Uncorrelated Uncertainty Condition
-When there is no systematic error [1][2] among the inputs, the uncertainties of the inputs are uncorrelated to each other, even though the inputs may have significant correlations among them.
+2.1 Uncorrelated Uncertainty Condition
+When inputs contain no systematic error [1][2], their uncertainties are uncorrelated, even though their values can have significant correlations.
 This condition can be characterized quantitatively in statistics as the uncorrelated uncertainty condition [17].
@@
 The horizontal axis is scaled as $\sqrt{\tilde{y}}$.
-Figure 2: Probability density function for $\tilde{y}=\sqrt{\tilde{x}}$, various values of $\mu$ as indicated in the legend.
+Figure 2: Probability density function for $\tilde{y}=\sqrt{\tilde{x}}$, for various values of $\mu$ as indicated in the legend.
 The variable $\tilde{x}$ follows a Gaussian distribution with the distributional mean $\mu$ and deviation $1$.
@@
 For example $\sqrt{\mu\pm 1}$ has a distributional zero at $\tilde{x}=0$, as shown in Figure 2.
-In both Figure 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely representation resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, such as $(5\pm 1)^{2}$ in Figure 1 and $\sqrt{5\pm 1}$ in Figure 2, thereby allowing for characterization of the output $f(x)$ using mean $\overline{f(x)}$ and deviation $\delta f(x)$ only.
+In both Figures 1 and 2, $\rho(\tilde{y},\mu_{y},\sigma_{y})$ closely resembles $\rho(\tilde{x},\mu,\sigma)$ when the mode of $\rho(\tilde{x},\mu,\sigma)$ lies sufficiently far away from either a distributional pole or a distributional zero, such as $(5\pm 1)^{2}$ in Figure 1 and $\sqrt{5\pm 1}$ in Figure 2, thereby allowing for characterization of the output $f(x)$ using mean $\overline{f(x)}$ and deviation $\delta f(x)$ only.
 2.3 Statistical Taylor Expansion
 […]
-$\zeta(n)$ $\equiv\int_{\varrho}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.2)
-$\zeta(1)$ $=0;$ (2.3)
-Let $\tilde{z}\in[\varrho,\kappa]$ where $\varrho,\kappa$ specify the bounding ranges.
-Formula (2.2) defines the corresponding bound moment $\zeta(n)$, which further satisfies the mean-reverting condition of Formula (2.3) so that $\kappa$ determines $\varrho$.
-For any symmetric $\rho(-\tilde{z})=\rho(\tilde{z}):\varrho=-\kappa,\zeta(2n+1)=0$.
-The probability of $\tilde{z}\not\in[\varrho,\kappa]$ is defined as the bounding leakage $\epsilon(\kappa)=1-\zeta(0,\kappa)$.
+$\zeta(n,\kappa)$ $\equiv\int_{\varrho}^{\kappa}\tilde{z}^{n}\rho(\tilde{z})d\tilde{z}$; (2.2)
+$\zeta(1,\kappa)$ $=0;$ (2.3)
 […]
-$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{x})\rho(\tilde{x},\mu,\sigma)d\tilde{x}=f(x)+\sum_{n=1}^{\infty}\sigma^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n);$ (2.5)
+[…]
+$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{z})\rho(\tilde{z})d\tilde{z}=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
 $\delta^{2}f(x)$ $=\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$
-$=\sum_{n=1}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.6)
-An analytic function $f(x)$ can be accurately evaluated over in a range using the Taylor series as shown in Formula (2.4).
-Formula (2.5) and Formula (2.3) yield the mean $\overline{f(x)}$ and the variance $\delta^{2}f(x)$ of $f(x)$, respectively.
+$=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
+An analytic function $f(x)$ can be accurately evaluated in a range using the Taylor series as shown in Formula (2.4).
+Formulas (2.5) and (2.3) yield the mean $\overline{f(x)}$ and variance $\delta^{2}f(x)$ of $f(x)$, respectively.
 The difference $\overline{f(x)}-f(x)$ is defined as the uncertainty bias, representing the effect of input uncertainty on the resulting value.
 […]
-$\overline{f(x,y)}$ $=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m)\zeta_{y}(n);$ (2.8)
-$\delta^{2}f(x,y)$ $=\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\sigma_{x})^{m}(\sigma_{y})^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
-$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\big(\zeta_{x}(m)\zeta_{y}(n)-\zeta_{x}(i)\zeta_{x}(m-i)\;\zeta_{y}(j)\zeta_{y}(n-j)\big);$ (2.9)
-Under the uncorrelated uncertainty condition, Formula (2.8) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.7), where $\zeta_{x}(m)$ and $\zeta_{y}(n)$ denote the variance moments for $x$ and $y$, respectively.
-Although Formula (2.3) is only for 2-dimensional, it can be extended easily to any dimension.
+$\overline{f(x,y)}$ $=\sum_{m=0}^{\infty}\sum_{n=0}^{\infty}(\delta x)^{m}(\delta y)^{n}\frac{f^{(m,n)}_{(x,y)}}{m!\;n!}\zeta_{x}(m,\kappa_{x})\zeta_{y}(n,\kappa_{y});$ (2.8)
+$\delta^{2}f(x,y)$ $=\sum_{m=1}^{\infty}\sum_{n=1}^{\infty}(\delta x)^{m}(\delta y)^{n}\sum_{i=0}^{m}\sum_{j=0}^{n}\frac{f^{(i,j)}_{(x,y)}}{i!\;j!}\frac{f^{(m-i,n-j)}_{(x,y)}}{(m-i)!\;(n-j)!}$
+$\big(\zeta_{x}(m,\kappa_{x})\zeta_{y}(n,\kappa_{y})-\zeta_{x}(i,\kappa_{x})\zeta_{x}(m-i,\kappa_{x})\;\zeta_{y}(j,\kappa_{y})\zeta_{y}(n-j,\kappa_{y})\big);$ (2.9)
+Under the uncorrelated uncertainty condition, Formulas (2.8) and (2.3) compute the mean and variance of the Taylor expansion given in Formula (2.7), where $\zeta_{x}(m,\kappa_{x})$ and $\zeta_{y}(n,\kappa_{y})$ denote the bound moments for $x$ and $y$, respectively.
+Although Formula (2.3) is only 2-dimensional problem space, it can be extended easily to any dimension.
 With the mean reverting condition of Formula (2.3):
 $\overline{x\pm y}$ $=\zeta(0,\kappa_{x})x\pm\zeta(0,\kappa_{x})y;$ (2.10)
-$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}+\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.11)
+$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.11)
 $\overline{xy}$ $=\zeta(0,\kappa_{x})x\;\zeta(0,\kappa_{y})y;$ (2.12)
-$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\sigma_{x})^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\sigma_{y})^{2}+\zeta(2,\kappa_{x})(\sigma_{x})^{2}\;\zeta(2,\kappa_{y})(\sigma_{y})^{2};$ (2.13)
-When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formula (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formula (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
+$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\delta y)^{2}+\zeta(2,\kappa_{x})(\delta x)^{2}\;\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.13)
+When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formulas (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formulas (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
 2.4 One-Dimensional Examples
-Formula (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
-$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}\sigma^{n}\zeta(n)\frac{1}{n!};$ (2.14)
-$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j!\;(n-j)!};$ (2.15)
-Formula (2.16) and (2.17) give the mean and variance for $\log(x)$, respectively:
-$\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n)}{n};$ (2.16)
-$\delta^{2}\log(x)$ $=\sum_{n=2}^{+\infty}P(x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n)-\zeta(j)\zeta(n-j)}{j(n-j)};$ (2.17)
-Formula (2.19) and (2.20) give the mean and variance for $\sin(x)$, respectively:
+Formulas (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
+$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\zeta(n,\kappa)\frac{1}{n!};$ (2.14)
+$\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!\;(n-j)!};$ (2.15)
+Formulas (2.16) and (2.17) give the mean and variance for $\log(x)$, respectively:
+$\overline{\log(x)}$ $=\log(x)+\sum_{n=1}^{+\infty}P(x)^{n}\frac{(-1)^{n+1}\zeta(n,\kappa)}{n};$ (2.16)
+$\delta^{2}\log(x)$ $=\sum_{n=2}^{+\infty}P(x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j(n-j)};$ (2.17)
+Formulas (2.19) and (2.20) give the mean and variance for $\sin(x)$, respectively:
 […]
-$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}\sigma^{n}\eta(n,x)\frac{\zeta(n)}{n!};$ (2.19)
-$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}\sigma^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.20)
-Formula (2.21) and (2.22) give the mean and variance for $x^{c}$, respectively:
-$\frac{\overline{x^{c}}}{x^{c}}$ $=1+\sum_{n=1}^{\infty}P(x)^{n}\zeta(n)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.21)
-$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big(\zeta(n)-\zeta(j)\zeta(n-j)\big);$ (2.22)
-The input and output in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\sigma\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\sigma\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
+$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n,\kappa)}{n!};$ (2.19)
+$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.20)
+Formulas (2.21) and (2.22) give the mean and variance for $x^{c}$, respectively:
+$\frac{\overline{x^{c}}}{x^{c}}$ $=1+\sum_{n=1}^{\infty}P(x)^{n}\zeta(n,\kappa)\begin{pmatrix}c\\ n\end{pmatrix}$; (2.21)
+$\frac{\delta^{2}x^{c}}{(x^{c})^{2}}$ $=\sum_{n=2}^{\infty}P(x)^{n}\sum_{j=1}^{n-1}\begin{pmatrix}c\\ j\end{pmatrix}\begin{pmatrix}c\\ n-j\end{pmatrix}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.22)
+The input and output in statistical Taylor expansion reflects the inherent characteristics of the calculation, such as $\delta x\rightarrow P(e^{x})$, $P(x)\rightarrow\delta\log(x)$, $\delta x\rightarrow\delta\sin(x)$, and $P(x)\rightarrow P(x^{c})$.
 2.5 Convergence of Variance
@@
 $\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.24)
-$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n)-\zeta(j)\zeta(n-j)=\frac{\kappa^{n}}{n};$ (2.25)
+$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)=\frac{\kappa^{n}}{n};$ (2.25)
+[…]
+  • Formula (2.15) for $e^{x\pm\delta x}$ and Formula (2.20) for $\sin(x\pm\delta x)$ both converge unconditionally.
+  • Formula (2.17) for $\log(x\pm\delta x)$ can be approximated by Formula (2.26), which converges when $P(x)<1/\kappa$.
+  • Formula (2.22) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.27), which converges when $P(x)\lesssim 1/\kappa$ although the precise upper bound for $P(x)$ varies with $c$.
 […]
 […]
-[…]
-  • Formula (2.15) for $e^{x\pm\delta x}$ and Formula (2.20) for $\sin(x\pm\delta x)$ both converge unconditionally.
-[…]
-  • Formula (2.22) for $(x\pm\delta x)^{c}$ can be approximated by Formula (2.27). which converges when $P(x)\lesssim 1/\kappa$ although the precise upper bound for $P(x)$ varies with $c$.
+Statistical Taylor expansion rejects the distributional zero of $\log(x)$ or $(x\pm\delta x)^{c}$ in the range of $P(x)>1/\kappa$ statistically because of the divergence of Formulas (2.26) and (2.27) mathematically, with $\zeta(2n,\kappa)$ providing the connection between these two perspectives.
 2.6 Statistical Bounding
-Figure 4: Measured bounding leakage $\epsilon(\kappa_{s},N)$ (y-axis) for varying measuring bounding range $\kappa_{s}$ (x-axis) and sample count NN (legend).
-Figure 5: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range $\kappa_{s}$ for the normal distribution.
+Figure 4: Measured linear leakage $\omega(\kappa_{s},N)$ (y-axis) for varying measuring bound range $\kappa_{s}$ (x-axis) and sample count NN (legend).
+Figure 5: Measured linear leakage $\omega(\kappa,N)$ (left y-axis) and corresponding measured bound range $\kappa$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is Gaussian, with different measuring bound range $\kappa_{s}$ (legend).
+Figure 6: Ideal leakage (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
+Figure 7: Ideal leakage (y-axis) for varying bound range $\kappa$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=6$ for Gaussian input uncertainties.
+$f(x)-\overline{f(x)}$ $\simeq(\delta x)^{2}\zeta(2,\kappa)f_{x}^{(2)};$ (2.28)
+$\delta^{2}f(x)$ $\simeq(\delta x)^{2}\zeta(2,\kappa)(f_{x}^{(1)})^{2};$ (2.29)
+Figures (2.28) and (2.29) show that the first-order approximation of Formulas (2.5) and (2.3) both contain the term $(\delta x)^{2}\zeta(2,\kappa)$.
+As $\kappa\rightarrow+\infty$, $\zeta(2,\kappa)\rightarrow 1$ such that $\delta^{2}f(x)$ reaches a stable value but the convergence range of $\delta^{2}f(x)$ reduces toward zero for Formulas (2.17) and (2.22).
+As a trade-off, the choice of $\kappa$ to calculate the stable $f(x)-\overline{f(x)}$ and $\delta^{2}f(x)$ is defined as the ideal bound range $\hat{\kappa}$, and the result is the ideal uncertainty bias $\widehat{\overline{f(x)}}-f(x)$ and ideal variance $\hat{\delta^{2}}f(x)$.
+For a Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$.
+For a Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$ by default.
+Define linear leakage $\omega(\kappa)\equiv 1-\zeta(2,\kappa)$.
 When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [4].
-This yields the sample bounding leakage $\epsilon(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional bounding leakage $\epsilon(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$.
-Because $\epsilon(\kappa)\neq\epsilon(\kappa,N)$ for finite $N$, let $\epsilon(\kappa)=\epsilon(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bonding range, and $\kappa(\kappa_{s},N)$ is the measured bounding range.
+This yields the sample linear leakage $\omega(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional linear leakage $\omega(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$.
+Because $\omega(\kappa)\neq\omega(\kappa,N)$ for finite $N$, let $\omega(\kappa)=\omega(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bound range, and $\kappa(\kappa_{s},N)$ is the measured bound range, with the latter used in calculating $\zeta(n,\kappa)$.
+$\epsilon(\kappa)$ $=\omega(\kappa)+2N(\kappa)\kappa;$ (2.30)
 $\epsilon(\kappa)$ $=1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.31)
 $\epsilon(\kappa_{s},N)$ $=1-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x-\overline{x}|}{\sqrt{2}})-\frac{1}{2}\xi(\frac{|\kappa_{s}\delta x+\overline{x}|}{\sqrt{2}});$ (2.32)
-When the underlying distribution is Normal, Formula (2.28) and (2.29) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
-Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction.
-It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
-Figure 5 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases, but converges to a larger stable bounding leakage $\epsilon(\kappa_{s})$.
-When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to the bounding leakage $\epsilon(N)$.
-Figure 5 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically.
-The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$.
-2.7 Ideal Statistics
-Figure 6: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Gaussian.
-Figure 7: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Uniform.
-$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.30)
-$\alpha(x\pm y)=$ $\frac{\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.31)
-When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}$ in Formula (2.30).
-Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
-  • When the underlying uncertainty distribution is Gaussian and $\kappa_{s}=5$, Figure 6 shows that the resulting variances $\delta^{2}f$ for a selected functions rise with $N$ until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 50$.
-  • When the underlying uncertainty distribution is Uniform, $\kappa_{s}=\sqrt{3}$, Figure 7 shows the same trend, however the stable variances $\widehat{\delta^{2}}f$ are reached only when $N\geq 10^{4}$.
-The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 6 and 7 respectively, except that $\overline{x}-x=0$.
-This similarity is expected because the first-order approximation in both Formula (2.5) and Formula (2.3) contains $\sigma^{2}\zeta(2,\kappa)$.
-In both Figure 6 and 7, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases.
-The stable resulting variances $\widehat{\delta^{2}}f$ is reached only when $\epsilon(s,N)<10^{-3}$.
-All functions approach their corresponding stable variances with exactly the same trend in each figure.
-These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer the confidential interval range $\kappa$ from bounding leakage $\epsilon(\kappa_{s})=1-\zeta(0,\kappa_{s})$ in Figure 4, or more precisely $1-\zeta(2,\kappa_{s})$ for $\delta^{2}f$.
-Define the stable variance when the sample count $N$ is sufficiently large as the ideal variance $\widehat{\delta^{2}}f$.
-The choice of $\kappa_{s}$ to calculate $\widehat{\delta^{2}}f$ is defined as the ideal bounding range $\hat{\kappa}$.
-For Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$.
-When the input distribution is unbounded, the choice of $\hat{\kappa}$ is a trade-off: a larger $\hat{\kappa}$ means a smaller bounding leakage for the result but a narrower convergence range for the input.
-For Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$.
-When the sample count $N$ is insufficient, $\widehat{\delta^{2}}f$ is calculated using $\hat{\kappa}$, to assume that $N$ is sufficient.
-$\delta^{2}f$ is calculated with the actual $N$.
-Define the ideal ratio as $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f$, which is 0 when $N=1$ and $1$ when $N$ is sufficiently large.
-The ideal ratio $\alpha$ quantifies the reliability of $\widehat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.31) as an example for $\alpha(x\pm y)$ .
+Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
+When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, while Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
+Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
+Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable result bounding leakage is smaller for larger bounding range $\kappa$.
+It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\hat{\kappa}$ is larger.
+$\varepsilon(x\pm y)$ $=\frac{\omega_{x}(\delta x)^{2}+\omega_{y}(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.33)
+Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example.
+Figure 6 shows that the stable ideal leakages for the selected functions are less than $10^{-4}$ for $\hat{\delta^{2}}f$, which is small enough for most applications.
+Figure 7 compares the ideal leakages for $\hat{\delta^{2}}f$ for the selected functions when different $\hat{\kappa}$ is used.
 2.7 Dependency Tracing
@@
   • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$.
-For a reversible transformation such as matrix reversion or FFT, after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
+For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
   • Formula (2.35) gives the variance of the linear transformation of a function, which can be applied to other formulas for more general dependency tracing.
@@
 2.8 Traditional Execution and Dependency Problem
-Dependency tracing requires an analytic form of the function to apply statistical Taylor expansion for the result mean and variance, as in Formula (2.5), (2.3), (2.8), and (2.3).
+Dependency tracing requires applying statistical Taylor expansion on an analytic form of the function in one step.
 This requirement often conflicts with conventional numerical methods for analytic functions:
@@
 For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 6.
-  • Furthermore, traditional approaches frequently apply approximations to result values during execution.
-Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than value in statistical Taylor expansion.
-Consequently, approximation strategies should prioritize variances than values.
-  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last digit.
-As demonstrated in Section 10, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty deviations explicitly incorporated into the output.
-  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls.
+  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls.
 However, this decomposition introduces dependency problems.
@@
 It represents an imprecise value $x\pm\delta x$ using a pair of 64-bit standard floating-point numbers and uses floating-point arithmetic for computation.
-Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms.
+Because of the finite precision and limited range of the floating-point representation, $\zeta(n,\kappa\leq 5)$ can be computed only to limited terms.
 Consequently, the following numerical rules are introduced:
@@
   • reliable: At every order, the deviation of the variance must be less than $1/5$ times the value of the variance.
-For simplicity of discussion, the Taylor coefficients in Formula (2.4) and (2.7) are assumed to be precise.
+For simplicity of discussion, the Taylor coefficients in Formulas (2.4) and (2.7) are assumed to be precise.
 3.1 Monotonic
@@
 The corresponding uncertainty bias and uncertainty are also shown (right y-axis).
-The x-axis is expressed in units of $\pi$.
 When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
@@
 The corresponding uncertainty bias and uncertainty are also shown (right y-axis).
-The x-axis is expressed in units of $\pi$.
 When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
 When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
-Beyond an upper bound $\delta x$, the expansion is no longer monotonic for $e^{x\pm\delta x}$, $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c}$.
-For Gaussian input uncertainty with $\hat{\kappa}=5$:
+Beyond an upper bound $\delta x$, the expansion is no longer monotonic for $e^{x\pm\delta x}$, $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c}$. For Gaussian input uncertainty with $\hat{\kappa}=5$:
   • For $e^{x\pm\delta x}$, $\delta x\lesssim 19.864$ and $P(e^{x\pm\delta x})\lesssim 1681.767$ regardless of $x$.
@@
 These limits follow directly from the relationship $P(x)\rightarrow\delta\log(x)$, as indicated in Formula (2.17).
-  • For $(x\pm\delta x)^{c}$, except when $c$ is a natural number, the upper bound $P(x)$ is close to $1/5$ but increasing with $c$. This trend is displayed in Figure 8.
-Similar trends holds when the input uncertainty is uniform but with different values.
-For example, after the respective upper bound $P(x)$ for $(x\pm\delta x)^{c}$ are normalized by the corresponding $\hat{\kappa}$, they show almost the same trends of increasing with $c$.
-Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$.
+  • For $(x\pm\delta x)^{c}$, except when $c$ is a natural number, the upper bound $P(x)$ is close to $1/5$ but increasing with $c$. This trend is shown in Figure 8.
+Similar trends holds when the input uncertainty follows other distribution.
+Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$ when the input uncertainty distribution is Uniform.
+It has almost the same trend of upper bounds $P(x)$ increasing with $c$ as shown in Figure 9 but just with different y-scaling: $1/\sqrt{3}$ instead of by $1/5$.
 3.2 Positive
@@
 The x-axis is expressed in units of $\pi$.
-Besides convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$.
+In addition to convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$.
 Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty.
 Beyond this upper bound, the result deviation is no longer positive at some expansion order.
-Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which Figure 10 satisfies.
+Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which is satisfied as shown in Figure 10.
 Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@
 Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$.
-Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP of $x$, where ULP refers to the Unit in the Last Place in conventional floating-point representation [8], because rounding errors are shown to be uniformly distributed within ULP [17].
+Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP (Unit in the Last Place [8]) of $x$, because rounding errors are shown to be uniformly distributed within ULP [17].
 3.4 Comparison
@@
   • Error distribution: the histogram of the normalized errors.
-Once input error from every source is accounted precisely, ideal coverage is achieved in either contexts:
-  • Distribution Test: When comparing the calculated mean and deviation with the result data set, the error deviation is exactly $1$ and the error distribution is Normal.
-Empirically, error distributions are close to Normal when error deviation are close to 1, regardless of input uncertainty distribution, and such convergence occurs rapidly [17].
-This is due to central limit theorem [4].
-  • Value Test: When comparing the calculated value and the result value in a corresponding data set, the error deviation is much less than $1$ and the error distribution is Delta.
+Once input error from every source is accounted precisely, ideal coverage is achieved in either context:
+  • Distribution Test: When comparing the calculated mean and deviation with the result data set, the error deviation is exactly $1$ and the error distribution is Normal, regardless of input uncertainty distribution.
+Such convergence to Normal distribution occurs rapidly [17] because of the central limit theorem [4].
+  • Value Test: When comparing values one-by-one between a calculated data set and the corresponding result data set, the error deviation is much less than $1$ and the error distribution is Delta.
 For example, a round-trip test is a value test.
@@
 $\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}$ $=\sum_{j=0}^{N}\tilde{x}^{j}P_{j},\;\;\;P_{j}\equiv\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (5.1)
-Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can only reach half of the maximal expansion order of Formula (2.3), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian.
+Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can reach only half of the maximal expansion order of Formula (2.3), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian, and $\hat{\kappa}=5$.
 5.1 Residual Error
 Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x-axis).
-The y-axis to the left shows both the value and the uncertainty of the residual errors.
-The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
-Figure 12 shows the residual error of $\sum_{j=0}^{224}x^{j}-1(1-x)$.
-It also displays the required expansion orders for 1/(1x)1/(1-x), which are all less than $224$.
+The left y-axis shows both the value and the uncertainty of the residual errors.
+The right y-axis indicates the expansion order needed to reach stable value for each $x$.
+Figure 12 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$.
+It also displays the required expansion orders for $\frac{1}{1-x}$, which are all less than $224$.
 Therefore, the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$.
-A detailed analysis indicates that the maximal residual error is $4$ times the ULP of 1/(1x)1/(1-x).
+A detailed analysis indicates that the maximal residual error is four times the ULP of $\frac{1}{1-x}$.
 In all cases, the calculated uncertainty bounds the residual error effectively for all $x$, with an error deviation of $2.60$ when the expansion order is less than $224$.
@@
 5.2 Continuity
-In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space.
-For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number $n$.
-The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$.
-5.3 Distributional Hole
-Figure 13: Histograms of normalized errors for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
-Unlike a distributional zero, a statistical bounding range in variance arithmetic can include a distributional pole, such as around $(0\pm\delta x)^{c},c>1$.
-The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram.
+In variance arithmetic, the result mean, variance and error distribution are generally continuous across parameter space.
+For example, when $c$ is a natural number $n$, $(x\pm\delta x)^{c}$ becomes a polynomial which has no upper bound for $\delta x$, in contrast to when $c$ is not a natural number as shown in Figure (8).
+However, the result mean, variance and error distribution of $(x\pm\delta x)^{c}$ remain continuous across $c=n$.
+5.3 Distributional Pole
+Figure 13: Error distributions for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
+A statistical bounding range in variance arithmetic can include a distributional pole, such as around $(0\pm\delta x)^{c},c>1$.
+The presence of such poles does not disrupt the continuity of the result mean, variance, or error distribution.
 Figure 13 illustrates the error distributions of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.
@@
   • When the second derivative is positive, the resulting distribution is right-sided Delta-like, such as the distribution when $n=2,x=0$, or when $n=2,x=\pm 0.2$, or when $n=3,x=0.2$.
-  • When the second derivative is negative, the resulted distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
-In each case, the transition from $x=0$ to $x=0.2$ is continuous for the resulting mean and deviation.
+  • When the second derivative is negative, the resulting distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
+In each case, the transition from $x=0$ to $x=\pm 0.2$ is continuous for the resulting mean, deviation and error distribution.
 6 Matrix Calculations
-Figure 14: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
+Figure 14: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of input noise precision (x-axis) and matrix size (y-axis).
 Figure 15: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (6.5).
@@
 The legend shows the size of the matrix, as well as the type of the matrix as Random for randomly generated matrix, and Hilbert as the Hilbert matrix.
+[…]
+$=\sum_{m=1\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$ (6.2)
 $\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (6.3)
@@
 Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20].
-Formula (6.1) and (6.2) gives the statistical Taylor expansion of the determinant of a square matrix $\mathbf{M}$ of size $n$, when the uncertainties of matrix elements are all independent of each other.
-In Formula (6.2), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices.
-The square matrix whose element is $(-1)^{i+j}|m_{j,i}|$ is defined as the adjugate matrix [20] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$, in which $m_{j,i}$ is the element of $\mathbf{M}$ at row index $j$ and column index $i$.
-Each $m_{j,i}$ is an integer between $2^{8},+2^{8}]$ so that $\mathbf{M}^{A}$ can be calculated precisely using integer arithmetic.
-By adding Gaussian noise of deviation $\delta x$ to each $m_{j,i}$, $\widetilde{\mathbf{M}}$ is created.
-The value error of $\mathbf{M}^{A}$ is the difference between $\widetilde{\mathbf{M}^{A}}$ and $\mathbf{M}^{A}$, while the result deviation is calculated by Formula (6.2), so that $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ is a distribution test.
-Figure 14 shows that the error deviations of $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ are very close to $1$ except when the matrix size is $8$ and the input uncertainty is $\delta x=10^{-17}$.
-This abnormality is caused by floating-point rounding errors only at this point, when the element values exceed the significand range of 64-bit floating-point representation, to result in a error deviation of $3.2$.
-Figure 14 demonstrates a typical distribution test of error deviation versus input noise deviation $\delta x$ and another specific dimension, with ideal coverage for all region except when $\delta x\rightarrow 0$.
+Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$.
+Fourmula (6.2) presents the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$ [20].
+In Formula (6.4), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [20].
+[…]
 $\mathbf{M}\times\mathbf{M}^{A}$ $=\mathbf{M}^{A}\times\mathbf{M}=|\mathbf{M}|\mathbf{I};$ (6.5)
@@
 Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [20].
-Formula (6.3) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ which leads to the definition of inverse matrix $\mathbf{M}^{-1}$ in Formula (6.4) [20].
-$\tilde{\mathbf{M}}\times\tilde{\mathbf{M}}^{A}-|\tilde{\mathbf{M}}|\mathbf{I}$ tests each element value so that it is a value test.
-Figure 15 shows the error deviation of a typical value test: the error deviation decreases linearly with increasing input uncertainty deviation $\delta x$, because the value should match unless at $\delta x\rightarrow 0$ when floating-point rounding errors dominate.
-The matrix size is the specific dimension in Figure 15.
-Because an element of the original matrix $\mathbf{M}$ appears multiple times in Formula (6.4), its result variance is very complicated using Formula (2.3).
+Formula (6.5) shows the relationship between $\mathbf{M}^{A}$ and $\mathbf{M}$ which leads to the definition of inverse matrix $\mathbf{M}^{-1}$ in Formula (6.6) [20].
+$\tilde{\mathbf{M}}\times\tilde{\mathbf{M}}^{A}-|\tilde{\mathbf{M}}|\mathbf{I}$ tests each element value such that it is a value test.
+Figure 15 shows the error deviation of a typical value test: the error deviation decreases linearly with increasing input uncertainty deviation $\delta x$, because the values should match unless at $\delta x\rightarrow 0$ when floating-point rounding errors dominate.
+The matrix size is the algorithm specific dimension in Figure 15.
+Because an element of the original matrix $\mathbf{M}$ appears multiple times in Formula (6.6), its variance, as obtained using Formula (2.3), is very complicated.
 For example, Formula (6.7) shows the simplest case for Formula (6.6): the first-order approximation of a 2x2 matrix.
-Contrary to traditional approach, statistical Taylor expansion uses Formula (6.4) for matrix inversion instead of Gaussian elimination [5], because logically, the result should be symmetric for all matrix elements, as what Formula (6.5) has demonstrated.
+Contrary to conventional approach, statistical Taylor expansion uses Formula (6.6) for matrix inversion instead of Gaussian elimination [5] because logically, the result should be symmetric for all matrix elements, as demonstrated by Formula (6.7).
 In Formula (6.6), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$.
-Figure 16 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices.
+Figure 16 shows that a strong linear correlation exists between conditional numbers [20] and the corresponding determinant precision of the matrices.
 As a reference, Figure 16 presents the Hilbert matrix [20] for each matrix size and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number.
-Thus, determinant precision can replace matrix condition number.
+Thus, determinant precision can replace matrix condition number to quantify overall matrix stability [20], and Formula (6.7) further shows the stability of each matrix element.
 7 Mathematical Library Functions
@@
 Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
+Uncertainty $\delta x\in[10^{-15},1]$ $\delta x\in[10^{-15},0.2x]$ $\delta x\in[10^{-15},0.2]$ $\delta x\in[10^{-15},1]$
 Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.978\pm 0.150$
-Table 1: The result error deviation of selected basic functions using variance arithmetic when $\delta x>10^{-15}$.
-The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverge regions are excluded.
+Table 1: The result error deviations of selected basic functions with Gaussian input noise $x\pm\delta x$ when $\delta x>10^{-15}$, obtained by variance arithmetic.
+The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverging regions are excluded.
 The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
 Figure 17: Error deviation for $\sin(x\pm\delta x)$ as a function of $x$ and $\delta x$.
-The x-axis represents $x$ value between $\pi$ and $\pi$.
-The y-axis represents $\delta x$ value between $10^{-16}$ and $1$.
-The z-axis shows the corresponding error deviation.
-Figure 18: Values and uncertainties of $\log(e^{x})-x$ and $e^{\log(x)}-x$ as functions of $x$, evaluated at $0.1$ increment.
-When $x$ is 2’s fractional such as $1/2$ or $1$, the result uncertainties are significantly smaller.
+The x-axis represents $x$ values between $\pi$ and $\pi$.
+The y-axis represents $\delta x$ values between $10^{-16}$ and $1$.
+The z-axis shows the corresponding error deviations.
+Figure 18: Values and uncertainties of $\log(e^{x})-x$ and $e^{\log(x)}-x$ as functions of $x$, evaluated in increment of $0.1$.
+When $x$ is 2’s fractional such as $1/2$ or $1$, the resulting uncertainties are significantly smaller because of floating-point representation.
 Figure 19: Normalized errors of $(x^{p})^{\frac{1}{p}}-x$ as functions of $x$ and $p$.
-Table 1 shows that Formula (2.15), (2.17), (2.20), and (2.22) provide almost perfect deviation for the result values which are calculated using corresponding math library functions by sampling from $x\pm\delta x$ in which $\delta x$ is Gaussian input noise deviation.
+Table 1 shows that by sampling random input $\tilde{x}$ from Gaussian noise $x+\delta x$, Formulas (2.15), (2.17), (2.20), and (2.22) provide almost perfect characterization $f(x)\pm\delta f$ for the result distributions of f(x~)f($\tilde${x}) in which f(x~)f($\tilde${x}) is the corresponding math library functions.
 Table 1 shows the results from Python, which differs slightly from those from either C++ or Java.
-The result error deviation is similar to Figure 14 but with the specific dimension as $x$ in $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$, or $c$ in $(1+\delta x)^{c}$.
-The coverage is only proper when $\delta x<10^{-15}$.
+The result error deviation is similar to that shown in Figure 14 but with the specific dimension as $x$ in $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$, or $c$ in $(1+\delta x)^{c}$.
+The coverage is proper when $\delta x<10^{-15}$.
 Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$.
@@
 As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$.
-The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ will be examined in greater detail in Section 10.
+The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ are examined in greater detail in Section 9.
 To test $f^{-1}(f(x))-x=0$ when $\delta x=0$ for the library functions:
   • Figure 18 shows that the value errors in $e^{\log(x)}-x$ are much less than those in $\log(e^{x})-x$. For $\log(e^{x})-x$, the error deviation is $0.41$ when $|x|\leq 1$, or 0 otherwise.
-  • Figure 19 shows that the error deviation for $(x^{p})^{1/p}-x$ is $0.56$, independent on neither $x$ nor $p$.
-The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has almost no value error, are not clear.
+  • Figure 19 shows that the error deviation for $(x^{p})^{1/p}-x$ is $0.56$, dependent on neither $x$ nor $p$.
+The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has nearly no value error, are not clear.
+It could relate to floating-point representation which stores each value $x$ as $\log_{2}(x)$.
+Figure 5 also suggests that using linear approximation $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$ have less error than $(x\pm\delta x)^{p}$.
 8 Moving-Window Linear Regression
@@
 The uncertainty for $Y$ is fixed at $0.2$.
-In the legend, Unadjusted refers to results obtained by directly applying Formula (8.5) and (8.6) using variance arithmetic, whereas Adjusted refers to using Formula (8.5) and (8.6) for $\alpha$ and $\beta$ values but Formula (8.7) and (8.8) for their variances.
+In the legend, Unadjusted refers to results obtained by directly applying Formulas (8.5) and (8.6) using variance arithmetic, whereas Adjusted refers to using Formulas (8.5) and (8.6) for $\alpha$ and $\beta$ values but Formulas (8.7) and (8.8) for their variances.
 Figure 21: Error deviations of the $\alpha+\beta\;Y$ fit vs time index.
@@
 For reference, the input time-series signal $Y$ is also plotted, with its values indicated on the y-axis on the right.
-Formula (8.1) and (8.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [5].
+Formulas (8.1) and (8.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [5].
 $\alpha$ $=\frac{\sum_{j}Y_{j}}{\sum_{j}1};$ (8.1)
@@
 For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window.
-This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (8.1) and (8.2) into Formula (8.3) and (8.4), respectively [17].
+This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formulas (8.1) and (8.2) into Formulas (8.3) and (8.4), respectively [17].
 $\alpha_{j}$ $=\alpha\;2H=\sum_{X=-H+1}^{H}Y_{j-H+X};$ (8.3)
 $\beta_{j}$ $=\beta\;\frac{H(H+1)(2H+1)}{3}=\sum_{X=-H}^{H}XY_{j-H+X};$ (8.4)
-The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [17].
+The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formulas (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formulas (8.5) and (8.6), respectively [17].
 $\beta_{j}$ $=\beta_{j-1}-\alpha_{j-1}+H\left(Y_{j-2H-1}+Y_{j}\right);$ (8.5)
@@
 $\delta^{2}\beta_{j}$ $=\sum_{X=-H}^{H}X^{2}(\delta Y_{j-H+X})^{2};$ (8.8)
-When the time series contains uncertainty, directly applying Formula (8.5) and (8.6) results in a loss of precision since both formulas reuse each input multiple times, thereby accumulating the variance of that input with every reuse.
-To prevent this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (8.6) and (8.5), respectively, while the variances should instead be computed using Formula (8.7) and (8.8), respectively.
+When the time series contains uncertainty, directly applying Formulas (8.5) and (8.6) results in a loss of precision since both formulas reuse each input multiple times, thereby accumulating the variance of that input with every reuse.
+To prevent this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formulas (8.6) and (8.5), respectively, while the variances should instead be computed using Formulas (8.7) and (8.8), respectively.
 Formula (8.8) is not progressive because the progressive form of $\delta^{2}\beta_{j}$ is more expensive in computation than Formula (8.8).
@@
 When (8.3) and (8.4) are applied to compute the uncertainties of $\alpha$ and $\beta$, both uncertainties increase exponentially with the time index $j$.
-In contrast, when Formula (8.3) and (8.4) are used exclusively for value calculation, while Formula (8.7) and (8.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$.
-Both are less than the input uncertainty $\delta Y$, due to the averaging effect of the moving window.
+In contrast, when Formulas (8.3) and (8.4) are used exclusively for value calculation, while Formulas (8.7) and (8.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$.
+Both are less than the input uncertainty $\delta Y$, because of the averaging effect of the moving window.
 8.3 Unspecified Input Error
@@
 This observation suggests that an error deviation exceeding 1 may indicate the presence of unspecified additional input errors beyond rounding errors, such as numerical errors in mathematical library functions.
-9 Regressive Generation of Sin and Cos
-Figure 22: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ versus $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
-Formula (9.2) and Formula (9.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (9.1).
-Formula (9.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, so that value errors will not accumulate when the regression order increases.
-$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (9.1)
-$\sin(\frac{\alpha+\beta}{2})=\sqrt{\frac{1-\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1-\cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)}{2}};$ (9.2)
-$\cos(\frac{\alpha+\beta}{2})=\sqrt{\frac{1+\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1+\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)}{2}};$ (9.3)
-$\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (9.4)
-Formula (9.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation, resulting in excessive uncertainty.
-In Figure 22, the Quart sine function exhibits uncertainties proportional to $\sin(\frac{\pi}{2^{L}})$, whereas the regression sine function shows increasing uncertainties as $x\rightarrow 0$.
-Unlike catastrophic cancellation in floating-point arithmetic, variance arithmetic uses very coarse precision to signal that the regression algorithm is unfit to compute $\sin(x\rightarrow 0)$.
 9 FFT (Fast Fourier Transformation)
@@
 $h[k]$ $=\frac{1}{N}\sum_{n=0}^{N-1}H[n]\;e^{-\frac{i2\pi}{N}nk};$ (9.2)
-For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (10.1) and (10.2), respectively [5].
-As a convention, $k$ denotes time index for $h[k]$ as a waveform, while $n$ represents frequency index for $H[n]$ as a spectrum.
+For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formulas (9.1) and (9.2), respectively [5].
+As a convention, $k$ denotes the time index for $h[k]$ as a waveform, whereas $n$ represents the frequency index for $H[n]$ as a spectrum.
 Figure 22: The DFT spectrum $H[n]$ of signal $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, as intensity (y-axis) and phase (embedded y-axis) versus frequency index $n\in[0,18]$ (x-axis and embedded x-axis) for different signal frequency $f$ (legend).
@@
 Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics
-To avoid the modeling errors inherent in DFT, only Formula (10.1) and (10.2) are used in this paper.
+To avoid the modeling errors inherent in DFT, only Formulas (9.1) and (9.2) are used in this study.
 9.2 FFT (Fast Fourier Transformation)
-When $N=2^{L}$, where $L$ is a natural number named FFT order, the generalized Danielson-Lanczos lemma can be applied to DFT to produce FFT [5].
-  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations such as Formula (2.10), (2.11), (2.12), and (2.13).
+When $N=2^{L}$, where $L$ is a natural number called the FFT order, the generalized Danielson-Lanczos lemma can be applied to DFT to produce FFT [5].
+  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations such as Formulas (2.10), (2.11), (2.12), and (2.13).
   • When $L$ is large, the substantial volume of input and output data enables high-quality statistical analysis.
-  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of two multiplications.
-  • Each step in the forward transformation doubles the variance, so the uncertainty deviation increases with the FFT order $L$ as $\sqrt{2}^{L}$.
+  • The computational complexity is proportional to $L$, because increasing $L$ by 1 adds an additional step involving a sum of two multiplications.
+  • Each step in the forward transformation doubles the variance; hence the uncertainty deviation increases with the FFT order $L$ as $\sqrt{2}^{L}$.
 Because the reverse transformation divides the result by $2^{L}$, its uncertainty deviation decreases with $L$ as $\sqrt{1/2}^{L}$.
-Consequently, the uncertainty deviation for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
-  • The forward and reverse transformations are identical except for a sign difference, meaning that they are essentially the same algorithm, and any observed difference arises solely from the input data.
-In normal usage, forward and reverse FFT transforms differ in their data prospective of time domain versus frequency domain:
-  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly by cancellation.
-  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly by spreading.
-The question is whether variance arithmetic can work properly in these two contrary cases.
+Consequently, the uncertainty deviation for the round-trip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
+  • The forward and reverse transformations are identical except for a sign difference, implying that they are essentially the same algorithm, and any observed difference arises solely from the input data.
 9.3 Testing Signals
 The following signals are used for testing:
-  • Sin: $h[k]=\sin(2\pi kf/N),f=1,2,...N/2-1$.
-  • Cos: $h[k]=\cos(2\pi kf/N),f=1,2,...N/2-1$.
+  • Sin: $h[k]=\sin(2\pi kf/N),f=1,2,...\frac{N}{2}-1$.
+  • Cos: $h[k]=\cos(2\pi kf/N),f=1,2,...\frac{N}{2}-1$.
   • Linear: $h[k]=k$, whose DFT is given by Formula (9.3).
@@
     $H[n]$ $=\frac{dG}{dy}=\begin{cases}n=0:\;\;\;\frac{N(N-1)}{2}\\ n\neq 0:\;\;\;-\frac{N}{2}(1+i\frac{\cos(n\frac{\pi}{N})}{\sin(n\frac{\pi}{N})})\end{cases};$ (9.3)
+For Sin or Cos signals, forward and reverse DFT transforms differ in their data prospective of time domain versus frequency domain:
+  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly during cancellation.
+  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
+The question is whether variance arithmetic can work properly in these two contrary cases.
 9.4 Trigonometric Library Errors
 Figure 23: Difference between Library and Quart $\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend).
-The uncertainties of the Quart $\sin(x)$ is $\sin(x)$ ULP, which shows a periodicity of $\pi$.
+The uncertainty of the Quart $\sin(x)$ is $\sin(x)$ ULP, which shows a periodicity of $\pi$.
 Figure 24: Difference between the Library and the Quart $\cos(x)/\sin(x)$ (y-axis) for $x=2\pi j/2^{L},j=0,1\dots 2^{L+2}$ (x-axis), and $L=5,6$ (legend).
-Formula (10.1) and (10.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, where $L$ is the FFT order.
+Formulas (9.1) and (9.2) restrict the use of $\sin(x)$ and $\cos(x)$ to $x=2\pi j/2^{L}$, where $L$ is the FFT order.
 To minimize numerical errors in computing $\sin(x)$, the following indexed sine can be used in place of standard library sine functions:
@@
   2. 2.
-    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are library sine directly, while the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from library $\cos(\pi(2^{L-1}-j)/2^{L})$.
+    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are the library sine directly, whereas the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from the library $\cos(\pi(2^{L-1}-j)/2^{L})$.
   3. 3.
@@
 The constructed indexed $\sin(x)$ is referred to as the Quart sine function. In contrast, the direct use of the standard library $\sin(x)$ is referred to as the Library sine function.
-Because the Quart sine function strictly preserves the symmetry and periodicity of sine function, it provides superior numerical accuracy compared to the Library sine function.
+Because the Quart sine function strictly preserves the symmetry and periodicity of the sine function, it provides better numerical accuracy than the Library sine function.
   • Figure 23 shows that the absolute value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
-  • Figure 25 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$ but are $10^{2}$ times larger than those observed for $\sin(x)$.
-Therefore, the linear spectrum in Formula (10.3) contains significantly larger numerical errors when computed using library sine functions.
+  • Figure 24 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but is $10^{2}$ times larger than that observed for $\sin(x)$.
+Therefore, the linear spectrum in Formula (9.3) contains significantly larger numerical errors when computed using the Library sine functions.
 9.5 Using Quart Sine for Sin/Cos Signals
-Figure 26: Histograms of normalized errors of Sin/Cos signals for forward, reverse and roundtrip transformations (legend) using the Quart sine function. The FFT order is $18$.
-With FFT order as the specific dimension, the error deviations for forward and reverse transformations resemble Figure 14, while those for round-trip transformation is almost identical to Figure 15, independent of Sin or Cos signals, or frequency of the signals.
+Figure 25: Error distributions of normalized errors of Sin/Cos signals for forward, reverse and round-trip transformations (legend) using the Quart sine function. The FFT order is $18$.
+With the FFT order as the specific dimension, the error deviations obtained using the Quart sine function for forward and reverse transformations resemble those shown in Figure 14, whereas those for round-trip transformation are nearly identical to those shown in Figure 15, independent of the signal frequency or whether a Sin or Cos signals is used.
 Therefore, the results for Sin and Cos signals across all frequencies are pooled together for statistical analysis, under the unified category Sin/Cos signals.
-When the FFT order $L$ is less than 8, the error deviations oscillate around $1$ due to insufficient sample count of $2^{L}$.
+When the FFT order $L$ is less than 8, the error deviations oscillate around $1$ because of an insufficient sample count of $2^{L}$.
 Even data for forward and reverse transformations are drastically different, variance arithmetic works effectively in all cases.
-When $L=18$ and $\delta x=0$, Figure 26 shows that the error distributions of Sin/Cos signal resemble Normal distribution, with an additional Delta-like distribution at $\tilde{z}=0$.
-The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in sine function.
+When $L=18$ and $\delta x=0$, Figure 25 shows that the error distributions of Sin/Cos signal resemble Normal distributions, with an additional Delta-like distribution at $\tilde{z}=0$.
+The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in the sine function.
 9.6 Using Library Sine for Sin/Cos Signals
-Figure 27: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
+Figure 26: Error distributions of normalized errors of Sin/Cos signal for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
 Figure 27: FFT value error spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
 The legend distinguishes between uncertainty and value error.
-The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
+The x-axis represents the frequency index, and the y-axis represents both uncertainty and value error.
 Figure 28: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation.
 The legend distinguishes between uncertainty and value error.
-The x-axis represents the time index, while the y-axis represents both uncertainty and value error.
-Figure 30: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the roundtrip transformation.
+The x-axis represents the time index, and the y-axis represents both uncertainty and value error.
+Figure 29: FFT value error waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the round-trip transformation.
 The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the frequency index, while the y-axis represents both uncertainty and value error.
-Figure 31: Error deviation (z-axis) of FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
-With FFT order as the specific dimension, the error deviations for the forward and the reverse transformations resemble Figure 14, while those for the round-trip transformation is almost identical to Figure 15.
+The x-axis represents the frequency index, and the y-axis represents both uncertainty and value error.
+Figure 30: Error deviation (z-axis) of the FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT order $L$ (y-axis).
+With FFT order as the specific dimension, the error deviations obtained using the Library sine for the forward and reverse transformations resemble those shown in Figure 14, whereas those for the round-trip transformation are nearly identical to those shown in Figure 15.
 In addition, the error deviations are larger than $1$ when $\delta x<10^{-15}$ for the forward transformation and $\delta x<10^{-14}$ for the reverse transformation.
-When $\delta x=0$, the error deviations for the reverse transformation increase with FFT order, to $6.2$ at FFT order $18$.
-As shown in Figure 24, the Library sine function contains more numerical errors, so the error distribution for the reverse transformation using Library sine function in Figure 27 is more structured and broader than that in Figure 26 using Quart sine function, while the error distribution for the forward transformations is more similar.
-This difference is consistent with a larger error deviation $6.2>1$ for the reverse transformation, but a comparable error deviation $1.1>1$ for the forward transformation when $\delta x=0$.
-Using the Library sine function, for a sine wave with a frequency of $3$, Figure 28, 29, and 30 presents the value errors for the forward, the reverse, and the round-trip transformations, respectively.
-In the reverse transformation, value errors exhibit a clear trend of increasing with the time index.
+When $\delta x=0$, the error deviations for the reverse transformation increase with the FFT order, to $6.2$ at FFT order $18$.
+As shown in Figure 23, the Library sine function contains more numerical errors, as a result, the error distribution for the reverse transformation obtained using the Library sine function in Figure 26 is more structured and broader than that shown in Figure 25 obtained using the Quart sine function, whereas the error distributions for the forward transformations are more similar.
+This difference is consistent with a larger error deviation $6.2>1$ for the reverse transformation, compared with the comparable error deviation $1.1>1$ for the forward transformation when $\delta x=0$.
+When $\delta x=0$, using the Library sine function, for a sine wave with a frequency of $3$, Figures 27, 28, and 29 present the value errors for the forward, reverse, and round-trip transformations, respectively.
+In the reverse transformation, the value errors exhibit a clear trend of increasing with the time index.
 These large value errors appear systematic rather than random and visually resemble a resonant pattern.
 Similar increases are observed at other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy.
-In contrast, such resonance is absent for the roundtrip transformation as shown in Figure 30, as well as when using the Quart sine function.
-Figure 31 demonstrates that the error deviations increase with sine or cosine frequency.
-Figure 24 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern in Figure 29.
+In contrast, such resonance is absent from the round-trip transformation shown in Figure 29, as well as when using the Quart sine function.
+Figure 30 demonstrates that the error deviations increase with sine or cosine frequency, regardless of FFT order $L$ when $L>8$.
+Figure 23 indicates that the numerical errors of $cos(x)/sin(x)$ obtained using the Library sine function increase with a periodicity of $\pi$, which can resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern shown in Figure 28.
 At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
-To suppress these numerical error resonances, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals.
+To suppress this numerical error resonance, an input noise of $\delta x=10^{-14}$ must be added to the sine or cosine signals.
 Such resonance of numerical errors can easily and mistakenly be taken as signals.
-10.7 Using Quart Sine for Linear Signal
+9.7 Using Quart Sine for Linear Signals
 Figure 31: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Quart sine function.
 Figure 32: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Quart sine function.
-Figure 34: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
-Figure 32 and 33 show the error deviation for the forward and the reverse transformations, respectively.
+Figure 33: Error distributions of normalized errors of Linear signals for forward, reverse and round-trip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
+Figures 31 and 32 show the error deviations for the forward and the reverse transformations, respectively.
 The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation.
 In other areas, both transformations achieve proper coverage with error values around $1$.
-When $L=18$ and $\delta x=0$, the error distribution of the reverse transformation in Figure 34 is narrower than that in Figure 27. The corresponding error deviations are $1.5<6.2$, respectively.
-10.8 Using Library Sine for Linear Signal
-Figure 35: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
+When $L=18$ and $\delta x=0$, the error distribution of the reverse transformation in Figure 33 is narrower than that shown in Figure 26.
+The corresponding error deviations are $1.5<6.2$, respectively.
+The error deviations for the round-trip transformation resemble those in Figure 15 but with the FFT order as the specific dimension.
+9.8 Using Library Sine for Linear Signals
+Figure 34: Error distributions of normalized errors of Linear signals for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
 Figure 35: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Library sine function.
 Figure 36: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Library sine function.
-Figure 38: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the roundtrip transformations of Linear signals computed using the Library sine function.
+Figure 37: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the round-trip transformations of Linear signals computed using the Library sine function.
 Figure 34 shows that the reverse error distribution when $\delta x=0$ is no longer bound.
-The difference between Figure 35 and Figure 34 is consistent with the large numerical errors as demonstrated in Figure 25.
-Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine.
-Figure 36 and 37 shows a much smaller ideal coverage areas than those in Figure 32 and 33.
-Because result deviations grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region.
-Outside the ideal coverage region, proper coverage cannot be achieved for the reverse transformation.
+The difference between Figures 34 and 33 is consistent with the large numerical errors as demonstrated in Figure 24.
+Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine function.
+Figures 35 and 36 show much smaller ideal coverage areas than those shown in Figures 31 and 32.
+Because uncertainty deviations grow more slowly in the reverse transformation than in the forward transformation, the former exhibits a smaller ideal coverage region.
+Outside of the ideal coverage region, proper coverage cannot be achieved for the reverse transformation.
 Furthermore, the range of input noise that produces ideal coverage decreases with increasing FFT order.
-At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable.
-Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function.
-Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
-Figure 38 shows that, even when variance arithmetic can no longer effectively track the value errors for the forward and the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as the plateau region at high $L$ and low $\delta x$ in Figure 38.
-Such error cancellation is due to dependence tracing in statistical Taylor expansion.
+At sufficiently high FFT orders (visually beyond FFT order $25$ for the reverse transformation), ideal coverage may no longer be achievable.
+Although FFT is widely regarded as one of the most robust numerical algorithms [5][10], and is generally insensitive to input errors, it can still fail because heavy calculation can amplify numerical errors in the Library sine function to completely obscure true signals.
+Such deterioration in calculation is not easily detectable when using conventional floating-point arithmetic.
+Figure 37 shows that, even when variance arithmetic can no longer effectively track the value errors for either the forward or the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as shown by the plateau region at high $L$ and low $\delta x$.
+Such error cancellation arises from dependence tracing in statistical Taylor expansion.
 9.9 Ideal Coverage
-Figure 39: Result error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
+Figure 38: Error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
 Adding noise to the input can suppress unspecified input errors.
-After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, while the error distribution for the round-trip transformation becomes Delta.
+After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, whereas the error distribution for the round-trip transformation becomes Delta.
 Figure 38 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
@@
   • As expected, the resulting uncertainty deviations for the reverse transformations decrease with FFT order $L$ as $1/\sqrt{2}^{L}$.
-  • As expected, the resulting uncertainty deviations for the round-trip transformations remains equal to the corresponding input uncertainties of $10^{-3}$.
+  • As expected, the resulting uncertainty deviations for the round-trip transformations remain equal to the corresponding input uncertainties of $10^{-3}$.
   • As expected, the resulting error deviations for the forward and the reverse transformations remain constant at $1$.
-  • As expected, the resulting error deviations for the round-trip transformations are far less than $1$ but increase with FFT order $L$ exponentially due to increasing calculation.
+  • As expected, the resulting error deviations for the round-trip transformations are far less than $1$ but increase exponentially with FFT order $L$ because of increasing calculation.
 Signal Sine function Forward Reverse Roundtrip
@@
 Table 2: The measured minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different signals and sine functions.
-Table 2 shows the minimal required noise to achieve ideal coverage for FFT transformations at FFT order $L=18$ for different signals and sine functions, which is consistent with the corresponding error distributions in Figure 26, 27, 34, and 35.
-This shows that an error distribution can qualify its input uncertainty quantification.
-Without knowing precise result, similar histogram can be constructed from the calculated mean $\overline{f}$ and deviation $\delta f$, and the result data set for $f$.
-It is worth investigating if such empirical histogram has similar power to reveal tracking of value errors by variance arithmetic.
+Table 2 shows the minimal required noise to achieve ideal coverage for FFT transformations at FFT order $L=18$ for different signals and sine functions, which is consistent with the corresponding error distributions shown in Figures 25, 26, 33, and 34.
+An error distribution can reflect its input uncertainty coverage.
+Without knowing the precise result, similar histogram can be constructed from the result data set for $f$ using the calculated mean $\overline{f}$ and deviation $\delta f$.
+It is worth investigating if such an empirical histogram has similar power to reveal input uncertainty coverage.
+10 Regressive Generation of Sine
+Figure 39: The resulting precision (left y-axis) and uncertainties (right y-axis) of $\sin(\pi j/2^{18})$ versus index $j$ (x-axis) using either Quart or regressive sine function (legend).
+$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (10.1)
+$\sin(\frac{\alpha+\beta}{2})=\sqrt{\frac{1-\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1-\cos(\alpha)\cos(\beta)+\sin(\alpha)\sin(\beta)}{2}};$ (10.2)
+$\cos(\frac{\alpha+\beta}{2})=\sqrt{\frac{1+\cos(\alpha+\beta)}{2}}=\sqrt{\frac{1+\cos(\alpha)\cos(\beta)-\sin(\alpha)\sin(\beta)}{2}};$ (10.3)
+$\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (10.4)
+Formulas (10.2) and (10.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1).
+Formula (10.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, such that value errors will not accumulate when the regression order $L$ increases.
+Formula (10.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation [9][10].
+As shown in Figure 39, the Quart sine function exhibits a constant precision around $10^{-16}$, whereas the regression sine function shows increasing precision up to $10^{-7}$ as $x\rightarrow 0$.
+Unlike the sneaky catastrophic cancellation in floating-point arithmetic, variance arithmetic uses coarser precision to demonstrate where and by how much the regression algorithm becomes unfit to compute $\sin(x)$.
 11 Conclusion and Discussion
 11.1 Summary
-When uncorrelated uncertainty condition is met, statistical Taylor expansion outputs the mean, deviation, and reliability of an analytic expression.
-It tracks the variable dependencies in intermediate steps, rejects invalid calculations, and explicitly incorporates the sample counts and the input uncertainty distributions into the statistical analysis.
-While statistical Taylor expansion eliminates the dependency problem, it also removes most execution flexibility.
-The presence of ideal coverage is a necessary condition for a numerical algorithm using statistical Taylor expansion to be considered correct.
-The ideal coverage defines the optimal applicable range for an algorithm.
-Variance arithmetic simplifies statistical Taylor expansion by introducing numerical rules to elimination invalid results, such as divergent, negative variance, unstable, infinite, or unreliable results.
-It provides proper coverage for floating-point rounding errors.
+When the uncorrelated uncertainty condition is satisfied, statistical Taylor expansion produces the mean, deviation, and reliability of an analytic expression.
+It tracks the variable dependencies in intermediate steps and rejects invalid calculations.
+Unlike conventional approaches, it explicitly incorporates the sample counts and uncertainty distributions into its result.
+Although statistical Taylor expansion eliminates the dependency problem, it also reduces execution flexibility.
+The presence of ideal coverage is a necessary condition for a numerical algorithm based on statistical Taylor expansion to be considered correct.
+Ideal coverage defines the optimal range of applicability for an algorithm.
+Variance arithmetic simplifies statistical Taylor expansion by introducing numerical rules that eliminate invalid results, such as divergent, negative variance, unstable, infinite, or unreliable results.
+It also provides proper coverage for floating-point rounding errors.
 The applicability of variance arithmetic has been demonstrated across a wide range of computational scenarios.
-Library mathematical functions should be recalculated using variance arithmetic, so that each output value is accompanied by its corresponding uncertainty.
-Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant effects on numerical results.
 The code and analysis framework for variance arithmetic are available as an open-source project at https://github.com/Chengpu0707/VarianceArithmetic.
-A more detailed description of this paper is presented at https://arxiv.org/abs/2410.01223.
+A more detailed description of this study is presented at https://arxiv.org/abs/2410.01223.
 11.2 Improvements Needed
-This paper presents statistical Taylor expansion and variance arithmetic, which is still in early stage of development, with several important questions remaining.
-Bound momentum $\zeta(2n)$ needs to be extended to all probability distributions.
-The statistical bounding to find the bounding range $\kappa$ for a sample count needs to be extended to discrete distributions.
-The choice of ideal bounding range $\hat{\kappa}$ needs to extend to other distributions.
-The statistical meaning for ideal ratio $\alpha$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\alpha$ the possibility for the result to be bound by $\hat{\kappa}$?
+This study presents statistical Taylor expansion and variance arithmetic, which are still in early stages of development. Accordingly, several important questions remain.
+Library mathematical functions should be recalculated using variance arithmetic, to ensure that each output value is accompanied by its corresponding uncertainty.
+Without this refinement, the value errors in the library functions can produce unpredictable and potentially significant result errors.
+Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions.
+The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of $1-\zeta(2,\kappa)$.
+The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions.
+The statistical meaning for ideal leakage $\varepsilon\in[0,1]$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\varepsilon$ the possibility for the result to be not bound by $\hat{\kappa}$?
 The performance of variance arithmetic must be improved for broader practical adoption.
-The fundamental formulas of statistical Taylor expansion, Formula (2.5), (2.3), (2.8), and (2.3), contain a large number of independent summations, making them excellent candidates for parallel processing.
+The fundamental formulas of statistical Taylor expansion, Formulas (2.5), (2.3), (2.8), and (2.3), contain many independent summations, making them excellent candidates for parallel processing.
 Moreover, the inherently procedural nature of these formulas allows statistical Taylor expansion to be implemented efficiently at the hardware level.
 A key open question is whether variance arithmetic can be adapted to achieve ideal coverage for floating-point rounding errors, because many theoretical calculations lack explicit input uncertainties.
-Variance arithmetic does not adjust uncertainty variance when floating-point rounding error occurs during calculation.
-This leads to error deviations larger than 1.
-Detecting floating-point rounding error also needs hardware implementation for efficiency.
-In variance arithmetic, deviations are comparable with values, but variances are used in calculation.
-This effectively limits the range of deviations to the square root of that of values.
-If the sign bit of the floating type can be re-purposed as an exponent bit in a new unsigned floating-point representation, the range of the deviation will be identical to that of values.
+Variance arithmetic does not adjust uncertainty characterization when floating-point rounding error occurs during calculation, which leads to error deviations larger than 1.
+Detecting floating-point rounding error and adjusting uncertainty characterization in real time needs hardware implementation for efficiency.
+In variance arithmetic, deviations are comparable to values; however variances are used in calculation.
+This approach effectively limits the range of deviations to the square root of that of the values.
+If the sign bit of the floating type can be re-purposed as an exponent bit in a new unsigned floating-point representation, the range of the deviations will be identical to that of the values.
 When an analytic expression undergoes statistical Taylor expansion, the resulting expression can become highly complex, as in the case of matrix inversion.
-However, modern symbolic computation tools such as SymPy and Mathematica can greatly facilitate these calculations.
+Modern symbolic computation tools such as SymPy and Mathematica can significantly facilitate such calculations.
 This observation suggests that it may be time to shift from purely numerical programming toward analytic programming, particularly for problems that possess inherently analytic formulations.
 As an enhancement to dependency tracing, source tracing identifies each input’s contribution to the overall result uncertainty.
-This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and in turn guide targeted improvements in data acquisition and processing strategies.
-For example, Formula (2.31) can guide how to improve the ideal ratio of $x\pm y$.
-Because traditional numerical approaches are based on floating-point arithmetic, they must be reexamined or even reinvented within the framework of variance arithmetic.
+This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and guides them to targeted improvements in data acquisition and processing strategies.
+For example, Formula (2.33) can provide guidance on improving the ideal leakage of $x\pm y$.
+Because conventional numerical approaches are based on floating-point arithmetic and path dependent in general, they must be reexamined or even reinvented within the framework of variance arithmetic.
 For instance, most conventional numerical algorithms aim to identify optimal computational paths, whereas statistical Taylor expansion conceptually rejects all path-dependent calculations.
@@
 11.3 Possible Connections to Quantum Physics
-For input distribution without bounding, Formula (2.3) may not converge, for example, when the input uncertainty is Laplace [4] so that $\zeta(2n)=(2n)!$.
-Formula (2.3) converges in all cases only when the input uncertainty is bounded due to sampling, which resembles the need for re-normalization in quantum field theory [21].
-The quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: what it is depends on how it is measured.
-For example, because mathematical divergence seems to correspond to distributional zero, it is even tempting to postulate if a localized experiment changes $\hat{\kappa}$ to produce distributional zero and the quantum result as “waveform collapse” [21].
-Both resemblances are worth further investigation.
-It is also worthwhile investigating if uncertainty bias produces Lamb shift [21].
+For input distribution without bounding, Formula (2.3) may not converge at all, for example, when the input uncertainty is Laplace [4], $\zeta(2n,\infty)=(2n)!$.
+Formula (2.3) converges in all cases only when the input uncertainty is bounded because of sampling, which resembles the need for re-normalization in quantum field theory [21].
+Furthermore, the quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: What it is depends on how it is measured.
+For example, it is possible that an experimental measurement reduces $\hat{\kappa}$ for a wave function, to cause the wave function to collapse.
+Both resemblances are worthy of further investigation.
 12 Statements and Declarations

### v19

@@
   • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this study.
-  • Calculating the reliability $\alpha\in[0,1]$ of $f\pm\delta f$ seems completely new.
+  • Calculating the reliability $[0,1]$ of $f\pm\delta f$ from input uncertainty distributions and sample counts seems completely new.
 As a new approach, statistical Taylor expansion is superior to all existing numerical arithmetic.
 Conventional floating-point arithmetic [6][7][8] computes only the result value $f$.
-Because $f$ contains unknown amount of rounding error [9][10][11], a floating-point representation with $10^{-16}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision.
+Because $f$ contains unknown amount of rounding error [9][10][11], a 32-bit floating-point representation with $10^{-7}$ resolution may not be good enough for inputs with $10^{-2}$ to $10^{-6}$ precision.
 However, statistical Taylor expansion can account for rounding errors as part of the result deviation $\delta f$.
@@
   • Section 6 describes the applications of variance arithmetic to matrix inversion, distinguishing between distribution tests and value tests.
-  • Section 7 discusses the evaluation of variance arithmetic on common mathematical library functions, showing the effect of distribution pole.
+  • Section 7 discusses the evaluation of variance arithmetic on common mathematical library functions, showing the effect of distributional pole.
   • Section 8 reveals the accumulation of numerical errors by reusing an input multiple time in a moving-window progressive algorithm.
@@
 […]
-$\overline{f(x)}$ $=\int_{-\varrho}^{+\kappa}f(x+\tilde{z})\rho(\tilde{z})d\tilde{z}=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
-$\delta^{2}f(x)$ $=\overline{(f(x)-\overline{f(x)})^{2}}=\overline{f(x)^{2}}-\overline{f(x)}^{2}$
-$=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
+$\overline{f(x)}$ $=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
+$\delta^{2}f(x)$ $=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
 An analytic function $f(x)$ can be accurately evaluated in a range using the Taylor series as shown in Formula (2.4).
@@
 With the mean reverting condition of Formula (2.3):
-$\overline{x\pm y}$ $=\zeta(0,\kappa_{x})x\pm\zeta(0,\kappa_{x})y;$ (2.10)
-$\delta^{2}(x\pm y)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}+\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.11)
-$\overline{xy}$ $=\zeta(0,\kappa_{x})x\;\zeta(0,\kappa_{y})y;$ (2.12)
-$\delta^{2}(xy)$ $=\zeta(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta(2,\kappa_{y})(\delta y)^{2}+\zeta(2,\kappa_{x})(\delta x)^{2}\;\zeta(2,\kappa_{y})(\delta y)^{2};$ (2.13)
+$\overline{x\pm y}$ $=\zeta_{x}(0,\kappa_{x})x\pm\zeta_{y}(0,\kappa_{x})y;$ (2.10)
+$\delta^{2}(x\pm y)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.11)
+$\overline{xy}$ $=\zeta_{x}(0,\kappa_{x})x\;\zeta_{y}(0,\kappa_{y})y;$ (2.12)
+$\delta^{2}(xy)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta_{y}(2,\kappa_{y})(\delta y)^{2}+\zeta_{x}(2,\kappa_{x})(\delta x)^{2}\;\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.13)
 When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formulas (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formulas (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
@@
 Formulas (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
-$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\zeta(n,\kappa)\frac{1}{n!};$ (2.14)
+$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{\zeta(n,\kappa)}{n!};$ (2.14)
 $\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!\;(n-j)!};$ (2.15)
@@
 $\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n,\kappa)}{n!};$ (2.19)
-$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\eta(j,x)\eta(n-j,x)}{j!(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.20)
+$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\eta(j,x)\eta(n-j,x)\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!(n-j)!};$ (2.20)
 Formulas (2.21) and (2.22) give the mean and variance for $x^{c}$, respectively:
@@
 Figure 3: When $\kappa=5$, the value of bound moment $\zeta(2n,\kappa)$ (left y-axis) for Gaussian distribution with order $2n$ (x-axis) versus $2N(\kappa)\frac{\kappa^{2n+1}}{2n+1}$ (left y-axis), and their ratio (right y-axis).
+The value of bound moment $\zeta(2n,\kappa)$ for Uniform distribution is also provided as an comparison.
 $\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.23)
-$\lim_{n\rightarrow+\infty}$ $\zeta(2n,\kappa)=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1};$ (2.24)
+[…]
 $\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)=\frac{\kappa^{n}}{n};$ (2.25)
@@
 Figure 5: Measured linear leakage $\omega(\kappa,N)$ (left y-axis) and corresponding measured bound range $\kappa$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is Gaussian, with different measuring bound range $\kappa_{s}$ (legend).
-Figure 6: Ideal leakage (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
-Figure 7: Ideal leakage (y-axis) for varying bound range $\kappa$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=6$ for Gaussian input uncertainties.
+Figure 6: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
+When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
+Figure 7: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) for Uniform input uncertainties.
+When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
 $f(x)-\overline{f(x)}$ $\simeq(\delta x)^{2}\zeta(2,\kappa)f_{x}^{(2)};$ (2.28)
@@
 Because $\omega(\kappa)\neq\omega(\kappa,N)$ for finite $N$, let $\omega(\kappa)=\omega(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bound range, and $\kappa(\kappa_{s},N)$ is the measured bound range, with the latter used in calculating $\zeta(n,\kappa)$.
-$\epsilon(\kappa)$ $=\omega(\kappa)+2N(\kappa)\kappa;$ (2.30)
+$\epsilon(\kappa)$ $=\omega(\kappa)+2\rho(\kappa)\kappa;$ (2.30)
 $\epsilon(\kappa)$ $=1-\xi(\frac{\kappa}{\sqrt{2}});$ (2.31)
@@
 Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$.
-When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, while Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
-Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
-Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable result bounding leakage is smaller for larger bounding range $\kappa$.
-It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\hat{\kappa}$ is larger.
+When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, whereas Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
+Figure 4 shows that $\omega(\kappa_{s})<\omega(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\omega(\kappa_{s},N)=\omega(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
+Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable resulting linear leakage $\omega$ is smaller for larger measuring bounding range $\kappa_{s}$.
+It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\kappa_{s}$ is larger.
 $\varepsilon(x\pm y)$ $=\frac{\omega_{x}(\delta x)^{2}+\omega_{y}(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.33)
 Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example.
-Figure 6 shows that the stable ideal leakages for the selected functions are less than $10^{-4}$ for $\hat{\delta^{2}}f$, which is small enough for most applications.
-Figure 7 compares the ideal leakages for $\hat{\delta^{2}}f$ for the selected functions when different $\hat{\kappa}$ is used.
+Figure 6 compares ideal leakage $\varepsilon$ and linear leakage $\omega$ for the selected functions when the input uncertainty distribution is Gaussian, and shows that the example stable ideal leakages are less than $10^{-4}$ when $\hat{\kappa}=5$, which is small enough for most applications.
+The unbiased bounding range $\kappa_{min},\kappa_{max}]$ of Uniform distribution is estimated as $1+\frac{1}{N}$ times of the minimal and maximal values in sampled data set.
+The bonding leakage $\epsilon$ is the percentage of $\kappa_{min},\kappa_{max}]$ outside the true bounding range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$.
+Because $\zeta(2n)$ for Uniform distribution increases much slower than $\zeta(2n,5)$ for Gaussian distribution, Figure 7 shows that for the selected functions, ideal leakage equals linear leakage for Uniform input uncertainty, and both decrease linearly with sample count $N$.
+The difference between Figures 6 and 7 shows the effect of input uncertainty distribution.
 2.7 Dependency Tracing
@@
 $\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.35)
-When all inputs satisfy the uncorrelated uncertainty assumption, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
+When all inputs satisfy the uncorrelated uncertainty condition, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
   • Formula (2.34) expresses $\delta^{2}(f+g)$, The dependency tracing of is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
@@
   • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$.
-For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
+For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original inputs should be restored.
   • Formula (2.35) gives the variance of the linear transformation of a function, which can be applied to other formulas for more general dependency tracing.
@@
 This requirement often conflicts with conventional numerical methods for analytic functions:
-  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships among the original input variables.
-  • Similarly, conditional executions are often employed to optimize performance and minimize rounding errors, for example, using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [20].
-For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 6.
   • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls.
@@
 For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
-  • Similarly, large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$.
-This approach also introduces the dependency problem by ignoring dependency tracing within $g(x)$ affecting $f(g(x))$, such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}$.
+  • Large calculations are often divided into sequential steps, such as computing $f(g(x))$ as $f(y)|_{y=g(x)}$.
+This approach also introduces the dependency problem by ignoring dependency tracing within $g(x)$ affecting $f(g(x))$, such as $\overline{(\sqrt{x})^{2}}>\overline{\sqrt{x^{2}}}>\overline{x}$ and $\delta^{2}(\sqrt{x})^{2}>\delta^{2}\sqrt{x^{2}}>\delta^{2}x$.
+  • Conditional executions are often employed to optimize performance and minimize rounding errors, for example, using Gaussian elimination to minimize floating-point rounding errors in matrix inversion [20].
+For dependency tracing, such conditional executions should instead be replaced by direct matrix inversion as described in Section 6.
+  • Traditionally, intermediate variables are widely used in computations; however, this practice disrupts dependency tracing by obscuring the relationships among the original input variables.
 Dependency tracing therefore removes nearly all flexibility from traditional numerical executions, effectively eliminating the associated dependency problems.
@@
 Figure 8: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty.
-The corresponding uncertainty bias and uncertainty are also shown (right y-axis).
+The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
 When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
 Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty.
-The corresponding uncertainty bias and uncertainty are also shown (right y-axis).
+The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
 When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
@@
 Figure 10: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for Gaussian uncertainty.
-The corresponding uncertainty bias and uncertainty are also shown (right y-axis).
+The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
 The x-axis is expressed in units of $\pi$.
 Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty.
-The corresponding uncertainty bias and uncertainty are also shown (right y-axis).
+The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
 The x-axis is expressed in units of $\pi$.
-In addition to convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$.
+In addition to convergence, the variance expansion may yield negative results when the expansion is terminated arbitrarily, as in Formula (2.20) for $sin(x\pm\delta x)$.
 Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty.
@@
 3.3 Floating-Point Rounding Errors
-Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$.
+Variance arithmetic incorporates floating-point rounding errors as $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$.
 Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP (Unit in the Last Place [8]) of $x$, because rounding errors are shown to be uniformly distributed within ULP [17].
@@
 However, if the input uncertainty is known only to order of magnitude, proper coverage is achieved when the error deviations fall within the range $[0.1,10]$.
-When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added, until ideal coverage is attained.
+When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added to inputs, until ideal coverage is attained.
 The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations.
@@
   • When the second derivative is negative, the resulting distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
-In each case, the transition from $x=0$ to $x=\pm 0.2$ is continuous for the resulting mean, deviation and error distribution.
 6 Matrix Calculations
 Figure 14: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of input noise precision (x-axis) and matrix size (y-axis).
-Figure 15: Error deviations (z-axis) as a function of matrix size (x-axis) and input noise precision (y-axis) for the difference of the two sides of Formula (6.5).
+Figure 15: Error deviations (z-axis) as a function of input noise precision (x-axis) and matrix size (y-axis) for the difference of the two sides of Formula (6.5).
 Figure 16: Linear correlation between the precision of a matrix determinant (y-axis) to its condition number (x-axis).
@@
 Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$.
-Fourmula (6.2) presents the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$ [20].
+Fourmulas (6.1) and (6.2) present the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$.
 In Formula (6.4), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [20].
-[…]
+The square matrix whose element is $(-1)^{i+j}|M_{j,i}|$ is defined as the adjugate matrix [20] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$.
+By assigning a random integer between $2^{8},+2^{8}]$ to each element, $\mathbf{M}$ is created such that $\mathbf{M}^{A}$ can be calculated precisely using floating-point arithmetic when $n\leq 7$.
+By adding Gaussian noise $\tilde{x}_{i,j}$ of deviation $\delta x$ to each element of $\mathbf{M}$, $\widetilde{\mathbf{M}}$ is created as shown in Formula (6.1).
+The value error of $\mathbf{M}^{A}$ is the difference between $\widetilde{\mathbf{M}^{A}}$ and $\mathbf{M}^{A}$, whereas the result deviation is calculated using Formula (6.4), such that $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ is a distribution test.
+Figure 14 shows that the error deviations of $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ are very close to $1$ except when the matrix size is $8$ and the input uncertainty is $\delta x=10^{-17}$.
+This abnormality is caused by floating-point rounding errors only at this point, to result in a error deviation of $3.2$.
+Figure 14 demonstrates a typical distribution test of error deviation versus input noise deviation $\delta x$ and another algorithm specific dimension, with ideal coverage for all region except when $\delta x\rightarrow 0$.
 $\mathbf{M}\times\mathbf{M}^{A}$ $=\mathbf{M}^{A}\times\mathbf{M}=|\mathbf{M}|\mathbf{I};$ (6.5)
@@
 The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has nearly no value error, are not clear.
-It could relate to floating-point representation which stores each value $x$ as $\log_{2}(x)$.
+It could relate to floating-point representation.
 Figure 5 also suggests that using linear approximation $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$ have less error than $(x\pm\delta x)^{p}$.
@@
 Because the Quart sine function strictly preserves the symmetry and periodicity of the sine function, it provides better numerical accuracy than the Library sine function.
-  • Figure 23 shows that the absolute value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
+  • Figure 23 shows that the value difference between the Library $\sin(x)$ and the Quart $\sin(x)$ increases approximately linearly with $|x|$.
   • Figure 24 shows the value difference between the Quart and Library $\cos(x)/\sin(x)$ also increases roughly linearly with $|x|$, but is $10^{2}$ times larger than that observed for $\sin(x)$.
@@
 Even data for forward and reverse transformations are drastically different, variance arithmetic works effectively in all cases.
-When $L=18$ and $\delta x=0$, Figure 25 shows that the error distributions of Sin/Cos signal resemble Normal distributions, with an additional Delta-like distribution at $\tilde{z}=0$.
+When $L=18$ and $\delta x=0$, Figure 25 shows that the error distributions of Sin/Cos signal resemble Normal distributions, with an additional Delta-like distribution at $\tilde{z}=0$ for the forward transformation.
 The error distribution of the reverse transformation is structured on top of the Normal distribution, suggesting that the reverse transformation is more sensitive to numerical errors in the sine function.
@@
 Figure 30 demonstrates that the error deviations increase with sine or cosine frequency, regardless of FFT order $L$ when $L>8$.
-Figure 23 indicates that the numerical errors of $cos(x)/sin(x)$ obtained using the Library sine function increase with a periodicity of $\pi$, which can resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern shown in Figure 28.
+Figure 23 indicates that the numerical errors of $sin(x)$ obtained using the Library sine function increase with a periodicity of $\pi$, which can resonate with a signal whose periodicity is an integer multiply of $\pi$, producing the resonant pattern shown in Figure 28.
 At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger.
@@
 9.9 Ideal Coverage
+Figure 38: Error distributions for Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) computed using the Library sine function.
+The input noise is either Gaussian or white (legend).
+The FFT order is $18$.
 Figure 39: Error deviation (left y-axis) and uncertainty deviation (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
+Figure 40: The resulting precision (left y-axis) and uncertainties (right y-axis) of $\sin(\pi j/2^{18})$ versus index $j$ (x-axis) using either Quart or regressive sine function (legend).
 Adding noise to the input can suppress unspecified input errors.
-After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, whereas the error distribution for the round-trip transformation becomes Delta.
+After adding a Gaussian input noise of $\delta x=10^{-3}$ to a Linear signal when using the Library sine function, the error distributions for both the forward and the reverse transformations become Normal, whereas the error distribution for the round-trip transformation becomes Delta, as shown in Figure 38.
 Figure 39 illustrates the corresponding error deviations and uncertainty deviations versus FFT order:
@@
 10 Regressive Generation of Sine
-Figure 39: The resulting precision (left y-axis) and uncertainties (right y-axis) of $\sin(\pi j/2^{18})$ versus index $j$ (x-axis) using either Quart or regressive sine function (legend).
 $\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (10.1)
@@
 Bound momentum $\zeta(n,\kappa)$ should be extended to all probability distributions.
-The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of $1-\zeta(2,\kappa)$.
+The choice of ideal bounding range $\hat{\kappa}$ should be extended to other distributions, perhaps by using the same value of linear leakage $\varepsilon\equiv 1-\zeta(2,\kappa)$.
 The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions.

### v20

@@
 Abstract
-As a rigorous extension of conventional Taylor expansion, statistical Taylor expansion replaces precise input variables with random input variables of known distributions and sample counts, to compute the mean, standard deviation, and reliability of each result.
+As a rigorous extension of conventional Taylor expansion, statistical Taylor expansion replaces precise input variables with random input variables of known distributions and sample counts, to compute the mean, deviation, and reliability of each result.
 This approach tracks the propagation of input uncertainties through intermediate steps, making the final result path independent.
@@
   • The traditional variance-covariance framework accounts only for linear interactions between random variables through an analytic function [3][4][5], whereas statistical Taylor expansion extends this framework to include higher-order interactions as expressed in Formula (2.3) in this study.
+A full expansion illuminates the convergence of the result.
   • Calculating the reliability $[0,1]$ of $f\pm\delta f$ from input uncertainty distributions and sample counts seems completely new.
@@
 […]
-  • A distributional pole occurs when $f^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$.
+  • A distributional pole occurs when $y^{(1)}_{x}=0\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=\infty$.
 For example, $(\mu\pm 1)^{2}$ has a distributional pole at $\tilde{x}=0$, as shown in Figure 1.
-  • A distributional zero occurs when $f^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$.
+  • A distributional zero occurs when $y^{(1)}_{x}=\infty\rightarrow\rho(\tilde{y},\mu_{y},\sigma_{y})=0$.
 For example $\sqrt{\mu\pm 1}$ has a distributional zero at $\tilde{x}=0$, as shown in Figure 2.
@@
 […]
-$\overline{f(x)}$ $=f(x)+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
-$\delta^{2}f(x)$ $=\sum_{n=1}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
+$\overline{f(x)}$ $=f(x)+\sum_{n=2}^{\infty}(\delta x)^{n}\frac{f^{(n)}_{x}}{n!}\zeta(n,\kappa);$ (2.5)
+$\delta^{2}f(x)$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{f^{(j)}_{x}}{j!}\frac{f^{(n-j)}_{x}}{(n-j)!}\big(\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)\big);$ (2.6)
 An analytic function $f(x)$ can be accurately evaluated in a range using the Taylor series as shown in Formula (2.4).
@@
 $\delta^{2}(x\pm y)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.11)
-$\overline{xy}$ $=\zeta_{x}(0,\kappa_{x})x\;\zeta_{y}(0,\kappa_{y})y;$ (2.12)
-$\delta^{2}(xy)$ $=\zeta_{x}(2,\kappa_{x})(\delta x)^{2}y^{2}+x^{2}\zeta_{y}(2,\kappa_{y})(\delta y)^{2}+\zeta_{x}(2,\kappa_{x})(\delta x)^{2}\;\zeta_{y}(2,\kappa_{y})(\delta y)^{2};$ (2.13)
+$\frac{\overline{xy}}{xy}$ $=\zeta_{x}(0,\kappa_{x})\zeta_{y}(0,\kappa_{y});$ (2.12)
+$P(xy)^{2}$ $=\zeta_{x}(2,\kappa_{x})P(x)^{2}+\zeta_{y}(2,\kappa_{y})P(y)^{2}+\zeta_{x}(2,\kappa_{x})P(x)^{2}\zeta_{y}(2,\kappa_{y})P(y)^{2};$ (2.13)
 When $N\rightarrow\infty$, $\zeta(0,\kappa)\rightarrow 1$ and $\zeta(2,\kappa)\rightarrow 1$, making Formulas (2.10) and (2.11) the convolution results for $x\pm y$ [4], and Formulas (2.12) and (2.13) the corresponding results of the product distribution for xyxy [4].
@@
 Formulas (2.14) and (2.15) give the mean and variance for $e^{x}$, respectively:
-$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=1}^{\infty}(\delta x)^{n}\frac{\zeta(n,\kappa)}{n!};$ (2.14)
+$\frac{\overline{e^{x}}}{e^{x}}$ $=1+\sum_{n=2}^{\infty}(\delta x)^{n}\frac{\zeta(n,\kappa)}{n!};$ (2.14)
 $\frac{\delta^{2}e^{x}}{(e^{x})^{2}}$ $=\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!\;(n-j)!};$ (2.15)
@@
 Formulas (2.18) and (2.19) give the mean and variance for $\sin(x)$, respectively:
-[…]
-$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\eta(n,x)\frac{\zeta(n,\kappa)}{n!};$ (2.19)
-$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\eta(j,x)\eta(n-j,x)\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!(n-j)!};$ (2.20)
+$\overline{\sin(x)}=$ $\sum_{n=0}^{\infty}(\delta x)^{n}\sin(x)_{x}^{(n)}\frac{\zeta(n,\kappa)}{n!};$ (2.18)
+$\delta^{2}\sin(x)=$ $\sum_{n=2}^{\infty}(\delta x)^{n}\sum_{j=1}^{n-1}\sin(x)_{x}^{(j)}\sin(x)_{x}^{(n-j)}\frac{\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)}{j!(n-j)!};$ (2.19)
 Formulas (2.20) and (2.21) give the mean and variance for $x^{c}$, respectively:
@@
 The value of bound moment $\zeta(2n,\kappa)$ for Uniform distribution is also provided as an comparison.
+$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)=\frac{\kappa^{n}}{n};$ (2.22)
 $\zeta(2n,\kappa)$ $=\int_{-\kappa}^{+\kappa}\tilde{z}^{2n}\frac{1}{2\sqrt{3}}d\tilde{z}=2\rho(\kappa)\frac{\kappa^{2n+1}}{2n+1},\;\;0\leq\kappa\leq\sqrt{3};$ (2.23)
 […]
-$\forall j$ $:\lim_{n\rightarrow+\infty}\zeta(n,\kappa)-\zeta(j,\kappa)\zeta(n-j,\kappa)=\frac{\kappa^{n}}{n};$ (2.25)
 […]
@@
 Figure 5: Measured linear leakage $\omega(\kappa,N)$ (left y-axis) and corresponding measured bound range $\kappa$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is Gaussian, with different measuring bound range $\kappa_{s}$ (legend).
-Figure 6: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
+Figure 6: Ideal leakage $\varepsilon$ (y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
 When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
-Figure 7: Ideal leakage $\varepsilon$ (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) for Uniform input uncertainties.
+Figure 7: Ideal leakage $\varepsilon$ (y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) for Uniform input uncertainties.
 When $f(x)=x$, ideal leakage equals linear leakage $\omega$.
-$f(x)-\overline{f(x)}$ $\simeq(\delta x)^{2}\zeta(2,\kappa)f_{x}^{(2)};$ (2.28)
-$\delta^{2}f(x)$ $\simeq(\delta x)^{2}\zeta(2,\kappa)(f_{x}^{(1)})^{2};$ (2.29)
-Figures (2.28) and (2.29) show that the first-order approximation of Formulas (2.5) and (2.6) both contain the term $(\delta x)^{2}\zeta(2,\kappa)$.
+The first-order approximation of Formulas (2.5) and (2.6) both contain the term $(\delta x)^{2}\zeta(2,\kappa)$.
 As $\kappa\rightarrow+\infty$, $\zeta(2,\kappa)\rightarrow 1$ such that $\delta^{2}f(x)$ reaches a stable value but the convergence range of $\delta^{2}f(x)$ reduces toward zero for Formulas (2.17) and (2.21).
-As a trade-off, the choice of $\kappa$ to calculate the stable $f(x)-\overline{f(x)}$ and $\delta^{2}f(x)$ is defined as the ideal bound range $\hat{\kappa}$, and the result is the ideal uncertainty bias $\widehat{\overline{f(x)}}-f(x)$ and ideal variance $\hat{\delta^{2}}f(x)$.
+As a trade-off, the choice of $\kappa$ to calculate the stable $f(x)-\overline{f(x)}$ and $\delta^{2}f(x)$ is defined as the ideal bound range $\hat{\kappa}$, and the result is the ideal variance $\hat{\delta^{2}}f(x)$.
 For a Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$.
@@
 Define linear leakage $\omega(\kappa)\equiv 1-\zeta(2,\kappa)$.
+In $\delta^{x}x=\zeta(2,\kappa)(\delta x)^{2}$, the result variance $\delta^{x}$ is less than input variance $(\delta x)^{2}$ due to linear leakage.
 When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [4].
@@
 When the underlying distribution is Normal, Formula (2.27) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, whereas Formulas (2.28) and (2.29) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4].
-Figure 4 shows that $\omega(\kappa_{s})<\omega(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\omega(\kappa_{s},N)=\omega(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
+Figure 4 shows that $\omega(\kappa_{s},N+1)<\omega(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\omega(\kappa_{s},N)=\omega(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N+1)<\kappa(\kappa_{s},N)$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction.
 Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable resulting linear leakage $\omega$ is smaller for larger measuring bounding range $\kappa_{s}$.
-It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\kappa_{s}$ is larger.
+It also shows that linear leakage converges toward 0 with increasing sample count $N$ faster when $\kappa_{s}$ is larger.
 $\varepsilon(x\pm y)$ $=\frac{\omega_{x}(\delta x)^{2}+\omega_{y}(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.30)
-Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example.
+$\varepsilon(xy)$ $=\frac{2\omega_{x}P(x)^{2}+2\omega_{y}P(y)^{2}-\omega_{x}P(x)^{2}\omega_{y}P(y)^{2}}{P(x)^{2}+P(y)^{2}+P(x)^{2}P(y)^{2}};$ (2.31)
+Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formulas (2.30) and 2.31 as two examples.
 Figure 6 compares ideal leakage $\varepsilon$ and linear leakage $\omega$ for the selected functions when the input uncertainty distribution is Gaussian, and shows that the example stable ideal leakages are less than $10^{-4}$ when $\hat{\kappa}=5$, which is small enough for most applications.
-The unbiased bounding range $\kappa_{min},\kappa_{max}]$ of Uniform distribution is estimated as $1+\frac{1}{N}$ times of the minimal and maximal values in sampled data set.
-The bonding leakage $\epsilon$ is the percentage of $\kappa_{min},\kappa_{max}]$ outside the true bounding range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$.
+The unbiased bounding range $\kappa_{min},\kappa_{max}]$ of Uniform distribution is estimated as $1+\frac{1}{N}$ times of the minimal and maximal values in sampled data set [20].
+The bounding leakage $\epsilon$ is the percentage of $\kappa_{min},\kappa_{max}]$ outside the true bounding range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$.
 Because $\zeta(2n)$ for Uniform distribution increases much slower than $\zeta(2n,5)$ for Gaussian distribution, Figure 7 shows that for the selected functions, ideal leakage equals linear leakage for Uniform input uncertainty, and both decrease linearly with sample count $N$.
-The difference between Figures 6 and 7 shows the effect of input uncertainty distribution.
+The difference between Figures 6 and 7 shows the effect of input uncertainty distribution to the output deviation $\delta f$.
 2.7 Dependency Tracing
-$\delta^{2}(f+g)=$ $\;\delta^{2}f+\delta^{2}g+2(\overline{fg}-\overline{f}\overline{g});$ (2.34)
-$\delta^{2}(c_{1}f+c_{0})=$ $\;c_{1}^{2}\delta^{2}f;$ (2.35)
-When all inputs satisfy the uncorrelated uncertainty condition, statistical Taylor expansion traces dependencies through the intermediate steps. For example:
-  • Formula (2.34) expresses $\delta^{2}(f+g)$, The dependency tracing of is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
+When all inputs satisfy the uncorrelated uncertainty condition, statistical Taylor expansion traces dependencies through the intermediate steps.
+For example, when the sample count $N$ is sufficiently large:
+  • The dependency tracing of $\delta^{2}(f+g)$ is illustrated by $\delta^{2}(f-f)=0$, and $\delta^{2}(f(x)+g(y))=\delta^{2}f+\delta^{2}g$, with the latter corresponding to Formula (2.11).
   • The dependence tracing of $\delta^{2}(fg)$ is illustrated by $\delta^{2}(f/f)=0$, $\delta^{2}(ff)=\delta^{2}f^{2}$, and $\delta^{2}(f(x)g(y))=\overline{f}^{2}(\delta^{2}g)+(\delta^{2}f)\overline{g}^{2}+(\delta^{2}f)(\delta^{2}g)$, with the latter corresponding to Formula (2.13).
@@
 For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original inputs should be restored.
-  • Formula (2.35) gives the variance of the linear transformation of a function, which can be applied to other formulas for more general dependency tracing.
 Statistical Taylor expansion employs dependency tracing to ensure that the calculated mean and variance satisfy statistics rigorously.
@@
 Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$ when the input uncertainty distribution is Uniform.
-It has almost the same trend of upper bounds $P(x)$ increasing with $c$ as shown in Figure 9 but just with different y-scaling: $1/\sqrt{3}$ instead of by $1/5$.
+It has almost the same trend of upper bounds $P(x)$ increasing with $c$ as shown in Figure 9 but just with different y-scaling: $1/\sqrt{3}$ instead of $1/5$.
 3.2 Positive
@@
 The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
-The x-axis is expressed in units of $\pi$.
 Figure 11: Measured upper bound $\delta x$ (left y-axis) for $\sin(x\pm\delta x)$ across different values of $x$ (x-axis) for uniform uncertainty.
 The corresponding resulting uncertainty bias and deviation are also shown (right y-axis).
-The x-axis is expressed in units of $\pi$.
-In addition to convergence, the variance expansion may yield negative results when the expansion is terminated arbitrarily, as in Formula (2.20) for $sin(x\pm\delta x)$.
+In addition to convergence, the variance expansion may yield negative results when the expansion is terminated arbitrarily, as in Formula (2.19) for $sin(x\pm\delta x)$, such that an upper bound for $\delta x$ is need to guarantee non-negative variance.
 Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty.
-Beyond this upper bound, the result deviation is no longer positive at some expansion order.
-Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which is satisfied as shown in Figure 10.
+Because $\sin(x)\in[-1,+1]$, $\overline{\sin(x)}\pm\delta\sin(x)$ should also follow this range, as indicated in 10.
 Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@
 In statistics, such two values have a $50\%$ possibility of being either less than or greater to each other but zero probability of being exactly equal [4].
-In variance arithmetic, however, they are treated as neither less nor greater than each other and therefore are considered equal.
+In variance arithmetic, however, they are considered equal.
 Otherwise, the standard z-statistic method [4] is applied to determine whether two imprecise values are statistically equal, less than, or greater than each other.
@@
 If the threshold probability for inequality is set at $50\%$, then $1.000\pm 0.002<1.002\pm 0.001$.
-Alternatively, an equivalent bounding range for z can be used, such as $|z|\leq 0.67448975$ for an equal probability threshold of $50\%$.
 Because the result of comparison depends on threshold probability which is application specific, comparison is not part of variance arithmetic.
@@
 When an input contains unspecified errors, such as numerical errors in library functions or floating-point rounding errors, Gaussian noise with progressively increasing deviations can be added to inputs, until ideal coverage is attained.
-The minimal noise deviation required provides a good estimate of the magnitude of the unspecified input uncertainty deviations.
+The minimal noise deviation required for the ideal coverage provides a good estimate of the magnitude of the unspecified input uncertainty deviations.
 Achieving ideal coverage serves as a necessary verification step to ensure that statistical Taylor expansion has been applied correctly within the given context.
@@
 $\sum_{j=0}^{N}c_{j}(x+\tilde{x})^{j}$ $=\sum_{j=0}^{N}\tilde{x}^{j}P_{j},\;\;\;P_{j}\equiv\sum_{k=0}^{N-j}x^{k-j}c_{j+k}\begin{pmatrix}j+k\\ j\end{pmatrix}$; (5.1)
-Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can reach only half of the maximal expansion order of Formula (2.6), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian, and $\hat{\kappa}=5$.
+Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can reach only half of the maximal expansion order of Formula (2.6), for example, $2j=224=448/2$ for $\zeta(2n,5)$ of Gaussian input uncertainty.
 5.1 Residual Error
@@
 […]
-$=\sum_{m=1\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$ (6.2)
 $\overline{|\mathbf{M}|}$ $=|\mathbf{M}|;$ (6.2)
-$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}|^{2}\prod_{k=1\dots n}^{i_{k}\in\{i_{1}\dots i_{m}\}}(\delta x_{i_{k},j_{k}})^{2};$ (6.4)
-Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20].
-Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$.
-Fourmulas (6.1) and (6.2) present the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$.
+$\delta^{2}|\mathbf{M}|$ $=\sum_{m=1}^{n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}|\mathbf{M}_{<i_{1}\dots i_{m}>_{n},<[j_{1}\dots j_{m}]>_{n}}|^{2}\prod_{k=1\dots m}\zeta_{i,j}(2,\kappa_{i,j})(\delta x_{i_{k},j_{k}})^{2};$ (6.3)
+Let $j_{1},j_{2}\dots j_{m}>_{n}$ denote a selection of $m$ numbers from the sequence of natural numbers $1,2\dots n$ in sequence, which is named as the ordered permutation.
+Let $j_{1},j_{2}\dots j_{m}]_{n}$ and $(-1)^{[j_{1},j_{2}\dots j_{m}]_{n}}$ denote a permutation of $j_{1},j_{2}\dots j_{m}]>_{n}$ and the corresponding sign [21], in which $j_{1},j_{2}\dots j_{m}]>_{n}$ denote the corresponding ordered permutation.
+Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2\dots n$ at row index $i$ and column index $j$.
+Fourmula (6.1) presents the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.2) and (6.3) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$.
 In Formula (6.3), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [21].
@@
 7 Mathematical Library Functions
-Basic Function $e^{x\pm\delta x}$ $\log(x\pm\delta x)$ $(1\pm\delta x)^{c}$ $\sin(x\pm\delta x)$
-Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
-Uncertainty $\delta x\in[10^{-15},1]$ $\delta x\in[10^{-15},0.2x]$ $\delta x\in[10^{-15},0.2]$ $\delta x\in[10^{-15},1]$
-Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.978\pm 0.150$
-Table 1: The result error deviations of selected basic functions with Gaussian input noise $x\pm\delta x$ when $\delta x>10^{-15}$, obtained by variance arithmetic.
-The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverging regions are excluded.
-The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
 Figure 17: Error deviation for $\sin(x\pm\delta x)$ as a function of $x$ and $\delta x$.
@@
 Figure 19: Normalized errors of $(x^{p})^{\frac{1}{p}}-x$ as functions of $x$ and $p$.
+Basic Function $e^{x\pm\delta x}$ $\log(x\pm\delta x)$ $(1\pm\delta x)^{c}$ $\sin(x\pm\delta x)$
+Range $x\in[-100,+100]$ $x\in[1/32,32]$ $c\in[-3,+3]$ $x\in[-\pi,+\pi]$
+Uncertainty $\delta x\in[10^{-15},1]$ $P(x)\in[10^{-15}x,0.2x]$ $P(x)\in[10^{-15},0.2]$ $\delta x\in[10^{-15},1]$
+Error Deviation $1.000\pm 0.010$ $0.999\pm 0.011$ $0.989\pm 0.104$ $0.997\pm 0.027$
+Table 1: The result error deviations of selected basic functions with Gaussian input noise $x\pm\delta x$ when $\delta x>10^{-15}$, obtained by variance arithmetic.
+The error deviation for $(1\pm\delta x)^{c}$ can be improved to $1.000\pm 0.007$ if diverging regions are excluded.
+The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
 Table 1 shows that by sampling random input $\tilde{x}$ from Gaussian noise $x+\delta x$, Formulas (2.15), (2.17), (2.19), and (2.21) provide almost perfect characterization $f(x)\pm\delta f$ for the result distributions of f(x~)f($\tilde${x}) in which f(x~)f($\tilde${x}) is the corresponding math library functions.
@@
 Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation.
-As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$.
 The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ are examined in greater detail in Section 9.
@@
 For reference, the input time-series signal $Y$ is also plotted, with its values indicated on the y-axis on the right.
-Formulas (8.1) and (8.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [5].
-$\alpha$ $=\frac{\sum_{j}Y_{j}}{\sum_{j}1};$ (8.1)
-$\beta$ $=\frac{\sum_{j}X_{j}Y_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}Y_{j}}{\sum_{j}X_{j}X_{j}\;\sum_{j}1-\sum_{j}X_{j}\;\sum_{j}X_{j}};$ (8.2)
-In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index.
-$Y_{j}$} is thus referred to as a time-series input, with $j$ corresponding to $X_{j}$.
-A moving window algorithm [5] is applied within a small window centered on each $j$.
-For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window.
-This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formulas (8.1) and (8.2) into Formulas (8.3) and (8.4), respectively [17].
 $\alpha_{j}$ $=\alpha\;2H=\sum_{X=-H+1}^{H}Y_{j-H+X};$ (8.1)
 $\beta_{j}$ $=\beta\;\frac{H(H+1)(2H+1)}{3}=\sum_{X=-H}^{H}XY_{j-H+X};$ (8.2)
-The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formulas (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formulas (8.5) and (8.6), respectively [17].
+In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index.
+$Y_{j}$} is thus referred to as a time-series input.
+A moving window algorithm [5] is applied within a small window centered on each $j$.
+For each calculation window, Formulas (8.1) and (8.2) present the least-square line-fit of $Y=\alpha+\beta j$ [17].
 $\beta_{j}$ $=\beta_{j-1}-\alpha_{j-1}+H\left(Y_{j-2H-1}+Y_{j}\right);$ (8.3)
 $\alpha_{j}$ $=\alpha_{j-1}-Y_{j-2H-1}+Y_{j};$ (8.4)
+The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formulas (8.1) and (8.2) to be reformulated into the progressive moving-window calculation given by Formulas (8.3) and (8.4), respectively [17].
 8.2 Variance Adjustment
@@
 9.1 DFT (Discrete Fourier Transformation)
+Figure 22: The DFT spectrum $H[n]$ of signal $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, as intensity (y-axis) and phase (embedded y-axis) versus frequency index $n\in[0,18]$ (x-axis and embedded x-axis) for different signal frequency $f$ (legend).
+This result agrees with both theoretical formula [17] and numerical computation from any mathematical libraries such as SciPy.
 $H[n]$ $=\sum_{k=0}^{N-1}h[k]\;e^{\frac{i2\pi}{N}kn};$ (9.1)
@@
 As a convention, $k$ denotes the time index for $h[k]$ as a waveform, whereas $n$ represents the frequency index for $H[n]$ as a spectrum.
-Figure 22: The DFT spectrum $H[n]$ of signal $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, as intensity (y-axis) and phase (embedded y-axis) versus frequency index $n\in[0,18]$ (x-axis and embedded x-axis) for different signal frequency $f$ (legend).
-This result agrees with both theoretical formula [17] and numerical computation from any mathematical libraries such as SciPy.
+However, DFT if not a faithful digital implementation of continuous Fourier transform (FT) [5].
 For example, the FT spectrum of a sine function is a Delta function at the signal frequency $f$ with a phase $\pi/2$ [5].
 Figure 22 shows the DFT spectra of the sine function $h[k]=\sin(f\frac{2\pi}{128}k),k\in[0,127]$, where $f$ is its signal frequency.
-If DFT is regarded as the digital implementation of FT, the spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise.
-Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics
+The DFT spectra exhibit no modeling error only when the input signal frequency $f$ is an integer, and display varying degrees of modeling errors otherwise.
+Because of these modeling errors, the use of DFT as the digital implementation of FT is questionable, even though such usage is ubiquitous, and fundamental to many areas of applied mathematics [5].
 To avoid the modeling errors inherent in DFT, only Formulas (9.1) and (9.2) are used in this study.
@@
 Consequently, the uncertainty deviation for the round-trip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
-  • The forward and reverse transformations are identical except for a sign difference, implying that they are essentially the same algorithm, and any observed difference arises solely from the input data.
 9.3 Testing Signals
@@
     $H[n]$ $=\frac{dG}{dy}=\begin{cases}n=0:\;\;\;\frac{N(N-1)}{2}\\ n\neq 0:\;\;\;-\frac{N}{2}(1+i\frac{\cos(n\frac{\pi}{N})}{\sin(n\frac{\pi}{N})})\end{cases};$ (9.3)
-For Sin or Cos signals, forward and reverse DFT transforms differ in their data prospective of time domain versus frequency domain:
-  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly during cancellation.
-  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
+The forward and reverse transformations are identical except for a sign difference, implying that they are essentially the same algorithm, and any observed difference arises mainly from the input data.
+For Sin or Cos signals:
+  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly during mutual cancellation of input data.
+  • The reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly.
 The question is whether variance arithmetic can work properly in these two contrary cases.
@@
   2. 2.
-    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are the library sine directly, whereas the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from the library $\cos(\pi(2^{L-1}-j)/2^{L})$.
+    The values of $\sin(\pi j/2^{L}),j\in[0,2^{L-2}]$ are the library sine directly, whereas the values of $\sin(\pi j/2^{L}),j\in[2^{L-2},2^{L-1}]$ are computed from the library $\cos(\pi(2^{L-1}-j)/2^{L}),j\in[0,2^{L-2}]$.
   3. 3.
@@
 9.5 Using Quart Sine for Sin/Cos Signals
-Figure 25: Error distributions of normalized errors of Sin/Cos signals for forward, reverse and round-trip transformations (legend) using the Quart sine function. The FFT order is $18$.
+Figure 25: Error distributions of normalized errors of Sin/Cos signals without added input noise for forward, reverse and round-trip transformations (legend) using the Quart sine function.
+The FFT order is $18$.
 With the FFT order as the specific dimension, the error deviations obtained using the Quart sine function for forward and reverse transformations resemble those shown in Figure 14, whereas those for round-trip transformation are nearly identical to those shown in Figure 15, independent of the signal frequency or whether a Sin or Cos signals is used.
@@
 9.6 Using Library Sine for Sin/Cos Signals
-Figure 26: Error distributions of normalized errors of Sin/Cos signal for forward, reverse and round-trip transformations (legend) computed using the Library sine function. The FFT order is $18$.
+Figure 26: Error distributions of normalized errors of Sin/Cos signal without added input noise for forward, reverse and round-trip transformations (legend) computed using the Library sine function.
+The FFT order is $18$.
 Figure 27: FFT value error spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation.
@@
 The legend distinguishes between the uncertainty and the value error.
-The x-axis represents the frequency index, and the y-axis represents both uncertainty and value error.
+The x-axis represents the time index, and the y-axis represents both uncertainty and value error.
 Figure 30: Error deviation (z-axis) of the FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT order $L$ (y-axis).
@@
 Figure 33: Error distributions of normalized errors of Linear signals for forward, reverse and round-trip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
-Figures 31 and 32 show the error deviations for the forward and the reverse transformations, respectively.
+Figures 31 and 32 show the error deviations for the forward and the reverse transformations, respectively, both of which resemble Figure 14.
 The forward transformation exhibits a larger ideal coverage area than the reverse transformations: $\delta x>10^{-12}$ for the forward transformation, and $\delta x>10^{-8}$ for the reverse transformation.
@@
 Variance arithmetic fails because of the large amount of unspecified numerical errors from the Library sine function.
-Figures 35 and 36 show much smaller ideal coverage areas than those shown in Figures 31 and 32.
+Figures 35 and 36 show much smaller ideal coverage areas than those shown in Figures 31 and 32, respectively.
 Because uncertainty deviations grow more slowly in the reverse transformation than in the forward transformation, the former exhibits a smaller ideal coverage region.
@@
 Such deterioration in calculation is not easily detectable when using conventional floating-point arithmetic.
-Figure 37 shows that, even when variance arithmetic can no longer effectively track the value errors for either the forward or the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as shown by the plateau region at high $L$ and low $\delta x$.
+Figure 37 shows that, even when variance arithmetic can no longer effectively track the value errors for either the forward or the reverse transformations, it can still effectively track the value errors for the round-trip transformation, as shown by the plateau region of error deviations at high $L$ and low $\delta x$.
 Such error cancellation arises from dependence tracing in statistical Taylor expansion.
@@
 $\sin(\alpha+\beta)=2\sin(\frac{\alpha+\beta}{2})\cos(\frac{\alpha+\beta}{2})=\sqrt{1-\cos(\frac{\alpha+\beta}{2})^{2}};$ (10.4)
-Formulas (10.2) and (10.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1).
+Formulas (10.2) and (10.3) calculate $\sin(j\pi/2^{L}),\cos(j\pi/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (10.1).
 Formula (10.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, such that value errors will not accumulate when the regression order $L$ increases.
@@
 The procedure to find the bound moments $\kappa$ for each sample count $N$ should be developed for discrete distributions.
-The statistical meaning for ideal leakage $\varepsilon\in[0,1]$ in the context of its choice of ideal bounding range $\hat{\kappa}$ needs further investigation: Is $\varepsilon$ the possibility for the result to be not bound by $\hat{\kappa}$?
 The performance of variance arithmetic must be improved for broader practical adoption.
@@
 This capability enables engineers to pinpoint the primary sources of measurement inaccuracy and guides them to targeted improvements in data acquisition and processing strategies.
-For example, Formula (2.33) can provide guidance on improving the ideal leakage of $x\pm y$.
+For example, Formulas (2.30) and (2.31) can provide guidance on improving the ideal leakage of $x\pm y$ and $xy$ respectively.
 Because conventional numerical approaches are based on floating-point arithmetic and path dependent in general, they must be reexamined or even reinvented within the framework of variance arithmetic.
@@
 Furthermore, the quantitative dependence of convergence on the choice of ideal bounding range $\hat{\kappa}$ is similar to the principle of quantum physics: What it is depends on how it is measured.
-For example, it is possible that an experimental measurement reduces $\hat{\kappa}$ for a wave function, to cause the wave function to collapse.
+For example, it is possible that an experimental measurement reduces $\hat{\kappa}$ for a wave function, to cause the wave function to collapse when the uncertainty deviation reduces from infinite to finite.
 Both resemblances are worthy of further investigation.